### v2

# Concurrent Composition for Differentially Private Continual Mechanisms

Abstract.
Many intended uses of differential privacy involve a continual mechanism that is set up to run continuously over a long period of time, making more statistical releases as either queries come in or the dataset is updated. In this paper, we give the first general treatment of privacy against adaptive adversaries for mechanisms that support dataset updates and a variety of queries, all arbitrarily interleaved. It also models a very general notion of neighboring, that includes both event-level and user-level privacy.
We prove several concurrent composition theorems for continual mechanisms, which ensure privacy even when an adversary can interleave its queries and dataset updates to the different composed mechanisms. Previous concurrent composition theorems for differential privacy were only for the case when the dataset is static, with no adaptive updates. Moreover, we also give the first interactive and continual generalizations of the “parallel composition theorem” for noninteractive differential privacy. Specifically, we show that the analogue of the noninteractive parallel composition theorem holds if either there are no adaptive dataset updates or each of the composed mechanisms satisfies pure differential privacy, but it fails to hold for composing approximately differentially private mechanisms with dataset updates.
We then formalize a set of general conditions on a continual mechanism $\mathcal{M}$ that runs multiple continual sub-mechanisms such that the privacy guarantees of $\mathcal{M}$ follow directly using the above concurrent composition theorems on the sub-mechanisms, without further privacy loss. This enables us to give a simpler and more modular privacy analysis of a recent continual histogram mechanism of Henzinger, Sricharan, and Steiner. In the case of approximate DP, ours is the first proof that shows that its privacy holds against adaptive adversaries.
11footnotetext: Institute of Science and Technology, Klosterneuburg, Austria (email: monika.henzinger@ista.ac.at).22footnotetext: Institute of Science and Technology, Klosterneuburg, Austria (email: roodabeh.safavi@ista.ac.at).33footnotetext: Harvard University (email: salil_vadhan@harvard.edu).
1. Introduction
Differential privacy [6] is the now-standard framework for protecting privacy when performing statistical analysis or machine learning on sensitive datasets about individuals. In addition to having a rich and rapidly growing scientific literature, differential privacy has also seen large-scale adoption by technology companies and government agencies.
The familiar formulation of differential privacy is for a noninteractive mechanism $\mathcal{M}:\mathcal{X}\rightarrow\mathcal{Y}$, which takes a dataset $x\in\mathcal{X}$ and produces a statistical release (e.g. a collection of statistics or the parameters of a machine learning model) $y\in\mathcal{Y}$. Differential privacy requires that every two datasets $x$ and $x^{\prime}$ that “differ on one individual’s data,” the output distributions $\mathcal{M}(x)$ and $\mathcal{M}(x^{\prime})$ are “close” to each other. Formally, we have a neighboring relation $\sim$ on $\mathcal{X}$, which specifies which datasets “differ on one individual’s data,” or more generally what information requires privacy protection. We say that $\mathcal{M}$ is $(\epsilon,\delta)$-DP with respect to $\sim$ if for all $x\sim x^{\prime}$, we have:
(1) $\forall S\subseteq\mathcal{Y}\qquad\Pr[\mathcal{M}(x)\in S]\leq e^{\epsilon}\cdot\Pr[\mathcal{M}(x^{\prime})\in S]+\delta.$
[…]
1.1. Differential Privacy Over Time
Many intended uses of differential privacy involve a system that is set up to run continuously over a long period of time, making more statistical releases as either queries come in or the dataset is updated, for example with people leaving and entering. (The latter case, with dataset updates, is known as differential privacy under continual observation, and was introduced in the concurrent works of Dwork, Naor, Pitassi, and Rothblum [7] and Chan, Shi, and Song [2].) These settings introduce a new attack surface, where an adversary may be adaptive, issuing queries or influencing dataset updates based on the responses it receives.
When $\delta>0$, there are mechanisms that are differentially private against oblivious adversaries but not against adaptive ones. [^1] A simple artificial example is the following: the mechanism’s first release can include a uniformly random number $r$ from $\{1,2,\ldots,\lceil 1/\delta\rceil\}$, and then if the adversary’s next query is $r$, the mechanism publishes the entire dataset in the clear. One natural example is the “advanced composition theorem” for differentially privacy [9], which is known to fail if the adversary can select the privacy-loss parameters $(\epsilon_{i},\delta_{i})$ of the composed mechanisms adaptively [19].
Until recently, adaptivity has been treated in an ad hoc manner in the differential privacy literature. For example, the aforementioned advanced composition theorem of [9] was formalized in a way that allows an adversary to adaptively choose the mechanisms and pairs of adjacent datasets, just not the privacy-loss parameters, and new composition theorems that allow for adaptive choices of privacy-loss parameters were given by [19]. Other early examples where adaptive queries were analyzed include the Sparse Vector Technique (SVT) [8, 11, 17] and the Private Multiplicative Weights algorithm [11].
A few years ago, Vadhan and Wang [20] initiated a systematic study of differentially private mechanisms $\mathcal{M}$ that support adaptive queries. Specifically, they formalized the notion of an interactive mechanism, which begins with a dataset as input and maintains a potentially secret state as it answers queries that come in over time. (Both SVT and Private Multiplicative Weights utilize a secret state.) They defined $\mathcal{M}$ to be an $(\epsilon,\delta)$-DP interactive mechanism if for every pair of adjacent datasets $x\sim x^{\prime}$ and every adaptive adversary $\mathcal{A}$, the view of $\mathcal{A}$ when interacting with $\mathcal{M}(x)$ is $(\epsilon,\delta)$-indistinguishable (in the sense of Inequality (1)) from its view when interacting with $\mathcal{M}(x^{\prime})$, where the view of $\mathcal{A}$ consists of the random coin tosses of $\mathcal{A}$ and the transcript of the interaction.
Vadhan and Wang [20] asked whether our existing composition theorems for differential privacy, where the mechanisms being composed are noninteractive mechanisms, extend to the concurrent composition of interactive mechanisms. In concurrent composition of interactive mechanisms, an adaptive adversary can interleave its queries to the different interactive mechanisms being composed. This is a realistic attack scenario when multiple interactive differentially private mechanisms are deployed in practice, and is also a useful primitive for analyzing differentially private algorithms that use several interactive mechanisms as building blocks. Vadhan and Wang showed that the basic composition theorem for pure differential privacy [6] does extend to concurrent composition, but for approximate differential privacy, they obtained a much weaker bound for the concurrent case. A tight bound for concurrent composition of approximate differential privacy was given by Lyu [18]. A series of works gave concurrent composition theorems for other flavors of differential privacy and composition, such as Rényi DP [18], $f$-DP [21], and adaptive choices of privacy parameters [10].
All of the above works are for the case where the queries to the mechanisms are adaptive, but the dataset $x$ is static, given as input to the mechanism at the start of the interaction. Jain, Raskhodnikova, Sivakumar, and Smith [13] recently gave a formalization of adaptive adversaries under continual observation (where there are dataset updates but the queries are fixed) for the special case of event-level privacy, where two streams of updates are considered adjacent if they differ on only one update. The lack of a more general formalism and toolkit for reasoning about adaptivity in dataset updates has led to complicated and ad hoc proofs of privacy of continual mechanisms (see e.g. [12]). Rectifying this situation is what motivated our work.
1.2. Our contributions
In this paper, we:
  • Give a formalism that captures privacy against adaptive adversaries for continual mechanisms, which support both queries and dataset updates. This generalizes the previously studied cases of interactive mechanisms (which have static datasets) and continual observation (which answer fixed queries). See Figure 1 for an illustration. We note that specific mechanisms, like variants of SVT, have been proposed to handle both dataset updates and queries [14], but without offering a definition of privacy against adversaries that can adaptively select both the updates and queries.
[…]
    Figure 1. Comparison between interactive mechanism (IMs), continual observation (CO), and continual mechanisms (CMs)
  • We prove concurrent composition theorems for continual mechanisms. (1) The first shows that, when the privacy-loss parameters are fixed, the standard composition theorems for noninteractive approximate differential privacy extend to the concurrent composition of continual mechanisms. (2) We extend the “parallel composition theorem” for noninteractive mechanisms to the concurrent parallel composition of a finite or infinite number of adaptively chosen interactive mechanisms with an adaptively chosen pair of neighboring datasets. (3) Then we extend this further to the concurrent composition of a finite or infinite number of adaptively chosen continual mechanisms when the composed mechanisms satisfy pure differential privacy. (4) We complement the latter with a counterexample for the case of approximate differential privacy.
  • We formalize a set of general conditions on a continual mechanism $\mathcal{M}$ that runs multiple continual sub-mechanisms such that the privacy guarantees of $\mathcal{M}$ follow directly using the above concurrent composition theorems on the sub-mechanisms, without further privacy loss.
  • We use our concurrent composition theorems for continual mechanisms, with the aforementioned general conditions, to give a simpler and more modular privacy analysis of a recent continual histogram mechanism [12]. In the case of approximate DP, ours is the first proof that the privacy of the mechanism holds against adaptive adversaries.
We elaborate on all of these contributions below.
A general formalism for privacy against adaptive adversaries.
Our basic object of study is a continual mechanism $\mathcal{M}$, which is an interactive mechanism that is not given any input at the start, but simply receives and sends messages in a potentially randomized and stateful manner. These messages can represent, for instance, queries or dataset updates. To model static datasets (as in standard interactive differential privacy), the entire dataset can be given to $\mathcal{M}$ as its first message.
Towards defining privacy, we consider adaptive adversaries $\mathcal{A}$ that send messages $m_{i}$ for $i=1,2,3,\dots$ that always are parsed as pairs $m_{i}=(m_{i}[0],m_{i}[1])$. Privacy is defined with respect to a verification function $f$ that takes a sequence of such pairs $(m_{1},m_{2},\ldots,m_{t})$ and outputs either $\top$ or $\bot$. When $f(m_{1},m_{2},\ldots,m_{t})=\top$, we consider the two sequences $m[0]=(m_{1}[0],m_{2}[0],\dots,m_{t}[0])$ and $m[1]=(m_{1}[1],m_{2}[1],\dots,m_{t}[1])$ to be neighboring.
Now, by the choice of the verification function $f$, we can easily capture previously studied notions of privacy and more. For example, for a standard interactive mechanism, $f(m_{1},m_{2},\ldots,m_{t})$ will check that $m_{1}[0]\sim m_{1}[1]$ are a pair of neighboring datasets, and that $m_{i}[0]=m_{i}[1]$ for all $i>1$ (these represent the adaptive queries). For continual observation under event-level privacy, we interpret all of the $m_{i}[b]$’s as dataset updates, and $f$ will require that $m[0]$ and $m[1]$ differ in at most one coordinate $i\in[k]$. For continual observation under “user-level” privacy, $f$ will require that the entries in which $m[0]$ and $m[1]$ differ correspond to only one user. For continual mechanisms, we can interleave adaptive queries among the dataset updates by having $f$ require that all $m_{i}[0]=m_{i}[1]$ whenever either one is a query (rather than a dataset update).
To go from a verification function $f$ to a definition of privacy against adaptive adversaries, we utilize the notion of an interactive post-processing mechanism (IPM) [10], which is a stateful procedure that can stand between an interactive mechanism and an adversary, or two interactive or continual mechanisms, or even two IPMs, and interact with both of them adaptively, sending and receiving messages to the mechanisms on its left and right. Specifically, we place two IPMs between the adversary $\mathcal{A}$ and the continual mechanism $\mathcal{M}$. The first is a verifier $\mathcal{V}[f]$ that, every time it receives a new message pair $m_{k}$ from $\mathcal{A}$, applies $f$ to the entire history of message pairs $(m_{1},\ldots,m_{k})$ sent by $\mathcal{A}$, and forwards $m_{k}$ onward only if $f$ returns $\top$. Otherwise $\mathcal{V}[f]$ halts the interaction. The second IPM is the identifier $\mathcal{I}$, which starts out with a secret bit $b\in\{0,1\}$, and every time it receives a message pair $(m_{i}[0],m_{i}[1])$ from $\mathcal{V}[f]$, it forwards $m_{i}[b]$ on to the mechanism $\mathcal{M}$. When $\mathcal{M}$ returns an answer $a_{i}$, $\mathcal{I}$ and $\mathcal{V}[f]$ just forward it back to $\mathcal{A}$. See Figure 2 for an illustration of an interaction between $\mathcal{A}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}[1]\!\circ^{*}\!\mathcal{M}$ consisting of three messages from $\mathcal{A}$ to $\mathcal{V}[f]$.
[…]
Figure 2. Illustration of the interactions between $\mathcal{A}$, $\mathcal{V}[f]$, $\mathcal{I}(b)$, and $\mathcal{M}$
This leads to the following definition of differential privacy for continual mechanisms:
Definition 1.1 (DP for continual mechanisms).
Let $f$ be a verification function. A continual mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP with respect to $f$ if for every adaptive adversary $\mathcal{A}$, the view of $\mathcal{A}$ when interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}[0]\!\circ^{*}\!\mathcal{M}$ is $(\epsilon,\delta)$-indistinguishable from its view when interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}[1]\!\circ^{*}\!\mathcal{M}$.
The idea of defining security against adaptive adversaries by having the adversary send pairs of messages and trying to distinguish whether the first message in each pair or the second message in each pair is being used goes back to the notion of left-or-right indistinguishability in crytography [1]. What is different here is its combination with a verification function $f$ to enforce the neighboring condition, which is necessary to capture differential privacy.
As far as we know, this definition captures all of the previous definitions of privacy against adaptive adversaries as special cases (by appropriate choices of $f$), as well as modeling adaptive adversaries in even more scenarios (such as continual observation with user-level privacy, and mixtures of adaptively-chosen dataset updates and queries).
Furthermore, it allows us to reduce the analysis of general continual mechanisms to that of the previously studied notion of interactive mechanisms. Indeed, by inspection, we observe that a continual mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP with respect to $f$ if and only if the interactive mechanism $\mathcal{M}^{\prime}(\cdot)=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(\cdot)\!\circ^{*}\!\mathcal{M}$ is $(\epsilon,\delta)$-DP on the dataset space $\{0,1\}$ (where $0\sim 1$).
Concurrent Composition Theorems for Continual Mechanisms.
With the above definitions in hand, we can formulate and prove new concurrent composition theorems for continual mechanisms. Recall that with concurrent composition of interactive mechanisms, an adaptive adversary (or honest analyst) can interleave its queries to the different interactive mechanisms being composed; for concurrent composition of continual mechanisms, we also allow the dataset updates to be interleaved with the queries and each other. (See Figure 3.)
Our first concurrent composition theorem is as follows:
Theorem 1.2 (concurrent composition of continual mechanisms, informally stated).
For every fixed sequence $(\epsilon_{1},\delta_{1}),\ldots,(\epsilon_{k},\delta_{k})$ of privacy-loss parameters and verification functions $f_{1},\ldots,f_{k}$, the concurrent composition of continual mechanisms $\mathcal{M}_{i}$ that are $(\epsilon_{i},\delta_{i})$-DP wrt $f_{i}$ is $(\epsilon,\delta)$-DP for the same $(\epsilon,\delta)$ that holds for composing noninteractive mechanisms that are $(\epsilon_{i},\delta_{i})$-DP.
[…]
Figure 3. Illustration of the interactions between an honest adversary $\mathcal{A}$ and the concurrent composition of $k$ continual mechanisms
Notably, this result allows the adversary to adaptively and concurrently choose the neighboring sequences fed to each of the $\mathcal{M}_{i}$’s based on the responses received from the other mechanisms.
Theorem 1.2 is proven by using the aforementioned reduction from continual to interactive mechanism. Then we use Lyu’s result  [18] for $(\epsilon,\delta)$-DP interactive mechanisms and extend it to hold for infinite (but discrete) answer spaces and an unbounded number of rounds of interaction as follows. In Lyu’s proof the restriction on the number of interactions comes from a backward definition of a so-called control function. We define a function which is the limit over this control function and, using monotone convergence, show that this “limit function” satisfies all the necessary conditions to be used in the rest of Lyu’s proof. This extends the result to an unbounded number of rounds of interaction. Furthermore, Lyu’s proof used an algorithmic construction iterating over the answer set, implying that the answer set must be finite. Instead with a mathematical approach, we extend the proof to work with a countable, infinite answer set.
While we focus on composing approximate DP with fixed privacy loss parameters $(\epsilon_{i},\delta_{i})$ in Theorem 1.2, we expect that there should also be analogues of it for other settings in which concurrent composition of interactive mechanisms have been studied, such as Rényi DP [18], $f$-DP [21], and adaptively chosen privacy-loss parameters [10].
Next, we explore the concurrent and continual analogue of parallel composition of differential privacy. To review parallel composition, suppose that $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$ are each $(\epsilon_{0},\delta_{0})$-DP noninteractive mechanisms with respect to $\sim$, where each $\mathcal{M}_{i}$ expects a dataset $x_{i}$ as input, and for a parameter $k\in\{0,1,\ldots,\ell\}$, we define a neighboring relation $\sim_{k}$ on $\ell$-tuples as follows: $(x_{1}[0],\ldots,x_{\ell}[0])\sim_{k}(x_{1}[1],\ldots,x_{\ell}[1])$ if:
  1. (1)
    For every $i\in[\ell]$, $x_{i}[0]\sim x_{i}[1]$ or $x_{i}[0]=x_{i}[1]$.
  2. (2)
    For at most $k$ values of $i\in[\ell]$, $x_{i}[0]\neq x_{i}[1]$.
We refer to the mechanism $\mathcal{M}^{\prime}(\cdot)=(\mathcal{M}_{1}(\cdot),\ldots,\mathcal{M}_{\ell}(\cdot))$ with neighbor relation $\sim_{k}$ as the $k$-sparse parallel composition of $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$. If $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$ are noninteractive mechanism, then it is known that $\mathcal{M}^{\prime}$ satisfies the same level of $(\epsilon,\delta)$-DP as the (standard) composition of $k$ (rather than $\ell$) $(\epsilon_{0},\delta_{0})$-DP mechanisms.
We study the extent to which parallel composition extends to the concurrent composition of continual mechanisms, meaning the neighboring relation is as above, but the adversary can interact concurrently with the mechanisms, and in particular decide adaptively on which $k$ of them to have different update sequences. Moreover, we also consider the case where $\ell$ is unbounded, meaning the adversary (or honest analyst) can adaptively decide how many mechanisms to interact with. We refer to this as concurrent $k$-sparse parallel composition.
First, we extend the parallel composition of noninteractive mechanisms to concurrent composition of interactive mechanisms (where there are adaptive queries but datasets $x_{i}$ are static, rather than dynamically updated).
Theorem 1.3 (concurrent parallel composition of interactive mechanisms).
For every $\epsilon_{0},\delta_{0}\geq 0$ and $k\in\mathbb{N}$, and every $(\epsilon_{0},\delta_{0})$-DP interactive mechanism $\mathcal{M}_{0}$, the concurrent $k$-sparse parallel composition of an unbounded number of copies of $\mathcal{M}_{0}$ satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},\delta_{0})$-DP mechanisms.
As discussed later in the introduction, we also prove a generalization of Theorem 1.3 for the case of composing interactive mechanisms $\mathcal{M}_{i}$ with differing $(\epsilon_{i},\delta_{i})$ parameters.
Then we turn to the case of continual mechanisms, where there can also be adaptive dataset updates. We prove a negative result for approximate DP ($\delta_{0}>0$):
Theorem 1.4 (concurrent parallel composition of approximate DP continual mechanisms, informally stated).
For every $\delta_{0}>0$, there is an $(0,\delta_{0})$-DP continual mechanism $\mathcal{M}_{0}$ such that for every $\ell\in\mathbb{N}$, the concurrent 1-sparse parallel composition of $\ell$ copies of $\mathcal{M}_{0}$ is not $(0,\delta)$-DP for any $\delta<1-(1-\delta_{0})^{\ell}$.
This theorem says that, at least in terms of the $\delta$ parameter, concurrent parallel composition of continual mechanisms does not provide any benefit over non-parallel composition, since the bound of $1-(1-\delta_{0})^{\ell}$, is the bound that we would get by just applying composition over all $\ell$ mechanisms.
However, we prove that if $\delta_{0}=0$, concurrent parallel composition behaves just like non-interactive parallel composition:
Theorem 1.5 (concurrent parallel composition of pure DP continual mechanisms, informally stated).
For every $\epsilon_{0}\geq 0$ and integers $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},0)$-DP mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms.
We note that the constraint of pure DP is only on the mechanisms being composed; we allow $\delta>0$ in the final privacy guarantee, so we can still, for example, recover the advanced composition bound applied to pure DP mechanisms. Like Theorem 1.3, we also prove a generalization of Theorem 1.5 for the case of composing mechanisms $\mathcal{M}_{i}$ with differing $(\epsilon_{i},\delta_{i})$ parameters.
To prove this theorem, we add a twist to the concurrent composition theorems for interactive mechanisms. Specifically, the papers [20, 18] proved their composition theorems for pure and approximate DP by showing that for every $(\epsilon_{0},\delta_{0})$-DP interactive mechanism $\mathcal{M}$ and pair $x_{0}\sim x_{1}$ of adjacent datasets, $\mathcal{M}$ can be simulated as an interactive post-processing of the noninteractive $(\epsilon_{0},\delta_{0})$-DP randomized response mechanism $\mathrm{RR}_{\epsilon_{0},\delta_{0}}$. That is, there is an IPM $\mathcal{T}$ (with no input) such that for each $b\in\{0,1\}$, $\mathcal{M}(x_{b})$ is equivalent to $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon_{0},\delta_{0}}(b)$.
We show that when $\delta_{0}=0$, the use of $\mathrm{RR}_{\epsilon_{0},\delta_{0}}$ by $\mathcal{T}$ can be delayed until the time that it receives a query that would have different answer distributions under $\mathcal{M}(x_{0})$ and $\mathcal{M}(x_{1})$. While intuitive, Theorem 1.4 tells us that this fact fails when $\delta_{0}>0$.
It is an interesting open problem to see if Theorem 1.5 can be extended to composing approximate DP mechanisms ($\delta_{0}>0$), if we allow an additional slackness in the final $\delta$, as shown necessary by Theorem 1.4.
An Algorithmic Application.
Composition theorems for differential privacy have two main uses. One is to bound the accumulated privacy loss when DP systems are used repeatedly over time. Another is to facilitate the design of complex DP algorithms from simpler ones. The concurrent composition theorems for interactive mechanisms have already found an application in the privacy analysis of matching algorithms in local edge differential privacy model [4]. (This work also has a result in the continual observation model, but the privacy is only proven for oblivious adversaries.)
Our concurrent composition theorem allows for a simpler and more modular privacy analysis of continual mechanisms built out of other continual mechanisms. In particular, we give a new privacy proof for the continual monotone histogram mechanism of Henzinger, Sricharan, and Steiner [12]. This mechanism involves concurrent executions of a simpler continual histogram mechanism and multiple instantations of the Sparse Vector Technique mechanism, and the original privacy analysis was rather intricate even for oblivious adversaries (because the mechanism itself is adaptive in its queries and dataset updates to the building-block mechanisms). With our concurrent composition theorems (both Theorem 1.2 and Theorem 1.5), we obtain a much simpler privacy proof. Moreover, for the case of approximate DP, the original proof was only for oblivious adversaries, whereas our proof also applies to adaptive adversaries.
[…]
Figure 4. Illustration of the interactions between an honest adversary $\mathcal{A}$ and a mechanism with $k$ sub-mechanisms
General Conditions for Privacy Analysis of Continual Mechanisms.
In the above application, the adversary does not directly interact with the sub-mechanisms, but rather it is the “main” HSS mechanism $\mathcal{M}$ that issues messages to its sub-mechanisms, based on the messages it receives. (See Figure 4 for an illustration.) These messages are not necessarily identical to the messages that $\mathcal{M}$ receives. Instead based on the input that $\mathcal{M}$ receives from an analyst or an adversary (passed through $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(\cdot)$), it generates an input sequence for its sub-mechanisms $\mathcal{M}_{i}$ that, on neighboring inputs to $\mathcal{M}$, must be identical or neighboring for their respective verification functions $f_{i}$. Note that $\mathcal{M}$ may fix the set of sub-mechanisms at initialization or select these mechanisms adaptively over time. For example, in the application above, the continual monotone histogram mechanism maintains $d$ fixed binary counters while adaptively instantiating an unbounded number of SVT and Laplace mechanisms as need.
The privacy analysis of such mechanisms $\mathcal{M}$ is challenging as the creation of sub-mechanisms and their inputs is adaptive and dependent on each other. Moreover, in contrast to concurrent composition of interactive mechanisms, $\mathcal{M}$ has direct access to the dataset updates, which cannot be revealed without differentially private processing (which needs to be done by the sub-mechanisms). We give a set of conditions for $\mathcal{M}$ and prove a theorem that shows that under these conditions the privacy of $\mathcal{M}$ can directly be computed by applying our concurrent composition theorems to the continual sub-mechanisms. The general idea is that $\mathcal{M}$ should only use its input sequence to generate inputs for the differentially private sub-mechanisms and any other action of $\mathcal{M}$, such as creating new sub-mechanism or generating a response, should be taken based only on the privatized outputs of the sub-mechanisms. This will ensure that $\mathcal{M}$ does not use any “privacy budget” in addition to the privacy budget used by the sub-mechanisms. For example, if $\mathcal{M}$ executes $k$ continual sub-mechanisms and, upon receiving neighboring inputs, issues neighboring message sequences for the sub-mechanisms, then the privacy of $\mathcal{M}$ equals the privacy of the concurrent composition of the sub-mechanism, as computed by Theorem 1.2. The same holds for the concurrent $k$-sparse parallel composition of sub-mechanisms, as computed in Theorems 1.3 and  1.5.
Our Variants of Concurrent Composition.
Here we elaborate on the different types of concurrent composition that we study in this paper, with more general and more precise formulations than described above.
  1. (1)
    Concurrent Composition of Continual Mechanisms: In the concurrent composition of fixed continual mechanism, $k$ continual mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ are fixed at the beginning, and an adversary adaptively issues pairs of messages for them such that for each $\mathcal{M}_{i}$, the sequences formed by the first messages of all message pairs and the sequence formed by the second messages of all message pairs are neighboring. A secret bit $b\in\{0,1\}$ determines whether each mechanism receives the first or second sequence from each pair. The adversary’s goal is to guess $b$ using the answers of all $k$ mechanisms. Theorem 4.10, which is the formal statement of Theorem 1.2 above, analyzes the privacy of this composition.
    A natural extension is the concurrent composition of continual mechanism with fixed parameters, where instead of fixing the mechanisms themselves, $k$ privacy parameter pairs $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed in advance. The adversary then adaptively selects mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ over time, ensuring that, for each $1\leq i\leq k$, the mechanism $\mathcal{M}_{i}$ is $(\epsilon_{i},\delta_{i})$-DP. The privacy guarantee for this variant is the same as Theorem 4.10 and is stated in Theorem 4.12.
  2. (2)
    Concurrent $k$-Sparse Parallel Composition of Continual Mechanisms: This composition generalizes the concurrent composition of continual mechanisms with $k$ fixed parameters by permitting the creation of an arbitrary number of continual mechanisms. As before, $k$ privacy parameters $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed at the beginning. The adversary adaptively creates a potentially unlimited number of continual mechanisms and issues pairs of messages to each mechanism such that for all but at most $k$ of these mechanisms, each message pair consists of identical messages. Let $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ denote the exceptional mechanisms, which may receive differing message sequences. For each $1\leq i\leq k$, the mechanism $\mathcal{M}_{i}$ must be $(\epsilon_{i},\delta_{i})$-DP, and the sequences of the first and second messages for $\mathcal{M}_{i}$ must be neighboring.
    As in previous compositions, a secret bit $b$ determines whether each mechanism receives the first or second message from each pair, and the adversary seeks to guess $b$ based on the outputs. Theorem 1.5 analyzes this composition in the special case where $\delta_{1}=\dots=\delta_{k}=0$ and $\epsilon_{1}=\dots=\epsilon_{k}=\epsilon_{0}$. The general version for non-identical $\epsilon_{i}$ parameters is given in Theorem 5.2.
  3. (3)
    Concurrent $k$-Sparse Parallel Composition of Interactive Mechanisms: In this composition, the adversary creates an unbounded number of interactive mechanisms over time, instead of continual ones. As before, $k$ privacy parameter pairs $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed in advance. Each time the adversary creates an interactive mechanism, they also select a pair of neighboring datasets for it. Instead of message pairs, the adversary adaptively sends (single) queries to each mechanism. A secret bit $b$ determines whether the mechanism is run on the first or second dataset of its respective pair. The adversary then tries to infer $b$ from the responses of all mechanisms. Theorem 1.3 analyzes the privacy of this composition when all $\mathcal{M}_{i}$ are $(\epsilon_{0},\delta_{0})$-DP. The more general case for mechanisms with possibly different $(\epsilon_{i},\delta_{i})$ parameters is stated in Theorem 5.5.
The paper is organized as follows. Section 2 introduces all necessary notation, including non-interactive mechanisms, interactive mechanisms, and interactive post-processing mechanism. Section 3 defines continual mechanisms, identifiers $\mathcal{I}$, verification functions $f$, verifiers $\mathcal{V}$ and differential privacy for continual mechanisms. Concurrent composition of continual mechanisms is defined in Section 4 and our first concurrent composition theorem is proven. Section 5 introduces concurrent $k$-sparse parallel composition of continual mechanisms and interactive mechanisms, proves the theorems for this type of composition and shows the counterexample. In Section 6, we state the conditions and the theorem for analyzing the privacy of complex continual mechanisms that use non-interactive, interactive and continual sub-mechanisms. Finally, in Section 7, we apply our theorem to analyze the privacy of the continual histogram mechanism of [12]. Appendix A presents our extension of Lyu’s result [18] to infinite answer spaces and an unbounded number of rounds of interaction.
2. Preliminaries
2.1. Non-interactive Mechanisms (NIM)
Let $\mathcal{X}$ be a family of datasets. A randomized mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is a randomized function that maps a dataset $x\in\mathcal{X}$ to an element of $\mathcal{Y}$. We call $\mathcal{M}$ a non-interactive mechanism (NIM) since it halts immediately after returning an output.
To define differential privacy, we require the definition of indistinguishability for random variables:
Definition 2.1 ($(\epsilon,\delta)$-indistinguishability [16]).
For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, two random variables $X_{1}$ and $X_{2}$ over the same domain $\mathcal{Y}$ are $(\epsilon,\delta)$-indistinguishable if for any subset $Y\subseteq\mathcal{Y}$,
$\Pr[X_{1}\in Y]\leq e^{\epsilon}\cdot\Pr[X_{2}\in Y]+\delta,$
and
$\Pr[X_{2}\in Y]\leq e^{\epsilon}\cdot\Pr[X_{1}\in Y]+\delta.$
Differential privacy is defined relative to a neighbor relation, which is a binary relation on the family of datasets $\mathcal{X}$. For example, we often consider two tabular datasets $x,x^{\prime}$ to be neighboring if they differ only by the presence or absence of a single record. Differential privacy for NIMs is defined as follows:
Definition 2.2 ($(\epsilon,\delta)$-DP for NIMs [6]).
For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, a randomized mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-differentially private (or $(\epsilon,\delta)$-DP for short) with respect to (w.r.t.) a neighbor relation $\sim$ if for every two datasets $x_{0},x_{1}\in\mathcal{X}$ satisfying $x_{0}\sim x_{1}$, the random variables $\mathcal{M}(x_{0})$ and $\mathcal{M}(x_{1})$ are $(\epsilon,\delta)$-indistinguishable.
An example of an $(\epsilon,\delta)$-DP NIM is the approximate randomized response, defined as follows:
Definition 2.3 ($\mathrm{RR}_{\epsilon,\delta}$).
For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, the approx randomized response $\mathrm{RR}_{\epsilon,\delta}$ is a non-interactive mechanism that takes a bit $b\in\{0,1\}$ as input and returns a pair according to the following distribution:
$\mathrm{RR}_{\epsilon,\delta}(b)=\begin{cases}(b,\top)&\quad\text{w.p. $\delta$}\\ (b,\bot)&\quad\text{w.p. $(1-\delta)\frac{e^{\epsilon}}{1+e^{\epsilon}}$}\\ (1-b,\bot)&\quad\text{w.p. $(1-\delta)\frac{1}{1+e^{\epsilon}}$}\end{cases}$
Lemma 2.4 ([15]).
For every $\epsilon\geq 0$ and $0\leq\delta\leq 1$, the approx randomized response $\mathrm{RR}_{\epsilon,\delta}$ is $(\epsilon,\delta)$-DP.
The reason [15] introduced $\mathrm{RR}_{\epsilon,\delta}$ is that it is ”universal” among $(\epsilon,\delta)$-DP NIMs in the following sense:
Theorem 2.5 ([15]).
A mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-DP w.r.t. a neighbor relation $\sim$ if and only if for any $x_{0}\sim x_{1}$, there exists a randomized ”post-processor” $\mathcal{P}:\{0,1\}\times\{\top,\bot\}\to\mathcal{Y}$ such that for both $b\in\{0,1\}$, $\mathcal{M}(x_{b})$ is identically distributed to $\mathcal{P}(\mathrm{RR}_{\epsilon,\delta})$.
2.2. Interactive Mechanisms (IM)
Unlike an NIM, an interactive mechanism (IM) continues interacting with the analyst and answers multiple adaptively asked queries. An IM is defined formally as follows.
Definition 2.6 (Interactive Mechanism (IM)).
An interactive mechanism (IM) is a randomized state function $\mathcal{M}:S_{\mathcal{M}}\times Q_{\mathcal{M}}\rightarrow S_{\mathcal{M}}\times A_{\mathcal{M}}$, where
  • $S_{\mathcal{M}}$ is the state space of $\mathcal{M}$,
  • $Q_{\mathcal{M}}$ is the set of possible incoming messages or queries for $\mathcal{M}$,
  • $A_{\mathcal{M}}$ is the set of possible outgoing messages or answers generated by $\mathcal{M}$.
Upon receiving a query $m\in Q_{\mathcal{M}}$, according to some distribution, $\mathcal{M}$ updates its current state $s\in S_{\mathcal{M}}$ to a new state $s^{\prime}\in S_{\mathcal{M}}$ and generates an answer $m^{\prime}\in A_{\mathcal{M}}$, which is described by $(s^{\prime},m^{\prime})\leftarrow\mathcal{M}(s,m)$. The set $S_{\mathcal{M}}^{\text{init}}\subseteq S_{\mathcal{M}}$ denotes the set of possible initial states for $\mathcal{M}$. For $s^{\text{init}}\in S_{\mathcal{M}}^{\text{init}}$, $\mathcal{M}(s^{\text{init}})$ represents the interactive mechanism $\mathcal{M}$ with the initial state $s^{\text{init}}$.
Remark 2.7.
In the differential privacy literature, a query typically refers to a question about the dataset held by a mechanism. However, in this paper, we use query as a shorthand for query message, meaning an input message to a mechanism that requires a response.
The state of an IM $\mathcal{M}$ serves as its memory. The initial state of $\mathcal{M}$ can include secret information, such as a dataset $x\in\mathcal{X}$, or it can simply be an empty memory. Thus, in some cases, the size of the initial state space $S_{\mathcal{M}}^{\text{init}}$ is as large as the family of datasets $\mathcal{X}$, while in some other cases, $S_{\mathcal{M}}^{\text{init}}$ is a singleton consisting of a single state representing the empty memory.
Notation 2.8.
If $S_{\mathcal{M}}^{\text{init}}$ is a singleton $\{s\}$, we denote $\mathcal{M}(s)$ as $\mathcal{M}$.
2.2.1. Interaction of IMs.
Two interactive mechanisms $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ satisfying $Q_{\mathcal{M}_{1}}=A_{\mathcal{M}_{2}}$ and $Q_{\mathcal{M}_{2}}=A_{\mathcal{M}_{1}}$ can interact with each other through message passing. For initial states $s_{1}\in S_{\mathcal{M}_{1}}^{\text{init}}$ and $s_{2}\in S_{\mathcal{M}_{2}}^{\text{init}}$, the interaction between $\mathcal{M}_{1}(s_{1})$ and $\mathcal{M}_{2}(s_{2})$ is defined as follows: It starts with round 1. Mechanism $\mathcal{M}_{1}(s_{1})$ sends a message to $\mathcal{M}_{2}(s_{2})$ in the odd rounds, and $\mathcal{M}_{2}(s_{2})$ responds in the subsequent even rounds. The interaction continues until one of the mechanisms halts the communication. A mechanism halts the communication by sending the specific message halt at its round.
In Definition 2.6, we defined an IM as a stateful randomized algorithm mapping its current state and an input message to a new state and an output message. To initiate the interaction and send the first message $m_{1}$, $\mathcal{M}_{1}(s_{1})$ invokes $\mathcal{M}_{1}(s_{1},\lambda)$, where $\lambda$ is the empty message.
Definition 2.9 (View of an IM).
Let $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ be two interactive mechanisms such that $Q_{\mathcal{M}_{1}}=A_{\mathcal{M}_{2}}$ and $Q_{\mathcal{M}_{2}}=A_{\mathcal{M}_{1}}$. Let $s_{1}$ and $s_{2}$ be initial states of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, respectively. Consider the interaction between $\mathcal{M}_{1}(s_{1})$ and $\mathcal{M}_{2}(s_{2})$, initiated by a message from $\mathcal{M}_{1}(s_{1})$. The view of $\mathcal{M}_{1}(s_{1})$ in this interaction, denoted $\mathrm{View}(\mathcal{M}_{1}(s_{1}),\mathcal{M}_{2}(s_{2}))$, is defined as
$\mathrm{View}(\mathcal{M}_{1}(s_{1}),\mathcal{M}_{2}(s_{2}))=(r_{\mathcal{M}_{1}},m_{1},m_{2},m_{3},\dots),$
where $r_{\mathcal{M}_{1}}$ represents the internal randomness of $\mathcal{M}_{1}$, and $(m_{1},m_{2},\dots)$ is the transcript of messages exchanged between $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. We define $\Pi_{\mathcal{M}_{1},\mathcal{M}_{2}}$ as the family of all possible views over all initial states $s_{1}\in S_{\mathcal{M}_{1}}^{\text{init}}$ and $s_{2}\in S_{\mathcal{M}_{2}}^{\text{init}}$.
Throughout the paper, whenever discussing the interaction between two IMs or the view of a mechanism in an interaction, we implicitly assume that the query and answer sets of the involved mechanisms match, i.e., $Q_{\mathcal{M}_{1}}=A_{\mathcal{M}_{2}}$ and $Q_{\mathcal{M}_{2}}=A_{\mathcal{M}_{1}}$.
Definition 2.10 (Equivalent IMs).
Let $\mathcal{M}$ and $\mathcal{M}^{\prime}$ be two interactive mechanisms with $Q_{\mathcal{M}}=Q_{\mathcal{M}^{\prime}}$ and $A_{\mathcal{M}}=A_{\mathcal{M}^{\prime}}$. Let $s$ and $s^{\prime}$ be initial states for $\mathcal{M}$ and $\mathcal{M}^{\prime}$, respectively. The mechanisms $\mathcal{M}(s)$ and $\mathcal{M}^{\prime}(s^{\prime})$ are equivalent if for any adversary $\mathcal{A}$ with $A_{\mathcal{A}}=Q_{\mathcal{M}}$ and $Q_{\mathcal{A}}=A_{\mathcal{M}}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}^{\prime}(s^{\prime}))$ are identically distributed.
2.2.2. Differential Privacy for IMs.
To define differential privacy for an interactive mechanism $\mathcal{M}$, we compare the views of a so-called adversary interacting with $\mathcal{M}(s)$ and $\mathcal{M}(s^{\prime})$ for neighboring initial states $s,s^{\prime}$.
Definition 2.11 (Adversary).
An (adaptive) adversary $\mathcal{A}$ is an interactive mechanism with a singleton initial state space. Let $s^{\text{init}}$ be the only initial state of $\mathcal{A}$. By Notation 2.8, we write $\mathcal{A}$ instead of $\mathcal{A}(s^{\text{init}})$.
Throughout this paper, adversaries always send the first message. Moreover, since adversaries are stateful algorithms, they can choose their messages adaptively based on the history of messages exchanged, meaning that all adversaries in this paper are adaptive.
Existing works [20, 21, 18, 10] view the initial state of $\mathcal{M}$ is a dataset, i.e., $S_{\mathcal{M}}^{\text{init}}=\mathcal{X}$, and (as is usual) they define neighbor relations as binary relations on $\mathcal{X}$. However, for us it will be convenient to conceptually separate $S_{\mathcal{M}}^{\text{init}}$ from the data (which is changing continuously over time), but still define DP w.r.t. neighbor relations on $S_{\mathcal{M}}^{\text{init}}$. This leads to the following definition of differential privacy for IMs.
Definition 2.12 ($(\epsilon,\delta)$-Differential Privacy for IMs).
Let $\mathcal{M}$ be an interactive mechanism, and let $\sim$ be a neighbor relation on $S_{\mathcal{M}}^{\text{init}}$. For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private (or $(\epsilon,\delta)$-DP for short) w.r.t. $\sim$ against adaptive adversaries if for every pair of neighboring initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$ satisfying $s_{0}\sim s_{1}$ and every adaptive adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s_{0}))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}(s_{1}))$ are $(\epsilon,\delta)$-indistinguishable.
2.3. Interactive Post-Processing (IPM)
For NIMs, a randomized post-processing function $\mathcal{T}:\mathcal{Y}\to\mathcal{Z}$ maps the output of a mechanism $\mathcal{N}:\mathcal{X}\to\mathcal{Y}$ to an element of $\mathcal{Z}$. Differential privacy for NIMs is known to be preserved under post-processing.
Haney et al. [10] generalize post-processing functions to interactive post-processing mechanisms (IPMs) as follows and show that interactive post-processing preserves privacy guarantees for IMs.
Definition 2.13 (Interactive Post-Processing Mechanism (IPM) [10]).
An interactive post-processing mechanism (IPM) is a stateful randomized algorithm $\mathcal{P}:S_{\mathcal{P}}\times\left((\{{\texttt{L}}\}\times Q_{\mathcal{P}}^{L})\cup(\{{\texttt{R}}\}\times Q_{\mathcal{P}}^{R})\right)\rightarrow S_{\mathcal{P}}\times\left((\{{\texttt{L}}\}\times A_{\mathcal{P}}^{L})\cup(\{{\texttt{R}}\}\times A_{\mathcal{P}}^{R})\right)$ that stands between two interactive mechanisms, referred to as the left and right mechanisms, communicating with them via message passing. The state space of $\mathcal{P}$ is denoted by $S_{\mathcal{P}}$, while $Q_{\mathcal{P}}^{L}$ and $Q_{\mathcal{P}}^{R}$ represent the sets of possible messages received by $\mathcal{P}$ from the left and right mechanisms, respectively. Similarly, $A_{\mathcal{P}}^{L}$ and $A_{\mathcal{P}}^{R}$ denote the sets of possible messages $\mathcal{P}$ sends to the left and right mechanisms. $\mathcal{P}$ randomly maps its current state $s\in S_{\mathcal{P}}$, an indicator $v\in\{{\texttt{L}},{\texttt{R}}\}$ specifying whether the incoming message is from the left or right mechanism, and a message $m\in Q_{\mathcal{P}}^{L}\cup Q_{\mathcal{P}}^{R}$ to a new state $s^{\prime}\in S_{\mathcal{P}}$, an indicator $v^{\prime}\in\{{\texttt{L}},{\texttt{R}}\}$ specifying whether the outgoing message is for the left or right mechanism, and a message $m^{\prime}\in A_{\mathcal{P}}^{L}\cup A_{\mathcal{P}}^{R}$. This is described by $(s^{\prime},v^{\prime},m^{\prime})\leftarrow\mathcal{P}(s,v,m)$.
The post-processing of an NIM $\mathcal{N}$ by a non-interactive post-processing function $\mathcal{T}$, denoted as $\mathcal{T}\circ\mathcal{N}$, is an NIM that first runs $\mathcal{N}$ on the input dataset and then applies $\mathcal{T}$ to $\mathcal{N}$’s output, returning $\mathcal{T}$’s output. In contrast, the post-processing of an IM $\mathcal{M}$ by an IPM $\mathcal{P}$, denoted by $\mathcal{P}\!\circ^{*}\!\mathcal{M}$, is an IM that runs $\mathcal{P}$ and $\mathcal{M}$ internally. In an interaction with another IM $\mathcal{A}$, $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ forwards its input messages from $\mathcal{A}$ to $\mathcal{P}$ as left messages. Then, $\mathcal{P}$ and $\mathcal{M}$ communicate multiple times, with $\mathcal{M}$ acting as the right mechanism, until $\mathcal{P}$ returns a response for its left mechanism, at which point $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ returns this response to $\mathcal{A}$. Note that $\mathcal{P}$ may return an answer for its left mechanism immediately after receiving a left message, without interacting with $\mathcal{M}$. The formal definition of $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ is given below:
Definition 2.14 (Post-Processing of an IM $\mathcal{M}$ by an IPM $\mathcal{P}$ ($\mathcal{P}\!\circ^{*}\!\mathcal{M}$)[10]).
Let $\mathcal{M}$ be an IM and $\mathcal{P}$ an IPM such that $Q_{\mathcal{P}}^{R}=A_{\mathcal{M}}$ and $A_{\mathcal{P}}^{R}=Q_{\mathcal{M}}$. The post-processing of $\mathcal{M}$ by $\mathcal{P}$ is an interactive mechanism $\mathcal{P}\!\circ^{*}\!\mathcal{M}:S_{\mathcal{P}\!\circ^{*}\!\mathcal{M}}\times Q_{\mathcal{P}}^{L}\to S_{\mathcal{P}\!\circ^{*}\!\mathcal{M}}\times A_{\mathcal{P}}^{L}$, defined in Algorithm 1. The initial state space of $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ equals $S_{\mathcal{P}\!\circ^{*}\!\mathcal{M}}^{\text{init}}=S_{\mathcal{M}}^{\text{init}}\times S_{\mathcal{P}}^{\text{init}}$. For readability, when $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ is initialized with the state $(s^{\mathcal{M}},s^{\mathcal{P}})\in S_{\mathcal{P}\!\circ^{*}\!\mathcal{M}}^{\text{init}}$, we denote it as $\mathcal{P}(s^{\mathcal{P}})\!\circ^{*}\!\mathcal{M}(s^{\mathcal{M}})$ instead of $\left(\mathcal{P}\!\circ^{*}\!\mathcal{M}\right)\left((s^{\mathcal{M}},s^{\mathcal{P}})\right)$.
Algorithm 1 Post-Processing of $\mathcal{M}$ by $\mathcal{P}$, denoted by $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ [10].
procedure $\left(\mathcal{P}\!\circ^{*}\!\mathcal{M}\right)$($s,m$):
     $(s^{\mathcal{M}},s^{\mathcal{P}})\leftarrow s$
     $(s^{\mathcal{P}},v,m^{\prime})\leftarrow\mathcal{P}(s^{\mathcal{P}},{\texttt{L}},m)$
     while $v={\texttt{R}}$ do
         $(s^{\mathcal{M}},m)\leftarrow\mathcal{M}(s^{\mathcal{M}},m^{\prime})$
         $(s^{\mathcal{P}},v,m^{\prime})\leftarrow\mathcal{P}(s^{\mathcal{P}},{\texttt{R}},m)$
     end while
     return $((s^{\mathcal{M}},s^{\mathcal{P}}),m^{\prime})$
end procedure
Haney et al. [10] shows that if an IM $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private, then for any IPM $\mathcal{P}$ with a singleton initial state space, $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ is also $(\epsilon,\delta)$-differentially private. This result is formally stated below.
Lemma 2.15 (Post-Processing Lemma for IMs [10]).
Let $\mathcal{M}$ be an IM $\mathcal{M}$ with an initial state space $S_{\mathcal{M}}^{\text{init}}$ and $\mathcal{P}$ be an IPM with a singleton initial state space $S_{\mathcal{P}}^{\text{init}}=\{s\}$. Let $\sim$ be a neighbor relation on $S_{\mathcal{M}}^{\text{init}}$. Define a binary relation $\sim_{*}$ on $S_{\mathcal{P}\!\circ^{*}\!\mathcal{M}}^{\text{init}}$ such that two initial states $(s_{\mathcal{M}}^{1},s)$ and $(s_{\mathcal{M}}^{2},s)$ satisfy $(s_{\mathcal{M}}^{1},s)\sim_{*}(s_{\mathcal{M}}^{2},s)$ if and only if $s_{\mathcal{M}}^{1}\sim s_{\mathcal{M}}^{2}$. If $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. the neighbor relation $\sim$ against adaptive adversaries, then $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ is also $(\epsilon,\delta)$-DP w.r.t. the neighbor relation $\sim_{*}$ against adaptive adversaries.  [^2] In [10], the definitions of IMs and IPMs are more restrictive. Specifically, for any IM $\mathcal{M}$, $S_{\mathcal{M}}=Q_{\mathcal{M}}=A_{\mathcal{M}}=\{0,1\}^{*}$, and for any IPM $\mathcal{P}$, $S_{\mathcal{P}}=M_{\mathcal{P}}=\{0,1\}^{*}$; However, their results do not rely on these restrictions, and the interactive post-processing lemma holds for arbitrary sets $S_{\mathcal{M}}$, $Q_{\mathcal{M}}$, $A_{\mathcal{M}}$, $S_{\mathcal{P}}$, and $M_{\mathcal{P}}$.
IPMs are not limited to interacting with IMs. Two IPMs can interact with each other as well, as follows:
Definition 2.16 (Chain of IPMs).
Let $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ be two IPMs satisfying $Q_{\mathcal{P}_{1}}^{R}=A_{\mathcal{P}_{2}}^{L}$ and $A_{\mathcal{P}_{1}}^{R}=Q_{\mathcal{P}_{2}}^{L}$. $\mathcal{P}_{1}\!\circ^{*}\!\mathcal{P}_{2}$ is an IPM with state space $S_{\mathcal{P}_{1}\!\circ^{*}\!\mathcal{P}_{2}}=S_{\mathcal{P}_{1}}\times S_{\mathcal{P}_{2}}$, left query space $Q_{\mathcal{P}_{1}\!\circ^{*}\!\mathcal{P}_{2}}^{L}=Q_{\mathcal{P}_{1}}^{L}$, left answer space $A_{\mathcal{P}_{1}\!\circ^{*}\!\mathcal{P}_{2}}^{L}=A_{\mathcal{P}_{1}}^{L}$, right query space $Q_{\mathcal{P}_{1}\!\circ^{*}\!\mathcal{P}_{2}}^{R}=Q_{\mathcal{P}_{2}}^{R}$, and right answer space $A_{\mathcal{P}_{1}\!\circ^{*}\!\mathcal{P}_{2}}^{R}=A_{\mathcal{P}_{2}}^{R}$. This mechanism is formally defined in Algorithm 2. For readability, when $\mathcal{P}_{1}\!\circ^{*}\!\mathcal{P}_{2}$ is initialized with $(s^{1},s^{1})\in S_{\mathcal{P}_{1}\!\circ^{*}\!\mathcal{P}_{2}}^{\text{init}}$, we denote it as $\mathcal{P}_{1}(s^{1})\!\circ^{*}\!\mathcal{P}_{2}(s^{2})$ instead of $\left(\mathcal{P}_{1}\!\circ^{*}\!\mathcal{P}_{2}\right)\left((s^{1},s^{2})\right)$.
Algorithm 2 Interactive Post-Processing Mechanism $\mathcal{P}_{1}\!\circ^{*}\!\mathcal{P}_{2}$.
procedure $\left(\mathcal{P}_{1}\!\circ^{*}\!\mathcal{P}_{2}\right)$($s,v,m$):
     $(s^{1},s^{2})\leftarrow s$
     while True do
         if $v={\texttt{L}}$ then
              $(s^{1},v,m)\leftarrow\mathcal{P}_{1}(s^{1},v,m)$
              if $v={\texttt{L}}$ then
                  Break the while loop
              end if
         else
              $(s^{2},v,m)\leftarrow\mathcal{P}_{2}(s^{2},v,m)$
              if $v={\texttt{R}}$ then
                  Break the while loop
              end if
         end if
     end while
     return $((s^{1},s^{2}),v,m)$
end procedure
3. Continual Mechanisms
In this section, we introduce continual mechanisms (CMs) and propose a general definition of differential privacy (DP) for them. This definition unifies and extends DP for IMs and DP under continual observation. Its generality and simplicity allow us to prove (filter) concurrent composition theorems for CMs without requiring complicated proofs. We later apply these theorems to simplify existing analyses and strengthen results in the continual observation setting.
This section is based on two crucial observations: (i) A message, as sent or received by IMs and IPMs, is a general and powerful concept capable of encoding various types of instructions, and (ii) the essence of DP is to protect the privacy of a single bit that distinguishes neighboring states, sequences, etc. The second observation has been used many times in the DP literature and was specifically used in [13] to model adaptivity in continual observation. Using these key observations, we first reformulate the definition of DP for IMs and continual observation. Then, inspired by the new formulations, we introduce continual mechanisms (CMs) and formally define DP in this setting.
3.1. An Alternative Definition of DP for IMs.
By Definition 2.12, an interactive mechanism $\mathcal{M}$ with initial state space $S_{\mathcal{M}}^{\text{init}}$ is $(\epsilon,\delta)$-DP w.r.t. a neighboring relation $\sim$, if for any two initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$ such that $s_{0}\sim s_{1}$ and any adversary $\mathcal{A}$, the view of $\mathcal{A}$ in the interaction with $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ is $(\epsilon,\delta)$-indistinguishable. This means that for the worst-case (i.e., adversarial) choice of neighboring initial states, an adversary by adaptively asking queries should not be able to determine whether $\mathcal{M}$ is running on $s_{0}$ or $s_{1}$.
In an alternative definition, an adversary sends a pair of neighboring initial states $(s_{0},s_{1})$ as its first message and adaptively asks a sequence of queries $q_{1},q_{2},\dots$ in the subsequent messages. In this setting, $\mathcal{M}$ receives one of $s_{0}$ and $s_{1}$ as its first message and responds with an acknowledgment message $\top$. Thereafter, $\mathcal{M}$ answers queries based on the received state, and the goal of the adversary is to determine whether $\mathcal{M}$ was initialized with $s_{0}$ or $s_{1}$. In this definition, the actual initial state of $\mathcal{M}$ (not the one received as the first message) carries no secret information and serves merely as an empty memory. Therefore, the initial state space of $\mathcal{M}$ consists of a single fixed initial state.
To formalize this, we restrict attention to adversaries that send a pair of neighboring initial states $(s_{0},s_{1})$ as their first message and pairs of identical queries $(q_{i},q_{i})$ in its subsequent messages. We also define an IPM $\mathcal{I}$ with the initial state space $S_{\mathcal{I}}^{\text{init}}=\{0,1\}$ to stand between the adversary and $\mathcal{M}$, filtering the pairs based on its, i.e., $\mathcal{I}$’s, initial state. Specifically, for $b\in\{0,1\}$, in an interaction between an adversary $\mathcal{A}$ and $\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$, upon receiving a message $(m_{0},m_{1})$ from $\mathcal{A}$, $\mathcal{I}(b)$ forwards $m_{b}$ to $\mathcal{M}$ and returns $\mathcal{M}$’s answer unchanged to $\mathcal{A}$. With this formulation, an interactive mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP if for any adversary of the above form, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}[0]\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}[1]\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
Actually, we will enforce the restrictions on the adversary also by an IPM, which we call a verifier $\mathcal{V}$. The verifier checks whether the initial message is a pair of neighboring datasets and the subsequent messages of the adversary are of the form $(q,q)$; only if these checks pass, are the messages forwarded to $\mathcal{I}$.
3.2. An Alternative Definition of DP under Continual Observation.
Dwork, Naor, Pitassi, and Rothblum [7] introduced differential privacy under continual observation, where the dataset is not given at the initialization time and instead evolves over time. In this model, the mechanism receives either an empty record or a data record at each step. If the record is non-empty, the mechanism adds it to its dataset. Then, whether the input record is empty or not, the mechanism answers a predetermined question about its current dataset.
To capture ”event-level privacy” (where the privacy unit is a single record), Dwork et al. [7] define neighboring as a binary relation on sequences of data and empty records: two sequences are neighboring if they are identical except for the presence or absence of a record in a single step. They called a mechanism $(\epsilon,\delta)$-DP under continual observation if the distributions of answer sequences are $(\epsilon,\delta)$-indistinguishable for any two neighboring input sequences. This definition does not allow for data sequences that are adaptively chosen based on the answers released.
Jain, Raskhodnikova, Sivakumar, and Smith [13] formalized the definition of DP against adaptive adversaries in the continual observation model as follows: An adaptive adversary $\mathcal{A}$ chooses a single (data or empty) record at every step, except for one challenge step, selected by the adversary, where $\mathcal{A}$ sends a pair of records. They called a mechanism $\mathcal{M}$ $(\epsilon,\delta)$-differentially private against adaptive adversaries in the continual observation model if for any such adaptive adversary $\mathcal{A}$, the distributions of $\mathcal{M}$’s answer sequences are $(\epsilon,\delta)$-indistinguishable when the first or the second element of the pair in the challenge step is handed to $\mathcal{M}$.
A mechanism $\mathcal{M}$ in the continual observation model can be viewed as an IM with a single possible initial state, representing an empty dataset. The input message set of $\mathcal{M}$ consists of the empty record and all possible data records. Instead of having a challenge step, we can alternatively restrict to adversaries that send pairs of identical records at every step except for one. Under this formulation, the definition of differential privacy proposed by [13] can be restated as follows: A mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP differential privacy against adaptive adversaries if for every such adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}[0]\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}[1]\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable, where $\mathcal{I}$ is the interactive post-processing mechanism discussed previously. Again, below we will enforce the restriction on the adversary via another IPM $\mathcal{V}$.
In this alternative definition, all adversary messages are pairs of identical data or empty records, except for a single message that is a pair of different records. Recall that in the alternative definition of DP for IMs, the first message sent by an adversary must be a pair of neighboring initial datasets, while all subsequent adversary messages must be pairs of identical queries. Next, we generalize the definition of differential privacy by allowing arbitrary conditions on the sequence of pairs sent by adversaries, without restricting the content of the messages.
3.3. Introducing CMs.
As discussed in the alternative formulation of IMs, we can assume that the initial state space of IMs is a singleton and encode its initial state as a message. A continual mechanism is defined as follows:
Definition 3.1 (Continual Mechanism (CM)).
A continual mechanism (CM) is an interactive mechanism with a singleton initial state space.
To the best of our knowledge, all mechanisms in DP literature can be represented as CMs:
  • Any non-interactive mechanism $\mathcal{N}$, such as the Laplace mechanism, can be modeled as a CM that receives a dataset $x$ as its first message and returns $\mathcal{N}(x)$ as a response. Upon receiving any subsequent input message, the CM representing $\mathcal{N}$ halts the communication.
  • Any interactive mechanism $\mathcal{M}$ with an initial state space $S_{\mathcal{M}}^{\text{init}}$ can be modeled as a CM that receives an initial state $s\in S_{\mathcal{M}}^{\text{init}}$ as its first message and responds with an acknowledgment message $\top$. This CM then answers the next input messages the same as $\mathcal{M}(s)$.
  • Mechanisms such as the Sparse Vector Technique (SVT) [8, 11, 17] and the continual counter [6, 2, 3, 5] in the streaming settings can also be represented as CMs receiving two types of messages: question queries and data-update queries. Question queries request information about the current dataset, while data-update queries consist of instructions for modifying the dataset.
3.4. DP for CMs.
To formulate the general definition of DP for CMs, we introduce two IPMs: an identifier $\mathcal{I}$ and a verifier $\mathcal{V}$. As described earlier, an identifier receives a secret bit $b$ as its initial state and maps messages of the form $(m_{0},m_{1})$ to $m_{b}$. To avoid defining a new family of adversaries for each problem or model, we introduce verifiers.
Let $\mathcal{M}$ be a CM and $\mathcal{I}$ be an identifier for $\mathcal{M}$. A verifier $\mathcal{V}$ with a verification function $f$, denoted by $\mathcal{V}[f]$, stands between an adversary $\mathcal{A}$ and the interactive mechanism $\mathcal{I}\!\circ^{*}\!\mathcal{M}$. Upon receiving a message from $\mathcal{A}$, $\mathcal{V}[f]$ applies $f$ to verify whether the sequence of pairs sent by the adversary up to that point satisfies a set of properties. Comparing the sequences formed by the first and second elements of the pairs, $f$ can be used to encode a neighboring property. If the pair sequence satisfies the properties, $\mathcal{V}[f]$ forwards $\mathcal{A}$’s message to $\mathcal{I}\!\circ^{*}\!\mathcal{M}$; otherwise, it halts the communication. The formal definitions of identifiers, verification functions, and verifiers are given below.
Definition 3.2 (Identifier $\mathcal{I}$).
An identifier $\mathcal{I}$ is an IPM with state space $S_{\mathcal{I}}$ and initial state space $S_{\mathcal{I}}^{\text{init}}=\{0,1\}$. $\mathcal{I}$ stands between two IMs, referred to as the left and right mechanisms. While there is no restriction on the format of messages received from the right mechanism, $\mathcal{I}$ only accepts messages of the form $(m_{0},m_{1})$ from its left mechanism. Upon receiving a message $(m_{0},m_{1})$ from the left mechanism, an instance of $\mathcal{I}$ with an initial state $b\in\{0,1\}$ forwards $m_{b}$ to the right mechanism. Upon receiving a message from the right mechanism, $\mathcal{I}(b)$ forwards the message unchanged to the left mechanism.
Definition 3.3 (Verification Function).
A verification function $f$ takes a sequence of messages as input and returns either $\top$ or $\bot$. A message sequence $m_{1},\dots,m_{k}$ is $f$-valid if $f(m_{1},\dots,m_{k})=\top$.
Remark 3.4.
Since the set of all possible messages does not exist, we assume that all message spaces considered in this paper are subsets of some universal set $\mathcal{U}$. For instance, in [10], the universal set is taken to be $\{0,1\}^{*}$. The domain of verification functions is $\mathcal{U}^{*}$.
Definition 3.5 (Verifier $\mathcal{V}$).
For a verification function $f$, a verifier $\mathcal{V}[f]$ is an IPM with the initial state space $S_{\mathcal{V}[f]}^{\text{init}}=\{()\}$, where ()()( ) represents an empty sequence. Since the initial state space of $\mathcal{V}[f]$ is a singleton, we denote an instance of $\mathcal{V}[f]$ with the initial state ()()( ) as $\mathcal{V}[f]$. $\mathcal{V}[f]$ stands between two IMs, referred to as the left and right mechanisms. The state of $\mathcal{V}[f]$ consists of the sequence of messages it has received from the left mechanism so far. Upon receiving a message $m$ from the left mechanism, $\mathcal{V}[f]$ appends $m$ to its state and applies the verification function $f$. If the sequence is $f$-valid, then $\mathcal{V}[f]$ forwards the message $m$ to the right mechanism; otherwise, it halts the communication. $\mathcal{V}[f]$ forwards any message from the right mechanism unchanged to the left mechanism.
Remark 3.6.
By definition, for any continual mechanism $\mathcal{M}$ and any verification function $f$, $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ is an IM with two possible initial states. The mechanism $\mathcal{I}\!\circ^{*}\!\mathcal{M}$ accepts messages of the form $(m_{0},m_{1})$, where both $m_{0}$ and $m_{1}$ are in the query space $Q_{\mathcal{M}}$ of $\mathcal{M}$. Therefore, when analyzing $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$, we assume that the verification function $f$ always returns $\bot$ whenever a message in its input sequence is not in $Q_{\mathcal{M}}\times Q_{\mathcal{M}}$, even if not explicitly say so.
Remark 3.6 implies that $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ can interact with any other IM. Thus, in the following definition of DP for CMs, we can use the general definition of adversaries given in Definition 2.11, without making any assumptions about the family of adversaries. Instead of defining DP w.r.t. neighboring relations, we introduce a more general formulation in terms of verification functions, as follows:
Definition 3.7 ($(\epsilon,\delta)$-Differential Privacy for CMs).
Let $\mathcal{M}$ be a CM and $f$ a verification function. We say that $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private (or $(\epsilon,\delta)$-DP for short) w.r.t. $f$ against adaptive adversaries if for any adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
[…]
Corollary 3.8.
A continual mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. a verification function $f$ against adaptive adversaries if and only if the interactive mechanism $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. the neighboring relation $\sim_{01}$ against adaptive adversaries.
CMs vs. IMs. In principle, CMs are a special class of IMs with a singleton initial state space. Conversely, as described earlier, IMs can be represented by CMs. Both CMs and IMs are defined by a randomized function mapping a state and query message to another state and an answer message, where message is a general concept. The key distinction between CMs and IMs lies in how differential privacy is defined for each.
By Definition 2.6, the initial state space of IMs contains multiple states, and DP for IMs is defined based on the neighboring initial states. However, by Definition 3.7, the initial state space of CMs is singleton, and a verifier forces an adversary to generate pairs such that the sequences of the first and second elements of the pairs satisfy certain properties. While the adversary chooses these two mechanisms, the identifier selects one of them to forward. Thus, the definition of DP for CMs is based on a bit.
4. Concurrent Composition of Continual Mechanisms
Composition of NIMs. For $1\leq i\leq k$, let $\mathcal{N}_{i}$ be an NIM with input domain $\mathcal{X}_{i}$, and let $\sim_{i}$ be a neighboring relation on $\mathcal{X}_{i}$. The composition of the NIMs $\mathcal{N}_{1},\dots,\mathcal{N}_{k}$ is represented by a non-interactive mechanism ${\textrm{Comp}}[\mathcal{N}_{1},\dots,\mathcal{N}_{k}]$ with the domain $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{k}$. This mechanism takes a tuple of datasets $x=(x_{1},\dots,x_{k})$ as input and returns
${{\textrm{Comp}}[\mathcal{N}_{1},\dots,\mathcal{N}_{k}](x)=(\mathcal{N}_{1}(x_{1}),\dots,\mathcal{N}_{k}(x_{k}))}.$
Define $\sim$ as a neighboring relation on $\mathcal{X}$, where two tuples $x=(x_{1},\dots,x_{k})$ and $x^{\prime}=(x_{1}^{\prime},\dots,x_{k}^{\prime})$ satisfy $x\sim x^{\prime}$ if and only if $x_{i}\sim_{i}x_{i}^{\prime}$ for all $1\leq i\leq k$. The composition of NIMs $\mathcal{N}_{1},\dots,\mathcal{N}_{k}$ is said to be $(\epsilon,\delta)$-DP if the non-interactive mechanism ${\textrm{Comp}}[\mathcal{N}_{1},\dots,\mathcal{N}_{k}]$ is $(\epsilon,\delta)$-DP w.r.t. $\sim$ (see Definition 2.2).
In the proofs of this section, we consider ${\textrm{Comp}}[\mathcal{N}_{1},\dots,\mathcal{N}_{k}]$ not as a non-interactive, but instead as an interactive mechanism with initial state space $\mathcal{X}$ that receives $x=(x_{1},\dots,x_{k})$ not as input, but instead as its first message. Immediately after receiving this message it executes ${\textrm{Comp}}[\mathcal{N}_{1},\dots,\mathcal{N}_{k}]$ on its initial state and returns its output. Upon receiving a second message, this interactive mechanism halts the communication.
Concurrent Composition of IMs. Vadhan and Wang [20] extended the definition of the composition of NIMs to the concurrent composition of IMs, where an adversary interacts with $k$ interactive mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ at the same time. In concurrent composition, the adversary generates a message and determines the receiver of this message (one of the $\mathcal{M}_{i}$s) adaptively based on all the previous messages exchanged between the adversary and the mechanisms $M_{1},\dots,M_{k}$. To formalize this, Vadhan and Wang [20] introduced an interactive mechanism ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$, which internally runs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ and, upon receiving a message $(q,j)$, forwards the query $q$ to $\mathcal{M}_{j}$ and returns its response.
The initial state space of ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ is $S_{\mathcal{M}_{1}}^{\text{init}}\times\dots\times S_{\mathcal{M}_{k}}^{\text{init}}$, where $S_{\mathcal{M}_{i}}^{\text{init}}$ is the initial state space of $\mathcal{M}_{i}$. Let $\sim_{i}$ be a neighboring relation on each $S_{\mathcal{M}_{i}}^{\text{init}}$, and define the neighboring relation $\sim$ on $S_{\mathcal{M}_{1}}^{\text{init}}\times\dots\times S_{\mathcal{M}_{k}}^{\text{init}}$ as follows: for any two initial states $s=(s_{1},\dots,s_{k})$ and $s^{\prime}=(s_{1}^{\prime},\dots,s_{k}^{\prime})$, we have $s\sim s^{\prime}$ if and only if $s_{i}\sim_{i}s_{i}^{\prime}$ for every $i\in[k]$.
Vadhan and Wang [20] show that if each IM $\mathcal{M}_{i}$ is $\epsilon_{i}$-DP w.r.t $\sim_{i}$, and if the composition of NIMs $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ is $(\epsilon,\delta)$-DP, then ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ is also $(\epsilon,\delta)$-DP against an adaptive adversary. Lyu [18] extends this result for $(\epsilon_{i},\delta_{i})$-DP IMs.
In the concurrent composition of IMs, queries asked to each mechanism are chosen adaptively, not only based on that mechanism’s previous responses but also using the responses of the other mechanisms up to that point. The results of [20] and [18] show that this concurrency and additional adaptivity do not compromise privacy guarantees. As a result, all $(\epsilon,\delta)$-DP composition theorems for NIMs, such as the basic [6], advanced [9], and optimal [15] composition theorems, can also be applied to the concurrent composition of IMs.
Concurrent Composition of CMs. We extend the definition of the concurrent composition for IMs to the concurrent composition of CMs, where an adversary can create new CMs or ask pairs of queries from the existing ones. Depending on the type of composition, the created mechanisms are constrained, and the sequence of query pairs asked from each CM satisfies certain properties. Based on a secret bit $b\in\{0,1\}$, each CM either receives the first element of every query pair asked from it or the second element. In Section 4.1, we generalize ConComp to ExtConComp and give a formal definition for the concurrent composition of CMs.
To extend Lyu’s result on the privacy guarantee of the concurrent composition of IMs to the concurrent composition of CMs, we want to use the following intermediate lemma from their paper: For every $(\epsilon,\delta)$-DP interactive mechanism $\mathcal{M}$ and any pair of neighboring initial states $s$ and $s^{\prime}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the IMs $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(b)$ are equivalent. However, as we discuss in Section 4.2, there are hidden assumptions in the proof of this lemma. Additionally, we show there that the construction of $\mathcal{P}$ in [20], for $\epsilon_{i}$-DP IMs, satisfies a property, required for a later proof.
In Section 4.3, we analyze the privacy of the concurrent composition of $k$ fixed CMs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$. In this composition, the adversary first creates the mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ and then asks pairs of queries from each mechanism, ensuring that the sequence of query pairs for each mechanism is valid. In Section 4.4, we extend this composition by fixing $k$ privacy parameters $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ and allowing the adversary to adaptively choose the $(\epsilon_{i},\delta_{i})$-DP mechanism $\mathcal{M}_{i}$ over time.
4.1. Formalization.
We extend the definition of ConComp in [20] and introduce ExtConComp that allows the creation of new mechanisms in addition to queries to existing ones. A creation query $(\mathcal{M},s,\epsilon,\delta,f)$ is a message that requests the creation of a continual mechanism $\mathcal{M}$ with the initial state $s$ and asserts that $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. the verification function $f$. The mechanism ExtConComp accepts creation queries and messages of the form $(q,i)$, where $q$ is a query for the $i$-th created mechanism. The formal definition of ExtConComp is as follows:
Definition 4.1.
Extended concurrent composition, ExtConComp, is a continual mechanism that receives two types of messages:
  • A creation query $(\mathcal{M},s,\epsilon,\delta,f)$, where $\mathcal{M}$ is a CM with a single initial state $s$ that is $(\epsilon,\delta)$-DP w.r.t. a verification function $f$, and
  • a query-index pair $(q,i)$, where $i$ is a positive integer, and $q$ is any message.
Upon receiving $(\mathcal{M},s,\epsilon,\delta,f)$, ExtConComp creates an instance of $\mathcal{M}$ with the initial state $s$, and when receiving $(q,i)$, it forwards $q$ to the $i$-th created mechanism and returns its response. If less than $i$ mechanisms have been created, ExtConComp halts the communication. The state of ExtConComp is a sequence of pairs $(\mathcal{M}_{i},s_{i}^{\prime})$, where the $i$-th pair represents the $i$-th created mechanism $\mathcal{M}_{i}$ and its current state $s_{i}^{\prime}$. The single initial state space of ExtConComp is the empty sequence ()()( ). Both query set and answer set of ExtConComp are the universal message set $\mathcal{U}$ in Remark 3.4.
To define DP for the extended concurrent composition, we must impose constraints on the created mechanisms and their query pairs to limit the adversary’s power. We force a set of constraints using a verifier $\mathcal{V}[f]$, where $f$ is a verification function returning $\bot$ whenever the adversary misbehaves. We define variants of the concurrent composition of CMs by choosing different verification functions $f$:
Definition 4.2.
For any verification function $f$, the $f$-concurrent composition of CMs is the interactive mechanism $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$.
Differential privacy for the $f$-concurrent composition of CMs follows the standard definition for IMs:
Definition 4.3 (DP for $f$-Concurrent Composition of CMs).
For a verification function $f$ and privacy parameter $\epsilon\geq 0$ and $0\leq\delta\leq 1$, the $f$-concurrent composition of CMs is $(\epsilon,\delta)$-DP against adaptive adversaries if the interactive mechanism $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-DP w.r.t. $\sim_{01}$. Equivalently, this holds if the continual mechanism ExtConComp is $(\epsilon,\delta)$-DP w.r.t. $f$ against adaptive adversaries.
Although verification functions $f$ for ExtConComp vary for different compositions, they must all ensure that the format of the adversary’s messages is compatible with $\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$. An adversary is expected to either create a new CM or send a pair of queries to an existing one. The creation of a new mechanism is independent of the initial state of $\mathcal{I}$. Thus, when requesting the creation of a new mechanism, the adversary must send a pair of identical creation queries to $\mathcal{V}[f]$. Moreover, due to the definition of $\mathcal{I}$ and ExtConComp, to request asking one of the queries $q_{0}$ and $q_{1}$ from the $i$-th created mechanism according to $\mathcal{I}$’s initial state, the adversary must send the message $\left((q_{0},i),(q_{1},i)\right)$ to $\mathcal{V}[f]$.
Next, we formalize these shared constraints by defining suitable message sequences. Later, when defining specific verification functions $f$ for different concurrent compositions, we set $f(m_{1},\dots,m_{k})=\bot$ whenever the sequence of messages $m_{1},\dots,m_{k}$ is not suitable.
Definition 4.4 (Suitable Sequence of Messages).
A sequence of messages $m_{1},\dots,m_{k}$ is suitable if for every $1\leq j\leq k$, (i) $m_{j}$ is of the form $(m_{j}^{0},m_{j}^{1})$, where $m_{j}^{0}$ and $m_{j}^{1}$ are either identical creation queries or messages of the forms $(q_{0},\ell)$ and $(q_{1},\ell)$ with the same index $\ell$; (ii) for each $b\in\{0,1\}$, if $m_{j}^{b}=(q,\ell)$, then there are at least $\ell$ creation queries in the sequence $m_{1}^{b},\dots,m_{j-1}^{b}$; (iii) for each $b\in\{0,1\}$, if $m_{j}^{b}=(q,\ell)$ and $\mathcal{M}_{\ell}$ is the $\ell$-th mechanism requested to be created in $m_{1}^{b},\dots,m_{j-1}^{b}$, then $q$ is in the query space of $\mathcal{M}_{\ell}$, i.e., $q\in Q_{\mathcal{M}_{\ell}}$; and (iv) if $m_{j}^{0}=m_{j}^{1}=(\mathcal{M},s,\epsilon,\delta,f)$, then $\mathcal{M}$ is a CM with a single initial state $s$, and it is $(\epsilon,\delta)$-DP w.r.t. $f$.
4.2. $(\epsilon,\delta)$-DP IMs and $\mathrm{RR}_{\epsilon,\delta}$
Lyu [18] shows that for every $(\epsilon,\delta)$-DP IM $\mathcal{M}$ and every pair of neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{T}$ (with a single initial state) such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(s_{b})$ are identical. Vadhan and Wang [20] show the same result in the special case where $\delta=0$.
Although not explicitly stated, the constructions of $\mathcal{T}$ proposed in both [18] and [20] require the answer distributions of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to be discrete. This assumption is formally defined as:
Definition 4.5.
An IM $\mathcal{M}$ has discrete answer distributions if $A_{\mathcal{M}}$ is countable, and for every initial state $s$, integer $t\in\mathbb{N}$, history $(q_{1},a_{1},\dots,q_{t-1},a_{t-1})\in(Q_{\mathcal{M}}\times A_{\mathcal{M}})^{t-1}$, and new query $q_{t}\in Q_{\mathcal{M}}$, the distribution of $\mathcal{M}(s)$’s answer in respond to $q_{t}$ conditioned on having the history $(q_{1},\dots,a_{t-1})$ has a probability mass function (PMF).
This is generally not a restrictive assumption in differential privacy. In practice, the binary nature of computers and the fact that the total number of memory bits in the world is finite ensure that every distribution sampled computationally is inherently discrete. Theoretically, when dealing with real-valued outputs, one can enforce discrete answer distributions by rounding each output to the nearest multiple of a sufficiently small constant $\alpha$. This discretization post-processing preserves the privacy of the mechanism and has a negligible effect on the accuracy. An example of this approach will be provided in Section 7.
Lyu [18] additionally assumes that there exists a pre-fixed upper bound $c$ on the number of queries $\mathcal{M}$ answers. In Appendix A, we will modify their proof to remove this additional assumption. Let $\mathcal{M}^{\prime}$ be a continual mechanism that is $(\epsilon,\delta)$-DP w.r.t. a verification function $f$. By Corollary 3.8, $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}^{\prime}$ is an IM that is $(\epsilon,\delta)$-DP w.r.t. $\sim_{01}$. Replacing $\mathcal{M}$ in Lemma A.1 with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}^{\prime}$, we get:
Lemma 4.6 ([18]).
For $\epsilon>0$ and $0\leq\delta\leq 1$, let $\mathcal{M}$ be a CM that is $(\epsilon,\delta)$-DP w.r.t. a verification function $f$. Suppose $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ has discrete answer distributions and responds to at most $c$ queries. Then, there exists an IPM $\mathcal{T}$ with a single initial state ()()( ) such that for each $b\in\{0,1\}$, the IMs $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$ are equivalent.
Let $\mathcal{M}$ be an IM that is $\epsilon$-DP w.r.t. a neighbor relation $\sim$. For every pair of neighboring initial states $s_{0}$ and $s_{1}$, Vadhan and Wang [20] construct an IPM $\mathcal{T}$ (with a single initial state) such that for each bit $b\in\{0,1\}$, the mechanisms $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,0}(b)$ and $\mathcal{M}(s_{b})$ are identical. Appendix B, we will review their construction and show that for any query sequence $q_{1},\dots,q_{t}$, if the distributions of the $t$ answers of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to the queries $q_{1},\dots,q_{t}$ are identical, then $\mathcal{T}$ does not need to know the output of $\mathrm{RR}_{\epsilon,0}(b)$ in the first $t$ steps. This observation allows us to delay the interaction of $\mathcal{T}$ and $\mathrm{RR}_{\epsilon,0}(b)$ as much as possible.
Let $\mathcal{M}=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}^{\prime}$, where $\mathcal{M}^{\prime}$ is an $\epsilon$-DP CM w.r.t. $f$. By definition, upon receiving a query sequence $q_{1},\dots,q_{t}$ where each $q_{j}=(q_{j}^{\prime},q_{j}^{\prime})$ is a pair of identical queries, $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}^{\prime}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M}^{\prime}$ have the same answer distribution, implying the following lemma:
Lemma 4.7 ([20]).
For $\epsilon\geq 0$, let $\mathcal{M}$ be a CM that is $\epsilon$-DP w.r.t. a verification function $f$. Suppose $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ has discrete answer distributions. Then, there exists an IPM $\mathcal{T}$ with a single initial state ()()( ) such that for each $b\in\{0,1\}$,
  • the IMs $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,0}(b)$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$ are equivalent, and
  • as long as $\mathcal{T}$ receives pairs of identical queries, it does not interact with $\mathrm{RR}_{\epsilon,0}(b)$.
Remark 4.8.
Since $\mathrm{RR}_{\epsilon,\delta}$ is a NIM, in both Lemma 4.6 and Lemma 4.7, the IPM $\mathcal{T}$ interacts with $\mathrm{RR}_{\epsilon,\delta}$ at most once.
4.3. Concurrent Composition of CMs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$
For $1\leq i\leq k$, let $\mathcal{M}_{i}$ be a CM with the initial state space $S_{i}=\{s_{i}\}$ such that $\mathcal{M}_{i}$ is $(\epsilon_{i},\delta_{i})$-DP w.r.t. a verification function $f_{i}$, and $\mathcal{V}[f_{i}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}_{i}$ has discrete answer distributions. In the concurrent composition of $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$, an adversary is restricted to the following steps: (1) It first creates the fixed mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ and (2) then adaptively issues pairs of queries for them such that the sequences of query pairs for each $\mathcal{M}_{i}$ is $f_{i}$-valid. Note that creating the CM $\mathcal{M}_{i}$ means creating an instance of $\mathcal{M}_{i}(s_{i})$, where $s_{i}$ is the unique initial state of $\mathcal{M}_{i}$. Based on a secret bit $b\in\{0,1\}$, each CM either receives the first element of every query pair addressed to it or the second element, and the adversary tries to guess $b$ using the answers of all $k$ mechanisms.
To formalize this type of concurrent composition of CMs, we introduce the verification function $f^{\alpha_{1},\dots,\alpha_{k}}$. Intuitively, this verification function enforces that the adversary creates the mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ in their first $k$ messages and then only sends messages containing pairs of queries for $\mathcal{M}_{i}$s. Additionally, it ensures that the sequence of query pairs for each $\mathcal{M}_{i}$ is $f_{i}$-valid. The formal definition of $f^{\alpha_{1},\dots,\alpha_{k}}$ is as follows:
Definition 4.9 ($f^{\alpha_{1},\dots,\alpha_{k}}$).
For $i\in[k]$, let $\alpha_{i}=(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})$ denote a creation query. For any input sequence of messages $m_{1},\dots,m_{t}$, the outcome of $f^{\alpha_{1},\dots,\alpha_{k}}$ returns $\top$ if and only if all of the following conditions hold:
  1. (i)
    The sequence $m_{1},\dots,m_{k}$ is suitable. For $j\in[t]$, let $m_{j}=(m_{j}^{0},m_{j}^{1})$.
  2. (ii)
    For $j\leq\min\{t,k\}$, the messages $m_{j}^{0}=m_{j}^{1}=\alpha_{j}$.
  3. (iii)
    For $i\in[k]$, the sequence of $(q^{0},q^{1})$s in the messages of the form $\left((q^{0},i),(q^{1},i)\right)$ is $f_{i}$-valid.
With the above definition of $f^{\alpha_{1},\dots,\alpha_{k}}$, we refer to the $f^{\alpha_{1},\dots,\alpha_{k}}$-concurrent composition of CMs as the concurrent composition of the continual mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$. Therefore, the concurrent composition of $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ is $(\epsilon,\delta)$-DP against adaptive adversaries if ExtConComp is $(\epsilon,\delta)$-DP w.r.t. $f^{\alpha_{1},\dots,\alpha_{k}}$ against adaptive adversaries.
Recall that in Section 3, we generalized the notion of neighboring relations for interactive mechanisms to verification functions for continual mechanisms. The neighboring initial states for an interactive mechanism are chosen non-adaptively at initialization. However, the choice of messages forming a valid sequence w.r.t. a verification function is adaptive and happens over time. That is, an adversary generates two different query sequences for a CM, and the choice of differing sequences is made adaptively throughout the interaction.
In the concurrent composition of IMs, queries to each mechanism are chosen adaptively, not only based on that mechanism’s previous responses but also using the responses of the other mechanisms up to that point. By definition, the choice of neighboring initial states for the interactive mechanisms is non-adaptive. However, in the concurrent composition of CMs, the choice of differences between query sequences, i.e., the neighboring query sequences is adaptive, based on not only the responses of the corresponding mechanism but also the responses of other mechanisms. (Note that, unlike IMs, CMs inherently have a single fixed initial state; hence, no choice of initial state is involved.) Despite this additional adaptivity, in the following theorem, we show that the privacy results of [18] for ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ also hold for $\mathcal{V}[f^{\alpha_{1},\dots,\alpha_{k}}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$.
Theorem 4.10.
For $1\leq i\leq k$, let $\mathcal{M}_{i}$ be a continual mechanism with initial state space $S_{\mathcal{M}_{i}}^{\text{init}}=\{s_{i}\}$ such that $\mathcal{M}_{i}$ is $(\epsilon_{i},\delta_{i})$-DP w.r.t. a verification function $f_{i}$ against adaptive adversaries, and $\mathcal{V}[f_{i}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}_{i}$ has discrete answer distributions. For $\epsilon>0$ and $0\leq\delta\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the concurrent composition of $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ is also $(\epsilon,\delta)$-DP against adaptive adversaries.
Proof.
Let $\alpha_{i}=(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})$ for each $i\in[k]$, and let $f$ denote $f^{\alpha_{1},\dots,\alpha_{k}}$ for simplicity. Our goal is to show that ExtConComp is $(\epsilon,\delta)$-DP w.r.t. $f$. To prove so, we construct an IPM $\mathcal{P}$ with a single initial state and show that for each $b\in\{0,1\}$, the mechanisms $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((b)_{i=1}^{k}\right)$ are equivalent.
Since $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}$ is an IPM with a single initial state, Lemma 2.15 implies that if ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$ is $(\epsilon,\delta)$-DP, then $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$ is also $(\epsilon,\delta)$-DP. Hence, for any adversary $\mathcal{A}$, the views of $\mathcal{A}$ when interacting with the mechanisms $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((0)_{i=1}^{k}\right)$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((1)_{i=1}^{k}\right)$ are $(\epsilon,\delta)$-indistinguishable. As we will show below, by construction, these views are identical to those obtained by interacting with the original mechanisms $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, implying that the concurrent composition of $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ is $(\epsilon,\delta)$-DP.
In $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, when $\mathcal{I}(b)$ receives pairs of identical creation queries from $\mathcal{V}[f]$, it forwards one of them to ExtConComp, which then creates an instance of $\mathcal{M}_{i}$ and returns $\top$. Subsequently, when $\mathcal{I}(b)$ receives two queries for $\mathcal{M}_{i}$, it forwards them to $\mathcal{M}_{i}$ through ExtConComp and returns its response. This procedure can be equivalently represented by an IM ${\textrm{IECC}}(b)$ as follows: when receiving pairs of identical creation queries from $\mathcal{V}[f]$, ${\textrm{IECC}}(b)$ creates an instance of $\mathcal{V}[f_{i}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{i}$ and returns $\top$. Subsequently, when receiving $\left((q_{0},i),(q_{1},i)\right)$, it forwards $(q_{0},q_{1})$ to $\mathcal{V}[f_{i}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{i}$ and returns its response. By definition of $f$ and $\mathcal{I}$, the mechanisms $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!{\textrm{IECC}}(b)$ are equivalent. Next, we define $\mathcal{P}$ explicitly and show $\mathcal{V}[f]\!\circ^{*}\!{\textrm{IECC}}(b)$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((b)_{i=1}^{k}\right)$ are equivalent.
Let $\mathcal{T}_{i}$ be the IPM for $\mathcal{M}_{i}$ given by Lemma 4.6, and let $t_{i}$ be its single initial state. The initial state of $\mathcal{P}$ is a pair of empty sequences. Let $(\sigma_{1},\sigma_{2})$ denote the state of $\mathcal{P}$. When receiving the first left message, $\mathcal{P}$ interacts with its right mechanism, ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((b)_{i=1}^{k}\right)$, and sets $\sigma_{1}$ to its output sequence $(r_{1},\dots,r_{k})$. $\mathcal{P}$ never interacts with its right mechanism again.
Upon receiving a left message $(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})^{2}$, $\mathcal{P}$ adds $(\mathcal{T}_{i},t_{i})$ to $\sigma_{2}$ and returns $\top$. Moreover, upon receiving a left message $\left((q_{0},i),(q_{1},i)\right)$, $\mathcal{P}$ sends $(q_{0},q_{1})$ as a left message to $\mathcal{T}_{i}$ and returns its response. If $\mathcal{T}_{i}$ requests a sample from $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$, $\mathcal{P}$ provides it with the previously stored sample $r_{i}$. Formally, $\mathcal{P}$ executes $(t_{i},v,m)\leftarrow\mathcal{T}_{i}\left(t_{i},{\texttt{L}},(q_{0},q_{1})\right)$ and as long as $v={\texttt{R}}$, it runs $(t_{i},v,m)\leftarrow\mathcal{T}_{i}\left(t_{i},v,r_{i}\right)$. Once $v={\texttt{L}}$, $\mathcal{P}$ returns the message $m$ to $\mathcal{V}[f]$. (By Remark 4.8, $v={\texttt{R}}$ at most once.) By Lemma 4.6, each $\mathcal{T}_{i}$, provided with the sample $r_{i}$ from $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$, simulates $\mathcal{V}[f_{i}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{i}$ identically. Hence, by definition of IECC and $\mathcal{P}$, the mechanisms ${\textrm{IECC}}(b)$ and $\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((b)_{i=1}^{k}\right)$ are equivalent, finishing the proof. ∎
4.4. Concurrent Composition of CMs with Fixed Parameters
The concurrent composition of CMs with fixed parameters generalizes the concurrent composition of fixed CMs by allowing the adversary to select the CMs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ adaptively over time, instead of forcing them to create a fixed set of mechanisms in their first $k$ messages. Each selected mechanism $\mathcal{M}_{i}$ must be $(\epsilon_{i},\delta_{i})$-DP w.r.t. a verification function $f_{i}$, where the multiset of privacy parameters $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ is fixed in advance, but the verification functions $f_{i}$ are chosen by the adversary. Note that this implies that at most $k$ continual mechanism are created by the adversary. Moreover, due to the aforementioned technical reasons, each $\mathcal{V}[f_{i}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}_{i}$ must have discrete answer distributions.
For instance, an adversary may first create a continual mechanism $\mathcal{M}_{3}$ that is $(\epsilon_{3},\delta_{3})$-DP w.r.t. a verification function $f_{3}$. Then, after interacting with $\mathcal{M}_{3}$ for some time and analyzing its responses, the adversary may then select a mechanism $\mathcal{M}_{1}$ from the class of $(\epsilon_{1},\delta_{1})$-DP CMs. Similar to the concurrent composition of fixed CMs, the adversary’s sequence of query pairs for each $\mathcal{M}_{i}$ must be $f_{i}$-valid.
To formally define this composition, we introduce a verification function, $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$ as follows:
Definition 4.11 ($f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$).
For any input sequence of messages $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$ returns $\top$ if and only if all of the following conditions hold:
  1. (i)
    The sequence $m_{1},\dots,m_{t}$ is suitable.
  2. (ii)
    There are at most $k$ pairs of identical creation queries in the sequence $m_{1},\dots,m_{t}$. Let $\mathcal{M}_{i}^{\prime}$ denote the $i$-th created mechanism and $(\epsilon_{i}^{\prime},\delta_{i}^{\prime})$ and $f_{i}^{\prime}$ be its privacy parameters and verification function, respectively. $\mathcal{V}[f_{i}^{\prime}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}_{i}^{\prime}$ has discrete answer distributions. Moreover, the multiset of the privacy parameters $(\epsilon_{i}^{\prime},\delta_{i}^{\prime})$ of the created mechanisms is a sub-multiset of the fixed privacy parameters $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$. Note that $(\epsilon_{i}^{\prime},\delta_{i}^{\prime})$ is adaptively chosen by the adversary, while $(\epsilon_{i},\delta_{i})$ is a predetermined publicly-known parameter.
  3. (iii)
    For each $i\in[k]$, the sequence of $(q^{0},q^{1})$s in the messages of the form $\left((q^{0},i),(q^{1},i)\right)$ is $f_{i}^{\prime}$-valid.
We refer to the $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$-concurrent composition of CMs as the concurrent composition of CMs with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$. A slight adaptation of the proof of Theorem 4.10 implies the following theorem.
Theorem 4.12.
For $\epsilon,\epsilon_{1},\dots,\epsilon_{k}>0$ and $0\leq\delta,\delta_{1},\dots,\delta_{k}\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the concurrent composition of $k$ CMs with the privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$ is also $(\epsilon,\delta)$-DP against adaptive adversaries.
Proof.
The proof is identical to the proof of Theorem 4.10 with the following modification in the construction of $\mathcal{P}$: Recall that $(r_{1},\dots,r_{k})$ denotes the output of ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$. In the proof of Theorem 4.10, the $i$-th creation query was known to be $(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})$. Consequently, $\mathcal{T}_{i}$ expected $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$ as its right mechanism, and when $\mathcal{T}_{i}$ needed a response from its right mechanism, $\mathcal{P}$ sent $r_{i}$ to it. However, in the concurrent composition with fixed parameters, the IPM requesting the sample from $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$ need not necessarily be $\mathcal{T}_{i}$. To handle this, we modify $\mathcal{P}$ so that it maps each output sample $r_{i}$ to the appropriate IPM $\mathcal{T}_{j}$ needing a sample from $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$.
Let $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$ denote the $j$-th pair of identical creation queries received by $\mathcal{P}$ from the verifier, and let $\mathcal{T}_{j}$ be the corresponding IPM given by Lemma 4.6. When $\mathcal{T}_{j}$ requests a sample from $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, $\mathcal{P}$ iterates over $\sigma_{1}$ to find the smallest $i^{*}\in[k]$ such that the $i^{*}$-th element of $\sigma_{1}$ is not null and $\epsilon_{j}^{\prime}=\epsilon_{i^{*}}$, $\delta_{j}^{\prime}=\delta_{i^{*}}$. After sending $r_{i^{*}}$ to $\mathcal{T}_{j}$, $\mathcal{P}$ updates $\sigma_{1}$ and sets $r_{i}$ to null. This adjustment ensures that even when there are duplicate privacy parameters among $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$, each sample $r_{i}$ is sent to at most one $\mathcal{T}_{j}$. Note that by the definition of $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$, $i^{*}$ always exists. ∎
5. Concurrent Parallel Composition of CMs
The concurrent $k$-sparse parallel composition extends the concurrent composition of CMs with $k$ fixed parameters by permitting the creation of an arbitrary number of CMs. Specifically, given a multiset params of $k$ privacy parameters $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$, the adversary can create an unlimited number of CMs with arbitrary privacy parameters. However, the adversary must ask pairs of identical queries to all created mechanisms, except for at most $k$ of them. Let $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ be these exceptional mechanisms. Each $\mathcal{M}_{i}$, selected by the adversary, must be $(\epsilon_{i},\delta_{i})$-DP w.r.t. a verification function $f_{i}$ and with $(\epsilon_{i},\delta_{i})\in{\texttt{params}}$, and the sequence of query pairs for $\mathcal{M}_{i}$ must be $f_{i}$-valid.
In Section 5.1, we formally define this composition and extend the result of Theorem 4.12 to the concurrent $k$-sparse parallel composition of CMs when $\delta_{1}=\dots=\delta_{k}=0$. In Section 5.2, we provide a counter example that demonstrates why such an extension is impossible for $\delta_{j}>0$.
In Section 5.3, we restrict the verification functions $f_{i}$ of the $k$ mechanisms receiving neighboring input and we show that Theorem 4.12 holds for the concurrent $k$-sparse parallel composition of CM in this scenario without any constraints on the values of $\delta_{j}$.
5.1. Definition and Composition Theorem for Pure DP
To formally define the concurrent $k$-sparse parallel composition of CMs with parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$, we introduce the verification function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$, where the subscript $\infty$ emphasizes that an unbounded number of CMs can be created.
Definition 5.1 ($f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$).
For any input sequence of messages $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$ returns $\top$ if and only if all of the following conditions hold:
  1. (i)
    The sequence $m_{1},\dots,m_{t}$ is suitable.
  2. (ii)
    Let $\mathcal{M}_{i}^{\prime}$ denote the $i$-th created mechanism and $(\epsilon_{i}^{\prime},\delta_{i}^{\prime})$ and $f_{i}^{\prime}$ be its privacy parameters and verification function, respectively. Define $\mathcal{J}$ as the set of indices $j$ such that there exists a message $\left((q_{0},j),(q_{1},j)\right)$ where $q_{0}\neq q_{1}$. Let $\texttt{params}_{\mathcal{J}}$ denote the multiset consisting of privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ for each $j\in\mathcal{J}$. This condition requires that $\texttt{params}_{\mathcal{J}}$ be a sub-multiset of the fixed privacy parameters $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$. Note that this implies that $|\mathcal{J}|\leq k.$
  3. (iii)
    For every $j\in\mathcal{J}$, the sequence of $(q_{0},q_{1})$s in the messages of the form $\left((q_{0},j),(q_{1},j)\right)$ must be $f_{i}$-valid. Note that by definition, for every index $j\notin\mathcal{J}$ and every message of the form $\left((q_{0},j),(q_{1},j)\right)$, it holds that $q_{0}=q_{1}$.
We refer to the $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$-concurrent composition of CMs as the concurrent $k$-sparse parallel composition of CMs with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$. Therefore, this composition is $(\epsilon,\delta)$-DP against adaptive adversaries if ExtConComp is $(\epsilon,\delta)$-DP w.r.t. $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$ against adaptive adversaries.
Theorem 5.2.
For $\epsilon,\epsilon_{1},\dots,\epsilon_{k}\geq 0$ and $0\leq\delta\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ is $(\epsilon,\delta)$-DP, then the concurrent $k$-sparse parallel composition of CMs with the privacy parameters $(\epsilon_{1},0),\dots,(\epsilon_{k},0)$ is also $(\epsilon,\delta)$-DP against an adaptive adversary.
Proof.
The proof is identical to the proof of Theorem 4.12, except for the definition of IPMs $\mathcal{T}_{j}$ that are executed by the IPM $\mathcal{P}$. Let $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$ denote the $j$-th pair of identical creation queries received by $\mathcal{P}$ from $\mathcal{V}[f]$. To ensure that the proof of Theorem 4.12 remains valid, we must guarantee that the newly defined IPMs $\mathcal{T}_{j}$ satisfy the following properties needed by the proof of Theorem 4.12. These are the only properties that this proof requires for the $\mathcal{T}_{j}$s:
  • The mechanisms $\mathcal{T}_{j}\!\circ^{*}\!\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ and $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ are equivalent, and
  • the multiset of privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ corresponding to IMPs $\mathcal{T}_{j}$ that are interacting with their right mechanism $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ is a sub-multiset of $(\epsilon_{1},0),\dots,(\epsilon_{k},0)$, implying that the remaining $\mathcal{T}_{j}$ never interact with a $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$.
The new definition of $\mathcal{T}_{j}$ is as follows:
  • If $\delta_{j}^{\prime}>0$, then $\mathcal{T}_{j}$ acts identical to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$: $\mathcal{T}_{j}$ has a single initial state $\left((),0,s^{\text{init}}\right)$, where $s^{\text{init}}$ is the single initial state of $\mathcal{M}_{j}^{\prime}$. Upon receiving a right message, $\mathcal{T}_{j}$ halts the communication, and upon receiving a left message, it updates its state and responds identically to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$.
  • If $\delta_{j}^{\prime}=0$, then $\mathcal{T}_{j}$ is the IPM corresponding to $\mathcal{M}_{j}^{\prime}$ given by Lemma 4.7.
By Definition 5.1 (ii), if $\delta_{j}^{\prime}>0$, then all future messages $\left((q_{0},j),(q_{1},j)\right)$ satisfy $q_{0}=q_{1}$. Therefore, by the above definition, for each $b\in\{0,1\}$, the mechanisms $\mathcal{T}_{j}\!\circ^{*}\!\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ and $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ are identical, and $\mathcal{T}_{j}$ never interacts with $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$.
Moreover, when $\delta_{j}^{\prime}=0$, by Definition 5.1 (iii), Lemma 4.7, and Remark 4.8, for each $b\in\{0,1\}$, the mechanisms $\mathcal{T}_{j}\!\circ^{*}\!\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ and $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ are identical, and $\mathcal{T}_{j}$ interacts with $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ at most once after receiving a pair of non-identical queries.
In the concurrent $k$-sparse parallel composition, an unlimited number of CMs can be created, but only at most $k$ of them (here, with $\delta_{j}^{\prime}=0$) can receive non-identical query pairs. By Definition 5.1 (ii), the privacy parameters of these mechanisms is a sub-multiset of $(\epsilon_{1},0),\dots,(\epsilon_{k},0)$. By design, the IPMs $\mathcal{T}_{j}$ corresponding to these mechanisms may require a sample from $\mathrm{RR}_{\epsilon_{j}^{\prime},0}$. ∎
5.2. Counter Example.
In this section, we illustrate that extending the result of Theorem 5.2 to general $\delta_{j}>0$ is impossible. Specifically, let $k=1$, $\epsilon_{1}=0$, and $\delta_{1}=\delta\in(0,1]$. We construct, an adversary $\mathcal{A}_{\delta}$ that interacts with $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ and successfully determines the secret bit $b\in\{0,1\}$. Our construction proceeds in three steps: (1) introducing a verification function $g$, (2) constructing a CM $\mathcal{M}_{\delta}$ that is $(0,\delta)$-DP w.r.t. $g$, and (3) designing $\mathcal{A}_{\delta}$.
Step 1 (Introducing $g$): The verification function $g$ accepts a message sequence $(m_{1},\dots,m_{t})$ and returns $\top$ if and only if (i) $t\leq 2$, (ii) each $m_{j}$ is a pair of binary numbers $(b_{j},b_{j}^{\prime})\in\{0,1\}^{2}$, and (ii) the sequences $(b_{1},\dots,b_{t})$ and $(b_{1}^{\prime},\dots,b_{t}^{\prime})$ differ in at most a single element.
[…]
  • $\mathcal{M}_{\delta}$ maps state $\top$ and an input message $b\in\{0,1\}$ to state $\top$ and output message $\top$ with probability $1-\delta$ and to state $\bot$ and output message $\bot$ with probability $\delta$.
  • $\mathcal{M}_{\delta}$ maps state $\bot$ and an input message $b\in\{0,1\}$ to state $\bot$ and output message $b$ with probability $1$.
Thus, $\mathcal{M}_{\delta}$ has two possible states, $\top$ and $\bot$, and accepts binary queries. Initially, $\mathcal{M}_{\delta}$ starts in state $\top$. On the first query, independent of the value of $b$, it switches to state $\bot$ with probability $\delta$ and returns its final state as an output message. On the second query, if the state is $\top$, the same as the first query, it switches to state $\bot$ with probability $\delta$ and returns the final state as output. However, if the state is $\bot$, $\mathcal{M}_{\delta}$ outputs the exact input message $b$.
Since with probability $1-\delta$, the first two outputs of $\mathcal{M}_{\delta}$ are independent of the input, $\mathcal{M}_{\delta}$ is $(0,\delta)$-DP w.r.t. the verification function $g$.
Step 3 (Designing $\mathcal{A}_{\delta}$): In the interaction with $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, the adversary $\mathcal{A}_{\delta}$ must determine the secret bit $b$. $\mathcal{A}_{\delta}$ accepts messages in $\{0,1,\top,\bot\}$ and sends the following messages: (a) the message $(\mathcal{M}_{\delta},\top,0,\delta,g)^{2}$, which is a pair of identical creation queries requesting the creation of $\mathcal{M}_{\delta}$, and (b) messages of the form $\left((b_{0},j),(b_{1},j)\right)$, where $b_{0},b_{1}\in\{0,1\}$ and $j\in\mathbb{N}$.
$\mathcal{A}_{\delta}$ starts the communication by sending $(\mathcal{M}_{\delta},\top,0,\delta,g)^{2}$ and setting its state $s=1$. While receiving $\top$ as a response, it increments $s$ and alternates its queries:
  • If $s$ is odd, it sends $(\mathcal{M}_{\delta},\top,0,\delta,g)^{2}$.
  • If $s$ is even, it sends $\left((0,s/2),(0,s/2)\right)$.
In this iterative procedure, $\mathcal{A}_{\delta}$ keeps creating a new instance of $\mathcal{M}_{\delta}$ and asking the query pair $(0,0)$ from it until one instance returns $\bot$. Since ExtConComp always returns $\top$ when creating a mechanism, this loop stops upon receiving a respond to $\left((0,s/2),(0,s/2)\right)$.
Once $\mathcal{A}_{\delta}$ receives $\bot$, without increasing $s$, it sends $\left((0,s/2),(1,s/2)\right)$. That is, $\mathcal{A}_{\delta}$ asks the pair $(0,1)$ from the last created instance of $\mathcal{M}_{\delta}$, which has returned $\bot$. If the initial state of the identifier $\mathcal{I}$ in $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $b$, then $\mathcal{M}_{\delta}$ receives the query $b$, and by definition it will return $b$ as output, revealing the secret bit $b$.
Therefore, with probability $1-(1-\delta)^{t}$, the adversary $\mathcal{A}_{\delta}$ can detect the secret bit $b$ by sending at most $12t+$ 1 queries. Since $\delta>0$, as $t$ goes to infinity, this probability tends to $1$.
5.3. Composition Theorem for Restricted Verification Functions.
In Section 5.2, we showed that for any positive $\delta$, the concurrent $1$-sparse parallel composition of CMs with the privacy parameter $(0,\delta)$ fails to preserve privacy. In our counterexample, the adversary leveraged the first outcome of a $(0,\delta)$-DP CM to decide whether to spend its budget on issuing non-identical queries to that mechanism. Intuitively, when $\delta>0$, there might exist certain “bad first answers” that occur with low probability; however, once we condition on their occurrences, the distribution of the second answer ceases to be differentially private.
In this section, we force the adversary to decide whether it will always issue pairs of identical queries to a CM when it sends the first message to that mechanism. Specifically, if the first message sent to a CM includes identical queries, then all subsequent messages for that CM must also have identical queries. This leads to the following definition.
Definition 5.3 (First-Pair Consistent).
A verification function $f$ is first-pair consistent if the following holds: given an input message sequence $m_{1},\dots,m_{t}$, $f$ returns $\bot$ if $m_{1}$ is a pair of identical messages, i.e., $m_{1}=(m^{\prime},m^{\prime})$ for some $m^{\prime}$, while there exists $2\leq j\leq t$ such that $m_{j}$ is not a pair of identical messages.
For instance, let $\mathcal{M}$ be an IM that is $(\epsilon,\delta)$-DP w.r.t. a neighbor relation $\sim$. As discussed in Section 3.1, $\mathcal{M}$ can be represented by a CM $\mathcal{M}^{\prime}$ (with initial state null) that receives an initial state $s\in S_{\mathcal{M}}^{\text{init}}$ as the first message and behaves the same as $\mathcal{M}(s)$ afterwards. Define the verification function $f_{\sim}$ as follows: given an input message sequence $(m_{1},\dots,m_{t})$, $f_{\sim}$ returns $\top$ if and only if $m_{1}$ is a pair of neighboring initial states in $S_{\mathcal{M}}^{\text{init}}$, and for all $2\leq j\leq t$, $m_{j}$ is a pair of identical queries in $Q_{\mathcal{M}}$. By definition, $f_{\sim}$ is first-pair consistent.
To formally define the restricted variant of the concurrent parallel composition of CMs, we introduce the verification function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$, where $FPC$ stands for first-pair consistency.
Definition 5.4 ($f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$).
For any input message sequence $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$ returns $\top$ if and only if both of the following conditions hold:
  1. (i)
    $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}(m_{1},\dots,m_{t})=\top$
  2. (ii)
    For every message $m_{j}$ of the form $(\mathcal{M}^{\prime},s^{\prime},\epsilon^{\prime},\delta^{\prime},f^{\prime})^{2}$, the verification function $f^{\prime}$ is first-pair consistent.
We refer to the $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$-concurrent composition of CMs as the FPC concurrent $k$-sparse parallel composition of CMs with privacy parameters $\left((\epsilon_{i},\delta_{i}\right)_{i=1}^{k}$.
Theorem 5.5.
For $\epsilon,\epsilon_{1},\dots,\epsilon_{k}\geq 0$ and $0\leq\delta,\delta_{1},\dots,\delta_{k}\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the FPC concurrent $k$-sparse parallel composition of CMs with the privacy parameters $(\epsilon_{1},0),\dots,(\epsilon_{k},0)$ is also $(\epsilon,\delta)$-DP against an adaptive adversary.
Proof.
We denote $f$ as shorthand for $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$. This proof is identical to the proof of Theorem 5.2, except for the following modification in the definition of $\mathcal{P}$: In the proof of Theorem 5.2, each $\mathcal{T}_{j}$ behaved identically to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ if $\delta_{j}^{\prime}>0$ and was defined according to Lemma 4.7 otherwise. In this proof, each $\mathcal{T}_{j}$ behaves identically to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ if the first message of the form $\left((q_{0},j),(q_{1},j)\right)$ satisfies $q_{0}=q_{1}$, and it is defined by Lemma 4.6 otherwise.
To implement this, $\mathcal{P}$’s state must include an additional sequence $\sigma_{3}$, storing a flag for each created mechanism, indicating whether the mechanism has previously received any messages. When $\mathcal{P}$ receives the creation query pair $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$ from the verifier, it adds the pair $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime})$ to $\sigma_{2}$, appends a 00 to the end of $\sigma_{3}$, and returns $\top$.
Upon receiving a message of the form $\left((q_{0},j),(q_{1},j)\right)$, if the $j$-th element of $\sigma_{3}$ is 00, $\mathcal{P}$ sets this element to $1$ and replaces the entry $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime})$ in $\sigma_{2}$ with $(\mathcal{T}_{j},t_{j})$, defined as follows:
  • If $q_{0}=q_{1}$, then $\mathcal{T}_{j}$ is an IPM with the single initial state $s_{j}^{\prime}$ that never interacts with its right mechanism and interacts with its left mechanism exactly as $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$.
  • If $q_{0}\neq q_{1}$, then $\mathcal{T}_{j}$ is the IPM defined by Lemma 4.6 corresponding to $\mathcal{M}_{j}^{\prime}$.
$\mathcal{T}_{j}$ with the first definition never requires a sample from $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, while a $\mathcal{T}_{j}$ with the second definition may require at most one sample (see Remark 4.8). We will show that the privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ corresponding to $\mathcal{T}_{j}$s with the second definition form a sub-multiset of $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$:
By Definition 5.4 (ii), we know that for each index $j$, if $q_{0}=q_{1}$ in the first message of the form $\left((q_{0},j),(q_{1},j)\right)$, then $q_{0}=q_{1}$ in all subsequent messages of this form. Thus, by Definition 5.4 (i) and Definition 5.1 (ii), the privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ corresponding to $\mathcal{T}_{j}$s with the second definition (which receive at least one pair of non-identical queries) form a sub-multiset of $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$. ∎
5.3.1. Concurrent $k$-sparse Parallel Composition of IM
Consider a setting where an adversary creates (an unlimited number of) IMs over time while adaptively asking queries from the existing ones. When creating an IM $\mathcal{M}_{i}$ that is $(\epsilon_{i},\delta_{i})$-DP w.r.t. a neighbor relation $\sim_{i}$, the adversary also selects a pair of neighboring initial states for $\mathcal{M}_{i}$. This selection depends on the history of queries asked so far and their responses. Based on a secret bit $b$, either all created mechanisms are initialized with the first initial state in their respective pairs, or all are initialized with the second. The goal of the adversary is to guess $b$.
This scenario is an special case of FPC concurrent $k$-sparse parallel composition of CMs, where, instead of creating $\mathcal{M}_{i}$ and assigning a pair of initial states to it, the adversary sends $(\mathcal{M}_{i}^{\prime},\text{null},\epsilon_{i},\delta_{i},f_{\sim_{i}})^{2}$, where $\mathcal{M}_{i}^{\prime}$ is the CM simulating $\mathcal{M}_{i}$. Moreover, to ask a query $q$ from the $i$-th created IM, the adversary sends $\left((q,i),(q,i)\right)$. Since each $f_{\sim_{i}}$ is first-pair consistent, Theorem 5.5 applies to the concurrent $k$-sparse parallel composition of IM with privacy parameters $\left((\epsilon_{i},\delta_{i}\right)_{i=1}^{k}$ (that are differentially private w.r.t. neighbor relations).
NIMs are a special subclass of IMs that halt the communication upon receiving a second input message. Since Theorem 5.5 applies to all IMs, it also holds for the concurrent $k$-sparse parallel composition of NIMs with privacy parameters $\left((\epsilon_{1},\delta_{i}\right)_{i=1}^{k}$. In the next section, we compare this result to the classical parallel composition theorem.
5.3.2. Comparison with the Parallel Composition of NIMs.
Consider a set of $\ell$ NIMs, $\mathcal{N}_{1},\dots,\mathcal{N}_{\ell}$, where each $\mathcal{N}_{i}$ has a domain $\mathcal{X}_{i}$ and is $(\epsilon_{i},\delta_{i})$-DP w.r.t. a neighbor relation $\sim_{i}$. The $k$-sparse parallel composition of NIMs $\mathcal{N}_{1},\dots,\mathcal{N}_{\ell}$ is said to be $(\epsilon,\delta)$-DP if the non-interactive mechanism ${\textrm{Comp}}(\mathcal{N}_{1},\dots,\mathcal{N}_{\ell})$ with the domain $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{\ell}$ is $(\epsilon,\delta)$-DP w.r.t. the neighbor relation $\sim_{\text{par}}$, defined as follows: for any two $x=(x_{1},\dots x_{\ell})$ and $x^{\prime}=(x_{1}^{\prime},\dots x_{\ell}^{\prime})$ in $\mathcal{X}$, we have $x\sim_{\text{par}}x^{\prime}$ if and only if there exists a subset $\mathcal{J}\subseteq\{1,\dots,\ell\}$ of size at most $k$ such that $x_{j}\sim_{j}x_{j}^{\prime}$ for every $j\in\mathcal{J}$, and $x_{j}=x_{j}^{\prime}$ for every $j\in\{1,\dots,\ell\}\setminus\mathcal{J}$.
The concurrent parallel composition of IMs generalizes the parallel composition of NIMs in several key ways: Rather than composing a fixed set of $\ell$ NIMs, an adversary adaptively selects an unbounded number of IMs with arbitrary privacy parameters. Additionally, instead of non-adaptively selecting worst-case neighboring initial states for (at most) $k$ mechanisms and identical initial states for the other $\ell-k$ mechanisms, the adversary adaptively (1) chooses pairs of initial states based on prior interactions with other mechanisms, while (2) ensuring that the initial states in at most $k$ pairs are neighboring, and that they are identical in the other pairs. In the concurrent parallel composition of CMs, the adversary is even stronger: instead of choosing a single pair of neighboring initial states, it adaptively selects a sequence of pairs, satisfying the constraints enforced by a verification function.
The $1$-sparse parallel composition of NIMs is commonly used when a dataset is partitioned among $\ell$ NIMs such that for any two neighboring datasets differing by a single record, all partitions remain identical except for the presence of the differing record in one partition. Consider the case where an adversary creates CMs receiving data or empty records as query messages. In the concurrent parallel composition of CMs, the adversary adaptively determines the number of partitions, assigns records to them, and selects which partitions contain differing sets of records.
It is known that the $k$-sparse parallel composition of $\ell$ $(\epsilon_{0},\delta_{0})$-DP NIMs is $(\epsilon,\delta)$-DP if the composition of $\ell$ instances of $\mathrm{RR}_{\epsilon_{0},\delta_{0}}$ is $(\epsilon,\delta)$-DP. We showed that any NIM that is $(\epsilon_{0},\delta_{0})$-DP w.r.t. a neighbor relation $\sim$ can be modeled by a CM that is $(\epsilon_{0},\delta_{0})$-DP w.r.t. $f_{\sim}$, which is a first-pair consistent verification function. Thus, setting $\epsilon_{1}=\dots=\epsilon_{k}=\epsilon_{0}$, $\delta_{1}=\dots=\delta_{k}=\delta_{0}$ in Theorem 5.5 recovers and generalizes the classic parallel composition theorem. Our theorem extends differential privacy guarantees to settings where an adversary can adaptively create an unbounded number of NIMs.
6. Designing Complex Continual Mechanisms
So far we discussed the setting where the messages are issued by an adversary. However, in the design of differentially private mechanisms in the streaming setting, it is common to use existing differentially private interactive mechanisms, such as the Sparse Vector Technique (SVT) and continual counters, as sub-mechanisms. Thus, it is the main mechanism $\mathcal{M}$ that issues messages to its sub-mechanisms. The main mechanism may fix the set of sub-mechanisms at initialization or select these mechanisms adaptively over time. For example, in Section 7, we present a mechanism that maintains $d$ fixed binary counters while adaptively instantiating an unbounded number of SVT and Laplace mechanisms as required.
The privacy analysis of such mechanisms $\mathcal{M}$ is challenging as the creation of sub-mechanisms and their inputs is adaptive and dependent. The goal of this section is to give a set of conditions for $\mathcal{M}$ and prove a theorem that shows that under these conditions the privacy of $\mathcal{M}$ can directly be computed by applying the known concurrent composition theorems for continual mechanisms to the sub-mechanisms. Specifically, given that the $f^{\prime}$-concurrent composition of CMs is $(\epsilon,\delta)$-DP w.r.t. some verification function $f^{\prime}$, we would like to show that $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. the given verification function $f$.
As discussed in Section 3, these sub-mechanisms can be modeled as CMs. We formalize $\mathcal{M}$, which runs multiple CMs, as a continual mechanism of the form $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$, where $\mathcal{P}$ is an IPM and ExtConComp is the CM defined in Section 4.1. Since $\mathcal{M}$ is a CM, its initial state space is a singleton, implying that the initial state space of $\mathcal{P}$ is also a singleton. Upon receiving a message, $\mathcal{P}$ may create new differentially private CMs or ask queries from the existing ones by communicating with ExtConComp, possibly multiple times.
To prove that $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. $f$, we need to show that the view of every adversary $\mathcal{A}$ interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M}=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable. Intuitively, the messages that $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ receives from $\mathcal{I}(b)$ contains information about $b$. Therefore, to deduce that $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-DP w.r.t. $f$ from the fact that ExtConComp is $(\epsilon,\delta)$-DP w.r.t. $f^{\prime}$, the IPM $\mathcal{P}$ must not directly use the messages from $\mathcal{I}$ to respond. Instead, it should only use these messages to determine which queries it wants to send to its its sub-mechanisms via ExtConComp. More precisely, $\mathcal{P}$ must generate an answer for $\mathcal{I}$ only based on the privatized answers provided by ExtConComp and not dependent on the un-privatized input sequence (Response Property). Moreover, $\mathcal{P}$ must decide whether to continue interacting with ExtConComp or to return a response based on these privatized answers (Destination Property).
Finally, we need a property for $\mathcal{P}$ to convert $f$-valid message sequences to $f^{\prime}$-valid ones: Let $m_{1}^{0},m_{2}^{0},\dots$ and $m_{1}^{1},m_{2}^{1},\dots$ be two sequences such that the sequence $(m_{1}^{0},m_{1}^{1}),(m_{2}^{0},m_{2}^{1}),\dots$ is $f$-valid. Upon receiving each sequence $m_{1}^{b},m_{2}^{b},\dots$, the mechanism $\mathcal{P}$ must send the messages $m_{1}^{b\prime},m_{2}^{b\prime},\dots$ to ExtConComp such that $(m_{1}^{0\prime},m_{1}^{1\prime}),(m_{2}^{0\prime},m_{2}^{1\prime}),\dots$ is $f^{\prime}$-valid ($f\to f^{\prime}$ Property).
In Section 6.1, we formalize the above properties for $\mathcal{P}$, and in Section 6.2, we prove the desired privacy guarantees for $\mathcal{M}$. Combining this result with the concurrent composition theorems provided in the previous sections, we can avoid complicated privacy analysis for many CMs.
6.1. Properties.
To formally state that $\mathcal{P}$ maps $f$-valid sequences to $f^{\prime}$-valid ones, we assume that $\mathcal{P}$ is deterministic. This means $\mathcal{P}$ can be expressed as a deterministic function that maps the sequence of received messages to its next output message. This assumption is not restrictive, as randomness can be modeled by a 00-DP sub-mechanism $\mathcal{M}_{i}$ that accepts a single message $\top$, and upon receiving it, generates a sufficiently long sequence of random bits or samples from a fixed distribution.
Recall that the goal of this section is to show $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-DP w.r.t. $f^{\prime}$ if ExtConComp is $(\epsilon,\delta)$-DP w.r.t. $f$. That is, we need to show that the view of any adversary in the interactions with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable.
The messages received by $\mathcal{P}$ from $\mathcal{I}$ contain information about the bit used by $\mathcal{I}$ to filter the adversary’s pairs. Therefore, to directly use the existing compositions theorems for ExtConComp to analyze the privacy of $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$, we need to ensure that $\mathcal{P}$ only utilizes the messages from $\mathcal{I}$ to generate queries for the existing private sub-mechanisms executed by ExtConComp. Moreover, we need to guarantee that the sequence $(m_{1}^{0\prime},m_{1}^{1\prime}),(m_{2}^{0\prime},m_{2}^{1\prime}),\dots$ is $f^{\prime}$-valid, where $m_{1}^{b\prime},m_{2}^{b\prime},\dots$ is the sequence of messages that $\mathcal{P}$ sends to ExtConComp when it receives messages from $\mathcal{I}(b)\!\circ^{*}\!\mathcal{V}[f]$. The following properties formalize these:
  • Destination Property. The destination L or R of messages sent by $\mathcal{P}$ is determined solely based on messages received from its right mechanism (here, ExtConComp). This ensures that if two transcripts contain identical sequences of messages from the right mechanism but different sequences from the left mechanism, $\mathcal{P}$ will still interact with the same mechanism in the next round.
  • Response Property. The responses sent by $\mathcal{P}$ to the left mechanism depend only on the messages received from the right mechanism. That is, if two transcripts contain identical sequences of messages from the right mechanism and $\mathcal{P}$ sends its next message to the left mechanism, then this message is identical in both cases.
  • $f\to f^{\prime}$ Property. For $v\in\{{\texttt{L}},{\texttt{R}}\}$, let the pair $(m,v)$ denote a message $m$ sent to or received from the left or right mechanism of $\mathcal{P}$. For $b\in\{0,1\}$, let $\sigma^{b}=\left((m_{1}^{b},v_{1}^{b}),\dots,(m_{t}^{b},v_{t}^{b})\right)$ be a sequence of $t$ input messages to $\mathcal{P}$, and let $\tau^{b}=\left((m_{1}^{b\prime},v_{1}^{b\prime}),\dots,(m_{t}^{b\prime},v_{t}^{b\prime})\right)$ be the corresponding sequence of $t$ output messages from $\mathcal{P}$. Suppose the following conditions hold:
    1. (i)
      $v_{i}^{0}=v_{i}^{1}$ for each $i\in[t]$, i.e., the sender of messages is identical;
    2. (ii)
      $m_{i}^{0}=m_{i}^{1}$ for each $i\in[t]$ where $v_{i}^{0}=v_{i}^{1}={\texttt{R}}$, meaning that messages from the right mechanism are identical; and
    3. (iii)
      the sequence of $(m_{i}^{0},m_{i}^{1})$s for all $i\in[t]$ where $v_{i}^{0}=v_{i}^{1}={\texttt{L}}$ is $f$-valid.
    Then, by the destination property of $\mathcal{P}$, $v_{i}^{0\prime}=v_{i}^{1\prime}$ for each $i\in[t]$, and by the response property, $m_{i}^{0\prime}=m_{i}^{1\prime}$ for each $i\in[t]$ with $v_{i}^{0}=v_{i}^{1}={\texttt{L}}$. The $f\to f^{\prime}$ property states that under these conditions, the sequence $\left((m_{i}^{0\prime},m_{i}^{1\prime})\right)$ for all $i\in[t]$ where $v_{i}^{0\prime}=v_{i}^{1\prime}={\texttt{R}}$ must be $f^{\prime}$-valid.
6.2. Privacy Result.
In this section, we will prove the following theorem:
Theorem 6.1.
Let $\mathcal{P}$ be a deterministic interactive post-processing mechanism with a singleton initial state space satisfying the destination, response, and $f\to f^{\prime}$ properties. If ExtConComp is $(\epsilon,\delta)$-DP w.r.t. $f^{\prime}$ against adaptive adversaries, then the continual mechanism $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-DP w.r.t. $f$ against adaptive adversaries.
In Theorem 6.1, we assume that the view of any adversary interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable. The goal is to show the same result for $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$. The key idea of the proof is to show that the IMs $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{T}\!\circ^{*}\!\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are identical, where $\mathcal{T}$ is an IPM with a singleton initial state space. The proof consists of two steps.
Step 1 (Swapping $\mathcal{P}$ and $\mathcal{I}(b)$). We start by swapping the positions of $\mathcal{P}$ and $\mathcal{I}(b)$ in $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$. The verifier $\mathcal{V}[f]$ sends pairs of query messages, while $\mathcal{P}$ expects a single one. Also, the identifier $\mathcal{I}(b)$ expects pairs of messages, while $\mathcal{P}$ generates a single response. Therefore, when swapping $\mathcal{P}$ and $\mathcal{I}(b)$, we should replace $\mathcal{P}\!\circ^{*}\!\mathcal{I}(b)$ with $\mathcal{I}(b)\!\circ^{*}\!\mathcal{P}^{2}$, where $\mathcal{P}^{2}$ is an interactive post-processing mechanism that runs two instances of $\mathcal{P}$ in parallel. The right mechanism of $\mathcal{P}^{2}$ is $\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$, and its left mechanism is $\mathcal{V}[f]$. The exact definition of $\mathcal{P}^{2}$ is given below.
Definition 6.2 ($\mathcal{P}^{2}$).
Let $\mathcal{P}:S_{\mathcal{P}}\times\left((\{{\texttt{L}}\}\times Q_{\mathcal{P}}^{L})\cup(\{{\texttt{R}}\}\times Q_{\mathcal{P}}^{R})\right)\rightarrow S_{\mathcal{P}}\times\left((\{{\texttt{L}}\}\times A_{\mathcal{P}}^{L})\cup(\{{\texttt{R}}\}\times A_{\mathcal{P}}^{R})\right)$ be a deterministic interactive post-processing mechanism with a singleton initial state space $S_{\mathcal{P}}^{\text{init}}=\{s^{\text{init}}\}$, satisfying the destination and response properties. $\mathcal{P}^{2}:S_{\mathcal{P}}^{2}\times\left((\{{\texttt{L}}\}\times Q_{\mathcal{P}}^{L})\cup(\{{\texttt{R}}\}\times Q_{\mathcal{P}}^{R}\times Q_{\mathcal{P}}^{R})\right)\rightarrow S_{\mathcal{P}}\times\left((\{{\texttt{L}}\}\times A_{\mathcal{P}}^{L})\cup(\{{\texttt{R}}\}\times A_{\mathcal{P}}^{R})\right)$ is an interactive post-processing mechanism that executes two instances of $\mathcal{P}$, referred to as $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$. The state of $\mathcal{P}^{2}$ is a pair $(s_{0},s_{1})$ denoting the current states of $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$. The initial state space of $\mathcal{P}^{2}$ is $\{(s^{\text{init}},s^{\text{init}})\}$.
When $\mathcal{P}^{2}$ receives a message $m\in Q_{\mathcal{P}}^{L}$ from the left mechanism, it forwards two copies of $m$ to both $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ as left messages. Specifically, it executes $(s_{0},v_{0},m_{0}^{\prime})\leftarrow\mathcal{P}_{0}(s_{0},{\texttt{L}},m)$ and $(s_{1},v_{1},m_{1}^{\prime})\leftarrow\mathcal{P}_{1}(s_{1},{\texttt{L}},m)$. When $\mathcal{P}^{2}$ receives a pair of messages $(m_{0},m_{1})\in Q_{\mathcal{P}}^{R}\times Q_{\mathcal{P}}^{R}$ from the right mechanism, it passes each $m_{b}$ to $\mathcal{P}_{b}$ as a right message. Specifically, it executes $(s_{0},v_{0},m_{0}^{\prime})\leftarrow\mathcal{P}_{0}(s_{0},{\texttt{R}},m_{0})$ and $(s_{1},v_{1},m_{1}^{\prime})\leftarrow\mathcal{P}_{1}(s_{1},{\texttt{R}},m_{0})$. By the destination property of $\mathcal{P}$ and the fact that both $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ receive identical left messages, we have $v_{0}^{\prime}=v_{1}^{\prime}$. If $v_{0}^{\prime}=v_{1}^{\prime}={\texttt{R}}$, then $\mathcal{P}^{2}$ sends $(m_{0}^{\prime},m_{1}^{\prime})$ to its right mechanism. If $v_{0}^{\prime}=v_{1}^{\prime}={\texttt{L}}$, then by the response property, $m_{0}^{\prime}=m_{1}^{\prime}$, and $\mathcal{P}^{2}$ returns $m_{0}^{\prime}$ to its left mechanism.
The following claim follow directly from the definition of $\mathcal{P}^{2}$:
Claim 6.3.
For each $b\in\{0,1\}$, the interactive mechanisms $\mathcal{I}(b)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{P}^{2}\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are identical.
Step 2 (Inserting $\mathcal{V}[f^{\prime}]$). Due to the $f\to f^{\prime}$ property, when $\mathcal{P}^{2}$ receives an $f$-valid message sequence from its left mechanism, it sends an $f^{\prime}$-valid message sequence to its right mechanism. We insert $\mathcal{V}[f^{\prime}]$ between $\mathcal{P}^{2}$ and $\mathcal{I}$ in $\mathcal{P}^{2}\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$. By definition, when $\mathcal{V}[f^{\prime}]$ receives an $f^{\prime}$-valid message sequence from its left mechanism, it acts like it does not exist. Specifically, it forwards every message from its left and right mechanism unchanged to the opposite mechanism. Thus, we have:
Claim 6.4.
For each $b\in\{0,1\}$, the interactive mechanisms $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}^{2}\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}^{2}\!\circ^{*}\!\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are identical.
proof of Theorem 6.1.
$\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}^{2}$ in Claim 6.4 is an IPM with a singleton initial state space. Thus, by Lemma 2.15, that $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$, is $(\epsilon,\delta)$-DP w.r.t. $\sim_{01}$ implies the view of any adversary in the interactions with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable, which finishes the proof. ∎
7. Application (Continual Monotone Histogram Query Problem)
We next show an application of our technique to the continual monotone histogram query problem. The continual monotone histogram query problem is defined as follows.
Definition 7.1.
Let $d\in\mathbb{N}$ and let $q:\mathbb{N}^{d}\rightarrow\mathbb{N}$ be a function that is monotone in its input. The input is a finite stream of row vectors $x^{1},\dots,x^{T}$ with $x^{t}\in\{0,1\}^{d}$ for all $t\in[T]$. Let $\sim$ be a neighbor relation on ($\left(\{0,1\}^{d}\right)^{*}$ where two input streams are neighboring if they are identical except for a single step. The output of the (differentially private) continual monotone histogram query problem at each time step $t$ is (an additive approximation to) $q(t)=q((\sum_{\ell=1}^{t}x_{i}^{\ell})_{i\in[d]})$.
Next we define query sensitivity.
Definition 7.2.
The function $q:\{0,1\}^{d}\rightarrow\mathbb{R}$ is $1$-sensitive if $\max_{x\sim y}|q(x)-q(y)|$ is at most $1$.
Henzinger, Sricharan, and Steiner [12] presented the best known differentially private mechanism for the monotone histogram query problem, where where the query is $1$-sensitive. For $\epsilon$-differential privacy it has, with probability at least $1-\beta$, an additive error of $O((d\log^{2}(dq^{*}/\beta)+\log T)\epsilon^{-1})$, and for $(\epsilon,\delta)$-differential privacy, it has an additive error of $O((d\log^{1.5}(dq^{*}/\beta)+\log T)\epsilon^{-1}\log\frac{1}{\delta})$, where $q^{*}=\max_{t\in[T]}q(t)$. The privacy of the mechanism is proven “from scratch” using the concept of a privacy game [13] and the latter bound only holds against an oblivious adversary.
In Section 7.1, we give an overview of the mechanism in [12], which we will call HSS, and in Section 7.2, we give a simple privacy proof for it using our approach. With the new privacy analysis, the result for $(\epsilon,\delta)$-differential privacy is stronger and holds against an adaptive adversary.
7.1. Overview of Mechanism.
[…]
Algorithm 3 Sparse vector technique for histogram queries
procedure SVT-init($\epsilon,q,h$):
     $h_{i}\leftarrow h_{i}$ for all $i\in[d]$
     $\epsilon=\epsilon$
     $q=q$
     $\tau\leftarrow\mathrm{Lap}(1/\epsilon)$
end procedure
procedure SVT-call($x,\texttt{Thresh}$):
     $h_{i}=h_{i}+x_{i}$ for all $i\in[d]$
     if $q(h)+\mathrm{Lap}(2/\epsilon)>\texttt{Thresh}+\tau$ then
         return $\top$ and halt
     else
         return $\bot$
     end if
end procedure
Let $f_{\sim}$ be a verification function that given a message sequence $m_{1},\dots,m_{t}$, returns $\top$ if and only if for each $j\in[t]$, the message $m_{j}$ is of the form $(m_{j}^{0},m_{j}^{1})$ and the message sequences $\sigma_{0}=(m_{1}^{0},\dots,m_{t}^{0})$ and $\sigma_{1}=(m_{1}^{1},\dots,m_{t}^{1})$ satisfy $\sigma_{0}\sim\sigma_{1}$. As part of their privacy proof, the following is shown in [12], which we restate in our own language.
Lemma 7.3.
If $q$ is a $1$-sensitive function, then for every $\epsilon$ and $h$, the mechanism $\textrm{SVT}_{\epsilon/6,q,h}$ is $\epsilon/3$-DP w.r.t. $f_{\sim}$.
For positive integers $t$ and $j$, $\beta\in(0,1]$, and $\epsilon>0$, let $\gamma(t,j,\beta,\epsilon)$ and $\xi(t,j,\beta,\epsilon)$ are some deterministic functions. Given an input stream $x_{1},x_{2},\dots$ with $x_{i}\in\{0,1\}^{d}$, the algorithm partitions the input stream into subsequences, called intervals. At the end of each interval, it gives the $d$-dimensional sum of all the inputs of the interval to a differentially-private histogram mechanism H. Note that H can be used directly to solve the problem; however, the new mechanism HSS has better accuracy bounds in many cases.
[…]
Algorithm 4 Mechanism for answering a monotone histogram query
[…]
$h_{i}\leftarrow 0$ for all $i\in[d]$, $c_{i}\leftarrow 0$ for all $i\in[d]$
$out\leftarrow q(h)$
$j\leftarrow 1$
[…]
$\texttt{Thresh}=\gamma(1,j,\beta,\epsilon)$
for $t\in\mathbb{N}$ do
     $c_{i}\leftarrow c_{i}+x^{(t)}_{i}$ for all $i\in[d]$,
[…]
         $h=\texttt{H-add}(c)$
         $out\leftarrow q(h)$
[…]
         $c_{i}\leftarrow 0$ for all $i\in[d]$
         if $q(c+h)+\mathrm{Lap}(3/\epsilon)>\texttt{Thresh}-\xi(t,j,\beta,\epsilon)$ then
              $\texttt{Thresh}=\texttt{Thresh}+\gamma(t,j,\beta,\epsilon)$
         end if
         $j\leftarrow j+1$
         $\texttt{Thresh}=\texttt{Thresh}-\gamma(t,j-1,\beta,\epsilon)+\gamma(t,j,\beta,\epsilon)$
     end if
     output $out$
     $\texttt{Thresh}=\texttt{Thresh}-\gamma(t,j,\beta,\epsilon)+\gamma(t+1,j,\beta,\epsilon)$
end for
7.2. New Privacy Analysis.
Algorithm 4 describes a CM HSS with query set $\{0,1\}^{d}$ and answer set $\mathbb{N}^{d}$. The state of HSS consists of the variables $(h,c,j,out,\texttt{Thresh})$, initialized with $(0^{d},0^{d},1,q(0^{d}),\gamma(1,1,\beta,\epsilon))$. The summation $q(c+h)+\mathrm{Lap}(3/\epsilon)$ in this algorithm can be rewritten as an in instance of Laplace mechanism with privacy parameter $\epsilon/3$ and input $q(c+h)$. Consequently, HSS executes one instance of a histogram mechanism, H, an unbounded number of SVT instances, and an unbounded number of Laplace mechanisms. The Laplace mechanism with privacy parameter $\epsilon/3$ is a NIM that is known to be $\epsilon/3$-DP when receiving real values differing by at most $1$. In other words, the Laplace mechanism is a CM that is $\epsilon/3$-DP w.r.t. a verification function $f_{\mathrm{Lap}}$ that returns $\top$ if and only if it receives a single input message $m=(m_{0},m_{1})\in\mathbb{R}^{2}$ satisfying $|m_{0}-m_{1}|\leq 1$.
The output of H is a vector in $\mathbb{R}^{d}$, while the output of the Laplace mechanism is a scalar in $\mathbb{R}$. Rounding up each real-valued output to the nearest integer does not affect privacy guarantees (by the post-processing lemma) but increases the additive error of each mechanism by at most $1$. This extra error is negligible compared to their original error. Therefore, we assume that both H and the Laplace mechanism have discrete answer distributions. Since the answer set of SVT is finite, it also has discrete answer distributions.
By design, HSS uses the inputs $x^{(t)}$ only to determine the inputs of its sub-mechanisms, and every other action is taken solely based on the privatized outputs of the sub-mechanisms. This allows us to apply Theorem 6.1 to analyze the privacy of HSS based on the privacy parameters of its sub-mechanisms.
By Lemma 7.3, each instance of SVT created by HSS is $\epsilon/3$-DP. Furthermore, for any two neighboring input streams, fixing the history of privatized outputs of the sub-mechanisms, HSS provides identical inputs to all but one SVT instance, which receives neighboring inputs. By Theorem 5.2, the parallel concurrent composition of all SVT instances is $\epsilon/3$-DP.
Similarly, HSS gives identical real-valued inputs to all but one instance of the Laplace mechanism, where the differing inputs vary by at most $1$. This implies that the concurrent composition of all Laplace mechanisms is also $\epsilon/3$-DP. Each concurrent composition is itself a continual mechanism. Additionally, it is given that H is $\epsilon/3$-DP. By basic composition, the composition of three instances of $\mathrm{RR}_{\epsilon/3,0}$ is $\epsilon$-DP. Therefore, by Theorem 4.10, the concurrent composition of all mechanisms—H, SVT instances, and Laplace mechanisms—is $\epsilon$-DP. Applying Theorem 6.1, the mechanism HSS is $\epsilon$-DP.
In the case where H is $(\epsilon/3,\delta)$-DP, since the composition of $\mathrm{RR}_{\epsilon/3,0}$, $\mathrm{RR}_{\epsilon/3,0}$, and $\mathrm{RR}_{\epsilon/3,\delta}$ is $(\epsilon,\delta)$-DP, the mechanism HSS is also $(\epsilon,\delta)$-DP. Both privacy guarantees provided in this section hold against adaptive adversaries.
Acknowledgments
1Monika Henzinger and Roodabeh Safavi were supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101019564), and the Austrian Science Fund (FWF) under grants DOI 10.55776/Z422, DOI 10.55776/I5982, and DOI 10.55776/P33775, with additional funding from the netidee SCIENCE Stiftung, 2020–2024.
2Salil Vadhan was supported by NSF grant BCS-2218803, a grant from the Sloan Foundation, and a Simons Investigator Award. Work began while a Visiting Researcher at the Bocconi University Department of Computing Sciences, supported by Luca Trevisan’s ERC Project GA-834861.
Appendix A Overview and Modification of [18]
In this section, we overview and modify the results of [18]. The original proof relies on the existence of a constant $T$ that upper bounds the number of queries adversaries ask; however, our modified construction does not. Moreover, an algorithmic construction in [18]’s proof needs the answer set to be finite. Using mathematical tools, we formalize their construction and weaken this assumption to having discrete answer distributions. We show:
Lemma A.1.
For $\epsilon>0$ and $0\leq\delta\leq 1$, let $\mathcal{M}:S_{\mathcal{M}}\times Q_{\mathcal{M}}\to S_{\mathcal{M}}\times A_{\mathcal{M}}$ be an $(\epsilon,\delta)$-DP IM w.r.t. a neighboring relation $\sim$. Suppose that $\mathcal{M}$ has discrete answer distributions. Then, for any two neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,\delta}(b)$ are identical.
Notation A.2.
For short, let $Q=Q_{\mathcal{M}}$ and $A=A_{\mathcal{M}}$. For any $T\in\mathbb{N}$, let $[T]=\{1,\dots,T\}$. For each $t\in\mathbb{N}$ and $b\in\{0,1\}$, we define the function $\mu_{t}^{b}:(Q\times A)^{t}\to[0,1]$ such that $\mu_{t}^{b}(q_{1},\dots,a_{t})$ denotes the probability of $\mathcal{M}(s_{b})$ returning answers $a_{1},\dots,a_{t}$ to the queries $q_{1},\dots,q_{t}$.
Fixing an upper bound $T$ on the number of queries, Lyu [18] reduces the construction of $\mathcal{P}$ to defining a sequence of functions $\phi_{t}^{0}$ and $\phi_{t}^{1}$ for each $t\in[T]$, satisfying three conditions. They then inductively construct each $\phi_{t}^{b}$ and verify that it satisfies the required conditions. The definition of these conditions is based on a sequence of control function $\mathrm{L}_{t,T}^{0}$ and $\mathrm{L}_{t,T}^{1}$, which have backward recursive definitions for $t$ from $T$ to $1$.
In Section A.1.1, we overview the definition of $\mathrm{L}_{t,T}^{b}$ and its properties. In Section A.1.2, we introduce a new control function $\mathrm{L}_{t}^{b}$ with no dependency on $T$. The goal is to use $\mathrm{L}_{t}^{b}$ instead of $\mathrm{L}_{t,T}^{b}$ in [18]’s proofs. So, in Section A.1.3, we show some properties for $\mathrm{L}_{t}^{b}$, required in later proofs.
In Section A.2, we overview the reduction presented in [18], with the control functions $\mathrm{L}_{t,T}^{b}$ replaced by $\mathrm{L}_{t}^{b}$. In this section, we prove the reduction still holds with the new control functions. Finally, in Section A.3, we present a construction for $\phi_{t}$ and show it satisfies all three conditions. We note that this construction is identical in essence to one given in [18], but we extend the proof to work with weaker assumptions.
A.1. Preliminaries
A.1.1. Control Functions $\mathrm{Lower}_{t,T}$ and $\mathrm{L}_{t,T}$.
Lyu [18] introduces the following control function:
Definition A.3 ($\mathrm{Lower}_{t,T}$).
For $T\in\mathbb{N}$, $b\in\{0,1\}$, and $t\in[T]$, the function $\mathrm{Lower}_{t,T}^{b}:(Q\times A)^{t-1}\times Q\to[0,1]$ is defined recursively as follows: for each transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
$\mathrm{Lower}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t})=\begin{cases}\sum\limits_{a_{t}\in A}\sup_{q_{t+1}\in Q}\mathrm{Lower}_{t+1,T}^{b}(q_{1},\dots a_{t},q_{t+1}),&t<T\\ \sum\limits_{a_{t}\in A}\max\{0,\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})\},&t=T\end{cases}$
We introduce another control function $\mathrm{L}_{t,T}$ based on $\mathrm{Lower}_{t,T}$:
Definition A.4 ($\mathrm{L}_{t,T}^{b}$).
For $T\in\mathbb{N}$, $b\in\{0,1\}$, and $t\in[T]$, the function $\mathrm{L}_{t,T}^{b}:(Q\times A)^{t}\to[0,1]$ is defined recursively as follows: for each transcript $(q_{1},\dots a_{t})\in(Q\times A)^{t}$,
$\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})=\begin{cases}\sup_{q_{t+1}\in Q}\mathrm{Lower}_{t+1,T}^{b}(q_{1},\dots,a_{t},q_{t+1}),&t<T\\ \max\{0,\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})\},&t=T\end{cases}$
The following corollary follows directly from the definitions of $\mathrm{Lower}_{t,T}^{b}$ and $\mathrm{L}_{t,T}^{b}$:
Corollary A.5.
For every $T\in\mathbb{N}$, $b\in\{0,1\}$, $t\in[T]$, and $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
$\mathrm{Lower}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t})=\sum_{a_{t}\in A}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t}).$
Lyu [18] shows the following two claims about $\mathrm{Lower}_{t,T}^{b}$:
Claim A.6 ([18]).
For each $b\in\{0,1\}$, $T\in\mathbb{N}$, and $q_{1}\in Q$ $\mathrm{Lower}_{1,T}^{b}(q_{1})\leq\delta$.
Claim A.7 ([18]).
For each $b\in\{0,1\}$, $T\in\mathbb{N}$, $t\in[T]$ [^3] In the original claim in [18], $t$ is assumed to be at most $T-1$, which is a typo., and transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$, we have
$\mathrm{Lower}_{t,T}^{b}(q_{1},a_{1},\dots,q_{t})$ $-e^{-\epsilon}\mathrm{Lower}_{t,T}^{1-b}(q_{1},a_{1},\dots,q_{t})\leq\mu_{t-1}^{b}(q_{1},\dots,a_{t-1})-e^{-\epsilon}\mu_{t-1}^{1-b}(q_{1},\dots,a_{t-1}).$
Claim A.6 combined with the definition of $\mathrm{L}_{t,T}^{b}$ imply:
Corollary A.8.
For each $b\in\{0,1\}$, and $T\in\mathbb{N}$, $\mathrm{L}_{0,T}^{b}()\leq\delta$.
As we point out later, in some part of the proof, Lyu [18] wrongly refers to Claim A.7, while the following lemma is required. For completeness, we prove this lemma and use it in the next section.
Lemma A.9.
For each $b\in\{0,1\}$, $T\in\mathbb{N}$, $t\in[T]\cup\{0\}$, and transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, we have
$\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$ $-e^{-\epsilon}\mathrm{L}_{t,T}^{1-b}(q_{1},\dots,a_{t})\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}).$
Proof.
For $t=T$, we need to show that
$\max\{0,\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})\}-e^{-\epsilon}\max\{0,\mu_{t}^{1-b}(q_{1},\dots,a_{t})-e^{\epsilon}\mu_{t}^{b}(q_{1},\dots,a_{t})\}$
$\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$
For short, we refer to the LHS and RHS as $A$ and $B$, respectively. There are three cases:
  • If $\mu_{t}^{b}(q_{1},\dots,a_{t})\leq e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$, then $A=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}),$ which is at most $B$.
  • If $\mu_{t}^{b}(q_{1},\dots,a_{t})\geq e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$, then, $A=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}),$ which is at most $B$ as $e^{\epsilon}\geq e^{-\epsilon}$.
  • Otherwise, if $e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})<\mu_{t}^{b}(q_{1},\dots,a_{t})<e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$, $A=0$ is at most $B$ since by the condition of the case distinction, $B=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$ is positive.
For $t<T$, we have
$\mathrm{L}_{t,T}^{b}(q_{1},a_{1},\dots,q_{t})-e^{-\epsilon}\mathrm{L}_{t,T}^{1-b}(q_{1},a_{1},\dots,q_{t})$
$=\sup_{q_{t+1}\in Q}\mathrm{Lower}_{t+1,T}^{b}(q_{1},a_{1},\dots,q_{t+1})-e^{-\epsilon}\sup_{q_{t+1}\in Q}\mathrm{Lower}_{t+1,T}^{1-b}(q_{1},a_{1},\dots,q_{t+1})$
$\leq\sup_{q_{t+1}\in Q}\left(\mathrm{Lower}_{t,T}^{b}(q_{1},a_{1},\dots,q_{t+1})-e^{-\epsilon}\mathrm{Lower}_{t,T}^{1-b}(q_{1},a_{1},\dots,q_{t+1})\right)$
$\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{b}(q_{1},\dots,a_{t}),$
where the last inequality holds because for every $q_{t+1}\in Q$, by Claim A.7
$\mathrm{Lower}_{t+1,T}^{b}(q_{1},a_{1},\dots,q_{t+1})$ $-e^{-\epsilon}\mathrm{Lower}_{t+1,T}^{1-b}(q_{1},a_{1},\dots,q_{t+1})\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{b}(q_{1},\dots,a_{t}).$
A.1.2. Control Function $\mathrm{L}_{t}^{b}$ with is the Limit of $\mathrm{L}_{t,T}$ (Definition).
This section introduces our main new technical contribution. We introduce $\mathrm{L}_{t}^{b}$ as limit of $\mathrm{L}_{t,T}^{b}$ and show that it satisfies all properties needed by the proof of [18]. Thus, differently from [18], we will use $\mathrm{L}_{t}^{b}$ to define the functions $\phi_{t}^{0}$ and $\phi_{t}^{1}$.
However, we first need to argue that the limit of $\mathrm{L}_{t,T}^{b}$ exists. the value of $\mathrm{L}_{t,T}^{b}$ is upper bounded by $1$. We also show that this function is non-decreasing in $T$. As a result, for each each $t\in\mathbb{N}\cup\{0\}$ and $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the limit $\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$ exists. To remove the assumption of having an upper bound $T$ on the number of queries, we introduce the following control function:
Definition A.10 ($\mathrm{L}_{t}^{b}$).
For each $b\in\{0,1\}$ and $t\in\mathbb{N}$, the function $\mathrm{L}_{t}^{b}:(Q\times A)^{t}\to[0,1]$ is defined as follows: for each transcript $(q_{1},\dots a_{t})\in(Q\times A)^{t}$,
$\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t})=\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t}).$
We need the following lemma throughout the section:
Lemma A.11.
For every $t\in\mathbb{N}$, transcript $(q_{1},\dots a_{t-1})\in(Q\times A)^{t-1}$, and query $q_{t}\in Q$, we have
$\mu_{t-1}^{b}(q_{1},a_{1},\dots,q_{t-1},a_{t-1})=\sum_{a_{t}\in A}\mu_{t}^{b}(q_{1},a_{1},\dots,q_{t},a_{t}).$
Proof.
For $a_{t}\in A$, let $\mu_{t}^{b}(a_{t}\mid q_{1},a_{1},\dots,q_{t-1},a_{t-1},q_{t})$ denote the probability that $\mathcal{M}(s_{b})$ returns the answer $a_{t}$ in response to $q_{t}$ conditioned on $q_{1},a_{1},\dots,q_{t-1},a_{t-1}$ being the history of previous queries and answers. By the chain rule, we have
$\mu_{t}^{b}(q_{1},a_{1},\dots,q_{t},a_{t})$ $=\sum_{a_{t}\in A}\mu_{t-1}^{b}(q_{1},a_{1},\dots,q_{t-1},a_{t-1})\cdot\mu_{t}^{b}(a_{t}\mid q_{1},\dots,a_{t-1},q_{t})$
$=\mu_{t-1}^{b}(q_{1},a_{1},\dots,q_{t-1},a_{t-1})\sum_{a_{t}\in A}\mu_{t}^{b}(a_{t}\mid q_{1},\dots,a_{t-1},q_{t})$
$=\mu_{t-1}^{b}(q_{1},a_{1},\dots,q_{t-1},a_{t-1})$
To upper bound $\mathrm{L}_{t,T}$, we first need to upper bound $\mathrm{Lower}_{t,T}$:
Lemma A.12.
For every $T\in\mathbb{N}$, $b\in\{0,1\}$, $t\in[T]$, and transcript $(q_{1},\dots a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
$\mathrm{Lower}_{t,T}^{b}(q_{1},\dots a_{t-1},q_{t})\leq\mu_{t}^{b}(q_{1},\dots,a_{t-1}).$
Proof.
Fix $T\geq t$. The proof is by induction on $t$ from $T$ to $1$. For $t=T$, by definition and Lemma A.11, we have
$\mathrm{Lower}_{t,T}^{b}(q_{1},\dots a_{t-1},q_{t})$ $=\sum_{a_{t}\in A}\max\{0,\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})\}$
$\leq\sum_{a_{t}\in A}\max\{0,\mu_{t}^{b}(q_{1},\dots,a_{t})\}=\mu_{t-1}^{b}(q_{1},\dots,a_{t-1})$
For $t<T$, by the inductive hypothesis,
$\mathrm{Lower}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t})$ $\leq\sum_{a_{t}\in A}\sup_{q_{t+1}\in Q}\mu_{t-1}^{b}(q_{1},\dots a_{t})=\sum_{a_{t}\in A}\mu_{t-1}^{b}(q_{1},\dots a_{t})=\mu_{t-1}^{b}(q_{1},\dots,a_{t-1}),$
completing the induction. ∎
The following lemma implies that the values of $\mathrm{L}_{t,T}$ are upper bounded by $1$.
Lemma A.13.
For every $T\in\mathbb{N}$, $b\in\{0,1\}$, $t\in[T]\cup\{0\}$, and transcript $(q_{1},\dots a_{t})\in(Q\times A)^{t}$,
$\mathrm{L}_{t,T}^{b}(q_{1},\dots a_{t})\leq\mu_{t}^{b}(q_{1},\dots,a_{t}).$
Proof.
For $t=T$, by definition and Lemma A.11, we have
$\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$ $=\max\{0,\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})\}$
$\leq\max\{0,\mu_{t}^{b}(q_{1},\dots,a_{t})\}=\mu_{t}^{b}(q_{1},\dots,a_{t})$
For $t<T$, by Lemma A.12,
$\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$ $=\sup_{q_{t+1}\in Q}\mathrm{Lower}_{t+1,T}^{b}(q_{1},\dots,a_{t},q_{t+1})\leq\sup_{q_{t}\in Q}\mu_{t}^{b}(q_{1},\dots,a_{t})=\mu_{t}^{b}(q_{1},\dots,a_{t})$
To ensure the function $\mathrm{L}_{t}^{b}$ is well-defined, it remains to show that $\mathrm{L}_{t,T}^{b}$ is non-decreasing in $T$. To prove so, we first need to show:
Lemma A.14.
For each $b\in\{0,1\}$ and $t\in\mathbb{N}$, the function $\mathrm{Lower}_{t,T}^{b}$ is monotonically non-decreasing in $T$.
Proof.
We need to show that for every $T\in\mathbb{N}$, $b\in\{0,1\}$, $t\leq T$, and transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
$\mathrm{Lower}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t})\leq\mathrm{Lower}_{t,T+1}^{b}(q_{1},\dots,a_{t-1},q_{t}).$
Fix $T\in\mathbb{N}$ and $b\in\{0,1\}$. We show this by reverse induction on $t$.
The base case is $t=T$. By Lemma A.11, for every $(q_{1},\dots,a_{t-1},q_{t})$, we have
$\mathrm{Lower}_{T,T}^{b}(q_{1},\dots,a_{T-1},q_{T})=\sum\limits_{a_{T}\in A}\max\{0,\mu_{T}^{b}(q_{1},\dots,a_{T})-e^{\epsilon}\mu_{T}^{1-b}(q_{1},\dots,a_{T})\}$
$=\sum\limits_{a_{T}\in A}\sup_{q_{T+1}\in Q}\max\{0,\sum\limits_{a_{T+1}\in A}\mu_{T+1}^{b}(q_{1},\dots,,a_{T+1})-e^{\epsilon}\mu_{T+1}^{1-b}(q_{1},\dots,a_{T+1})\}$
$\leq\sum\limits_{a_{T}\in A}\sup_{q_{T+1}\in Q}\sum\limits_{a_{T+1}\in A}\max\{0,\mu_{T+1}^{b}(q_{1},\dots,a_{T+1})-e^{\epsilon}\mu_{T+1}^{1-b}(q_{1},\dots,a_{T+1})\}$
$=\sum\limits_{a_{T}\in A}\sup_{q_{T+1}\in Q}\mathrm{Lower}_{T+1,T+1}(q_{1},\dots,a_{T},q_{T+1})$
$=\mathrm{Lower}_{T,T+1}^{b}(q_{1},\dots,a_{T-1},q_{T})$
For the induction step, assume that $t<T$ and that for every $(q_{1},\dots,a_{t},q_{t+1})$, we have $\mathrm{Lower}_{t+1,T}^{b}(q_{1},\dots,a_{t},q_{t+1})\leq\mathrm{Lower}_{t+1,T+1}^{b}(q_{1},\dots,a_{t},q_{t+1})$. We need to show $\mathrm{Lower}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t})\leq\mathrm{Lower}_{t,T+1}^{b}(q_{1},\dots,a_{t-1},q_{t})$ for every $(q_{1},\dots,a_{t-1},q_{t})$. By definition,
$\mathrm{Lower}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t})$ $=\sum\limits_{a_{t}\in A}\sup_{q_{t+1}\in Q}\mathrm{Lower}_{t+1,T}^{b}(q_{1},\dots a_{t},q_{t+1})$
$\leq\sum\limits_{a_{t}\in A}\sup_{q_{t+1}\in Q}\mathrm{Lower}_{t+1,T+1}^{b}(q_{1},\dots a_{t},q_{t+1})$
$=\mathrm{Lower}_{t,T+1}^{b}(q_{1},\dots,a_{t-1},q_{t})$
Lemma A.15.
For each $b\in\{0,1\}$ and $t\in\mathbb{N}\cup\{0\}$, the function $\mathrm{L}_{t,T}^{b}$ is monotonically non-decreasing in $T$.
Proof.
We need to show that for every $T\in\mathbb{N}$, $t\in[T]\cup\{0\}$, and transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$,
$\mathrm{L}_{t,T}(q_{1},\dots,a_{t})\leq\mathrm{L}_{t,T+1}(q_{1},\dots,a_{t})$
Fix $T\in\mathbb{N}$ and $b\in\{0,1\}$. We show this by reverse induction on $t$ from $T$ to 00.
The base case is $t=T$. By Lemma A.11, for every $(q_{1},\dots,a_{T})$ we have
$\mathrm{L}_{T,T}^{b}(q_{1},\dots,a_{T})=\max\{0,\mu_{T}^{b}(q_{1},\dots,a_{T})-e^{\epsilon}\mu_{T}^{1-b}(q_{1},\dots,a_{T})\}$
$=\sup_{q_{T+1}\in Q}\max\{0,\sum\limits_{a_{T+1}\in A}\mu_{T+1}^{b}(q_{1},\dots,a_{T+1})-e^{\epsilon}\mu_{T+1}^{1-b}(q_{1},\dots,a_{T+1}))\}$
$\leq\sup_{q_{T+1}\in Q}\sum\limits_{a_{T+1}\in A}\max\{0,\mu_{T+1}^{b}(q_{1},\dots,a_{T+1})-e^{\epsilon}\mu_{T+1}^{1-b}(q_{1},\dots,a_{T+1})\}$
$=\sup_{q_{T+1}\in Q}\mathrm{Lower}_{T+1,T+1}(q_{1},\dots,a_{T},q_{T+1})$
$=\mathrm{L}_{T,T+1}^{b}(q_{1},\dots,a_{T-1},q_{T},a_{T})$
For the induction step, assume that $t<T$ and that for every $(q_{1},\dots,a_{t+1})$, we have $\mathrm{L}_{t+1,T}^{b}(q_{1},\dots,a_{t+1})\leq\mathrm{L}_{t+1,T+1}^{b}(q_{1},\dots,a_{t+1})$. We need to show $\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})\leq\mathrm{L}_{t,T+1}^{b}(q_{1},\dots,a_{t})$ for every $(q_{1},\dots,a_{t})$. By definition of $\mathrm{L}_{t,T+1}$ and Lemma A.14, we have
$\mathrm{L}_{t,T}^{b}(q_{1},\dots,q_{t},a_{t})$ $=\sup_{q_{t+1}\in Q}\mathrm{Lower}_{t+1,T}^{b}(q_{1},\dots a_{t},q_{t+1})$
$\leq\sup_{q_{t+1}\in Q}\mathrm{Lower}_{t+1,T+1}^{b}(q_{1},\dots a_{t},q_{t+1})$
$=\mathrm{L}_{t,T+1}^{b}(q_{1},\dots,q_{t},a_{t})$
A.1.3. Control Function $\mathrm{L}_{t}^{b}$ (Properties).
In this section, we present the properties of $\mathrm{L}_{t}^{b}$, required in the next sections.
The definition of $\mathrm{L}_{t}^{b}$ combined with Corollary A.8, Lemma A.9, and Lemma A.13 give the following corollaries, respectively:
Corollary A.16.
For each $b\in\{0,1\}$, $\mathrm{L}_{0}^{b}()\leq\delta$.
Corollary A.17.
For each $b\in\{0,1\}$, $t\in\mathbb{N}\cup\{0\}$, and transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, we have
$\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t})$ $-e^{-\epsilon}\mathrm{L}_{t}^{1-b}(q_{1},\dots,a_{t})\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}).$
Lemma A.18.
For each $b\in\{0,1\}$, $t\in\mathbb{N}$, and transcript $(q_{1},\dots a_{t})\in(Q\times A)^{t}$,
$\mathrm{L}_{t}^{b}(q_{1},\dots a_{t})\leq\mu_{t}^{b}(q_{1},\dots,a_{t}).$
To show the next property, we need the following known fact:
Lemma A.19 (Monotone Convergence).
Let $f:\mathbb{N}^{2}\to\mathbb{R}$ be a function such that $\sum_{m}f(m,1)>-\infty$ and that is non-decreasing in the second parameter. Then
$\lim_{n\to\infty}\sum_{m=1}^{\infty}f(m,n)=\sum_{m=1}^{\infty}\lim_{n\to\infty}f(m,n)$
Lemma A.19 combined with Lemma A.15 implies:
Corollary A.20.
For any $b\in\{0,1\}$, $t\in\mathbb{N}$, and $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
$\sum_{a_{t}\in A}\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})=\lim_{T\geq t,T\to\infty}\sum_{a_{t}\in A}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$
Lemma A.21.
For every $b\in\{0,1\}$, $t\in\mathbb{N}$, transcript $(q_{1},\dots a_{t-1})\in(Q\times A)^{t-1}$, and query $q_{t}\in Q$,
$\mathrm{L}_{t-1}^{b}(q_{1},\dots a_{t-1})\geq\sum_{a_{t}\in A}\mathrm{L}_{t}^{b}(q_{1},\dots a_{t})$
Proof.
First, we will show that for any $T>t-1$,
$\mathrm{L}_{t-1,T}^{b}(q_{1},\dots a_{t-1})\geq\sum_{a_{t}\in A}\mathrm{L}_{t,T}^{b}(q_{1},\dots a_{t})$
Since $t\neq T$, by definition of $\mathrm{L}_{t-1,T}^{b}$ and Corollary A.5, we have
$\mathrm{L}_{t-1,T}^{b}(q_{1},\dots a_{t-1})$ $=\sup_{q\in Q}\mathrm{Lower}_{t,T}(q_{1},\dots a_{t-1},q)\geq\mathrm{Lower}_{t,T}(q_{1},\dots a_{t-1},q_{t})=\sum_{a_{t}\in A}\mathrm{L}_{t,T}^{b}(q_{1},\dots a_{t}).$
Thus, by Corollary A.20, we have
$\mathrm{L}_{t-1}^{b}(q_{1},\dots a_{t-1})$ $=\lim_{T\geq t-1,T\to\infty}\mathrm{L}_{t-1,T}^{b}(q_{1},\dots a_{t-1})=\lim_{T\geq t,T\to\infty}\mathrm{L}_{t-1,T}^{b}(q_{1},\dots a_{t-1})$
$\geq\lim_{T\geq t,T\to\infty}\sum_{a_{t}\in A}\mathrm{L}_{t,T}^{b}(q_{1},\dots a_{t})=\sum_{a_{t}\in A}\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots a_{t})=\sum_{a_{t}\in A}\mathrm{L}_{t}^{b}(q_{1},\dots a_{t})$
A.2. Reduction.
To define $\mathcal{P}$, Lyu [18] constructs four families of PMFs $\mathcal{F}_{(0,\bot)}$, $\mathcal{F}_{(1,\bot)}$, $\mathcal{F}_{(0,\top)}$, and $\mathcal{F}_{(1,\top)}$, corresponding to each outcome of $\mathrm{RR}_{\epsilon,\delta}$. Intuitively, each family describes an IM. Specifically, each family consists of PMFs over the answer set $A$ for every $t\in[T]$ and every history $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t}\times Q$, representing the probability of returning each answer in response to $q_{t}$, conditioned on $q_{1},\dots,a_{t-1}$ being the history of queries and answers.
A.2.1. Constructing $\mathcal{P}$ using $\mathcal{F}_{r}$s.
Roughly speaking, in the interaction between $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,\delta}(b)$ and an adversary $\mathcal{A}$, $\mathcal{P}$ interacts with $\mathrm{RR}_{\epsilon,\delta}(b)$ once and stores its output $r$ as thee first parameter in its state. It then answers the adversary’s queries by itself using the PMFs provided by $\mathcal{F}_{r}$. Specifically, $\mathcal{P}$ has a single initial state $\left(\text{null},()\right)$, where null is an empty field and ()()( ) represents an empty sequence. Let $h$ denote the sequence of interactions between $\mathcal{P}$ and $\mathcal{A}$, which is stored by $\mathcal{P}$ as second parameter of its state. After interacting with $\mathrm{RR}_{\epsilon,\delta}(b)$, $\mathcal{P}$ updates null to $r$. Moreover, upon receiving a query $q$ and before returning an answer $a$, $\mathcal{P}$ appends $q$ and $a$ to the history sequence $h$. To generate a response $a$ to a query $q$, $\mathcal{P}$ samples $a$ from the PMF corresponding to $(h,q)$ in $\mathcal{F}_{r}$, where $h$ is the history of all previous queries and answers.
A.2.2. Constructing $\mathcal{F}_{(b,\bot)}$s.
Lyu [18] constructs a function $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for each $t\in[T]\cup\{0\}$. Let $\phi_{t}^{0}$ and $\phi_{t}^{1}$ denote the functions returning the first and second elements of the output pairs of $\phi_{t}$. For each $b\in\{0,1\}$, the family $\mathcal{F}_{(b,\bot)}$ is defined based on $\phi_{t}^{b}$s. Intuitively, $\phi_{t}^{b}(q_{1},\dots,a_{t})$ represents the probability that $\mathcal{M}(s_{b})$ returns the answers $a_{1},\dots,a_{t}$ to the fixed queries $q_{1},\dots,q_{t}$. That is, for each query sequence $q_{1},\dots,q_{t}$, $\phi_{t}^{b}(q_{1},\cdot,q_{2},\cdot,\dots,q_{t},\cdot)$ is a PMF over $A^{t}$.
We do not provide the construction of $\phi_{t}$ in this section. However, assuming the existence of $\phi_{t}$ satisfying the following three conditions, we show that for every $b\in\{0,1\}$, the mechanisms $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}$ and $\mathcal{M}(s_{b})$ are equivalent. These three conditions were originally stated in [18], with the control function $\mathrm{L}_{t,T}^{b}$ instead of $\mathrm{L}_{t}^{b}$.
For $t=0$, let $\phi_{t}()=(1,1)$. Suppose for every $t\in[T]$, there exists a function $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ satisfying the following conditions for each $b\in\{0,1\}$:
  1. (I)
    For any transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$,
    $\delta\phi_{t}^{b}(q_{1},\dots,a_{t})\geq\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t}).$
  2. (II)
    For any transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$,
    $\delta\phi_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\delta\phi_{t}^{1-b}(q_{1},\dots,a_{t})\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}).$
  3. (III)
    For any transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
    $\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)=\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
For each $b\in\{0,1\}$, the family $\mathcal{F}_{(b,\bot)}$ consists of functions $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}:A\to[0,1]$ for each $(q_{1},\dots,a_{t-1},q_{t})\in\bigcup_{t=1}^{T}(Q\times A)^{t-1}\times Q$, where $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ is defined as follows:
  • If $\phi_{t-1}(q_{1},\dots,a_{t-1})>0$, then for each answer $a_{t}\in A$,
    $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=\frac{\phi_{t}(q_{1},\dots,a_{t})}{\phi_{t-1}(q_{1},\dots,a_{t-1})}.$
  • Otherwise, if $\phi_{t-1}(q_{1},\dots,a_{t-1})=0$, $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=0$ for every answer, except for one fixed arbitrary answer $a^{*}\in A$, where $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a^{*})=1$. This case is included only for formality, and the corresponding PMFs are never used by $\mathcal{P}$.
Since $\phi_{t-1}$ and $\phi_{t}$ are non-negative, $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})$ is also non-negative. Furthermore, by (III), we have $\sum_{a_{t}\in A}f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a)=1$, implying that $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})$ is a PMF.
A.2.3. Constructing $\mathcal{F}_{(b,\top)}$s.
Similar to the previous section, $\mathcal{F}_{(0,\top)}$ and $\mathcal{F}_{(1,\top)}$ are defined based on a sequence of functions $\psi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for each $t\in[T]\cup\{0\}$. For $t=0$, $\psi_{t}()=(1,1)$. For $t\in[T]$, $\psi_{t}$ is defined as follows:
  • If $\delta=1$, then $\psi_{t}$ is identical to $\phi_{t}$.
  • Otherwise, for each $b\in\{0,1\}$ and $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, $\psi_{t}^{b}(q_{1},\dots,a_{t})$ equals
    $\frac{1}{(1-\delta)(e^{\epsilon}-1)}\Big{(}e^{\epsilon}\mu_{t}^{b}(q_{1},\dots,a_{t})-\delta e^{\epsilon}\phi_{t}^{b}(q_{1},\dots,a_{t})-\mu_{t}^{1-b}(q_{1},\dots,a_{t})+\delta\phi_{t}^{1-b}(q_{1},\dots,a_{t})\Big{)}.$
For each $b\in\{0,1\}$, the family $\mathcal{F}_{(b,\top)}$ consists of functions $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}:A\to[0,1]$ for each $(q_{1},\dots,a_{t-1},q_{t})\in\bigcup_{t=1}^{T}(Q\times A)^{t-1}\times Q$, where $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ is defined as follows:
  • If $\psi_{t-1}(q_{1},\dots,a_{t-1})>0$, then for each answer $a_{t}\in A$,
    $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=\frac{\psi_{t}(q_{1},\dots,a_{t})}{\psi_{t-1}(q_{1},\dots,a_{t-1})}.$
  • Otherwise, $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=0$ for every answer, except for one fixed arbitrary answer $a^{*}\in A$, where $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a^{*})=1$.
By condition (II), each $\psi_{t}^{b}$ is non-negative. Also, by Lemma A.11 and condition (III), for every $(q_{1},\dots,a_{t},q_{t})\in(Q\times A)^{t-1}\times Q$,
$\sum_{a\in A}\psi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)=\psi_{t-1}^{b}(q_{1},\dots,a_{t-1}),$
implying that each $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ forms a valid PMF.
A.2.4. Proof of Lemma A.1.
By definition of $\mathrm{RR}$ and of $\mathcal{P}$, for each $t\in\mathbb{N}$ and transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the probability of $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,\delta}(b)$ returning the answers $a_{1},\dots,a_{t}$ to the queries $q_{1},\dots,q_{t}$ is
$\delta\phi_{t}^{b}(q_{1},\dots,a_{t})+(1-\delta)\frac{e^{\epsilon}}{1+e^{\epsilon}}\psi_{t}^{b}(q_{1},\dots,a_{t})+(1-\delta)\frac{1}{1+e^{\epsilon}}\psi_{t}^{1-b}(q_{1},\dots,a_{t}).$
by the definition of $\psi_{t}$, this probability equals
$\delta\phi_{t}^{b}(q_{1},\dots,a_{t})+\frac{e^{\epsilon}}{(1+e^{\epsilon})(e^{\epsilon}-1)}\Big{(}e^{\epsilon}\mu_{t}^{b}(q_{1},\dots,a_{t})$
$\quad-\delta e^{\epsilon}\phi_{t}^{b}(q_{1},\dots,a_{t})-\mu_{t}^{1-b}(q_{1},\dots,a_{t})+\delta\phi_{t}^{1-b}(q_{1},\dots,a_{t})\Big{)}$
$\quad+\frac{1}{(1+e^{\epsilon})(e^{\epsilon}-1)}\Big{(}e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})-\delta e^{\epsilon}\phi_{t}^{1-b}(q_{1},\dots,a_{t})-\mu_{t}^{b}(q_{1},\dots,a_{t})+\delta\phi_{t}^{b}(q_{1},\dots,a_{t})\Big{)}$
$=\mu_{t}^{b}(q_{1},\dots,a_{t}).$
Thus, the views of any adversary interacting with $\mathcal{M}(s_{b})$ and $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,\delta}(b)$ are identically distributed, implying that these mechanisms are equivalent.
A.3. Construction of $\phi_{t}$
In this section, we formalize the construction of [18] differently to remove the finiteness assumption on $A$.
Roughly speaking, Lyu [18] initializes the values of $\phi_{t}$ with the ones from $\mathrm{L}_{t}$. Then, for each $q_{1},\dots,a_{t-1},q_{t}$, they iterate over all possible answers $a_{t}\in A$ in a fixed order and gradually increase the value of $\phi_{t}(q_{1},\dots,a_{t})$ to satisfy conditions (II) and (III), while ensuring that condition (I) remains valid. This iterative procedure requires $A$ to be finite; otherwise, the construction algorithm does not terminate. To extend this approach to infinite $A$, we define $\phi_{t}$ mathematically and apply the transfinite recursion theorem to prove its existence.
For each $t\in\mathbb{N}$, knowing that $\phi_{t-1}$ exists and satisfies condition (I), we construct $\phi_{t}$ and show it satisfies conditions (I), (II), and (III). For $t=0$, we set $\phi_{0}()=(1,1)$. By Corollary A.16, $\phi_{0}$ satisfies condition (I).
Remark A.22.
In the rest of this section, we assume $\phi_{t-1}$ exists and satisfies condition (I).
In Section A.3.1, we introduce the necessary preliminaries for formally defining $\phi_{t}$ when $A$ is infinite. In Section A.3.2, we construct $\phi_{t}$, and in Section A.3.3, we prove that $\phi_{t}$ satisfies conditions (I) to (III).
A.3.1. Preliminaries.
In this section, we introduce the fundamental concepts and notations needed for defining $\phi_{t}$ when $A$ is not finite.
Definition A.23 (Well-Order).
A well-order on a set $X$ is a total order (i.e., for every $x,y\in X$, either $x<y$, $x>y$, or $x=y$) such that every non-empty subset of $X$ has a least element.
Since $A$ is countable, we can inject $A$ into $\mathbb{N}$. By listing elements of $A$ in the order given by their images under this injection, we get a well‐order $<_{A}$ on $A$. Note that for any non-empty subset $B\subseteq A$, the image of $B$ in $\mathbb{N}$ has a least element, whose preimage is the least element of $B$ under $<_{A}$. We will later use the following remark when constructing $\xi_{t}$ and $\phi_{t}$.
Remark A.24.
By the definition of $<_{A}$, each element of $A$ is either the least element of $A$ or has an immediate predecessor in $A$.
The query set $Q$ is not necessarily countable. However, by the well‐ordering theorem (which is equivalent to the axiom of choice), there exists a well‐order $<_{Q}$ on $Q$, regardless of its cardinality.
Definition A.25 (Lexicographic Order $<_{t}$ on $(Q\times A)^{t}$).
For $t\in\mathbb{N}\cup\{0\}$, we define an order $<_{t}$ on $(Q\times A)^{t}$ by lexicographic comparison: For any two distinct tuples $(q_{1},\dots,a_{t})$ and $(q_{1}^{\prime},\dots,a_{t}^{\prime})$ in $(Q\times A)^{t}$, let $i$ be the smallest index such that $(q_{i},a_{i})\neq(q_{i}^{\prime},a_{i}^{\prime})$. Then:
$(q_{1},\dots,a_{t})<_{t}(q_{1}^{\prime},\dots,a_{t}^{\prime})\text{ if and only if }\left[q_{i}<_{Q}q_{i}^{\prime}\right]\text{ or }\left[q_{i}=q_{i}^{\prime}\text{ and }a_{i}<_{A}a_{i}^{\prime}\right]$
Since $<_{A}$ and $<_{Q}$ are well-orders, $<_{t}$ is also a well-order: $<_{t}$ is clearly a total order. We will show that every non-empty subset $B\subseteq(Q\times A)^{t}$ has a least element. Define
$B_{1}^{Q}$ =$=\{q\in Q\mid\exists a_{1},\dots,a_{t}\in A,\exists q_{2},\dots,q_{t}\in Q\text{ such that }(q,a_{1},q_{2},\dots,a_{t})\in B\}\text{ and }$
$B_{1}^{A}$ =$=\{a\in A\mid\exists a_{2},\dots,a_{t}\in A,\exists q_{2},\dots,q_{t}\in Q\text{ such that }(q_{1}^{*},a,q_{2},\dots,a_{t})\in B\}.$
Since $B$ is non-empty, $B_{1}^{Q}$ is also non-empty. By the well-ordering of $Q$, $B_{1}^{Q}$ has a least element $q_{1}^{*}$. Similarly, since $B_{1}^{A}$ is non-empty, it has a least element $a_{1}^{*}$.
Again, since $B_{2}^{Q}$ is non-empty, it has a least element $q_{2}^{*}$. Repeating this process $t$ times, we obtain the least element $(q_{1}^{*},\dots,a_{t}^{*})$ of $B$.
Lemma A.26 (Transfinite Recursion).
Let $f$ be a function recursively defined over $(Q\times A)^{t}$ according to the well-order $<_{t}$. If for each sequence $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the value of $f(q_{1},\dots,a_{t})$ is uniquely determined by the values $f(q_{1}^{\prime},\dots,a_{t}^{\prime})$ for every $(q_{1}^{\prime},\dots,a_{t}^{\prime})<_{t}(q_{1},\dots,a_{t})$, then $f$ exists and is unique.
In later sections, we define functions $f$ recursively over $(Q\times A)^{t}$ using the well-order $<_{t}$. Specifically, we determine $f(q_{1},\dots,a_{t})$ based on the values $f(q_{1},\dots,a_{t-1},q_{t},a)$ for every $a<_{A}a_{t}$. By the definition of $<_{t}$, we have:
$(q_{1},\dots,a_{t-1},q_{t},a)<_{t}(q_{1},\dots,a_{t}).$
Finally, we define two key sets that will be used in our constructions: for any $a_{t}\in A$,
$\Gamma_{<a_{t}}=\{a\in A\mid a<_{A}a_{t}\},$
and
$\Gamma_{>a_{t}}=\{a\in A\mid a_{t}<_{A}a\}.$
A.3.2. Defining $\phi_{t}$ and showing its existence.
We define the function $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ recursively for each $(q_{1},\dots,a_{t})$ based on the $<_{t}$ order. For each $(q_{1},\dots,a_{t})$, we determine $\phi_{t}(q_{1},\dots,a_{t})$ by solving an optimization problem with the variables $\phi_{t}^{0}(q_{1},\dots,a_{t})$ and $\phi_{t}^{1}(q_{1},\dots,a_{t})$. The constraints for this problem involve the known functions $\mu_{t}$, $\phi_{t-1}$, $\mathrm{L}_{t}$, and the known values $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a)$ for every $a\in\Gamma_{<a_{t}}$.
We define $\phi_{t}^{0}(q_{1},\dots,a_{t})$ and $\phi_{t}^{1}(q_{1},\dots,a_{t})$ as the largest possible real numbers in $[0,1]$ satisfying the following constraints:
  1. ($\phi$1)
    For each $b\in\{0,1\}$,
    $\delta\phi_{t}^{b}(q_{1},\dots,a_{t})\geq\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t}).$
  2. ($\phi$2)
    For each $b\in\{0,1\}$,
    $\sum_{a\in\Gamma_{<a_{t}}}$ $\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)+\phi_{t}^{b}(q_{1},\dots,a_{t})+\sum_{a\in\Gamma_{>a_{t}}}{\frac{1}{\delta}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$
  3. ($\phi$3)
    For each $b\in\{0,1\}$,
    $\delta\phi_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\delta\phi_{t}^{1-b}(q_{1},\dots,a_{t})\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$
Although $\Gamma_{<a_{t}}$ is always finite, $\Gamma_{>a_{t}}$ could be infinite. Thus, to ensure ($\phi$2) is well-defined, we need to show:
Claim A.27.
For any $t\in[T]$, $b\in\{0,1\}$, and $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$, the summation $\sum_{a\in\Gamma_{>a_{t}}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ exists.
Proof.
By Lemma A.18,
$\sum\limits_{a\in\Gamma_{>a_{t}}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\sum\limits_{a\in A}\mu_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)=\mu_{t-1}^{b}(q_{1},\dots,a_{t-1})\leq 1.$
Moreover each summand $\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ is non-negative, implying that the summation is within the range $[0,1]$, and thus it exists. ∎
Claim A.28.
The function $\phi_{t}$ is well-defined and exists.
Proof.
The proof consists of two steps: (i) ensuring that the optimization problems are feasible, and (ii) applying the transfinite recursion theorem (Lemma A.26) and proving that $\phi_{t}$ exists.
Fix any $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$. By Remark A.24, each answer in $A$ is either the smallest element of $A$ or has an immediate predecessor. We will inductively show that for every $a_{t}\in A$, the values $\frac{1}{\delta}\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})$ and $\frac{1}{\delta}\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})$ satisfy the constraints of the optimization problem defining $\phi_{t}^{0}(q_{1},\dots,a_{t})$ and $\phi_{t}^{1}(q_{1},\dots,a_{t})$.
$\frac{1}{\delta}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t})$ trivially satisfies constraint ($\phi$1). Moreover, by Corollary A.17, it satisfies constraint ($\phi$3). To prove constraints ($\phi$2) is satisfied, we need to show that
(2) $\sum_{a\in\Gamma_{<a_{t}}}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ $+\sum_{a\in\Gamma_{>a_{t}}\cup\{a_{t}\}}\frac{1}{\delta}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}),$
We prove this for the base case (i.e., when $a_{t}$ is the smallest element in $A$) and the inductive step (i.e., when $a_{t}$ has a predecessor $a^{*}$) separately:
Base Case: Suppose $a_{t}$ is the least element of $A$. Then, $\Gamma_{<a_{t}}$ is empty, reducing our goal to showing that
$\sum_{\begin{subarray}{c}a\in A\end{subarray}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\delta\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
We know that $\phi_{t-1}(q_{1},\dots,a_{t-1})$ satisfies condition (I). Thus, by Lemma A.21,
$\delta\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$ $\geq\mathrm{L}_{t-1}^{b}(q_{1},\dots,a_{t-1})\geq\sum_{\begin{subarray}{c}a\in A\end{subarray}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a),$
finishing the proof for the base case.
Inductive Step: Let $a_{*}$ be the immediate predecessor of $a_{t}$. Since $\Gamma_{<a_{t}}=\Gamma_{<a^{*}}\cup\{a^{*}\}$, constraints ($\phi$2) for defining $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a^{*})$ is identical to (2). By inductive hypothesis, $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a^{*})$ exists and satisfies these constraints. Thus, inequality (2) holds, completing the induction.
Since $<_{t}$ is a well-order and the value of $\phi_{t}(q_{1},\dots,a_{t})$ depends only on preceding values in this order, by the transfinite recursion theorem, the function $\phi_{t}$ is well-defined and exists. ∎
A.3.3. Satisfying Conditions (I) to (III)
Our main objective is to construct $\phi_{t}^{b}$ so that it fulfills conditions (I), (II), and (III). Constraints ($\phi$1) and ($\phi$3) directly imply conditions (I) and (II), respectively. It remains to show $\phi_{t}$ satisfies conditions (III).
First, we claim for each $b\in\{0,1\}$ and every transcript $(q_{1},\dots,a_{t-1},q_{t})$, $\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ is at most $\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$: By constraint ($\phi$2) and non-negativity of $\mathrm{L}^{b}$, for every $a_{t}$,
$\sum_{a\in\Gamma_{<a_{t}}\cup\{a_{t}\}}\phi_{t}^{b}(q_{1},\dots,a_{t})\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
Thus, when $A$ is finite and $a^{\prime}$ is the maximum element of $A$, we have
$\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)=\sum_{\mathclap{a\in\Gamma_{<a^{\prime}}\cup\{a^{\prime}\}}}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a^{\prime})\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
Moreover, when $A$ is infinite and $a_{1}^{\prime}<_{A}a_{2}^{\prime}<_{A}\dots$ are the sorted elements of $A$, we have
$\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ $=\lim_{k\to\mathbb{N}}\sum_{a\in\Gamma_{<a_{k}^{\prime}}\cup\{a_{k}^{\prime}\}}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a_{k}^{\prime})\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
The rest of proof is identical to the proof given in [18]. We repeat it for completeness. For the sake of contradiction, suppose there exists $b\in\{0,1\}$ and $(q_{1},\dots,a_{t-1},q_{t})$ such that
(3) $\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)<\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
Fix $b$. As $\phi_{t}^{b}$ is lower bounded by $\mathrm{L}_{t}^{b}/\delta$, this implies that for every $a_{t}\in A$, in the optimization problem determining the value of $\phi_{t}(q_{1},\dots,a_{t})$, constraint ($\phi$2) for $b$ is not tight. Consequently, for every $a_{t}\in A$, we must have had
$\phi_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\phi_{t}^{1-b}(q_{1},\dots,a_{t})=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$
otherwise, the value of $\phi_{t}^{b}(q_{1},\dots,a_{t})$ could have increased, which contradicts the maximality of $\phi^{t}_{b}$.
Taking a sum over all $a_{t}\in A$, we get
(4) $\sum_{a_{t}\in A}\delta\phi_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\delta\sum_{a_{t}\in A}\phi_{t}^{1-b}(q_{1},\dots,a_{t})=\mu_{t-1}^{b}(q_{1},\dots,a_{t-1})-e^{-\epsilon}\mu_{t-1}^{1-b}(q_{1},\dots,a_{t-1})$
By (II) and the inductive hypothesis, we know that the RHS is at least
$\delta\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})-e^{-\epsilon}\delta\phi_{t-1}^{1-b}(q_{1},\dots,a_{t-1})$
Plugging this to (4) and reformulating it, we get
$\sum_{a_{t}\in A}\phi_{t}^{b}(q_{1},\dots,a_{t})$ $-\phi_{t}^{b}(q_{1},\dots,a_{t-1})\geq e^{-\epsilon}(\sum_{a_{t}\in A}\phi_{t}^{1-b}(q_{1},\dots,a_{t})-\phi_{t}^{1-b}(q_{1},\dots,a_{t-1}))$
By (3), the LHS is negative, implying that the RHS is negative too. Thus, $\sum_{a\in A}\phi_{t}^{1-b}(q_{1},\dots,a_{t-1},q_{t},a)<\phi_{t-1}^{1-b}(q_{1},\dots,a_{t-1})$. Hence, by an identical argument, we have
$\sum_{a_{t}\in A}\phi_{t}^{1-b}(q_{1},\dots,a_{t})$ $-\phi_{t}^{1-b}(q_{1},\dots,a_{t-1})\geq e^{-\epsilon}(\sum_{a_{t}\in A}\phi_{t}^{b}(q_{1},\dots,a_{t})-\phi_{t}^{b}(q_{1},\dots,a_{t-1})),$
which is a contradiction when $\epsilon>0$, which implies that condition (III) holds.
Appendix B Overview of [20]
For $\epsilon\geq 0$, let $\mathcal{M}:S_{\mathcal{M}}\times Q_{\mathcal{M}}\to S_{\mathcal{M}}\times A_{\mathcal{M}}$ be an $\epsilon$-DP IM w.r.t. a neighboring relation $\sim$. Suppose that $\mathcal{M}$ has discrete answer distributions. Vadhan and Wang [20] shows that for any two neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,0}(b)$ are identical. In this section, we overview the construction of $\mathcal{P}$ and show it satisfies the following property:
Claim B.1.
For any $t\in\mathbb{N}$ and any query sequence $(q_{1},\dots,q_{t})\in Q_{\mathcal{M}}^{t}$, if the distributions of the first $t$ answers of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to the queries $(q_{1},\dots,q_{t})$ are identical, then for every $b\in\{0,1\}$, when $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,0}(b)$ receives the queries $q_{1},\dots,q_{t}$, $\mathcal{P}$ generates answers itself, without interacting with $\mathrm{RR}_{\epsilon,0}(b)$.
Since $\delta=0$, $\mathrm{RR}_{\epsilon,0}(b)$ returns $(b,\bot)$ with probability $\frac{e^{\epsilon}}{1+e^{\epsilon}}$ and $(1-b,\bot)$ otherwise. For each $b$ and query sequence $(q_{1},\dots,q_{t})\in Q_{\mathcal{M}}^{t}$, let $\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{b})}$ denote the distribution of the first $t$ answers of $\mathcal{M}(s_{b})$ in response to $q_{1},\dots,q_{t}$. If $\epsilon=0$, then define $\nu_{q_{1},\dots,q_{t}}^{b}$ to be identical to $\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{b})}$. Otherwise, set
$\nu_{q_{1},\dots,q_{t}}^{b}=\frac{e^{\epsilon}\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{b})}-\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{1-b})}}{e^{\epsilon}-1}$
For $t\in\mathbb{N}$, $b\in\{0,1\}$, and $(q_{1},\dots,a_{t-1},q_{t})\in(Q_{\mathcal{M}}\times A_{\mathcal{M}})^{t-1}\times Q$, if $\nu_{q_{1},\dots,q_{t-1}}^{b}(q_{1},\dots,a_{t-1})>0$, the function $T_{q_{1},\dots,a_{t-1},q_{t}}:A_{\mathcal{M}}\to[0,1]$ is defined as
$T_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=\frac{\nu_{q_{1},\dots,q_{t}}^{b}(q_{1},\dots,a_{t})}{\nu_{q_{1},\dots,q_{t-1}}^{b}(q_{1},\dots,a_{t-1})}$
Vadhan and Wang [20] define $\mathcal{P}$ as follows: The initial state of $\mathcal{P}$ is $\left(\text{null},()\right)$, where null is an empty field and ()()( ) represents an empty sequence. Let $h$ denote the sequence. Let $(r,\bot)$ denote the outcome of $\mathrm{RR}_{\epsilon,0}(b)$. After interacting with $\mathrm{RR}_{\epsilon,0}(b)$, $\mathcal{P}$ updates null to $r$. Moreover, upon receiving a query $q$ and before returning an answer $a$, $\mathcal{P}$ appends $q$ and $a$ to the history sequence $h$. To generate a response $a$ to a query $q$, $\mathcal{P}$ samples $a$ from the distribution $T_{h,q}^{r}$, where $h$ is the history of all previous queries and answers.
We observe that as long as $T_{h,q}^{0}$ and $T_{h,q}^{1}$ are identical, $\mathcal{P}$ does not need to know the value of $r$. Thus, to defer interaction with $\mathrm{RR}_{\epsilon,0}(b)$, we define $\mathcal{P}$ to interact with $\mathrm{RR}_{\epsilon,0}(b)$ only when these distributions are different for the first time.
Proof of Claim B.1.
That the distributions $\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{0})}$ and $\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{1})}$ are identical implies that for every $k\leq t$, $\mu_{q_{1},\dots,q_{k}}^{\mathcal{M}(s_{0})}$ and $\mu_{q_{1},\dots,q_{k}}^{\mathcal{M}(s_{1})}$ are identical. Consequently, by definition, both distributions $\nu_{q_{1},\dots,q_{k}}^{0}$ and $\nu_{q_{1},\dots,q_{k}}^{1}$ are identical to $\mu_{q_{1},\dots,q_{k}}^{0}$. Therefore, the conditional distributions $T_{q_{1},\dots,a_{t-1},q_{t}}^{0}$ and $T_{q_{1},\dots,a_{t-1},q_{t}}^{1}$ used by $\mathcal{P}$ are also identical, implying that $\mathcal{P}$ does not interact with $\mathrm{RR}_{\epsilon,0}$ to generate the first $t$ answers. ∎

### v3

@@
 Previous concurrent composition theorems for differential privacy were only for the case when the dataset is static, with no adaptive updates.
-Moreover, we also give the first interactive and continual generalizations of the “parallel composition theorem” for noninteractive differential privacy.
+We also give the first interactive and continual generalizations of the “parallel composition theorem” for noninteractive differential privacy.
 Specifically, we show that the analogue of the noninteractive parallel composition theorem holds if either there are no adaptive dataset updates or each of the composed mechanisms satisfies pure differential privacy, but it fails to hold for composing approximately differentially private mechanisms with dataset updates.
+Thus, we prove a tight new composition theorem for this case.
+In addition, we prove concurrent filter compositions theorems for the scenarios in which the privacy parameters are adaptively chosen.
 We then formalize a set of general conditions on a continual mechanism $\mathcal{M}$ that runs multiple continual sub-mechanisms such that the privacy guarantees of $\mathcal{M}$ follow directly using the above concurrent composition theorems on the sub-mechanisms, without further privacy loss.
@@
 (1) $\forall S\subseteq\mathcal{Y}\qquad\Pr[\mathcal{M}(x)\in S]\leq e^{\epsilon}\cdot\Pr[\mathcal{M}(x^{\prime})\in S]+\delta.$
-[…]
+Typically we think of $\epsilon$ as a small constant, e.g.
+$\epsilon=.1$, while $\delta$ is cryptographically negligible.
+The case that $\delta=0$ is referred to as pure differential privacy, and otherwise Inequality 1 is known as approximate differential privacy.
+This simple definition is very useful and easy to work with, but does not capture many scenarios in which we want to apply differential privacy.
 1.1. Differential Privacy Over Time
@@
 Until recently, adaptivity has been treated in an ad hoc manner in the differential privacy literature.
-For example, the aforementioned advanced composition theorem of [9] was formalized in a way that allows an adversary to adaptively choose the mechanisms and pairs of adjacent datasets, just not the privacy-loss parameters, and new composition theorems that allow for adaptive choices of privacy-loss parameters were given by [19].
-Other early examples where adaptive queries were analyzed include the Sparse Vector Technique (SVT) [8, 11, 17] and the Private Multiplicative Weights algorithm [11].
+For example, the aforementioned advanced composition theorem of [9] was formalized in a way that allows an adversary to adaptively choose the mechanisms and pairs of adjacent datasets, just not the privacy-loss parameters, and new composition theorems that allow for adaptive choices of privacy-loss parameters (known as privacy filters and privacy odometers) were given by [21, 25].
+Other early examples where adaptive queries were analyzed include the Sparse Vector Technique (SVT) [8, 12, 18] and the Private Multiplicative Weights (PMW) algorithm [12].
 A few years ago, Vadhan and Wang [23] initiated a systematic study of differentially private mechanisms $\mathcal{M}$ that support adaptive queries.
-Specifically, they formalized the notion of an interactive mechanism, which begins with a dataset as input and maintains a potentially secret state as it answers queries that come in over time.
-(Both SVT and Private Multiplicative Weights utilize a secret state.) They defined $\mathcal{M}$ to be an $(\epsilon,\delta)$-DP interactive mechanism if for every pair of adjacent datasets $x\sim x^{\prime}$ and every adaptive adversary $\mathcal{A}$, the view of $\mathcal{A}$ when interacting with $\mathcal{M}(x)$ is $(\epsilon,\delta)$-indistinguishable (in the sense of Inequality (1)) from its view when interacting with $\mathcal{M}(x^{\prime})$, where the view of $\mathcal{A}$ consists of the random coin tosses of $\mathcal{A}$ and the transcript of the interaction.
+Specifically, they formalized the notion of an interactive mechanism $\mathcal{M}$, which begins with a dataset as input and maintains a potentially secret state as it answers queries that come in over time.
+The use of a secret state, as utilized in both SVT and PMW, allows for correlating the answers to queries over time, and often allows for the privacy-loss budget to degrade much more slowly than composing noninteractive mechanisms would allow.
+They defined $\mathcal{M}$ to be an $(\epsilon,\delta)$-DP interactive mechanism if for every pair of adjacent datasets $x\sim x^{\prime}$ and every adaptive adversary $\mathcal{A}$, the view of $\mathcal{A}$ when interacting with $\mathcal{M}(x)$ is $(\epsilon,\delta)$-indistinguishable (in the sense of Inequality (1)) from its view when interacting with $\mathcal{M}(x^{\prime})$, where the view of $\mathcal{A}$ consists of the random coin tosses of $\mathcal{A}$ and the transcript of the interaction.
 Vadhan and Wang [23] asked whether our existing composition theorems for differential privacy, where the mechanisms being composed are noninteractive mechanisms, extend to the concurrent composition of interactive mechanisms.
@@
 This is a realistic attack scenario when multiple interactive differentially private mechanisms are deployed in practice, and is also a useful primitive for analyzing differentially private algorithms that use several interactive mechanisms as building blocks.
+The challenge in analyzing concurrent composition comes from the fact that the adversary can base its queries to $\mathcal{M}_{1}$ on answers given by $\mathcal{M}_{2}$ and vice-versa.
+The privacy guarantees on an interactive mechanism $\mathcal{M}$ only promise privacy against adversaries that do not get additional information about the sensitive dataset in the middle of its interaction with $\mathcal{M}$.
+Reasoning about the effect of such interleaved interactions is particularly tricky when the mechanisms $\mathcal{M}_{1}$ and/or $\mathcal{M}_{2}$ are not pure DP and can maintain secret state (like in the Sparse Vector Technique).
+Thus, it took several years before we had concurrent composition theorems that matched our composition theorems for noninteractive mechanisms.
+Note that the extension of composition theorems for noninteractive mechanism to concurrent composition theorems for interactive mechanisms is non-trivial: Let us call an interactive mechanism $M$ $M$-adaptive if it is differentially private against an adversary that was only interactive with $M$.
+Consider an adversary that first interacts adaptively with $M_{1}$, then $M_{2}$ and then with $M_{1}$ again.
+Note that for the second sequence of interactions with $M_{1}$ the adversary takes into account its prior interaction with $M_{1}$ as well as with $M_{2}$.
+Thus, the adversary has more information about the data than if it had only interacted with $M_{1}$ before and the fact that $M_{1}$ is $M_{1}$-adaptive does not mean that it is adaptive against this interleaved composition of $M_{1}$ and $M_{2}$.
 Vadhan and Wang showed that the basic composition theorem for pure differential privacy [6] does extend to concurrent composition, but for approximate differential privacy, they obtained a much weaker bound for the concurrent case.
@@
 A series of works gave concurrent composition theorems for other flavors of differential privacy and composition, such as Rényi DP [19], $f$-DP [24], and adaptive choices of privacy parameters [11].
-All of the above works are for the case where the queries to the mechanisms are adaptive, but the dataset $x$ is static, given as input to the mechanism at the start of the interaction.
+Haney, Shoemate, Tian, Vadhan, Vyrros, Wang, and Xu [11] study the concurrent extension of composition theorems for non-interactive mechanisms where the adversary instantiates arbitrary mechanisms with adaptive privacy parameters.
+This setting is referred to as filter composition, where a filter is a function $\mathrm{Filt}$ that maps the privacy parameters $(\epsilon_{i},\delta_{i})$ of the created mechanisms to an element of $\{\top,\bot\}$.
+Suppose an adversary adaptively creates noninteractive mechanisms such that $g$ applied to the privacy parameters of the created mechanisms equals $\top$ at every point in time.
+The $\mathrm{Filt}$-filter composition of noninteractive mechanisms is said to satisfy $(\epsilon,\delta)$-DP if the view of every such adversary is $(\epsilon,\delta)$-indistinguishable when all created mechanisms are run on dataset $x$ or its neighboring dataset $x^{\prime}$.
+Haney et al.
+extend this definition to the concurrent $\mathrm{Filt}$-filter composition of interactive mechanisms, where the adversary may not only create new mechanisms but also interact with existing interactive ones.
+They prove that if the $\mathrm{Filt}$-filter composition of non-interactive mechanisms is $(\epsilon,\delta)$-DP, then the concurrent $\mathrm{Filt}$-filter composition of interactive mechanisms is also $(\epsilon,\delta)$-DP.
+All of the above works are for the case where the queries to the mechanisms are adaptive, but the datasets are static, given as input to the mechanisms at the start of the interaction.
+[^2] One can extend the concurrent composition theorems to the case where, similarly to the advanced composition theorem [9], an adversary can adaptively select a pair of adjacent datasets $(x_{i},x_{i}^{\prime})$ for the $i$’th mechanism $\mathcal{M}_{i}$, but once the interactive mechanism $\mathcal{M}_{i}$ is started, its dataset is fixed and no further updates are allowed.
+Thus, whenever $(x_{i+1},x_{i+1}^{\prime})$ are given, a new interactive mechanism $\mathcal{M}_{i+1}$ has to be started that does not know the secret state of $\mathcal{M}_{i}$.
+However, in some real-world applications, both queries and dataset updates may be issued adaptively, i.e., the dynamic updates to the dataset might be chosen adaptively.
+For instance, Ghazi et al.
+[10] study an advertising problem in which the server receives both private data and queries over time.
+They design algorithms that accept dataset updates while applying non-interactive mechanisms—such as the discrete Laplace mechanism—to answer queries and analyze the privacy guarantees for their algorithm using existing results on privacy filters for composing non-interactive mechanisms with adaptive privacy-loss parameters.
+This independent work [^3] The first version of this work was available on arXiv prior to [10].
+shows that if a mechanism receiving both queries and dataset updates produces a differentially private output at each time step $i$ (with adaptive parameters $\epsilon_{i},\delta_{i}$ at step $i$) such that $\sum_{i}\epsilon_{i}\leq\epsilon$ and $\sum_{i}\delta_{i}\leq\delta$, then the overall mechanism is $(\epsilon,\delta)$-DP.
+This guarantee can be seen as a special case of our filter composition theorem, where the composition rule is simple summation and each composed mechanism is restricted to a non-interactive mechanism corresponding to a single step.
+While some continual mechanisms do indeed employ independent non-interactive mechanisms at each step, many important examples—such as the Sparse Vector Technique (SVT)—generate outputs that are correlated across time.
+For these mechanisms, privacy must be analyzed over the entire output sequence rather than on a step-by-step basis to get the desired guarantees.
+(If the $\epsilon$-DP SVT mechanism were analyzed on a step-by-step basis, then the privacy guarantee would degrade to $\Theta(q\cdot\epsilon)$-DP, where $q$ is the number of queries issued, defeating the entire point of SVT, which is to have a privacy guarantee and noise scale that is independent of the number of queries.) As a consequence, in the concurrent composition of continual mechanisms (or even interactive mechanisms), one cannot simply reduce the analysis to the composition of per-step non-interactive mechanisms.
+Indeed, this is why concurrent composition theorems for interactive mechanisms took several years to establish [23, 19, 24, 11].
+Thus, the results of Ghazi et al.
+[10] do not apply to the more general setting of (filter) concurrent composition of continual mechanisms.
+[^4] It is Ghazi et al.’s definition of Individual DP (IDP, their Definition 4.1) and their composition theorem that is framed in terms of IDP that restricts to step-by-step privacy analyses corresponding to composition of noninteractive mechanisms.
+Their definition of $(\epsilon,\delta)$-DP interactive mechanisms (their Definition 3.6) is somewhat more general than their definition of IDP, but it still does not allow for exploiting correlated randomness across queries to reduce privacy loss, as in SVT.
+(This is because their Definition 3.4 requires the next state of the mechanism to be a function of only the dataset and the query, and not of the randomness used to generate the answer.
 Jain, Raskhodnikova, Sivakumar, and Smith [14] recently gave a formalization of adaptive adversaries under continual observation (where there are dataset updates but the queries are fixed) for the special case of event-level privacy, where two streams of updates are considered adjacent if they differ on only one update.
@@
 Rectifying this situation is what motivated our work.
+Comparison with algorithms and data structures.
+We note that the different flavors of mechanisms discussed above are differential privacy analogues of familiar variants of randomized algorithms and data structures.
+A non-interactive mechanism is simply a static (or batch) algorithm.
+An interactive mechanism corresponds to static data structure.
+Our focus in this paper is on continual mechanisms, which allow both queries and updates, and these can be viewed as dynamic data structures.
+Adaptivity is a common concern for the correctness properties of randomized data structures; here we are concerned instead with its effect on privacy properties.
 1.2. Our contributions
-In this paper, we:
-  • Give a formalism that captures privacy against adaptive adversaries for continual mechanisms, which support both queries and dataset updates.
+(A) We give a formalism that captures privacy against adaptive adversaries for continual mechanisms, which support both queries and dataset updates.
 This generalizes the previously studied cases of interactive mechanisms (which have static datasets) and continual observation (which answer fixed queries).
@@
 Figure 1. Comparison between interactive mechanism (IMs), continual observation (CO), and continual mechanisms (CMs)
-  • We prove concurrent composition theorems for continual mechanisms.
+(B) We prove concurrent composition theorems for continual mechanisms, where an adaptive adversary may interleave queries and updates across mechanisms based on the entire history of observed outputs.
+As discussed for the concurrent composition of interactive mechanisms, the concurrent setting involves additional adaptivity that is not supported by composition theorems for noninteractive mechanisms.
+Existing results on the concurrent composition of interactive mechanisms do not apply to the concurrent composition of continual mechanisms, since in our model the adversary adaptively selects both (a) the mechanisms receiving neighboring inputs (including their privacy parameters and neighboring definition) and (b) the neighboring update sequences they receive over time, which is not the case for interactive mechanisms.
+We show tight privacy bounds for the following useful forms of concurrent composition of continual mechanisms:
 (1) The first shows that, when the privacy-loss parameters are fixed, the standard composition theorems for noninteractive approximate differential privacy extend to the concurrent composition of continual mechanisms.
-(2) We extend the “parallel composition theorem” for noninteractive mechanisms to the concurrent parallel composition of a finite or infinite number of adaptively chosen interactive mechanisms with an adaptively chosen pair of neighboring datasets.
-(3) Then we extend this further to the concurrent composition of a finite or infinite number of adaptively chosen continual mechanisms when the composed mechanisms satisfy pure differential privacy.
-(4) We complement the latter with a counterexample for the case of approximate differential privacy.
-  • We formalize a set of general conditions on a continual mechanism $\mathcal{M}$ that runs multiple continual sub-mechanisms such that the privacy guarantees of $\mathcal{M}$ follow directly using the above concurrent composition theorems on the sub-mechanisms, without further privacy loss.
-  • We use our concurrent composition theorems for continual mechanisms, with the aforementioned general conditions, to give a simpler and more modular privacy analysis of a recent continual histogram mechanism [12].
+(2) We extend the classical “parallel composition theorem” for noninteractive mechanisms to the concurrent parallel composition of an unbounded number of adaptively chosen interactive mechanisms with an adaptively chosen pair of neighboring datasets.
+(3) We construct a counterexample demonstrating that the concurrent parallel composition of an unbounded number of continual mechanisms with adaptively chosen dataset updates is not private.
+Then imposing an upper bound on the sum of the privacy parameter $\delta_{i}$ of the created mechanisms, we introduce a new composition theorem for concurrent parallel composition with this upper bound.
+(4) As a special case, our new composition theorem extends the “parallel composition theorem” for noninteractive mechanisms to the concurrent composition of an unbounded number of adaptively chosen continual mechanisms when the composed mechanisms satisfy pure differential privacy.
+(C) We formalize a set of general conditions on a continual mechanism $\mathcal{M}$ that runs multiple continual sub-mechanisms such that the privacy guarantees of $\mathcal{M}$ follow directly using the above concurrent composition theorems on the sub-mechanisms, without further privacy loss.
+These conditions are also reformulated as a set of properties for pseudocode describing continual mechanisms.
+(D) We use our concurrent composition theorems for continual mechanisms, with the aforementioned general conditions, to give a simpler and more modular privacy analysis of a recent continual histogram mechanism [13].
 In the case of approximate DP, ours is the first proof that the privacy of the mechanism holds against adaptive adversaries.
+(E) We how that, for every filter $\mathrm{Filt}$, the concurrent $\mathrm{Filt}$-filter composition of continual mechanisms enjoys the same privacy guarantees as the $\mathrm{Filt}$-filter composition of non-interactive mechanisms.
 We elaborate on all of these contributions below.
@@
 Theorem 1.2 (concurrent composition of continual mechanisms, informally stated).
-For every fixed sequence $(\epsilon_{1},\delta_{1}),\ldots,(\epsilon_{k},\delta_{k})$ of privacy-loss parameters and verification functions $f_{1},\ldots,f_{k}$, the concurrent composition of continual mechanisms $\mathcal{M}_{i}$ that are $(\epsilon_{i},\delta_{i})$-DP wrt $f_{i}$ is $(\epsilon,\delta)$-DP for the same $(\epsilon,\delta)$ that holds for composing noninteractive mechanisms that are $(\epsilon_{i},\delta_{i})$-DP.
+For every fixed sequence $(\epsilon_{1},\delta_{1}),\ldots,(\epsilon_{k},\delta_{k})$ of privacy-loss parameters and verification functions $f_{1},\ldots,f_{k}$, the concurrent composition of continual mechanisms $\mathcal{M}_{i}$ that are $(\epsilon_{i},\delta_{i})$-DP w.r.t.
+$f_{i}$ is $(\epsilon,\delta)$-DP for the same $(\epsilon,\delta)$ that holds for composing noninteractive mechanisms that are $(\epsilon_{i},\delta_{i})$-DP.
 […]
@@
 Notably, this result allows the adversary to adaptively and concurrently choose the neighboring sequences fed to each of the $\mathcal{M}_{i}$’s based on the responses received from the other mechanisms.
-Theorem 1.2 is proven by using the aforementioned reduction from continual to interactive mechanism.
-Then we use Lyu’s result  [18] for $(\epsilon,\delta)$-DP interactive mechanisms and extend it to hold for infinite (but discrete) answer spaces and an unbounded number of rounds of interaction as follows.
-In Lyu’s proof the restriction on the number of interactions comes from a backward definition of a so-called control function.
-We define a function which is the limit over this control function and, using monotone convergence, show that this “limit function” satisfies all the necessary conditions to be used in the rest of Lyu’s proof.
-This extends the result to an unbounded number of rounds of interaction.
-Furthermore, Lyu’s proof used an algorithmic construction iterating over the answer set, implying that the answer set must be finite.
-Instead with a mathematical approach, we extend the proof to work with a countable, infinite answer set.
+A key step in the proof of Theorem 1.2 is to show that for any $(\epsilon_{i},\delta_{i})$-DP continual mechanism $\mathcal{M}$, the interactive mechanism $\mathcal{M}^{\prime}(\cdot)=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(\cdot)\!\circ^{*}\!\mathcal{M}$ can be simulated by post-processing the randomized response mechanism $\mathrm{RR}_{\epsilon_{i},\delta_{i}}(\cdot)$.
+Lyu [19] proves such a result under the assumptions that the answer space of $\mathcal{M}$ is finite and that there exists a predetermined upper bound $T$ on the number of interactions.
+We remove the latter assumption and relax the former to having an infinite (but discrete) answer space.
 While we focus on composing approximate DP with fixed privacy loss parameters $(\epsilon_{i},\delta_{i})$ in Theorem 1.2, we expect that there should also be analogues of it for other settings in which concurrent composition of interactive mechanisms have been studied, such as Rényi DP [19], $f$-DP [24], and adaptively chosen privacy-loss parameters [11].
@@
 We refer to this as concurrent $k$-sparse parallel composition.
+A more restricted version of this composition, which is limited to noninteractive mechanisms, was previously studied in [20].
 First, we extend the parallel composition of noninteractive mechanisms to concurrent composition of interactive mechanisms (where there are adaptive queries but datasets $x_{i}$ are static, rather than dynamically updated).
 Theorem 1.3 (concurrent parallel composition of interactive mechanisms).
-For every $\epsilon_{0},\delta_{0}\geq 0$ and $k\in\mathbb{N}$, and every $(\epsilon_{0},\delta_{0})$-DP interactive mechanism $\mathcal{M}_{0}$, the concurrent $k$-sparse parallel composition of an unbounded number of copies of $\mathcal{M}_{0}$ satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},\delta_{0})$-DP mechanisms.
-As discussed later in the introduction, we also prove a generalization of Theorem 1.3 for the case of composing interactive mechanisms $\mathcal{M}_{i}$ with differing $(\epsilon_{i},\delta_{i})$ parameters.
+For every $\epsilon_{0}>0$ and $0\leq\delta_{0}\leq 1$ and $k\in\mathbb{N}$, and every $(\epsilon_{0},\delta_{0})$-DP interactive mechanism $\mathcal{M}_{0}$, the concurrent $k$-sparse parallel composition of an unbounded number of copies of $\mathcal{M}_{0}$ satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},\delta_{0})$-DP mechanisms.
+As discussed later in the paper, we also prove a generalization of Theorem 1.3 for the case of composing interactive mechanisms $\mathcal{M}_{i}$ with differing $(\epsilon_{i},\delta_{i})$ parameters.
 Then we turn to the case of continual mechanisms, where there can also be adaptive dataset updates. We prove a negative result for approximate DP ($\delta_{0}>0$):
-Theorem 1.4 (concurrent parallel composition of approximate DP continual mechanisms, informally stated).
+Theorem 1.4 (counterexample for the concurrent parallel composition of approximate DP continual mechanisms, informally stated).
 For every $\delta_{0}>0$, there is an $(0,\delta_{0})$-DP continual mechanism $\mathcal{M}_{0}$ such that for every $\ell\in\mathbb{N}$, the concurrent 1-sparse parallel composition of $\ell$ copies of $\mathcal{M}_{0}$ is not $(0,\delta)$-DP for any $\delta<1-(1-\delta_{0})^{\ell}$.
 This theorem says that, at least in terms of the $\delta$ parameter, concurrent parallel composition of continual mechanisms does not provide any benefit over non-parallel composition, since the bound of $1-(1-\delta_{0})^{\ell}$, is the bound that we would get by just applying composition over all $\ell$ mechanisms.
-However, we prove that if $\delta_{0}=0$, concurrent parallel composition behaves just like non-interactive parallel composition:
-Theorem 1.5 (concurrent parallel composition of pure DP continual mechanisms, informally stated).
-For every $\epsilon_{0}\geq 0$ and integers $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},0)$-DP mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms.
-We note that the constraint of pure DP is only on the mechanisms being composed; we allow $\delta>0$ in the final privacy guarantee, so we can still, for example, recover the advanced composition bound applied to pure DP mechanisms.
-Like Theorem 1.3, we also prove a generalization of Theorem 1.5 for the case of composing mechanisms $\mathcal{M}_{i}$ with differing $(\epsilon_{i},\delta_{i})$ parameters.
-To prove this theorem, we add a twist to the concurrent composition theorems for interactive mechanisms.
-Specifically, the papers [20, 18] proved their composition theorems for pure and approximate DP by showing that for every $(\epsilon_{0},\delta_{0})$-DP interactive mechanism $\mathcal{M}$ and pair $x_{0}\sim x_{1}$ of adjacent datasets, $\mathcal{M}$ can be simulated as an interactive post-processing of the noninteractive $(\epsilon_{0},\delta_{0})$-DP randomized response mechanism $\mathrm{RR}_{\epsilon_{0},\delta_{0}}$.
-That is, there is an IPM $\mathcal{T}$ (with no input) such that for each $b\in\{0,1\}$, $\mathcal{M}(x_{b})$ is equivalent to $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon_{0},\delta_{0}}(b)$.
-We show that when $\delta_{0}=0$, the use of $\mathrm{RR}_{\epsilon_{0},\delta_{0}}$ by $\mathcal{T}$ can be delayed until the time that it receives a query that would have different answer distributions under $\mathcal{M}(x_{0})$ and $\mathcal{M}(x_{1})$.
-While intuitive, Theorem 1.4 tells us that this fact fails when $\delta_{0}>0$.
-It is an interesting open problem to see if Theorem 1.5 can be extended to composing approximate DP mechanisms ($\delta_{0}>0$), if we allow an additional slackness in the final $\delta$, as shown necessary by Theorem 1.4.
-An Algorithmic Application.
+Inspired by this counterexample, we prove a weaker version of Theorem 1.3 for the concurrent parallel composition of continual mechanisms.
+Theorem 1.5 (concurrent parallel composition of approx DP continual mechanisms, informally stated).
+For every $\epsilon_{0}>0$, $0\leq\delta^{\prime}\leq 1$, and $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},\delta_{i})$-DP continual mechanisms, where the parameters $\delta_{i}$ are chosen adaptively such that $\sum_{i}\delta_{i}\leq\delta^{\prime}$, satisfies $(\epsilon,\delta+\delta^{\prime})$-DP if the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms is $(\epsilon,\delta)$-DP.
+We note that the formal version of this result applies to $(\epsilon_{i},\delta_{i})$-DP continual mechanisms with possibly different $(\epsilon_{i},\delta_{i})$ parameters.
+The proof of Theorem 1.5 relies on a key property of our new post-processing construction (recall that $\mathrm{RR}_{\epsilon,\delta}(b)$ denotes the randomized response mechanism with parameters $\epsilon$ and $\delta$ and input $b\in\{0,1\}$):
+Lemma 1.6 (post-processing of randomized response mechanisms, informal statement).
+For every $(\epsilon,\delta)$-DP interactive mechanism $\mathcal{M}$ and every pair of neighboring datasets $x_{0},x_{1}$, there exists an interactive post-processing mechanism $\mathcal{P}$ that receives the outputs of $\mathrm{RR}_{\epsilon,0}(b)$ and $\mathrm{RR}_{0,\delta}(b)$ and simulates $\mathcal{M}(x_{b})$ identically.
+As long as the answer distributions of $\mathcal{M}(x_{0})$ and $\mathcal{M}(x_{1})$ to the queries asked so far are identical, $\mathcal{P}$ does not access the outcome of $\mathrm{RR}_{\epsilon,0}(b)$.
+Setting $\delta^{\prime}=0$, Theorem 1.5 implies:
+Corollary 1.7 (concurrent parallel composition of pure DP continual mechanisms, informally stated).
+For every $\epsilon_{0}>0$ and integer $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},0)$-DP mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms.
+Like Theorem 1.5, we also prove a generalization of Corollary 1.7 for the case of composing mechanisms $\mathcal{M}_{i}$ with differing $\epsilon_{i}$ parameters.
+Concurrent filter composition theorems for continual mechanisms.
+Let $\mathrm{Filt}:\left([0,\infty)\times[0,1]\right)^{*}\to\{\top,\bot\}$ be a filter.
+Consider adversaries that create continual mechanisms with adaptive privacy parameters such that $\mathrm{Filt}(\sigma)=\top$ at every step, where $\sigma$ is the sequence of privacy parameters $(\epsilon_{i},\delta_{i})$ of the created mechanisms $\mathcal{M}_{i}$ up to that point.
+Suppose each $\mathcal{M}_{i}$ is $(\epsilon_{i},\delta_{i})$-DP with respect to a verification function $f_{i}$.
+In addition to creating new mechanisms, these adversaries may adaptively generate pairs of inputs for existing mechanisms such that $f_{i}$ applied to the sequence of input pairs for each $\mathcal{M}_{i}$ evaluates to $\top$.
+The concurrent $\mathrm{Filt}$-filter composition of continual mechanisms is said to be $(\epsilon,\delta)$-DP if the view of such adversaries is $(\epsilon,\delta)$-indistinguishable when all mechanisms receive either the first inputs from their respective pairs or all receive the second ones.
+Theorem 1.8 (Concurrent filter composition for continual mechanisms).
+Let $\mathrm{Filt}:\left([0,\infty)\times[0,1]\right)^{*}\to\{\top,\bot\}$ be a filter.
+The concurrent $\mathrm{Filt}$-filter composition of continual mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantee as the $\mathrm{Filt}$-filter composition of non-interactive mechanisms.
+General Conditions for Privacy Analysis of Continual Mechanisms.
 Composition theorems for differential privacy have two main uses.
@@
 The concurrent composition theorems for interactive mechanisms have already found an application in the privacy analysis of matching algorithms in local edge differential privacy model [4].
-(This work also has a result in the continual observation model, but the privacy is only proven for oblivious adversaries.)
-Our concurrent composition theorem allows for a simpler and more modular privacy analysis of continual mechanisms built out of other continual mechanisms.
-In particular, we give a new privacy proof for the continual monotone histogram mechanism of Henzinger, Sricharan, and Steiner [12].
-This mechanism involves concurrent executions of a simpler continual histogram mechanism and multiple instantations of the Sparse Vector Technique mechanism, and the original privacy analysis was rather intricate even for oblivious adversaries (because the mechanism itself is adaptive in its queries and dataset updates to the building-block mechanisms).
-With our concurrent composition theorems (both Theorem 1.2 and Theorem 1.5), we obtain a much simpler privacy proof.
-Moreover, for the case of approximate DP, the original proof was only for oblivious adversaries, whereas our proof also applies to adaptive adversaries.
+(This work also has a result in the continual observation model, but the privacy is only proven for oblivious adversaries, which allows it to be proven without using a composition theorem for continual mechanisms.) Our concurrent composition theorem allows for a simpler and more modular privacy analysis of continual mechanisms built out of other continual mechanisms.
+There are various applications, where a continual mechanism executes multiple differentially private mechanisms, such as the Sparse Vector Technique (SVT) and continual counters, as sub-mechanisms and concurrently interacts with them.
+Let $\mathcal{M}$ denote such a composite mechanism.
+Upon receiving an input, $\mathcal{M}$ may create new sub-mechanisms, interact multiple times with the existing ones, and then return an output.
+$\mathcal{M}$ generates inputs for its sub-mechanisms based on its own inputs and the previous outputs of all sub-mechanisms.(See Figure 4 for an illustration.) Analyzing the privacy of such mechanism designs is challenging because the creation and inputs of sub-mechanisms are interdependent and adaptive.
+Even if the inputs of $\mathcal{M}$ are non-adaptive, the inputs to its sub-mechanisms may become adaptive: Let $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ denote the sub-mechanisms of $\mathcal{M}$.
+If the inputs of $\mathcal{M}_{1}$ depend on the outputs of $\mathcal{M}_{2}$ and the inputs of $\mathcal{M}_{2}$ depend on the outputs of $\mathcal{M}_{1}$, then $\mathcal{M}_{1}$ receives inputs correlated with its own past outputs.
+Later, we see an example of this situation, where a histogram mechanism creates $d$ private continual counters at the initialization and an adaptive number of SVTs and Laplace mechanisms over time.
+In that design, the inputs of the counters depend on the output of the (active instance of the) SVT and vice versa.
+In such modular designs, we model $\mathcal{M}$ as the post-processing of a specific concurrent composition of its (continual) sub-mechanisms.
+Since $\mathcal{M}$ accesses the raw (secret) data itself, it could in principle incur additional privacy loss beyond that of the concurrent composition of the sub-mechanisms.
+To rule out such privacy loss, we introduce a set of conditions on $\mathcal{M}$ and prove that, under these conditions, the privacy guarantee of $\mathcal{M}$ is exactly that of the concurrent composition of its sub-mechanisms.
+Intuitively, $\mathcal{M}$ must restrict its use of raw inputs solely to generating inputs for differentially private sub-mechanisms, while all other actions, such as creating new sub-mechanisms or determining output values, must depend only on the privatized outputs of the sub-mechanisms.
+This will ensure that $\mathcal{M}$ does not use any “privacy budget” in addition to the “privacy budget” used by the sub-mechanisms.
+Continual mechanisms are typically presented using pseudocode (rather than as formal state machines exchanging messages, which are used by our conditions on $\mathcal{M}$), where the pseudocode receives an input, updates its variables, and returns an output.
+We therefore reformulate our conditions in this setting.
+We call a variable a decision variable if it is either the output of a private sub-mechanism or computed solely based on other decision variables.
+Our conditions are now as follows:
+  • Destination condition: Decisions about whether to produce an output, instantiate a new sub-mechanism (and of which type), or interact with an existing one (and which one) must be based only on decision variables.
+  • Response condition: The value of $\mathcal{M}$’s output must be determined solely by decision variables.
+  • Mapping condition: Fixing the values of decision variables, neighboring inputs to $\mathcal{M}$ must be mapped to input sequences for the sub-mechanisms with a structure for which concurrent composition privacy is private (This condition is referred to as $f\to f^{\prime}$ property in Section 6).
+We show that if these conditions are satisfied, then the privacy guarantee of $\mathcal{M}$ is the same as that of the underlying concurrent composition.
+For example, if $\mathcal{M}$ executes $k$ continual sub-mechanisms and ensures that its neighboring inputs yield neighboring input sequences for each sub-mechanism, then the privacy guarantee of $\mathcal{M}$ matches that of the concurrent composition of its sub-mechanisms as given in Theorem 1.2.
+The same conclusion holds for the concurrent $k$-sparse parallel composition of sub-mechanisms, as established in Theorem 1.3 and Corollary 1.7.
 […]
 Figure 4. Illustration of the interactions between an honest adversary $\mathcal{A}$ and a mechanism with $k$ sub-mechanisms
-General Conditions for Privacy Analysis of Continual Mechanisms.
-In the above application, the adversary does not directly interact with the sub-mechanisms, but rather it is the “main” HSS mechanism $\mathcal{M}$ that issues messages to its sub-mechanisms, based on the messages it receives.
-(See Figure 4 for an illustration.) These messages are not necessarily identical to the messages that $\mathcal{M}$ receives.
-Instead based on the input that $\mathcal{M}$ receives from an analyst or an adversary (passed through $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(\cdot)$), it generates an input sequence for its sub-mechanisms $\mathcal{M}_{i}$ that, on neighboring inputs to $\mathcal{M}$, must be identical or neighboring for their respective verification functions $f_{i}$.
-Note that $\mathcal{M}$ may fix the set of sub-mechanisms at initialization or select these mechanisms adaptively over time.
-For example, in the application above, the continual monotone histogram mechanism maintains $d$ fixed binary counters while adaptively instantiating an unbounded number of SVT and Laplace mechanisms as need.
-The privacy analysis of such mechanisms $\mathcal{M}$ is challenging as the creation of sub-mechanisms and their inputs is adaptive and dependent on each other.
-Moreover, in contrast to concurrent composition of interactive mechanisms, $\mathcal{M}$ has direct access to the dataset updates, which cannot be revealed without differentially private processing (which needs to be done by the sub-mechanisms).
-We give a set of conditions for $\mathcal{M}$ and prove a theorem that shows that under these conditions the privacy of $\mathcal{M}$ can directly be computed by applying our concurrent composition theorems to the continual sub-mechanisms.
-The general idea is that $\mathcal{M}$ should only use its input sequence to generate inputs for the differentially private sub-mechanisms and any other action of $\mathcal{M}$, such as creating new sub-mechanism or generating a response, should be taken based only on the privatized outputs of the sub-mechanisms.
-This will ensure that $\mathcal{M}$ does not use any “privacy budget” in addition to the privacy budget used by the sub-mechanisms.
-For example, if $\mathcal{M}$ executes $k$ continual sub-mechanisms and, upon receiving neighboring inputs, issues neighboring message sequences for the sub-mechanisms, then the privacy of $\mathcal{M}$ equals the privacy of the concurrent composition of the sub-mechanism, as computed by Theorem 1.2.
-The same holds for the concurrent $k$-sparse parallel composition of sub-mechanisms, as computed in Theorems 1.3 and  1.5.
+An Algorithmic Application.
+We give a new privacy proof for the continual monotone histogram mechanism of Henzinger, Sricharan, and Steiner (HSS) [13].
+This mechanism involves concurrent executions of a simpler continual histogram mechanism and multiple instantiations of the Sparse Vector Technique mechanism, and the original privacy analysis was rather intricate even for oblivious adversaries (because the mechanism itself is adaptive in its queries and dataset updates to the building-block mechanisms).
+Combining our concurrent composition theorems (both Theorem 1.2 and Corollary 1.7), we obtain a much simpler privacy proof.
+Moreover, for the case of approximate DP, the original proof was only for oblivious adversaries, whereas our proof also applies to adaptive adversaries.
 Our Variants of Concurrent Composition.
 Here we elaborate on the different types of concurrent composition that we study in this paper, with more general and more precise formulations than described above.
-  1. (1)
-    Concurrent Composition of Continual Mechanisms: In the concurrent composition of fixed continual mechanism, $k$ continual mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ are fixed at the beginning, and an adversary adaptively issues pairs of messages for them such that for each $\mathcal{M}_{i}$, the sequences formed by the first messages of all message pairs and the sequence formed by the second messages of all message pairs are neighboring.
+(1) Concurrent Composition of Continual Mechanisms: In the concurrent composition of fixed continual mechanism, $k$ continual mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ are fixed at the beginning, and an adversary adaptively issues pairs of messages for them such that for each $\mathcal{M}_{i}$, the sequences formed by the first messages of all message pairs and the sequence formed by the second messages of all message pairs are neighboring.
 A secret bit $b\in\{0,1\}$ determines whether each mechanism receives the first or second sequence from each pair.
@@
 The privacy guarantee for this variant is the same as Theorem 4.9 and is stated in Theorem 4.11.
-  2. (2)
-    Concurrent $k$-Sparse Parallel Composition of Continual Mechanisms: This composition generalizes the concurrent composition of continual mechanisms with $k$ fixed parameters by permitting the creation of an arbitrary number of continual mechanisms.
+(2) Concurrent $k$-Sparse Parallel Composition of Continual Mechanisms: This composition generalizes the concurrent composition of continual mechanisms with $k$ fixed parameters by permitting the creation of an arbitrary number of continual mechanisms.
 As before, $k$ privacy parameters $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed at the beginning.
@@
 For each $1\leq i\leq k$, the mechanism $\mathcal{M}_{i}$ must be $(\epsilon_{i},\delta_{i})$-DP, and the sequences of the first and second messages for $\mathcal{M}_{i}$ must be neighboring.
+We note the choice of mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ is made adaptively over time.
+Specifically, the adversary decides whether to issue a pair of differing messages for a mechanism based on the previous answers of all mechanisms, including the outputs of that mechanism.
 As in previous compositions, a secret bit $b$ determines whether each mechanism receives the first or second message from each pair, and the adversary seeks to guess $b$ based on the outputs.
-Theorem 1.5 analyzes this composition in the special case where $\delta_{1}=\dots=\delta_{k}=0$ and $\epsilon_{1}=\dots=\epsilon_{k}=\epsilon_{0}$.
-The general version for non-identical $\epsilon_{i}$ parameters is given in Theorem 5.2.
-  3. (3)
-    Concurrent $k$-Sparse Parallel Composition of Interactive Mechanisms: In this composition, the adversary creates an unbounded number of interactive mechanisms over time, instead of continual ones.
+Let $\mathcal{M}_{j}^{\prime}$ denote the $j$-th created mechanism, and let $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ be its privacy parameter.
+Theorem 5.4, informally stated in Theorem 1.5, analyzes the privacy of the concurrent $k$-sparse parallel composition of continual mechanisms when $\sum\delta_{j}^{\prime}$ is upper bounded by a predetermined value.
+Corollary 5.5, informally stated in Corollary 1.7, sets this upper bound to 0 and provides privacy guarantees for the concurrent $k$-sparse parallel composition of purely differentially private continual mechanisms.
+(3) Concurrent $k$-Sparse Parallel Composition of Interactive Mechanisms: In this composition, the adversary creates an unbounded number of interactive mechanisms over time, instead of continual ones.
 As before, $k$ privacy parameter pairs $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed in advance.
@@
 In Section 6, we state the conditions and the theorem for analyzing the privacy of complex continual mechanisms that use non-interactive, interactive and continual sub-mechanisms.
-Finally, in Section 7, we apply our theorem to analyze the privacy of the continual histogram mechanism of [12].
-Appendix A presents our extension of Lyu’s result [18] to infinite answer spaces and an unbounded number of rounds of interaction.
+In Section 7, we apply our theorem to analyze the privacy of the continual histogram mechanism of [13].
+Finally, in Section 8 and Section 9, we prove two technical intermediate lemmas, required for the previous results.
 2. Preliminaries
@@
 For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, the approx randomized response $\mathrm{RR}_{\epsilon,\delta}$ is a non-interactive mechanism that takes a bit $b\in\{0,1\}$ as input and returns a pair according to the following distribution:
-$\mathrm{RR}_{\epsilon,\delta}(b)=\begin{cases}(b,\top)&\quad\text{w.p.
-$\delta$}\\ (b,\bot)&\quad\text{w.p.
-$(1-\delta)\frac{e^{\epsilon}}{1+e^{\epsilon}}$}\\ (1-b,\bot)&\quad\text{w.p.
+$\mathrm{RR}_{\epsilon,\delta}(b)=\begin{cases}(\top,b)&\quad\text{w.p.
+$\delta$}\\ (\bot,b)&\quad\text{w.p.
+$(1-\delta)\frac{e^{\epsilon}}{1+e^{\epsilon}}$}\\ (\bot,1-b)&\quad\text{w.p.
 $(1-\delta)\frac{1}{1+e^{\epsilon}}$}\end{cases}$
@@
 A mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-DP w.r.t.
-a neighbor relation $\sim$ if and only if for any $x_{0}\sim x_{1}$, there exists a randomized ”post-processor” $\mathcal{P}:\{0,1\}\times\{\top,\bot\}\to\mathcal{Y}$ such that for both $b\in\{0,1\}$, $\mathcal{M}(x_{b})$ is identically distributed to $\mathcal{P}(\mathrm{RR}_{\epsilon,\delta})$.
+a neighbor relation $\sim$ if and only if for any $x_{0}\sim x_{1}$, there exists a randomized ”post-processor” $\mathcal{P}:\mathcal{Y}\to\{0,1\}\times\{\top,\bot\}$ such that for both $b\in\{0,1\}$, $\mathcal{M}(x_{b})$ is identically distributed to $\mathcal{P}(\mathrm{RR}_{\epsilon,\delta}(b))$.
 2.2. Interactive Mechanisms (IM)
@@
 If $S_{\mathcal{M}}^{\text{init}}$ is a singleton $\{s\}$, we denote $\mathcal{M}(s)$ as $\mathcal{M}$.
+An example of an IM is the interactive randomized response mechanism, denoted $\mathrm{IRR}_{\epsilon,\delta}$, which will repeatedly be used in our proofs.
+Ghazi et al.
+[10] define this mechanism to initially flip a coin and determine whether to answer all subsequent queries truthfully or in a privacy-preserving manner.
+Our definition of interactive randomized response mechanism is different but shares the same intuition.
+Recall that the standard randomized response mechanism $\mathrm{RR}_{\epsilon,\delta}$ is an NIM that outputs a pair from the space $\{\top,\bot\}\times\{0,1\}$.
+In contrast, we define the interactive variant $\mathrm{IRR}_{\epsilon,\delta}$ to output the pair in two separate steps: it first produces an element from $\{\top,\bot\}$, and later returns a bit in $\{0,1\}$, such that the joint distribution of the two outputs matches that of $\mathrm{RR}_{\epsilon,\delta}$.
+Specifically, $\mathrm{IRR}_{\epsilon,\delta}$ is an IM with a single query message $q^{*}$.
+Given an initial state $b\in\{0,1\}$, $\mathrm{IRR}_{\epsilon,\delta}(b)$ responds to the first query $q^{*}$ with an element in $\{\top,\bot\}$ and returns a bit in $\{0,1\}$ when receiving $q^{*}$ for the second time.
+$\mathrm{IRR}_{\epsilon,\delta}(b)$ halts the communication upon receiving $q^{*}$ for the third time.
+The formal definition is given below.
+Definition 2.9 ($\mathrm{IRR}_{\epsilon,\delta}$).
+For any $\epsilon\geq 0$ and $0\leq\delta\leq 1$, the interactive randomized response mechanism $\mathrm{IRR}_{\epsilon,\delta}$ is an IM with state space $S=\{\text{null}\}\cup\{0,1\}\cup\{\top,\bot\}\times\{0,1\}$, initial state space $S^{\text{init}}=\{0,1\}$, and query space $Q=\{q^{*}\}$.
+$\mathrm{IRR}_{\epsilon,\delta}$ is represented by a randomized function that maps a state $s\in S$ and query $q^{*}\in Q$ to a new state $s^{\prime}\in S$ and a message $m$ as follows:
+  • If $s$ is a bit $b\in\{0,1\}$ (i.e., $q^{*}$ is received for the first time), the output $m$ is set to $\bot$ with probability $\delta$ and to $\top$ with probability $1-\delta$.
+The new state is then set to $s^{\prime}=(m,b)$.
+  • If $s$ is a pair $(\tau,b)$ in $\{\top,\bot\}\times\{0,1\}$ (i.e., $q^{*}$ is received for the first time), $s^{\prime}$ is set to null and $m$ is determined as follows:
+    • If $\tau=\bot$ (i.e., the first response was $\bot$), then $m=b$ with probability $1$.
+    • If $\tau=\top$ (i.e., the first response was $\top$), $m=b$ with probability $\frac{e^{\epsilon}}{1+e^{\epsilon}}$ and m=1bm=1-b with probability $\frac{1}{1+e^{\epsilon}}$.
 2.2.1. Interaction of IMs.
@@
 Let $\mathcal{M}$ be an IM $\mathcal{M}$ with an initial state space $S_{\mathcal{M}}^{\text{init}}$ and $\mathcal{P}$ be an IPM with a singleton initial state space $S_{\mathcal{P}}^{\text{init}}=\{s\}$.
-Let $\sim$ be a neighbor relation on $S_{\mathcal{M}}^{\text{init}}$.
-Define a binary relation $\sim_{*}$ on $S_{\mathcal{P}\!\circ^{*}\!\mathcal{M}}^{\text{init}}$ such that two initial states $(s_{\mathcal{M}}^{1},s)$ and $(s_{\mathcal{M}}^{2},s)$ satisfy $(s_{\mathcal{M}}^{1},s)\sim_{*}(s_{\mathcal{M}}^{2},s)$ if and only if $s_{\mathcal{M}}^{1}\sim s_{\mathcal{M}}^{2}$.
-If $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t.
-the neighbor relation $\sim$ against adaptive adversaries, then $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ is also $(\epsilon,\delta)$-DP w.r.t.
-the neighbor relation $\sim_{*}$ against adaptive adversaries.
+For every pair of initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$, the following holds: If for every adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s_{0}))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}(s_{1}))$ are $(\epsilon,\delta)$-indistinguishable, then for every adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{P}\!\circ^{*}\!\mathcal{M}(s_{0}))$ and $\mathrm{View}(\mathcal{A},\mathcal{P}\!\circ^{*}\!\mathcal{M}(s_{1}))$ are $(\epsilon,\delta)$-indistinguishable.
 [^5] In [11], the definitions of IMs and IPMs are more restrictive.
@@
 By Definition 2.13, an interactive mechanism $\mathcal{M}$ with initial state space $S_{\mathcal{M}}^{\text{init}}$ is $(\epsilon,\delta)$-DP w.r.t.
-a neighboring relation $\sim$, if for any two initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$ such that $s_{0}\sim s_{1}$ and any adversary $\mathcal{A}$, the view of $\mathcal{A}$ in the interaction with $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ is $(\epsilon,\delta)$-indistinguishable.
+a neighboring relation $\sim$, if for every two initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$ such that $s_{0}\sim s_{1}$ and every adversary $\mathcal{A}$, the view of $\mathcal{A}$ in the interaction with $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ is $(\epsilon,\delta)$-indistinguishable.
 This means that for the worst-case (i.e., adversarial) choice of neighboring initial states, an adversary by adaptively asking queries should not be able to determine whether $\mathcal{M}$ is running on $s_{0}$ or $s_{1}$.
@@
 Specifically, for $b\in\{0,1\}$, in an interaction between an adversary $\mathcal{A}$ and $\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$, upon receiving a message $(m_{0},m_{1})$ from $\mathcal{A}$, $\mathcal{I}(b)$ forwards $m_{b}$ to $\mathcal{M}$ and returns $\mathcal{M}$’s answer unchanged to $\mathcal{A}$.
-With this formulation, an interactive mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP if for any adversary of the above form, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}[0]\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}[1]\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
+With this formulation, an interactive mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP if for every adversary of the above form, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}[0]\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}[1]\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
 Actually, we will enforce the restrictions on the adversary also by an IPM, which we call a verifier $\mathcal{V}$.
@@
 [7] define neighboring as a binary relation on sequences of data and empty records: two sequences are neighboring if they are identical except for the presence or absence of a record in a single step.
-They called a mechanism $(\epsilon,\delta)$-DP under continual observation if the distributions of answer sequences are $(\epsilon,\delta)$-indistinguishable for any two neighboring input sequences.
+They called a mechanism $(\epsilon,\delta)$-DP under continual observation if the distributions of answer sequences are $(\epsilon,\delta)$-indistinguishable for every two neighboring input sequences.
 This definition does not allow for data sequences that are adaptively chosen based on the answers released.
@@
 Instead of having a challenge step, we can alternatively restrict to adversaries that send pairs of identical records at every step except for one.
-Under this formulation, the definition of differential privacy proposed by [13] can be restated as follows: A mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP differential privacy against adaptive adversaries if for every such adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}[0]\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}[1]\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable, where $\mathcal{I}$ is the interactive post-processing mechanism discussed previously.
+Under this formulation, the definition of differential privacy proposed by [14] can be restated as follows: A mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP differential privacy against adaptive adversaries if for any such adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}[0]\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}[1]\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable, where $\mathcal{I}$ is the interactive post-processing mechanism discussed previously.
 Again, below we will enforce the restriction on the adversary via another IPM $\mathcal{V}$.
@@
 To the best of our knowledge, all mechanisms in DP literature can be represented as CMs:
-  • Any non-interactive mechanism $\mathcal{N}$, such as the Laplace mechanism, can be modeled as a CM that receives a dataset $x$ as its first message and returns $\mathcal{N}(x)$ as a response.
+  • Every non-interactive mechanism $\mathcal{N}$, such as the Laplace mechanism, can be modeled as a CM that receives a dataset $x$ as its first message and returns $\mathcal{N}(x)$ as a response.
 Upon receiving any subsequent input message, the CM representing $\mathcal{N}$ halts the communication.
-  • Any interactive mechanism $\mathcal{M}$ with an initial state space $S_{\mathcal{M}}^{\text{init}}$ can be modeled as a CM that receives an initial state $s\in S_{\mathcal{M}}^{\text{init}}$ as its first message and responds with an acknowledgment message $\top$.
+  • Every interactive mechanism $\mathcal{M}$ with an initial state space $S_{\mathcal{M}}^{\text{init}}$ can be modeled as a CM that receives an initial state $s\in S_{\mathcal{M}}^{\text{init}}$ as its first message and responds with an acknowledgment message $\top$.
 This CM then answers the next input messages the same as $\mathcal{M}(s)$.
@@
 Definition 3.5 (Verifier $\mathcal{V}$).
-For a verification function $f$, a verifier $\mathcal{V}[f]$ is an IPM with the initial state space $S_{\mathcal{V}[f]}^{\text{init}}=\{()\}$, where ()()( ) represents an empty sequence.
-Since the initial state space of $\mathcal{V}[f]$ is a singleton, we denote an instance of $\mathcal{V}[f]$ with the initial state ()()( ) as $\mathcal{V}[f]$.
+For a verification function $f$, a verifier $\mathcal{V}[f]$ is an IPM with the initial state space $S_{\mathcal{V}[f]}^{\text{init}}=\{()\}$, where $()$ represents an empty sequence.
+Since the initial state space of $\mathcal{V}[f]$ is a singleton, we denote an instance of $\mathcal{V}[f]$ with the initial state $()$ as $\mathcal{V}[f]$.
 $\mathcal{V}[f]$ stands between two IMs, referred to as the left and right mechanisms.
@@
 Remark 3.6.
-By definition, for any continual mechanism $\mathcal{M}$ and any verification function $f$, $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ is an IM with two possible initial states.
+By definition, for every continual mechanism $\mathcal{M}$ and every verification function $f$, $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ is an IM with two possible initial states.
 The mechanism $\mathcal{I}\!\circ^{*}\!\mathcal{M}$ accepts messages of the form $(m_{0},m_{1})$, where both $m_{0}$ and $m_{1}$ are in the query space $Q_{\mathcal{M}}$ of $\mathcal{M}$.
@@
 We say that $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private (or $(\epsilon,\delta)$-DP for short) w.r.t.
-$f$ against adaptive adversaries if for any adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
-[…]
+$f$ against adaptive adversaries if for every adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
+Let $S_{\mathcal{M}}^{\text{init}}=\{s^{\text{init}}\}$ be the initial state space of the continual mechanism $\mathcal{M}$ in Definition 3.7.
+By definition, the initial state spaces of $\mathcal{I}$ and $\mathcal{V}[f]$ are $\{0,1\}$ and $\{()\}$, respectively.
+Thus, the initial state space of $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ is $\left\{((),0,s^{\text{init}}),((),1,s^{\text{init}})\right\}$.
+Define $\sim_{01}$ as a neighboring on this set, where only $((),0,s^{\text{init}})\sim_{01}((),1,s^{\text{init}})$.
+Comparing Definition 3.7 and Definition 2.13 leads to the following corollary:
 Corollary 3.8.
@@
 The initial state space of ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ is $S_{\mathcal{M}_{1}}^{\text{init}}\times\dots\times S_{\mathcal{M}_{k}}^{\text{init}}$, where $S_{\mathcal{M}_{i}}^{\text{init}}$ is the initial state space of $\mathcal{M}_{i}$.
-Let $\sim_{i}$ be a neighboring relation on each $S_{\mathcal{M}_{i}}^{\text{init}}$, and define the neighboring relation $\sim$ on $S_{\mathcal{M}_{1}}^{\text{init}}\times\dots\times S_{\mathcal{M}_{k}}^{\text{init}}$ as follows: for any two initial states $s=(s_{1},\dots,s_{k})$ and $s^{\prime}=(s_{1}^{\prime},\dots,s_{k}^{\prime})$, we have $s\sim s^{\prime}$ if and only if $s_{i}\sim_{i}s_{i}^{\prime}$ for every $i\in[k]$.
+Let $\sim_{i}$ be a neighboring relation on each $S_{\mathcal{M}_{i}}^{\text{init}}$, and define the neighboring relation $\sim$ on $S_{\mathcal{M}_{1}}^{\text{init}}\times\dots\times S_{\mathcal{M}_{k}}^{\text{init}}$ as follows: for every two initial states $s=(s_{1},\dots,s_{k})$ and $s^{\prime}=(s_{1}^{\prime},\dots,s_{k}^{\prime})$, we have $s\sim s^{\prime}$ if and only if $s_{i}\sim_{i}s_{i}^{\prime}$ for all $i\in[k]$.
 Vadhan and Wang [23] show that if each IM $\mathcal{M}_{i}$ is $\epsilon_{i}$-DP w.r.t $\sim_{i}$, and if the composition of NIMs $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ is $(\epsilon,\delta)$-DP, then ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ is also $(\epsilon,\delta)$-DP against an adaptive adversary.
@@
 In Section 4.1, we generalize ConComp to ExtConComp and give a formal definition for the concurrent composition of CMs.
-To extend Lyu’s result on the privacy guarantee of the concurrent composition of IMs to the concurrent composition of CMs, we want to use the following intermediate lemma from their paper: For every $(\epsilon,\delta)$-DP interactive mechanism $\mathcal{M}$ and any pair of neighboring initial states $s$ and $s^{\prime}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the IMs $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(b)$ are equivalent.
+To extend Lyu’s result on the privacy guarantee of the concurrent composition of IMs to the concurrent composition of CMs, we want to use the following intermediate lemma from their paper: For every $(\epsilon,\delta)$-DP interactive mechanism $\mathcal{M}$ and every pair of neighboring initial states $s$ and $s^{\prime}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the IMs $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(b)$ are equivalent.
 However, as we discuss in Section 4.2, there are hidden assumptions in the proof of this lemma.
@@
 The state of ExtConComp is a sequence of pairs $(\mathcal{M}_{i},s_{i}^{\prime})$, where the $i$-th pair represents the $i$-th created mechanism $\mathcal{M}_{i}$ and its current state $s_{i}^{\prime}$.
-The single initial state space of ExtConComp is the empty sequence ()()( ).
+The single initial state space of ExtConComp is the empty sequence $()$.
 Both query set and answer set of ExtConComp are the universal message set $\mathcal{U}$ in Remark 3.4.
@@
 Definition 4.2.
-For any verification function $f$, the $f$-concurrent composition of CMs is the interactive mechanism $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$.
+Let $f$ be a verification function. The $f$-concurrent composition of CMs is the interactive mechanism $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$.
 Differential privacy for the $f$-concurrent composition of CMs follows the standard definition for IMs:
@@
 Definition 4.4 (Suitable Sequence of Messages).
-A sequence of messages $m_{1},\dots,m_{k}$ is suitable if for every $1\leq j\leq k$, (i) $m_{j}$ is of the form $(m_{j}^{0},m_{j}^{1})$, where $m_{j}^{0}$ and $m_{j}^{1}$ are either identical creation queries or messages of the forms $(q_{0},\ell)$ and $(q_{1},\ell)$ with the same index $\ell$; (ii) for each $b\in\{0,1\}$, if $m_{j}^{b}=(q,\ell)$, then there are at least $\ell$ creation queries in the sequence $m_{1}^{b},\dots,m_{j-1}^{b}$; (iii) for each $b\in\{0,1\}$, if $m_{j}^{b}=(q,\ell)$ and $\mathcal{M}_{\ell}$ is the $\ell$-th mechanism requested to be created in $m_{1}^{b},\dots,m_{j-1}^{b}$, then $q$ is in the query space of $\mathcal{M}_{\ell}$, i.e., $q\in Q_{\mathcal{M}_{\ell}}$; and (iv) if $m_{j}^{0}=m_{j}^{1}=(\mathcal{M},s,\epsilon,\delta,f)$, then $\mathcal{M}$ is a CM with a single initial state $s$, and it is $(\epsilon,\delta)$-DP w.r.t.
-$f$.
+A sequence of messages $m_{1},\dots,m_{k}$ is said to be suitable if the following conditions hold for every $1\leq j\leq k$:
+  1. (i)
+    Each message $m_{j}$ is either
+    1. (a)
+      a pair of identical creation queries of the form $(\mathcal{M},s,\epsilon,\delta,f)^{2}$, or
+    2. (b)
+      a query pair message of the form $\left((q_{0},\ell),(q_{1},\ell)\right)$, where both queries $q_{0}$ and $q_{1}$ share the same index $\ell$.
+  2. (ii)
+    If $m_{j}=\left((q_{0},\ell),(q_{1},\ell)\right)$, then there are at least $\ell$ creation-query pairs in the sequence $m_{1},\dots,m_{j-1}$.
+  3. (iii)
+    If $m_{j}=\left((q_{0},\ell),(q_{1},\ell)\right)$ and $\mathcal{M}_{\ell}$ is the $\ell$-th mechanism requested to be created in the preceding messages $m_{1},\dots,m_{j-1}$, then both $q_{0}$ and $q_{1}$ belong to the query space of $\mathcal{M}_{\ell}$, i.e., $q_{0},q_{1}\in Q_{\mathcal{M}_{\ell}}$.
+  4. (iv)
+    If $m_{j}=(\mathcal{M},s,\epsilon,\delta,f)^{2}$, then $\mathcal{M}$ is a CM with a single initial state $s$, and it is $(\epsilon,\delta)$-DP w.r.t. $f$.
+  5. (v)
+    If $m_{j}=(\mathcal{M},s,\epsilon,\delta,f)^{2}$ is the $\ell$-th creation-query pair in the sequence $m_{1},\dots,m_{j-1}$, then the corresponding sequence of query pairs $(q_{0},q_{1})$ from the messages of the form $\left((q_{0},\ell),(q_{1},\ell)\right)$ in $m_{1},\dots,m_{k}$ is $f$-valid.
 4.2. $(\epsilon,\delta)$-DP IMs and $\mathrm{RR}_{\epsilon,\delta}$
 Lyu [19] shows that for every $(\epsilon,\delta)$-DP IM $\mathcal{M}$ and every pair of neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{T}$ (with a single initial state) such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(s_{b})$ are identical.
-Vadhan and Wang [20] show the same result in the special case where $\delta=0$.
-Although not explicitly stated, the constructions of $\mathcal{T}$ proposed in both [18] and [20] require the answer distributions of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to be discrete.
+Although not explicitly stated, the construction of $\mathcal{T}$ proposed in [19] requires the answer distributions of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to be discrete.
 This assumption is formally defined as:
@@
 Lyu [19] additionally assumes that there exists a pre-fixed upper bound $c$ on the number of queries $\mathcal{M}$ answers.
-In Appendix A, we will modify their proof to remove this additional assumption.
+This assumption stems from their use of a backward-defined control function.
+By employing the limit control functions which we introduce in Section 8, one can modify their proof to remove this additional assumption.
+Moreover, using the formalization of Section 8 to define recursive functions, we can relax the requirement that the answer set is finite to allow countably infinite answer sets.
 Let $\mathcal{M}^{\prime}$ be a continual mechanism that is $(\epsilon,\delta)$-DP w.r.t.
@@
 $\sim_{01}$.
-Replacing $\mathcal{M}$ in Lemma A.1 with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}^{\prime}$, we get:
+We have:
 Lemma 4.6 ([19]).
@@
 a verification function $f$.
-Suppose $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ has discrete answer distributions and responds to at most $c$ queries.
-Then, there exists an IPM $\mathcal{T}$ with a single initial state ()()( ) such that for each $b\in\{0,1\}$, the IMs $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$ are equivalent.
-Let $\mathcal{M}$ be an IM that is $\epsilon$-DP w.r.t.
-a neighbor relation $\sim$.
-For every pair of neighboring initial states $s_{0}$ and $s_{1}$, Vadhan and Wang [20] construct an IPM $\mathcal{T}$ (with a single initial state) such that for each bit $b\in\{0,1\}$, the mechanisms $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,0}(b)$ and $\mathcal{M}(s_{b})$ are identical.
-Appendix B, we will review their construction and show that for any query sequence $q_{1},\dots,q_{t}$, if the distributions of the $t$ answers of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to the queries $q_{1},\dots,q_{t}$ are identical, then $\mathcal{T}$ does not need to know the output of $\mathrm{RR}_{\epsilon,0}(b)$ in the first $t$ steps.
-This observation allows us to delay the interaction of $\mathcal{T}$ and $\mathrm{RR}_{\epsilon,0}(b)$ as much as possible.
-Let $\mathcal{M}=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}^{\prime}$, where $\mathcal{M}^{\prime}$ is an $\epsilon$-DP CM w.r.t.
-$f$.
-By definition, upon receiving a query sequence $q_{1},\dots,q_{t}$ where each $q_{j}=(q_{j}^{\prime},q_{j}^{\prime})$ is a pair of identical queries, $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}^{\prime}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M}^{\prime}$ have the same answer distribution, implying the following lemma:
-Lemma 4.7 ([20]).
-For $\epsilon\geq 0$, let $\mathcal{M}$ be a CM that is $\epsilon$-DP w.r.t.
-a verification function $f$.
 Suppose $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ has discrete answer distributions.
-Then, there exists an IPM $\mathcal{T}$ with a single initial state ()()( ) such that for each $b\in\{0,1\}$,
-  • the IMs $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,0}(b)$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$ are equivalent, and
-  • as long as $\mathcal{T}$ receives pairs of identical queries, it does not interact with $\mathrm{RR}_{\epsilon,0}(b)$.
+Then, there exists an IPM $\mathcal{T}$ with a single initial state such that for each $b\in\{0,1\}$, the IMs $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$ are equivalent.
 Remark 4.7.
-Since $\mathrm{RR}_{\epsilon,\delta}$ is a NIM, in both Lemma 4.6 and Lemma 4.7, the IPM $\mathcal{T}$ interacts with $\mathrm{RR}_{\epsilon,\delta}$ at most once.
+Since $\mathrm{RR}_{\epsilon,\delta}$ is a NIM, in Lemma 4.6, the IPM $\mathcal{T}$ interacts with $\mathrm{RR}_{\epsilon,\delta}$ at most once.
 4.3. Concurrent Composition of CMs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$
@@
 For $i\in[k]$, let $\alpha_{i}=(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})$ denote a creation query.
-For any input sequence of messages $m_{1},\dots,m_{t}$, the outcome of $f^{\alpha_{1},\dots,\alpha_{k}}$ returns $\top$ if and only if all of the following conditions hold:
+For every $t\in\mathbb{N}$ and every input sequence of messages $m_{1},\dots,m_{t}$, the verification function $f^{\alpha_{1},\dots,\alpha_{k}}$ returns $\top$ if and only if all of the following conditions hold:
   1. (i)
@@
     For $j\leq\min\{t,k\}$, the messages $m_{j}^{0}=m_{j}^{1}=\alpha_{j}$.
-  3. (iii)
-    For $i\in[k]$, the sequence of $(q^{0},q^{1})$s in the messages of the form $\left((q^{0},i),(q^{1},i)\right)$ is $f_{i}$-valid.
 With the above definition of $f^{\alpha_{1},\dots,\alpha_{k}}$, we refer to the $f^{\alpha_{1},\dots,\alpha_{k}}$-concurrent composition of CMs as the concurrent composition of the continual mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$.
@@
 Theorem 4.9.
-For $1\leq i\leq k$, let $\mathcal{M}_{i}$ be a continual mechanism with initial state space $S_{\mathcal{M}_{i}}^{\text{init}}=\{s_{i}\}$ such that $\mathcal{M}_{i}$ is $(\epsilon_{i},\delta_{i})$-DP w.r.t.
+For $1\leq i\leq k$, $\epsilon_{1},\dots,\epsilon_{k}>0$, and $0\leq\delta_{1},\dots,\delta_{k}\leq 1$, let $\mathcal{M}_{i}$ be a continual mechanism with initial state space $S_{\mathcal{M}_{i}}^{\text{init}}=\{s_{i}\}$ such that $\mathcal{M}_{i}$ is $(\epsilon_{i},\delta_{i})$-DP w.r.t.
 a verification function $f_{i}$ against adaptive adversaries, and $\mathcal{V}[f_{i}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}_{i}$ has discrete answer distributions.
-For $\epsilon>0$ and $0\leq\delta\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the concurrent composition of $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ is also $(\epsilon,\delta)$-DP against adaptive adversaries.
+For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the concurrent composition of $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ is also $(\epsilon,\delta)$-DP against adaptive adversaries.
 Proof.
@@
 Since $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}$ is an IPM with a single initial state, Lemma 2.16 implies that if ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$ is $(\epsilon,\delta)$-DP, then $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$ is also $(\epsilon,\delta)$-DP.
-Hence, for any adversary $\mathcal{A}$, the views of $\mathcal{A}$ when interacting with the mechanisms $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((0)_{i=1}^{k}\right)$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((1)_{i=1}^{k}\right)$ are $(\epsilon,\delta)$-indistinguishable.
+Hence, for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ when interacting with the mechanisms $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((0)_{i=1}^{k}\right)$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((1)_{i=1}^{k}\right)$ are $(\epsilon,\delta)$-indistinguishable.
 As we will show below, by construction, these views are identical to those obtained by interacting with the original mechanisms $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, implying that the concurrent composition of $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ is $(\epsilon,\delta)$-DP.
@@
 Definition 4.10 ($f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$).
-For any input sequence of messages $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$ returns $\top$ if and only if all of the following conditions hold:
+For every $t\in\mathbb{N}$ and every input sequence of messages $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$ returns $\top$ if and only if all of the following conditions hold:
   1. (i)
@@
 Note that $(\epsilon_{i}^{\prime},\delta_{i}^{\prime})$ is adaptively chosen by the adversary, while $(\epsilon_{i},\delta_{i})$ is a predetermined publicly-known parameter.
-  3. (iii)
-    For each $i\in[k]$, the sequence of $(q^{0},q^{1})$s in the messages of the form $\left((q^{0},i),(q^{1},i)\right)$ is $f_{i}^{\prime}$-valid.
 We refer to the $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$-concurrent composition of CMs as the concurrent composition of CMs with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$.
 A slight adaptation of the proof of Theorem 4.9 implies the following theorem.
+For convenience, in the rest of the paper, we represent a pair of identical creation queries $(\mathcal{M},s,\epsilon,\delta,f)$ by $(\mathcal{M},s,\epsilon,\delta,f)^{2}$.
 Theorem 4.11.
-For $\epsilon,\epsilon_{1},\dots,\epsilon_{k}>0$ and $0\leq\delta,\delta_{1},\dots,\delta_{k}\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the concurrent composition of $k$ CMs with the privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$ is also $(\epsilon,\delta)$-DP against adaptive adversaries.
+For $\epsilon_{1},\dots,\epsilon_{k}>0$, $\epsilon\geq 0$ and $0\leq\delta,\delta_{1},\dots,\delta_{k}\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the concurrent composition of $k$ CMs with the privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$ is also $(\epsilon,\delta)$-DP against adaptive adversaries.
 Proof.
-The proof is identical to the proof of Theorem 4.10 with the following modification in the construction of $\mathcal{P}$: Recall that $(r_{1},\dots,r_{k})$ denotes the output of ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$.
-In the proof of Theorem 4.10, the $i$-th creation query was known to be $(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})$.
+The proof is identical to the proof of Theorem 4.9 with a modification in the construction of $\mathcal{P}$: Recall that $(r_{1},\dots,r_{k})$ denotes the output of ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$.
+In the proof of Theorem 4.9, the $i$-th creation query was known in advance to be $(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})$.
 Consequently, $\mathcal{T}_{i}$ expected $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$ as its right mechanism, and when $\mathcal{T}_{i}$ needed a response from its right mechanism, $\mathcal{P}$ sent $r_{i}$ to it.
-However, in the concurrent composition with fixed parameters, the IPM requesting the sample from $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$ need not necessarily be $\mathcal{T}_{i}$.
+However, in the concurrent composition with fixed parameters, the IPM requesting the sample from $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$ need not necessarily be the $i$-th IPM, $\mathcal{T}_{i}$.
 To handle this, we modify $\mathcal{P}$ so that it maps each output sample $r_{i}$ to the appropriate IPM $\mathcal{T}_{j}$ needing a sample from $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$.
+The implementation details are as follows.
 Let $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$ denote the $j$-th pair of identical creation queries received by $\mathcal{P}$ from the verifier, and let $\mathcal{T}_{j}$ be the corresponding IPM given by Lemma 4.6.
-When $\mathcal{T}_{j}$ requests a sample from $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, $\mathcal{P}$ iterates over $\sigma_{1}$ to find the smallest $i^{*}\in[k]$ such that the $i^{*}$-th element of $\sigma_{1}$ is not null and $\epsilon_{j}^{\prime}=\epsilon_{i^{*}}$, $\delta_{j}^{\prime}=\delta_{i^{*}}$.
-After sending $r_{i^{*}}$ to $\mathcal{T}_{j}$, $\mathcal{P}$ updates $\sigma_{1}$ and sets $r_{i}$ to null.
+In the proof of Theorem 4.9, the state of $\mathcal{P}$ consists of a pair of two sequences $\sigma_{1}=(r_{1},\dots,r_{k})$ and $\sigma_{2}=\left((\mathcal{T}_{j},t_{j})\right)_{j=1}^{k}$, where $r_{j}$ is the randomized response defined above and $t_{j}$ denotes the current state of $\mathcal{T}_{j}$.
+We extend the state of $\mathcal{P}$ to additionally store an index $y_{i}$ with each $r_{i}$ in $\sigma_{1}$ and the privacy parameters $\epsilon_{j}^{\prime},\delta_{j}^{\prime}$ with each $\mathcal{T}_{j}$ in $\sigma_{2}$.
+Variable $y_{i}$ is initialized to 0 and used to record the index $j$ of the mechanism $\mathcal{T}_{j}$ associated with $r_{i}$
+Upon receiving $(r_{1},\dots,r_{k})$ from ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$, $\mathcal{P}$ sets $\sigma_{1}=\left((r_{1},y_{1}=0),\dots,(r_{k},y_{k}=0)\right)$.
+Also, upon receiving a pair of creation queries $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$, $\mathcal{P}$ appends the tuple $(\mathcal{T}_{j},t_{j},\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ to $\sigma_{2}$.
+When $\mathcal{T}_{j}$ requests a sample from $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, $\mathcal{P}$ iterates over $\sigma_{1}$ to find the smallest $i^{*}\in[k]$ such that $y_{i^{*}}=0$, $\epsilon_{j}^{\prime}=\epsilon_{i^{*}}$, and $\delta_{j}^{\prime}=\delta_{i^{*}}$.
+It then assigns $y_{i^{*}}=j$ and sends $r_{i^{*}}$ to $\mathcal{T}_{j}$ as a right response.
 This adjustment ensures that even when there are duplicate privacy parameters among $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$, each sample $r_{i}$ is sent to at most one $\mathcal{T}_{j}$.
@@
 a verification function $f_{i}$ and with $(\epsilon_{i},\delta_{i})\in{\texttt{params}}$, and the sequence of query pairs for $\mathcal{M}_{i}$ must be $f_{i}$-valid.
-In Section 5.1, we formally define this composition and extend the result of Theorem 4.12 to the concurrent $k$-sparse parallel composition of CMs when $\delta_{1}=\dots=\delta_{k}=0$.
-In Section 5.2, we provide a counter example that demonstrates why such an extension is impossible for $\delta_{j}>0$.
-In Section 5.3, we restrict the verification functions $f_{i}$ of the $k$ mechanisms receiving neighboring input and we show that Theorem 4.12 holds for the concurrent $k$-sparse parallel composition of CM in this scenario without any constraints on the values of $\delta_{j}$.
-5.1. Definition and Composition Theorem for Pure DP
+In Section 5.1, we formally define this composition and show that the result of Theorem 4.11 cannot be extended to the concurrent $k$-sparse parallel composition of CMs when any $\delta_{j}>0$.
+In Section 5.2, we prove a parallel composition theorem that matches the privacy guarantee of Theorem 4.11 when only purely differentially privacy CMs are created.
+Finally, in Section 5.3, we consider a restricted class of verification functions $f_{i}$ for the $k$ mechanisms receiving neighboring inputs.
+Under certain assumptions about these functions, we show that the privacy guarantee of Theorem 4.11 still holds even if the created mechanism are approximately dp.
+5.1. Definition and Counterexample
 To formally define the concurrent $k$-sparse parallel composition of CMs with parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$, we introduce the verification function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$, where the subscript $\infty$ emphasizes that an unbounded number of CMs can be created.
 Definition 5.1 ($f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$).
-For any input sequence of messages $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$ returns $\top$ if and only if all of the following conditions hold:
+For every $t\in\mathbb{N}$ and every input sequence of messages $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$ returns $\top$ if and only if all of the following conditions hold:
   1. (i)
@@
 Note that this implies that $|\mathcal{J}|\leq k.$
-  3. (iii)
-    For every $j\in\mathcal{J}$, the sequence of $(q_{0},q_{1})$s in the messages of the form $\left((q_{0},j),(q_{1},j)\right)$ must be $f_{i}$-valid.
-Note that by definition, for every index $j\notin\mathcal{J}$ and every message of the form $\left((q_{0},j),(q_{1},j)\right)$, it holds that $q_{0}=q_{1}$.
 We refer to the $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$-concurrent composition of CMs as the concurrent $k$-sparse parallel composition of CMs with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$.
@@
 $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$ against adaptive adversaries.
+The following theorem shows that for $k=1$, $\epsilon_{1}=0$, and any $\delta_{1}\in(0,1]$, the $f^{0,\delta_{1}}_{\infty}$-concurrent composition of CMs is not differentially private:
 Theorem 5.2.
-For $\epsilon,\epsilon_{1},\dots,\epsilon_{k}\geq 0$ and $0\leq\delta\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ is $(\epsilon,\delta)$-DP, then the concurrent $k$-sparse parallel composition of CMs with the privacy parameters $(\epsilon_{1},0),\dots,(\epsilon_{k},0)$ is also $(\epsilon,\delta)$-DP against an adaptive adversary.
+For every $\delta\in(0,1]$, there exists an adversary $\mathcal{A}_{\delta}$ such that the views of $\mathcal{A}_{\delta}$ interacting with $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ have disjoint supports.
 Proof.
-The proof is identical to the proof of Theorem 4.12, except for the definition of IPMs $\mathcal{T}_{j}$ that are executed by the IPM $\mathcal{P}$.
-Let $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$ denote the $j$-th pair of identical creation queries received by $\mathcal{P}$ from $\mathcal{V}[f]$.
-To ensure that the proof of Theorem 4.12 remains valid, we must guarantee that the newly defined IPMs $\mathcal{T}_{j}$ satisfy the following properties needed by the proof of Theorem 4.12.
-These are the only properties that this proof requires for the $\mathcal{T}_{j}$s:
-  • The mechanisms $\mathcal{T}_{j}\!\circ^{*}\!\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ and $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ are equivalent, and
-  • the multiset of privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ corresponding to IMPs $\mathcal{T}_{j}$ that are interacting with their right mechanism $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ is a sub-multiset of $(\epsilon_{1},0),\dots,(\epsilon_{k},0)$, implying that the remaining $\mathcal{T}_{j}$ never interact with a $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$.
-The new definition of $\mathcal{T}_{j}$ is as follows:
-  • If $\delta_{j}^{\prime}>0$, then $\mathcal{T}_{j}$ acts identical to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$: $\mathcal{T}_{j}$ has a single initial state $\left((),0,s^{\text{init}}\right)$, where $s^{\text{init}}$ is the single initial state of $\mathcal{M}_{j}^{\prime}$.
-Upon receiving a right message, $\mathcal{T}_{j}$ halts the communication, and upon receiving a left message, it updates its state and responds identically to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$.
-  • If $\delta_{j}^{\prime}=0$, then $\mathcal{T}_{j}$ is the IPM corresponding to $\mathcal{M}_{j}^{\prime}$ given by Lemma 4.7.
-By Definition 5.1 (ii), if $\delta_{j}^{\prime}>0$, then all future messages $\left((q_{0},j),(q_{1},j)\right)$ satisfy $q_{0}=q_{1}$.
-Therefore, by the above definition, for each $b\in\{0,1\}$, the mechanisms $\mathcal{T}_{j}\!\circ^{*}\!\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ and $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ are identical, and $\mathcal{T}_{j}$ never interacts with $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$.
-Moreover, when $\delta_{j}^{\prime}=0$, by Definition 5.1 (iii), Lemma 4.7, and Remark 4.8, for each $b\in\{0,1\}$, the mechanisms $\mathcal{T}_{j}\!\circ^{*}\!\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ and $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ are identical, and $\mathcal{T}_{j}$ interacts with $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ at most once after receiving a pair of non-identical queries.
-In the concurrent $k$-sparse parallel composition, an unlimited number of CMs can be created, but only at most $k$ of them (here, with $\delta_{j}^{\prime}=0$) can receive non-identical query pairs.
-By Definition 5.1 (ii), the privacy parameters of these mechanisms is a sub-multiset of $(\epsilon_{1},0),\dots,(\epsilon_{k},0)$.
-By design, the IPMs $\mathcal{T}_{j}$ corresponding to these mechanisms may require a sample from $\mathrm{RR}_{\epsilon_{j}^{\prime},0}$.
-∎
-5.2. Counter Example.
-In this section, we illustrate that extending the result of Theorem 5.2 to general $\delta_{j}>0$ is impossible.
-Specifically, let $k=1$, $\epsilon_{1}=0$, and $\delta_{1}=\delta\in(0,1]$.
-We construct, an adversary $\mathcal{A}_{\delta}$ that interacts with $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ and successfully determines the secret bit $b\in\{0,1\}$.
-Our construction proceeds in three steps: (1) introducing a verification function $g$, (2) constructing a CM $\mathcal{M}_{\delta}$ that is $(0,\delta)$-DP w.r.t.
+We construct $\mathcal{A}_{\delta}$ in three steps: (1) introducing a verification function $g$, (2) constructing a CM $\mathcal{M}_{\delta}$ that is $(0,\delta)$-DP w.r.t.
 $g$, and (3) designing $\mathcal{A}_{\delta}$.
+We then prove that for every $b\in\{0,1\}$, with probability $1$, $\mathcal{A}_{\delta}$ can determine the secret initial state $b$ by interacting with $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ finitely many times.
+Specifically, we show that with probability $1$, the view of $\mathcal{A}_{\delta}$ interacting with $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ is finite and includes $b$ as the last message.
+This immediately implies that the views of $\mathcal{A}_{\delta}$ interacting with $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ have disjoint supports.
 Step 1 (Introducing $g$): The verification function $g$ accepts a message sequence $(m_{1},\dots,m_{t})$ and returns $\top$ if and only if (i) $t\leq 2$, (ii) each $m_{j}$ is a pair of binary numbers $(b_{j},b_{j}^{\prime})\in\{0,1\}^{2}$, and (ii) the sequences $(b_{1},\dots,b_{t})$ and $(b_{1}^{\prime},\dots,b_{t}^{\prime})$ differ in at most a single element.
-[…]
+Step 2 (Constructing $\mathcal{M}_{\delta}$): The CM $\mathcal{M}_{\delta}:\{\bot,\top\}\times\{0,1\}\to\{\bot,\top\}\times\{0,1,\top,\bot\}$ with the initial state space $\{\top\}$ is defined as follows:
   • $\mathcal{M}_{\delta}$ maps state $\top$ and an input message $b\in\{0,1\}$ to state $\top$ and output message $\top$ with probability $1-\delta$ and to state $\bot$ and output message $\bot$ with probability $\delta$.
@@
 If the initial state of the identifier $\mathcal{I}$ in $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $b$, then $\mathcal{M}_{\delta}$ receives the query $b$, and by definition it will return $b$ as output, revealing the secret bit $b$.
-Therefore, with probability $1-(1-\delta)^{t}$, the adversary $\mathcal{A}_{\delta}$ can detect the secret bit $b$ by sending at most $12t+$ 1 queries.
+Upon receiving the response $b$, $\mathcal{A}_{\delta}$ halts the communication.
+Therefore, with probability $1-(1-\delta)^{t}$, the adversary $\mathcal{A}_{\delta}$ can detect the secret bit $b$ by sending at most $2t+1$ queries.
 Since $\delta>0$, as $t$ goes to infinity, this probability tends to $1$.
+Thus, with probability $1$, the view of $\mathcal{A}_{\delta}$ interacting with $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ is finite and includes $b$ as the last message.
+∎
+5.2. Composition Theorem
+In the counterexample of Section 5.1, the adversary runs an unbounded number of $(0,\delta)$-DP CMs but sends a pair of non-identical queries to only one of them, leading to the exposure of the secret bit with probability $1$.
+In that construction, each individual mechanism may fail to preserve privacy with probability $\delta$, and these failures accumulate, ultimately leading to complete exposure.
+To prevent such leakage, we restrict the second privacy parameters $\delta_{j}^{\prime}$ of all created mechanisms to satisfy a certain condition.
+Definition 5.3 ($f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta^{\prime}}$).
+For every $t\in\mathbb{N}$ and every input sequence of messages $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta}$ returns $\top$ if and only if all of the following conditions hold:
+  1. (i)
+    The sequence $m_{1},\dots,m_{t}$ is suitable.
+  2. (ii)
+    Let $\mathcal{M}_{i}^{\prime}$ denote the $i$-th created mechanism and $(\epsilon_{i}^{\prime},\delta_{i}^{\prime})$ be its privacy parameters.
+Define $\mathcal{J}$ as the set of indices $j$ such that there exists a message $\left((q_{0},j),(q_{1},j)\right)$ where $q_{0}\neq q_{1}$.
+Let $\texttt{params}_{\mathcal{J}}$ denote the multiset consisting of the first privacy parameters $\epsilon_{j}^{\prime}$ for each $j\in\mathcal{J}$.
+This condition requires that $\texttt{params}_{\mathcal{J}}$ be a sub-multiset of the fixed privacy parameters $\epsilon_{1},\dots,\epsilon_{k}$.
+  3. (iii)
+    The inequality $1-\prod_{j}(1-\delta_{j}^{\prime})\leq\delta^{\prime}$ holds, where the product is taken over the indices of the created mechanisms.
+We refer to the $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta}$-concurrent composition of CMs as the concurrent $k$-sparse parallel composition of CMs with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$ and upper bound $\delta$.
+We show next the following theorem.
+Theorem 5.4.
+For $\epsilon_{1},\dots,\epsilon_{k}>0$, $\epsilon\geq 0$ and $0\leq\delta,\delta^{\prime}\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ is $(\epsilon,\delta)$-DP, then the $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta^{\prime}}$-concurrent composition of CMs is $(\epsilon,\delta+\delta^{\prime})$-DP against adaptive adversaries.
+It is known that the composition of $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ is $\sum_{i=1}^{k}\epsilon_{i}$-DP.
+Thus, by converging all parameters $\epsilon_{1},\dots,\epsilon_{k}$ to zero, we conclude that the $f^{0,\dots,0}_{\infty,\delta^{\prime}}$-concurrent composition of CMs is $(0,\delta^{\prime})$-DP.
+The proof of Theorem 5.2 shows that this composition cannot be $(0,\delta^{\prime\prime})$-DP for $\delta^{\prime\prime}<\delta^{\prime}$, and thus the result of Theorem 5.4 is tight.
+Setting $\delta^{\prime}=0$ in Theorem 5.4 means that all created mechanisms must be purely differentially private:
+Corollary 5.5.
+For $\epsilon_{1},\dots,\epsilon_{k}>0$, $\epsilon\geq 0$ and $0\leq\delta\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ is $(\epsilon,\delta)$-DP, then the $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,0}$-concurrent composition of CMs is $(\epsilon,\delta)$-DP against adaptive adversaries.
+That is, the concurrent $k$-sparse parallel composition of purely differentially private CMs with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$ is $(\epsilon,\delta)$-DP against adaptive adversaries.
+In the proofs of the previous composition theorems (namely, Theorems 4.9 and 4.11), each created CM $\mathcal{M}_{j}^{\prime}$ with privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ is simulated by the IPM $\mathcal{T}_{j}$ from Lemma 4.6.
+For each $b\in\{0,1\}$, the mechanism $\mathcal{T}_{j}$ interacts with $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ once and simulates $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ identically.
+Next, we present a new construction for $\mathcal{T}_{j}$, where it interacts with $\mathrm{IRR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$ instead of $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$.
+This construction satisfies additional properties needed in the proof of Theorem 5.4.
+In Section 8, we prove the following lemma, which directly implies the desired IPM as stated in Corollary 5.7.
+Lemma 5.6.
+For $\epsilon>0$ and $0\leq\delta\leq 1$, let $\mathcal{M}:S_{\mathcal{M}}\times Q_{\mathcal{M}}\to S_{\mathcal{M}}\times A_{\mathcal{M}}$ be an $(\epsilon,\delta)$-DP IM w.r.t.
+a neighbor relation $\sim$.
+Suppose that $\mathcal{M}$ has discrete answer distributions.
+Then, for every two neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$:
+  • The mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\!\circ^{*}\!\mathrm{IRR}_{\epsilon,\delta}(b)$ are equivalent.
+  • For every $k\in\mathbb{N}$ and every query sequence $(q_{1},\dots,q_{k})\in Q_{\mathcal{M}}^{k}$, if the distributions of answers produced by $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to $(q_{1},\dots,q_{k})$ are identical, then when $\mathcal{P}\!\circ^{*}\!\mathrm{IRR}_{\epsilon,\delta}(b)$ receives the queries $q_{1},\dots,q_{k}$, the IPM $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}(b)$ only once.
+Otherwise, $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}(b)$ at most twice.
+To avoid confusion in later proofs, we replace the IPM $\mathcal{P}$ in Lemma 5.6 with $\mathcal{T}$ in the following corollary.
+Corollary 5.7.
+For $\epsilon>0$ and $0\leq\delta\leq 1$, let $\mathcal{M}$ be a CM that is $(\epsilon,\delta)$-DP w.r.t.
+a verification function $f$.
+Suppose $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ has discrete answer distributions.
+Then there exists an IPM $\mathcal{T}$ such that for each $b\in\{0,1\}$:
+  • The IMs $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$ and $\mathcal{T}\!\circ^{*}\!\mathrm{IRR}_{\epsilon,\delta}(b)$ are equivalent.
+  • For every $k\in\mathbb{N}$ and every message sequence $m_{1},\dots,m_{k}$ such that each message $m_{j}$ is a pair of identical messages, if $\mathcal{T}\!\circ^{*}\!\mathrm{IRR}_{\epsilon,\delta}(b)$ receives $m_{1},\dots,m_{k}$ as input, then the IPM $\mathcal{T}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}(b)$ at most once.
+  • $\mathcal{T}$ sends at most two messages to its right mechanism.
+Proof.
+By Corollary 3.8, $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ is an IM that is $(\epsilon,\delta)$-DP w.r.t.
+$\sim_{01}$.
+By definition of the identifier $\mathcal{I}$, for every $k\in\mathbb{N}$ and every message sequence $m_{1},\dots,m_{k}$, if each message $m_{j}$ is a pair of identical messages, then the distributions of answers produced by $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M}$ in respond to $(m_{1},\dots,m_{k})$ are identical.
+Applying Lemma 5.6 to the IM $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ then yields the desired result.
+∎
+Informally, by Definition 2.9, Corollary 5.7 states that each created CM $\mathcal{M}_{j}^{\prime}$, with the privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$, can be simulated by post-processing the outcomes of two randomized response mechanisms, namely $\mathrm{RR}_{\epsilon_{j}^{\prime},0}$ and $\mathrm{RR}_{0,\delta_{j}^{\prime}}$.
+While simulating $\mathcal{M}_{j}^{\prime}$ in general requires access to both outputs, by this lemma, as long as the adversary asks pairs of identical queries from $\mathcal{M}_{j}^{\prime}$, the output of $\mathrm{RR}_{\epsilon_{j}^{\prime},0}$ is unnecessary.
+By definition of $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta^{\prime}}$, the multiset of the parameters $\epsilon_{j}^{\prime}$ of the mechanisms receiving non-identical query pairs is a sub-multiset of $\epsilon_{1},\dots,\epsilon_{k}$.
+The high-level idea behind the proof of Theorem 5.4 is to separately compose $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ and $\mathrm{RR}_{0,\delta_{1}^{\prime}},\mathrm{RR}_{0,\delta_{2}^{\prime}},\dots$, and then apply basic composition to these two sets of composed mechanisms to conclude the final result.
+Since the parameters $\delta_{j}^{\prime}$ are chosen adaptively, the composition of $\mathrm{RR}_{0,\delta_{1}^{\prime}},\mathrm{RR}_{0,\delta_{2}^{\prime}},\dots$ is an instance of the filter composition of NIMs [21].
+We do not state this composition in terms of filters and define it using a verification function $f^{\delta^{\prime}}_{\mathrm{RR}}$ that only allows the creation of randomized response mechanisms $\mathrm{RR}_{0,\delta_{j}^{\prime}}$ and returns $\bot$ if $1-\prod(1-\delta_{j}^{\prime})>\delta^{\prime}$.
+Definition 5.8 ($f^{\delta}_{\mathrm{RR}}$).
+For every $t\in\mathbb{N}$ and every input sequence of messages $m_{1},\dots,m_{t}$, the function $f^{\delta}_{\mathrm{RR}}$ returns $\top$ if and only if all of the following conditions hold:
+  1. (i)
+    The sequence $m_{1},\dots,m_{t}$ is suitable.
+  2. (ii)
+    Let $\mathcal{M}_{j}^{\prime}$ denote the $j$-th mechanism requested to be created and $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ and $f_{j}^{\prime}$ be its privacy parameters and verification function, respectively.
+Each $\mathcal{M}_{j}^{\prime}$ is the randomized response mechanism $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$ (see Definition 2.3), $f_{j}^{\prime}$ is its corresponding verification function, and $\epsilon_{j}=0$.
+By definition of $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, $f_{j}^{\prime}$ returns $\top$ if and only if it receives a single message $m$ where $m$ is a pair of bits.
+  3. (iii)
+    The inequality $1-\prod_{j}(1-\delta_{j}^{\prime})\leq\delta$ holds, where the product is taken over the indices of the created mechanisms.
+The following lemma is proved in Section 9.
+[^6] By setting $\epsilon_{1}=\epsilon_{2}=\dots=0$ in Theorem 2 of [25], it follows that for every $\epsilon>0$, $\delta^{\prime}>0$, and $0\leq\delta^{\prime\prime}\leq 1$, the composition of randomized response mechanisms $\mathrm{RR}_{0,\delta_{1}},\mathrm{RR}_{0,\delta_{2}},\dots$ with adaptively chosen privacy parameters $\delta_{1},\delta_{2},\dots$ satisfying $\sum_{i}\delta_{i}\leq\delta^{\prime\prime}$ guarantees $(\epsilon,\delta^{\prime}+\delta^{\prime\prime})$-DP.
+Taking $\epsilon$ and $\delta^{\prime}$ to zero, we can conclude that this composition is $(0,\delta^{\prime\prime})$-DP.
+To get tighter bounds, we relax the requirement $\sum_{i}\delta_{i}\leq\delta^{\prime\prime}$ to $1-\prod_{i}(1-\delta_{i})\leq\delta^{\prime\prime}$ and prove the composition is still $(0,\delta^{\prime\prime})$-DP.
+Lemma 5.9.
+For every $0\leq\delta\leq 1$, the $f^{\delta}_{\mathrm{RR}}$-concurrent composition of continual mechanisms is $(0,\delta)$-differentially private.
+Before proving Theorem 5.4, we show one final technical lemma that formally composes the fixed-parameter mechanisms $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ and the adaptively chosen mechanisms $\mathrm{RR}_{0,\delta_{1}^{\prime}},\mathrm{RR}_{0,\delta_{2}^{\prime}},\dots$.
+To define the composition, we introduce the following verification function:
+Definition 5.10 ($f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta,RR}$).
+For every input sequence of messages $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta,RR}$ returns $\top$ if and only if all of the following conditions hold:
+  1. (i)
+    The sequence $m_{1},\dots,m_{t}$ is suitable.
+  2. (ii)
+    Let $\mathcal{M}_{j}^{\prime}$ denote the $j$-th created mechanism and $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ and $f_{j}^{\prime}$ be its privacy parameters and verification function, respectively.
+Each $\mathcal{M}_{j}^{\prime}$ equals the randomized response mechanism $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, and at least one of $\epsilon_{j}^{\prime}$ and $\delta_{j}^{\prime}$ is zero.
+  3. (iii)
+    The multiset of positive $\epsilon_{j}^{\prime}$ is a sub-multiset of the fixed parameters $\epsilon_{1},\dots,\epsilon_{k}$.
+  4. (iii)
+    The inequality $1-\prod_{j}(1-\delta_{j}^{\prime})\leq\delta^{\prime}$ holds, where the product is taken over the indices of the created mechanisms.
+Lemma 5.11.
+For every $\epsilon\geq 0$ and $0\leq\delta,\delta^{\prime}\leq 1$, if ${\textrm{Comp}}(\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0})$ is $(\epsilon,\delta)$-DP, then the $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta^{\prime},RR}$-concurrent composition of CMs is $(\epsilon,\delta+\delta^{\prime})$-DP.
+Proof.
+Let $f^{\mathrm{RR}}=f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta^{\prime},RR}$. We design an IPM $\mathcal{F}$ with a single initial state and a verification function $f^{*}$ and show:
+  1. (1)
+    For each $b\in\{0,1\}$, the IMs $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{F}\!\circ^{*}\!\mathcal{V}[f^{*}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are equivalent.
+  2. (2)
+    ExtConComp is $(\epsilon,\delta+\delta^{\prime})$-DP w.r.t. $f^{*}$.
+From property (2), it follows that for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ interacting with $\mathcal{V}[f^{*}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{*}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ are $(\epsilon,\delta+\delta^{\prime})$-indistinguishable.
+Then, by the post-processing lemma, for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ interacting with $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{F}\!\circ^{*}\!\mathcal{V}[f^{*}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{F}\!\circ^{*}\!\mathcal{V}[f^{*}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ are $(\epsilon,\delta+\delta^{\prime})$-indistinguishable.
+Thus, by property (1), the views of any adversary interacting with $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ are $(\epsilon,\delta+\delta^{\prime})$-indistinguishable, implying that the $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta^{\prime},RR}$-concurrent composition of CMs is $(\epsilon,\delta+\delta^{\prime})$-DP.
+Intuitively, $\mathcal{F}$ partitions the randomized responses mechanisms $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$ requested to be created into two groups: those with $\epsilon_{j}^{\prime}=0$ and those with $\delta_{j}^{\prime}=0$.
+Then, instead of directly requesting ExtConComp to execute the randomized response mechanisms, $\mathcal{F}$ requests ExtConComp, called the outer ExtConComp, to run two so-called inner instances of ExtConComp, each responsible for executing one group.
+Note that ExtConComp can create any continual sub-mechanism including another instance of ExtConComp.
+To implement this, $\mathcal{F}$ begins by creating two internal instances of ExtConComp.
+It then rewrites incoming messages to ensure they are routed to the appropriate internal ExtConComp.
+To formally define $\mathcal{F}$, we first define the verification function $f^{*}$: For each $i\in[k]$, let $\beta_{i}$ denote the creation query for $\mathrm{RR}_{\epsilon_{i},0}$.
+Let $\gamma_{1}=({\textrm{ExtConComp}},(),\epsilon,\delta,f^{\beta_{1},\dots,\beta_{k}})$ be a creation query for creating an instance of ExtConComp running $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ (see Definition 4.8).
+Also, let $\gamma_{2}=({\textrm{ExtConComp}},(),0,\delta,f^{\delta}_{\mathrm{RR}})$ be a creation query for creating another instance of ExtConComp executing adaptively chosen $\mathrm{RR}_{0,\delta_{1}^{\prime}},\mathrm{RR}_{0,\delta_{2}^{\prime}},\dots$ satisfying $1-\prod(1-\delta_{j}^{\prime})\leq\delta^{\prime}$ (see Definition 5.8).
+We define $f^{*}=f^{\gamma_{1},\gamma_{2}}$ as in Definition 4.8.
+We now define $\mathcal{F}$.
+Initially (i.e., upon receiving its first left message), $\mathcal{F}$ creates the two internal instances of ExtConComp by sending the pairs of identical creation queries $(\gamma_{1},\gamma_{1})$ and $(\gamma_{2},\gamma_{2})$ to its right mechanism $\mathcal{V}[f^{*}]$.
+$\mathcal{F}$ ignores the acknowledge responses $\top$.
+The left query set of $\mathcal{F}$ equals the set of messages the verifier $\mathcal{V}[f^{\mathrm{RR}}]$ sends to its right mechanism, which consists of (A) pairs of identical creation queries and (B) messages of the form $\left((c,j),(c^{\prime},j)\right)$ that requests giving the $j$-th created (randomized response) mechanism one of the bits $c$ and $c^{\prime}$ as input and returning its response.
+We denote the $j$-th pair of creation queries by $(\alpha_{j},\alpha_{j})$ and its corresponding randomized response mechanism by $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$.
+Note that the decision of whether to feed $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$ with $c$ or $c^{\prime}$ is made by the identifier.
+(A) Let $(\alpha_{j},\alpha_{j})$ denote the $j$-th pair of identical creation queries received by $\mathcal{F}$, and let $\epsilon_{j}^{\prime}$ and $\delta_{j}^{\prime}$ be the privacy parameters in $\alpha_{j}$.
+Upon receiving $(\alpha_{j},\alpha_{j})$ from the left mechanism $\mathcal{V}[f^{\mathrm{RR}}]$, $\mathcal{F}$ takes one of the following actions:
+  • If $\epsilon_{j}^{\prime}=0$, $\mathcal{F}$ sends the message $\left((\alpha_{j},2),(\alpha_{j},2)\right)$ to its right mechanism.
+This message passes through an identifier before reaching (the outer) ExtConComp.
+The outer mechanism ExtConComp receives $(\alpha_{j},2)$ and forwards the creation query $\alpha_{j}$ to its second inner instance of ExtConComp.
+  • Otherwise, by definition of $f^{\mathrm{RR}}$, $\delta_{j}^{\prime}=0$.
+In this case, $\mathcal{F}$ sends the message $\left((\alpha_{j},1),(\alpha_{j},1)\right)$ to its right mechanism, resulting in the first inner instance of ExtConComp to execute $\alpha_{j}$.
+$\mathcal{F}$ maintains an initially empty array in its state, initially empty.
+When processing $(\alpha_{j},\alpha_{j})$, it adds the pair $(w_{j},t_{j})$ to this array, where $w_{j}\in\{1,2\}$ identifies the inner ExtConComp that has created $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, and $t_{j}$ indicates the index of the created mechanism within that instance.
+(B) Upon receiving a left message of the form $\left((c,j),(c^{\prime},j)\right)$, $\mathcal{F}$ sends the message $\left(((c,t_{j}),w_{j}),((c^{\prime},t_{j}),w_{j})\right)$ to the right mechanism, using the stored values $w_{j}$ and $t_{j}$.
+Except for the first two acknowledgment messages for creating the internal instances of ExtConComp, upon receiving any message from the right, $\mathcal{F}$ forwards it to the left mechanism unchanged.
+By the definition of $f^{\mathrm{RR}}$, the privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ associated with the creation queries satisfy the following properties: (i) $\epsilon_{j}^{\prime}=0$ for all creation queries $\alpha_{j}$ except at most $k$ of them, which form a sub-multiset of the predetermined parameters $\epsilon_{1},\dots,\epsilon_{k}$, and (ii) the inequality $1-\prod(1-\delta_{j}^{\prime})$ holds.
+Therefore, by the construction of $\mathcal{F}$, the verifier $\mathcal{V}[f^{*}]$ in $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{F}\!\circ^{*}\!\mathcal{V}[f^{*}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$ always receives valid inputs and never halts the communication.
+Thus, for each $b\in\{0,1\}$, the IMs $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{F}\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{F}\!\circ^{*}\!\mathcal{V}[f^{*}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are equivalent.
+Furthermore, by the design of $\mathcal{F}$, for each $b\in\{0,1\}$, the IMs $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{F}\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are equivalent: They both use the same verification function to validate messages.
+Upon receiving a query to create $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, they both create the mechanism and return $\top$.
+And upon receiving the pair of input datasets $(c_{0},c_{1})$ for $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, they both give $c_{b}$ to $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$ and return its response.
+Thus, the IMs $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{F}\!\circ^{*}\!\mathcal{V}[f^{*}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are equivalent, and the property (1) is satisfied.
+It remains to prove that the property (2) holds.
+The assumption that ${\textrm{Comp}}(\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0})$ is $(\epsilon,\delta)$-DP is equivalent to that ExtConComp is $(\epsilon,\delta)$-DP w.r.t.
+the verification function $f^{\beta_{1},\dots,\beta_{k}}$.
+Moreover, by Lemma 5.9, ExtConComp is $(0,\delta^{\prime})$-DP w.r.t.
+$f^{\mathrm{RR}}$.
+Therefore, by Theorem 4.9 and the fact that the composition of $\mathrm{RR}_{\epsilon,\delta}$ and $\mathrm{RR}_{0,\delta^{\prime}}$ is $(\epsilon,\delta+\delta^{\prime})$-DP, ExtConComp is $(\epsilon,\delta+\delta^{\prime})$-DP w.r.t.
+the verification function $f^{*}=f^{\gamma_{1},\gamma_{2}}$, finishing the proof.
+∎
+Remark 5.12 (Improved Basic Composition).
+An immediate consequence of Lemma 5.11 is an improvement in the well-known basic composition theorem, stating that the composition of $k$ randomized response mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $\left(\sum_{i=1}^{k}\epsilon_{i},\sum_{i=1}^{k}\delta_{i}\right)$-DP.
+By separately composing $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ using the original basic composition and composing $\mathrm{RR}_{0,\delta_{1}},\dots,\mathrm{RR}_{0,\delta_{k}}$ using Lemma 5.9, we can conclude that the composition of $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $\left(\sum_{i=1}^{k}\epsilon_{i},1-\prod_{i=1}^{k}(1-\delta_{i})\right)$-DP, where $1-\prod_{i=1}^{k}(1-\delta_{i})\leq\sum_{i=1}^{k}\delta_{i}$ always holds.
+Proof of Theorem 5.4.
+Let $f=f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta^{\prime}}$ and $f^{\mathrm{RR}}=f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta^{\prime},RR}$.
+The proof proceeds in two steps: (1) We design an IPM $\mathcal{P}$ and prove that for each $b\in\{0,1\}$, the IMs $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are equivalent.
+(2) By Lemma 5.11, the views of any adversary interacting with $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta+\delta^{\prime})$-indistinguishable.
+Therefore, combining (1) and (2) with the post-processing lemma, it follows that the views of any adversary interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ are $(\epsilon,\delta+\delta^{\prime})$-indistinguishable.
+This in turn implies that the $f$-concurrent composition of continual mechanisms is $(\epsilon,\delta+\delta^{\prime})$-DP.
+We are left with showing (1).
+In $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$, the IPM $\mathcal{P}$ receives from the left either (i) pairs of identical creation queries, or (ii) messages of the form $((q,j),(q^{\prime},j))$, where $q$ and $q^{\prime}$ are queries to the $j$-th created mechanism.
+Let $(\alpha_{j},\alpha_{j})$ be the $j$-th pair of creation queries, with $\alpha_{j}=(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})$.
+Also, for any $\epsilon^{\prime\prime}$ and $\delta^{\prime\prime}$, let $\beta_{\epsilon^{\prime\prime},\delta^{\prime\prime}}$ denote the creation query for the mechanism $\mathrm{RR}_{\epsilon^{\prime\prime},\delta^{\prime\prime}}$.
+The state of $\mathcal{P}$ is a pair $(n,\sigma)$, where $n$ is a counter (initialized to 0) and $\sigma$ is a sequence (initialized to an empty sequence).
+Upon receiving $(\alpha_{i},\alpha_{i})$ from the left, $\mathcal{P}$ appends the tuple $(T_{j},t_{j},\epsilon_{j}^{\prime},\delta_{j}^{\prime},w_{j}=\text{null})$ to $\sigma$, where $\mathcal{T}_{j}$ is the IPM corresponding to $\mathcal{M}_{j}^{\prime}$ from Corollary 5.7, $t_{j}$ is $\mathcal{T}_{j}$’s initial state, $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ are the privacy parameters associated with $\mathcal{M}_{j}^{\prime}$, and $w_{j}$ is an initially null variable that will later be replaced by an outcome of a randomized response mechanism.
+$\mathcal{P}$ then returns the acknowledgment message $\top$ as response to its left mechanism.
+We note that although the privacy parameters $\epsilon_{1},\dots,\epsilon_{k}$ in the statement of Theorem 5.4 are assumed to be strictly positive, $\epsilon_{j}^{\prime}$ might be zero.
+However, Corollary 5.7 is only applicable when $\epsilon_{j}^{\prime}>0$.
+To fix this technical issue, if the privacy parameter $\epsilon_{j}^{\prime}$ of the $j$-th creation query is zero, we change it to a small positive constant, e.g., $0.1$.
+Since $\mathcal{M}_{j}^{\prime}$ satisfies $(0,\delta_{j}^{\prime})$-DP w.r.t.
+$f_{j}^{\prime}$, it also satisfies $(0.1,\delta_{j}^{\prime})$-DP w.r.t.
+$f_{j}^{\prime}$.
+Thus, by Corollary 5.7, there exists an IPM $\mathcal{T}_{j}$ interacting with $\mathrm{IRR}_{0.1,\delta_{j}^{\prime}}$ and simulating $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$.
+By Definition 5.3 (ii), since $0\not\in\{\epsilon_{1},\dots,\epsilon_{k}\}$, the adversary always asks pairs of identical queries from $\mathcal{M}_{j}^{\prime}$, implying that $\mathcal{T}_{j}$ interacts with its right mechanism at most once.
+The IPM $\mathcal{T}_{j}$ expects $\mathrm{IRR}_{0.1,\delta_{j}^{\prime}}$ as its right mechanism, and $\mathrm{IRR}_{0.1,\delta_{j}^{\prime}}$ generates its first response only based on $\delta_{j}^{\prime}$, and thus changing the value of $\epsilon_{j}^{\prime}$ from zero to the arbitrary constant $0.1$ is unimportant.
+Upon receiving a left message of the form $((q,j),(q^{\prime},j))$, the IPM $\mathcal{P}$ sends $(q,q^{\prime})$ as a left message to $\mathcal{T}_{j}$.
+Upon receipt, $\mathcal{T}_{j}$ may interact with its right mechanism.
+As described below, $\mathcal{P}$ simulates the interaction of $\mathcal{T}_{j}$ with its right mechanism until $\mathcal{T}_{j}$ sends a left message $m$.
+$\mathcal{P}$ then returns $m$ as a response to its left mechanism.
+By Corollary 5.7, $\mathcal{T}_{j}$ sends at most two right messages.
+When $\mathcal{T}_{j}$ sends a right message for the first time (i.e., when $w_{j}=\text{null}$), $\mathcal{P}$ requests the creation of $\mathrm{RR}_{0,\delta_{j}^{\prime}}$ by sending the message $(\beta_{0,\delta_{j}^{\prime}},\beta_{0,\delta_{j}^{\prime}})$ to its right mechanism $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$.
+$\mathcal{P}$ discards the acknowledgment response $\top$, increments $n$ by one, and sends the message $\left((0,n),(1,n)\right)$ to its right mechanism.
+We will later show that $\mathcal{V}[f^{\mathrm{RR}}]$ never halts the communication.
+Given the initial state $b\in\{0,1\}$, upon receiving $\left((0,n),(1,n)\right)$, the identifier $\mathcal{I}(b)$ forwards $(b,n)$ to ExtConComp, which in turn sends $b$ to $\mathrm{RR}_{0,\delta_{j}^{\prime}}$.
+The response of $\mathrm{RR}_{0,\delta_{j}^{\prime}}$ is then forwarded to $\mathcal{P}$ unchanged.
+Let $(\tau_{j},z_{j})\in\{\top,\bot\}\times\{0,1\}$ denote this response.
+$\mathcal{P}$ sets $w_{j}=(\tau_{j},z_{j})$ and gives $\tau$ as a right response to $\mathcal{T}_{j}$.
+Note that if $\tau_{j}=\bot$, the bit $b$ is exposed and $z_{j}=b$.
+When $\mathcal{T}_{j}$ sends a second right message (i.e., when $w_{j}\neq\text{null}$), $\mathcal{P}$ proceeds as follows:
+  • If $\tau_{j}=\bot$, then $\mathcal{P}$ directly returns $z_{j}$ as the right response to $\mathcal{T}_{j}$, thereby revealing the initial state of $\mathcal{I}$.
+  • Otherwise, $\mathcal{P}$ requests the creation of $\mathrm{RR}_{\epsilon_{j}^{\prime},0}$ by sending $(\beta_{\epsilon_{j}^{\prime},0},\beta_{\epsilon_{j}^{\prime},0})$ to its right mechanism $\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$.
+As before, $\mathcal{P}$ drops the acknowledgment response $\top$, increments the counter $n$, and sends $\left((0,n),(1,n)\right)$ to its right mechanism.
+In this case, $\mathrm{RR}_{\epsilon_{j}^{\prime},0}$ receives the initial state of $\mathcal{I}$ as input and returns a response $(\tau_{j}^{\prime},z_{j}^{\prime})$, which is forwarded unchanged to $\mathcal{P}$.
+By definition $\tau_{j}^{\prime}=\top$ with probability $1$.
+Finally, $\mathcal{P}$ sends $z_{j}^{\prime}$ as a right response to $\mathcal{T}_{j}$.
+By the definition of $f$, there are at most $k$ mechanisms among the creation queries sent by $\mathcal{V}[f]$ that receive messages of the form $((q,j),(q^{\prime},j))$ with $q\neq q^{\prime}$, and the parameter $\epsilon_{j}^{\prime}$ of these mechanisms form a sub-multiset of $\epsilon_{1},\dots,\epsilon_{k}$.
+By Corollary 5.7, only the IPM $\mathcal{T}_{j}$ corresponding to these mechanisms may send a second message to their right mechanisms.
+Moreover, by the definition of $f$, the parameters $\delta_{j}^{\prime}$ of all mechanisms satisfy $1-\prod_{j}(1-\delta_{j}^{\prime})\delta^{\prime}$.
+Thus, by the design of $\mathcal{P}$, the verifier $\mathcal{V}[f^{\mathrm{RR}}]$ never halts the communication and forwards all messages to the opposite side unchanged.
+Comparing Definition 2.9 and Definition 2.3, we observe that the left responses generated by each $\mathcal{T}_{j}$ are identically distributed to those that would be produced if it were interacting directly with the IM $\mathrm{IRR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$.
+Thus, by Corollary 5.7, each $\mathcal{T}_{j}$ simulates $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ identically, implying that the IMs $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!\mathcal{V}[f^{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are equivalent.
+∎
 5.3. Composition Theorem for Restricted Verification Functions.
-In Section 5.2, we showed that for any positive $\delta$, the concurrent $1$-sparse parallel composition of CMs with the privacy parameter $(0,\delta)$ fails to preserve privacy.
+In Theorem 5.2, we showed that for any positive $\delta$, the concurrent $1$-sparse parallel composition of CMs with the privacy parameter $(0,\delta)$ fails to preserve privacy.
 In our counterexample, the adversary leveraged the first outcome of a $(0,\delta)$-DP CM to decide whether to spend its budget on issuing non-identical queries to that mechanism.
@@
 Definition 5.13 (First-Pair Consistent).
-A verification function $f$ is first-pair consistent if the following holds: given an input message sequence $m_{1},\dots,m_{t}$, $f$ returns $\bot$ if $m_{1}$ is a pair of identical messages, i.e., $m_{1}=(m^{\prime},m^{\prime})$ for some $m^{\prime}$, while there exists $2\leq j\leq t$ such that $m_{j}$ is not a pair of identical messages.
+A verification function $f$ is first-pair consistent if the following holds: for every input message sequence $m_{1},\dots,m_{t}$, $f(m_{1},\dots,m_{t})=\bot$ whenever $m_{1}$ is a pair of identical messages, i.e., $m_{1}=(m^{\prime},m^{\prime})$ for some $m^{\prime}$, and there exists $2\leq j\leq t$ such that $m_{j}$ is not a pair of identical messages.
 For instance, let $\mathcal{M}$ be an IM that is $(\epsilon,\delta)$-DP w.r.t.
@@
 By definition, $f_{\sim}$ is first-pair consistent.
-To formally define the restricted variant of the concurrent parallel composition of CMs, we introduce the verification function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$, where $FPC$ stands for first-pair consistency.
-Definition 5.4 ($f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$).
-For any input message sequence $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$ returns $\top$ if and only if both of the following conditions hold:
+To formally define the restricted variant of the concurrent parallel composition of CMs, we introduce the verification function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,\mathit{FPC}}$, where $FPC$ stands for first-pair consistency.
+Definition 5.14 ($f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,\mathit{FPC}}$).
+For any $t\in\mathbb{N}$ and any input message sequence $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,\mathit{FPC}}$ returns $\top$ if and only if both of the following conditions hold:
   1. (i)
@@
     For every message $m_{j}$ of the form $(\mathcal{M}^{\prime},s^{\prime},\epsilon^{\prime},\delta^{\prime},f^{\prime})^{2}$, the verification function $f^{\prime}$ is first-pair consistent.
-We refer to the $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$-concurrent composition of CMs as the FPC concurrent $k$-sparse parallel composition of CMs with privacy parameters $\left((\epsilon_{i},\delta_{i}\right)_{i=1}^{k}$.
+We refer to the $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,\mathit{FPC}}$-concurrent composition of CMs as the FPC concurrent $k$-sparse parallel composition of CMs with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$.
 Theorem 5.15.
-For $\epsilon,\epsilon_{1},\dots,\epsilon_{k}\geq 0$ and $0\leq\delta,\delta_{1},\dots,\delta_{k}\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the FPC concurrent $k$-sparse parallel composition of CMs with the privacy parameters $(\epsilon_{1},0),\dots,(\epsilon_{k},0)$ is also $(\epsilon,\delta)$-DP against an adaptive adversary.
+For $\epsilon_{1},\dots,\epsilon_{k}>0$, $\epsilon\geq 0$ and $0\leq\delta,\delta_{1},\dots,\delta_{k}\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the FPC concurrent $k$-sparse parallel composition of CMs with the privacy parameters $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ is also $(\epsilon,\delta)$-DP against an adaptive adversary.
 Proof.
-We denote $f$ as shorthand for $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$.
-This proof is identical to the proof of Theorem 5.2, except for the following modification in the definition of $\mathcal{P}$: In the proof of Theorem 5.2, each $\mathcal{T}_{j}$ behaved identically to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ if $\delta_{j}^{\prime}>0$ and was defined according to Lemma 4.7 otherwise.
-In this proof, each $\mathcal{T}_{j}$ behaves identically to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ if the first message of the form $\left((q_{0},j),(q_{1},j)\right)$ satisfies $q_{0}=q_{1}$, and it is defined by Lemma 4.6 otherwise.
-To implement this, $\mathcal{P}$’s state must include an additional sequence $\sigma_{3}$, storing a flag for each created mechanism, indicating whether the mechanism has previously received any messages.
-When $\mathcal{P}$ receives the creation query pair $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$ from the verifier, it adds the pair $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime})$ to $\sigma_{2}$, appends a 00 to the end of $\sigma_{3}$, and returns $\top$.
-Upon receiving a message of the form $\left((q_{0},j),(q_{1},j)\right)$, if the $j$-th element of $\sigma_{3}$ is 00, $\mathcal{P}$ sets this element to $1$ and replaces the entry $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime})$ in $\sigma_{2}$ with $(\mathcal{T}_{j},t_{j})$, defined as follows:
-  • If $q_{0}=q_{1}$, then $\mathcal{T}_{j}$ is an IPM with the single initial state $s_{j}^{\prime}$ that never interacts with its right mechanism and interacts with its left mechanism exactly as $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$.
+We denote $f$ as shorthand for $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,\mathit{FPC}}$.
+The proof is identical to the proof of Theorem 4.11, except for the definition of IPMs $\mathcal{T}_{j}$ that are executed by the IPM $\mathcal{P}$: In the proof of Theorem 4.11, each $\mathcal{T}_{j}$ was defined according to Lemma 4.6.
+In this proof, each $\mathcal{T}_{j}$ is defined by Corollary 5.7 if the first message of the form $\left((q_{0},j),(q_{1},j)\right)$ satisfies $q_{0}\neq q_{1}$, and it behaves identically to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ otherwise.
+To implement this, the sequence $\sigma_{2}$ in $\mathcal{P}$’s state must include an additional boolean variable $u_{j}$ for each created mechanism $\mathcal{M}_{j}^{\prime}$, indicating whether the mechanism has previously received any messages.
+Specifically, when $\mathcal{P}$ receives the creation query pair $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$ from the verifier, instead of adding $(\mathcal{T}_{j},t_{j},\epsilon_{j}^{\prime},\delta_{j}^{\prime})$, it adds the tuple $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},u_{j}=0)$ to $\sigma_{2}$.
+Upon receiving a left message of the form $\left((q_{0},j),(q_{1},j)\right)$, if $u_{j}=0$, before responding, $\mathcal{P}$ replaces the $j$-th entry $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},u_{j}=0)$ in $\sigma_{2}$ with $(\mathcal{T}_{j},t_{j},\epsilon_{j}^{\prime},\delta_{j}^{\prime},1)$, where $T_{j}$ is an IPM defined as follows and $t_{j}$ is its initial state:
+  • If $q_{0}=q_{1}$, then $\mathcal{T}_{j}$ is an IPM with the single initial state $\left((),0,s_{j}^{\prime}\right)$ that never interacts with its right mechanism and interacts with its left mechanism exactly as $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$.
+That is, upon receiving a right message, $\mathcal{T}_{j}$ halts the communication, and upon receiving a left message, it updates its state and responds identically to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$.
   • If $q_{0}\neq q_{1}$, then $\mathcal{T}_{j}$ is the IPM defined by Lemma 4.6 corresponding to $\mathcal{M}_{j}^{\prime}$.
-$\mathcal{T}_{j}$ with the first definition never requires a sample from $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, while a $\mathcal{T}_{j}$ with the second definition may require at most one sample (see Remark 4.8).
-We will show that the privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ corresponding to $\mathcal{T}_{j}$s with the second definition form a sub-multiset of $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$:
+To generate a response to a left message of the form $\left((q_{0},j),(q_{1},j)\right)$, as before, $\mathcal{P}$ forwards $(q_{0},q_{1})$ to $\mathcal{T}_{j}$ as a left message and provide a properly chosen $r_{i}$ as a right response if needed (see the proof of Theorem 4.11).
 By Definition 5.14 (ii), we know that for each index $j$, if $q_{0}=q_{1}$ in the first message of the form $\left((q_{0},j),(q_{1},j)\right)$, then $q_{0}=q_{1}$ in all subsequent messages of this form.
-Thus, by Definition 5.4 (i) and Definition 5.1 (ii), the privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ corresponding to $\mathcal{T}_{j}$s with the second definition (which receive at least one pair of non-identical queries) form a sub-multiset of $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$.
+In this case, by the above definition, for each $b\in\{0,1\}$, the mechanism $\mathcal{T}_{j}$ never interacts with $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ and simulates $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ identically.
+Moreover, when $q_{0}\neq q_{1}$ in the first message, by Definition 5.14 (i) and Definition 5.1 (ii), the privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ corresponding to $\mathcal{T}_{j}$s with the second definition form a sub-multiset of $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$.
+Therefore, whenever an IPM $\mathcal{T}_{j}$ requires a sample from $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, $\mathcal{P}$ can find such sample in $\sigma_{1}$ that is unused.
 ∎
@@
 Moreover, to ask a query $q$ from the $i$-th created IM, the adversary sends $\left((q,i),(q,i)\right)$.
-Since each $f_{\sim_{i}}$ is first-pair consistent, Theorem 5.5 applies to the concurrent $k$-sparse parallel composition of IM with privacy parameters $\left((\epsilon_{i},\delta_{i}\right)_{i=1}^{k}$ (that are differentially private w.r.t.
+Since each $f_{\sim_{i}}$ is first-pair consistent, Theorem 5.15 applies to the concurrent $k$-sparse parallel composition of IM with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$ (that are differentially private w.r.t.
 neighbor relations).
@@
 Since Theorem 5.15 applies to all IMs, it also holds for the concurrent $k$-sparse parallel composition of NIMs with privacy parameters $\left((\epsilon_{1},\delta_{i}\right)_{i=1}^{k}$.
-In the next section, we compare this result to the classical parallel composition theorem.
+We note that the restricted version of Theorem 5.15 for the $k$-sparse parallel composition of (unbounded number of) NIMs was previously proved in [20] as an intermediate lemma.
+In the next section, we compare the $k$-sparse parallel composition of IMs to the classical parallel composition theorem for NIMs.
 5.3.2. Comparison with the Parallel Composition of NIMs.
@@
 a neighbor relation $\sim$ can be modeled by a CM that is $(\epsilon_{0},\delta_{0})$-DP w.r.t.
-$f_{\sim}$, which is a first-pair consistent verification function.
+$f_{\sim}$, which is a first-pair consistent verification function (see the example below Definition 5.13).
 Thus, setting $\epsilon_{1}=\dots=\epsilon_{k}=\epsilon_{0}$, $\delta_{1}=\dots=\delta_{k}=\delta_{0}$ in Theorem 5.15 recovers and generalizes the classic parallel composition theorem.
@@
 Combining this result with the concurrent composition theorems provided in the previous sections, we can avoid complicated privacy analysis for many CMs.
+To obtain rigorous results, our statements and proofs must be presented in a technical and somewhat complex form.
+In contrast, mechanisms are usually described more simply using pseudocode rather than message-passing state machines.
+In Section 6.3, we reformulate our results in terms of pseudocode, enabling new mechanisms to be analyzed directly without the need for reformulating to the setting of in this paper.
 6.1. Properties.
@@
 This means $\mathcal{P}$ can be expressed as a deterministic function that maps the sequence of received messages to its next output message.
-This assumption is not restrictive, as randomness can be modeled by a 00-DP sub-mechanism $\mathcal{M}_{i}$ that accepts a single message $\top$, and upon receiving it, generates a sufficiently long sequence of random bits or samples from a fixed distribution.
+This assumption is not restrictive, as randomness can be modeled by a 0-DP sub-mechanism $\mathcal{M}_{i}$ that accepts a single message $\top$, and upon receiving it, generates a sufficiently long sequence of random bits or samples from a fixed distribution.
 Recall that the goal of this section is to show $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-DP w.r.t.
@@
 $f$.
-That is, we need to show that the view of any adversary in the interactions with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable.
+That is, we need to show that the view of every adversary interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable.
 The messages received by $\mathcal{P}$ from $\mathcal{I}$ contain information about the bit used by $\mathcal{I}$ to filter the adversary’s pairs.
@@
 $f$ against adaptive adversaries.
-In Theorem 6.1, we assume that the view of any adversary interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable.
+In Theorem 6.1, we assume that the view of every adversary interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable.
 The goal is to show the same result for $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$.
@@
 $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}^{2}$ in Claim 6.4 is an IPM with a singleton initial state space.
-Thus, by Lemma 2.15, that $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$, is $(\epsilon,\delta)$-DP w.r.t.
-$\sim_{01}$ implies the view of any adversary in the interactions with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable, which finishes the proof.
+Thus, by Lemma 2.16, that for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable implies that for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable, which finishes the proof.
 ∎
+6.3. Simplified Formulation for Pseudocodes
+Continual mechanisms are typically not described as state machines exchanging messages.
+Instead, they are presented in the form of pseudocode: at each time step, the pseudocode receives an input, updates its internal variables, and produces an output.
+For clarity and practicality, we restate the results of this section in the pseudocode framework.
+Decision vs. Auxiliary Variables.
+We partition the variables of a continual mechanism $\mathcal{M}$ into two categories:
+  • Decision variables: These are either outputs of submechanisms or variables that are computed solely based on decision variables.
+Intuitively, it follows that they are post-processing of the privatized outputs.
+  • Auxiliary variables: These are all remaining variables. The values of these variables may depend on the mechanism’s inputs as well as on both decision and auxiliary variables.
+Note.
+For simplicity, we assume pseudocodes are deterministic as any randomness could be modeled as a submechanism call.
+For instance, sampling from the uniform distribution could be represented as invoking a 0-DP submechanism that outputs such a sample on a fixed input.
+Properties.
+For $t\in\mathbb{N}$, given two sequences $\sigma^{0}=(m_{1}^{0},\dots,m_{t}^{0})$ and $\sigma^{1}=(m_{1}^{1},\dots,m_{t}^{1})$, their combination is defined as $\big((m_{1}^{0},m_{1}^{1}),\dots,(m_{t}^{0},m_{t}^{1})\big)$.
+Recall the three properties from Section 6.1.
+We reformulate them for pseudocodes as follows:
+  1. (1)
+    Destination property.
+A pseudocode satisfies this property if (i) the decision of whether to produce an output, create a new submechanism, or interact with an existing one, (ii) the choice of the new submechanism in the second case, and (iii) the selection of the submechanism to interact with in the third case is made solely based on decision variables.
+  2. (2)
+    Response property. A pseudocode satisfies this property if its output is determined solely by decision variables.
+  3. (3)
+    $f\to f^{\prime}$ property.
+Let $f$ and $f^{\prime}$ be verification functions.
+A pseudocode satisfies this property if, for every $t\in\mathbb{N}$ and every pair of input sequences $(\sigma^{0},\sigma^{1})$ of length $t$ forming an $f$-valid combination, the following holds: Assume that all decision variables of $\mathcal{M}$ are updated identically through the end of time step $t$.
+For each $b\in\{0,1\}$, define $\sigma^{b}$ to denote the sequence of creation queries and inputs to submechanisms when processing $\sigma^{b}$.
+Here, a creation query is a tuple specifying a created submechanisms, its privacy parameters, and associated verification functions.
+Also, an input $m$ to the $j$-th submechanism is represented by the pair $(m,j)$.
+Then, the combination of $\sigma^{0}$ and $\sigma^{1}$ is required to be $f^{\prime}$-valid.
+Corollary 6.5.
+Let $f$ and $f^{\prime}$ be two verification functions.
+For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, suppose the $f^{\prime}$-concurrent composition of continual mechanisms is $(\epsilon,\delta)$-DP.
+Then, every continual mechanism whose pseudocode satisfies the destination, response, and $f\to f^{\prime}$ properties, defined above, is $(\epsilon,\delta)$-DP w.r.t.
+$f$.
 7. Application (Continual Monotone Histogram Query Problem)
-We next show an application of our technique to the continual monotone histogram query problem. The continual monotone histogram query problem is defined as follows.
+We next show an application of our technique to the continual monotone histogram query problem.
+So far, we defined differential privacy (DP) with respect to (w.r.t.) verification functions.
+To be consistent with the literature, in this section we use the notion of DP w.r.t.
+“neighboring input sequences” instead:
+Remark 7.1.
+Let $\sim$ be a binary relation on a family of message sequences.
+Throughout this section, we say that a mechanism is $(\epsilon,\delta)$-DP w.r.t.
+the neighbor relation $\sim$ (or simply it is $(\epsilon,\delta)$-DP) if the mechanism is $(\epsilon,\delta)$-DP w.r.t.
+to the verification function $f_{\sim}$ against adaptive adversaries, where $f_{\sim}$ is defined as follows: For every message sequence $m_{1},\dots,m_{t}$, $f_{\sim}(m_{1}\dots,m_{t})=\top$ if and only if the message sequence is suitable (implying $m_{j}=(m_{j}^{0},m_{j}^{1})$ for every $j\in\{1,\dots,t\}$) and the sequences $\sigma_{0}=(m_{1}^{0},\dots,m_{t}^{0})$ and $\sigma_{1}=(m_{1}^{1},\dots,m_{t}^{1})$ satisfy $\sigma_{0}\sim\sigma_{1}$.
 Definition 7.2.
-Let $d\in\mathbb{N}$ and let $q:\mathbb{N}^{d}\rightarrow\mathbb{N}$ be a function that is monotone in its input.
-The input is a finite stream of row vectors $x^{1},\dots,x^{T}$ with $x^{t}\in\{0,1\}^{d}$ for all $t\in[T]$.
-Let $\sim$ be a neighbor relation on ($\left(\{0,1\}^{d}\right)^{*}$ where two input streams are neighboring if they are identical except for a single step.
-The output of the (differentially private) continual monotone histogram query problem at each time step $t$ is (an additive approximation to) $q(t)=q((\sum_{\ell=1}^{t}x_{i}^{\ell})_{i\in[d]})$.
-Next we define query sensitivity.
+Let $d,T\in\mathbb{N}$ and let $q:\mathbb{N}^{d}\to\mathbb{N}$ be a function that is monotone in its input.
+In the (differentially private) continual monotone histogram query problem, a continual mechanism $\mathcal{M}$ receives $T$ inputs of row vectors $x^{1},\dots,x^{T}$, each in $\{0,1\}^{d}$.
+Define $\sim$ to be a neighbor relation on $\left(\{0,1\}^{d}\right)^{*}$ where two input streams are neighboring if they are identical except for a single step.
+Differential privacy for $\mathcal{M}$ is defined with respect to this neighbor relation (see Remark 7.1).
+The output of $\mathcal{M}$ at each time step $t$ is (an additive approximation to) $q(t)=q((\sum_{\ell=1}^{t}x_{i}^{\ell})_{i\in[d]})$.
 Definition 7.3.
-The function $q:\{0,1\}^{d}\rightarrow\mathbb{R}$ is $1$-sensitive if $\max_{x\sim y}|q(x)-q(y)|$ is at most $1$.
-Henzinger, Sricharan, and Steiner [12] presented the best known differentially private mechanism for the monotone histogram query problem, where where the query is $1$-sensitive.
+Recall function $q:\{0,1\}^{d}\to\mathbb{R}$ and neighbor relation $\sim$ from Definition 7.2. $q$ is said to be $1$-sensitive if $\max_{x^{0}\sim x^{1}}|q(x^{0})-q(x^{1})|$ is at most $1$.
+Henzinger, Sricharan, and Steiner [13] presented the best known differentially private mechanism for the monotone histogram query problem, where the query is $1$-sensitive.
 For $\epsilon$-differential privacy it has, with probability at least $1-\beta$, an additive error of $O((d\log^{2}(dq^{*}/\beta)+\log T)\epsilon^{-1})$, and for $(\epsilon,\delta)$-differential privacy, it has an additive error of $O((d\log^{1.5}(dq^{*}/\beta)+\log T)\epsilon^{-1}\log\frac{1}{\delta})$, where $q^{*}=\max_{t\in[T]}q(t)$.
@@
 7.1. Overview of Mechanism.
-[…]
-[…]
+The mechanism HSS, described in Algorithm 4, employs a variant of the Sparse Vector Technique mechanism, SVT, and a $d$-dimensional counter mechanism, d-counter, as submechanisms.
+We start by defining these mechanisms and then explain HSS.
+Sparse vector technique mechanism (SVT).
+The mechanism SVT is described in Algorithm 3 using two operations.
+The first, $\texttt{SVT-init}(\epsilon,q,h)$, initializes the mechanism with a privacy parameter $\epsilon$, a query $q$, and an initial $d$-dimensional vector $h$.
+This procedure indeed describes the creation of a continual mechanism $\textrm{SVT}_{h,\epsilon,q}$.
+The second operation, $\texttt{SVT-update}(x,\texttt{Thresh})$, maps inputs $(x,\texttt{Thresh})\in\{0,1\}^{d}\times\mathbb{R}$ to outputs in $\{\top,\bot\}$.
+The mechanism maintains a running sum of all inputs $x$ and, in a differentially private manner, checks whether applying $q$ to the running sum yields a value at least Thresh.
+If so, it outputs $\top$ and halts.
+Otherwise, it outputs $\bot$ and continues.
+These two procedures are given in detail in Algorithm 3.
+The command $\mathrm{Lap}(b)$ denotes a random sample from the Laplace distribution with scale parameter $b$.
+To clarify the connection between these pseudocodes and our definition of continual mechanisms (Definition 3.1), we note that the state of $\textrm{SVT}_{h,\epsilon,q}$ is the tuple of its variables.
+When $\texttt{SVT-init}(\epsilon,q,h)$ is invoked, the continual mechanism $\textrm{SVT}_{h,\epsilon,q}$ with the single initial state $(\epsilon,q,h,\tau=\text{null})$ is created.
+In the pseudocode of SVT-init, the variable $\tau$ takes value.
+In our setting, such initialization is considered to happen when $\textrm{SVT}_{h,\epsilon,q}$ receives its first input.
+The SVT-update operation then describes how $\textrm{SVT}_{h,\epsilon,q}$ maps its current state (i.e., set of variables) and an input message $m=(x,\texttt{Thresh})\in\{0,1\}^{d}\times\mathbb{R}$ to a new states (with updated variables) and a response $m^{\prime}\in\{\top,\bot\}$.
 Algorithm 3 Sparse vector technique for histogram queries
@@
 end procedure
-procedure SVT-call($x,\texttt{Thresh}$):
+procedure SVT-update($x,\texttt{Thresh}$):
   $h_{i}=h_{i}+x_{i}$ for all $i\in[d]$
@@
 end procedure
-Let $f_{\sim}$ be a verification function that given a message sequence $m_{1},\dots,m_{t}$, returns $\top$ if and only if for each $j\in[t]$, the message $m_{j}$ is of the form $(m_{j}^{0},m_{j}^{1})$ and the message sequences $\sigma_{0}=(m_{1}^{0},\dots,m_{t}^{0})$ and $\sigma_{1}=(m_{1}^{1},\dots,m_{t}^{1})$ satisfy $\sigma_{0}\sim\sigma_{1}$.
-As part of their privacy proof, the following is shown in [12], which we restate in our own language.
+As part of their privacy proof, the following is shown in [13]:
 Lemma 7.4.
-If $q$ is a $1$-sensitive function, then for every $\epsilon$ and $h$, the mechanism $\textrm{SVT}_{\epsilon/6,q,h}$ is $\epsilon/3$-DP w.r.t. $f_{\sim}$.
-For positive integers $t$ and $j$, $\beta\in(0,1]$, and $\epsilon>0$, let $\gamma(t,j,\beta,\epsilon)$ and $\xi(t,j,\beta,\epsilon)$ are some deterministic functions.
-Given an input stream $x_{1},x_{2},\dots$ with $x_{i}\in\{0,1\}^{d}$, the algorithm partitions the input stream into subsequences, called intervals.
-At the end of each interval, it gives the $d$-dimensional sum of all the inputs of the interval to a differentially-private histogram mechanism H.
-Note that H can be used directly to solve the problem; however, the new mechanism HSS has better accuracy bounds in many cases.
-[…]
-[…]
+If $q$ is a $1$-sensitive function, then for every $\epsilon$ and $h$, the mechanism $\textrm{SVT}_{\epsilon/6,q,h}$ is $\epsilon/3$-DP w.r.t. the neighbor relation $\sim$ in Definition 7.2.
+$d$-dimensional continual counter.
+The mechanism d-counter is described by two operations.
+The first, $\texttt{d-counter-init}(\epsilon)$, creates $d$ differentially private continual binary counters with privacy parameter $\epsilon/d$.
+The second, $\texttt{d-counter-update}(x)$, gives the elements of its $d$-dimensional input $x$ to their corresponding binary counters and returns the aggregated outputs as a $d$-dimensional vector.
+We denote this continual mechanism by $\texttt{d-counter}_{\epsilon}$.
+In another version, $\texttt{d-counter-init}(\epsilon,\delta)$ creates $d$ differentially private continual binary counters with privacy parameters $\epsilon/d$ and $\delta/d$.
+We represent this continual mechanism by $\texttt{d-counter}_{\epsilon,\delta}$.
+By definition, mechanisms $\texttt{d-counter}_{\epsilon}$ and $\texttt{d-counter}_{\epsilon,\delta}$ are the concurrent composition of $d$ many $\epsilon/d$- and $(\epsilon/d,\delta/d)$-DP fixed private continual counters.
+We can assume that these continual counters return an integer value at each point of time, and thus they have discrete answer distributions.
+If the output space is real, we use a counter with integer outputs that post-processes the original counter by rounding its output up to the nearest integer.
+By the post-processing lemma, the privacy guarantees still hold and the additive error is only negligibly affected by at most $1$.
+Therefore, Theorem 4.9 is applicable and by the basic composition, we have:
+Corollary 7.5.
+Mechanisms $\texttt{d-counter}_{\epsilon}$ and $\texttt{d-counter}_{\epsilon,\delta}$ are $\epsilon$-DP and $(\epsilon,\delta)$-DP, respectively.
+Continual histogram query mechanism.
+The mechanism HSS, described in Algorithm 4, is initialized with a privacy parameter $\epsilon>0$ and an accuracy parameter $0<\beta\leq 1$ and receives a vector $x\in\{0,1\}^{d}$ as input at each time step.
+This mechanism partitions the input stream into subsequences, called intervals.
+At the end of each interval, it forwards the $d$-dimensional sum of all the inputs of the interval to the differentially private $d$-dimensional continual counter d-counter.
+(Note that d-counter can be used directly to solve the problem; however, HSS achieves better accuracy bounds in many cases.) The mechanism HSS maintains two $d$-dimensional vectors: (1) variable $c$, where $c_{i}$ stores the exact sum of coordinate $i$ within the current interval, and (2) variable $h$, where $h_{i}$ stores the approximate sum of coordinate $i$ over all previous intervals computed by d-counter.
+Both vectors are initialized to $0^{d}$.
+At the beginning of each interval (including the initial time step), HSS creates a fresh instance of the Sparse Vector Technique mechanism SVT with privacy parameter $\epsilon/6$, query $q$, and initial vector $h=0^{d}$ (denoted by $\textrm{SVT}_{\epsilon/6,q,h}$).
+The mechanism discards the previous instance.
+HSS keeps track of the current time step and the current interval using variables $t$ and $j$, respectively.
+In Algorithm 4, $\gamma$ and $\xi$ are some deterministic functions, mapping $t,j,\beta,\epsilon$ to a real value.
+The mechanism HSS also stores a threshold variable Thresh, initially set to $\gamma(1,1,\beta,\epsilon)$.
+Upon receiving an input $x$ at step $t$, the mechanism updates $c$ and invokes $\texttt{SVT-update}(x,\texttt{Thresh})$.
+If the response is $\top$, the current interval ends: $c$ is passed to d-counter, $h$ is updated with the output of d-counter, and $c$ is reset to $0^{d}$.
+Since SVT halts after returning $\top$, a new instance $\textrm{SVT}_{\epsilon/6,q,h}$ is created.
+In this case, HSS privately evaluates $q(c+h)$ via the Laplace noise and compares the noisy result to $\texttt{Thresh}-\xi(t,j,\beta,\epsilon)$.
+If the inequality holds, Thresh is suitably increased.
+Moreover, independent of the outcome of this comparison, Thresh is updated again to reflect the increase in the number of intervals.
+Then, regardless of the output of $\texttt{SVT-update}(x,\texttt{Thresh})$, HSS updates Thresh once more to account for the increased number of time steps.
+Finally, HSS outputs $h$.
 Algorithm 4 Mechanism for answering a monotone histogram query
-[…]
+procedure Hist-init($\epsilon,q,\beta$):
+  $\texttt{d-counter-init}(\epsilon/3)$ (or $\texttt{d-counter-init}(\epsilon/3,\delta)$)
   $h_{i}\leftarrow 0$ for all $i\in[d]$, $c_{i}\leftarrow 0$ for all $i\in[d]$
@@
   $j\leftarrow 1$
-[…]
+  $\texttt{SVT-init}(\epsilon/6,q,h)$
   $\texttt{Thresh}=\gamma(1,j,\beta,\epsilon)$
-for $t\in\mathbb{N}$ do
-     $c_{i}\leftarrow c_{i}+x^{(t)}_{i}$ for all $i\in[d]$,
-[…]
-         $h=\texttt{H-add}(c)$
+  $t\leftarrow 0$
+end procedure
+procedure Hist-update($x$):
+  $t\leftarrow t+1$
+  $c_{i}\leftarrow c_{i}+x_{i}$ for all $i\in[d]$,
+  if $\texttt{SVT-update}(x,\texttt{Thresh})==\top$ then
+   $h=\texttt{d-counter-update}(c)$
    $out\leftarrow q(h)$
-[…]
+   $\texttt{SVT-init}(\epsilon/6,q,h)$
    $c_{i}\leftarrow 0$ for all $i\in[d]$
@@
   end if
-     output $out$
   $\texttt{Thresh}=\texttt{Thresh}-\gamma(t,j,\beta,\epsilon)+\gamma(t+1,j,\beta,\epsilon)$
-end for
+  return outout
+end procedure
 7.2. New Privacy Analysis.
-Algorithm 4 describes a CM HSS with query set $\{0,1\}^{d}$ and answer set $\mathbb{N}^{d}$.
-The state of HSS consists of the variables $(h,c,j,out,\texttt{Thresh})$, initialized with $(0^{d},0^{d},1,q(0^{d}),\gamma(1,1,\beta,\epsilon))$.
-The summation $q(c+h)+\mathrm{Lap}(3/\epsilon)$ in this algorithm can be rewritten as an in instance of Laplace mechanism with privacy parameter $\epsilon/3$ and input $q(c+h)$.
-Consequently, HSS executes one instance of a histogram mechanism, H, an unbounded number of SVT instances, and an unbounded number of Laplace mechanisms.
-The Laplace mechanism with privacy parameter $\epsilon/3$ is a NIM that is known to be $\epsilon/3$-DP when receiving real values differing by at most $1$.
-In other words, the Laplace mechanism is a CM that is $\epsilon/3$-DP w.r.t.
-a verification function $f_{\mathrm{Lap}}$ that returns $\top$ if and only if it receives a single input message $m=(m_{0},m_{1})\in\mathbb{R}^{2}$ satisfying $|m_{0}-m_{1}|\leq 1$.
-The output of H is a vector in $\mathbb{R}^{d}$, while the output of the Laplace mechanism is a scalar in $\mathbb{R}$.
-Rounding up each real-valued output to the nearest integer does not affect privacy guarantees (by the post-processing lemma) but increases the additive error of each mechanism by at most $1$.
-This extra error is negligible compared to their original error.
-Therefore, we assume that both H and the Laplace mechanism have discrete answer distributions.
-Since the answer set of SVT is finite, it also has discrete answer distributions.
-By design, HSS uses the inputs $x^{(t)}$ only to determine the inputs of its sub-mechanisms, and every other action is taken solely based on the privatized outputs of the sub-mechanisms.
-This allows us to apply Theorem 6.1 to analyze the privacy of HSS based on the privacy parameters of its sub-mechanisms.
-By Lemma 7.3, each instance of SVT created by HSS is $\epsilon/3$-DP.
-Furthermore, for any two neighboring input streams, fixing the history of privatized outputs of the sub-mechanisms, HSS provides identical inputs to all but one SVT instance, which receives neighboring inputs.
-By Theorem 5.2, the parallel concurrent composition of all SVT instances is $\epsilon/3$-DP.
-Similarly, HSS gives identical real-valued inputs to all but one instance of the Laplace mechanism, where the differing inputs vary by at most $1$.
-This implies that the concurrent composition of all Laplace mechanisms is also $\epsilon/3$-DP.
-Each concurrent composition is itself a continual mechanism.
-Additionally, it is given that H is $\epsilon/3$-DP.
-By basic composition, the composition of three instances of $\mathrm{RR}_{\epsilon/3,0}$ is $\epsilon$-DP.
-Therefore, by Theorem 4.10, the concurrent composition of all mechanisms—H, SVT instances, and Laplace mechanisms—is $\epsilon$-DP.
-Applying Theorem 6.1, the mechanism HSS is $\epsilon$-DP.
-In the case where H is $(\epsilon/3,\delta)$-DP, since the composition of $\mathrm{RR}_{\epsilon/3,0}$, $\mathrm{RR}_{\epsilon/3,0}$, and $\mathrm{RR}_{\epsilon/3,\delta}$ is $(\epsilon,\delta)$-DP, the mechanism HSS is also $(\epsilon,\delta)$-DP.
-Both privacy guarantees provided in this section hold against adaptive adversaries.
-Acknowledgments
-1Monika Henzinger and Roodabeh Safavi were supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No.
-101019564), and the Austrian Science Fund (FWF) under grants DOI 10.55776/Z422, DOI 10.55776/I5982, and DOI 10.55776/P33775, with additional funding from the netidee SCIENCE Stiftung, 2020–2024.
-2Salil Vadhan was supported by NSF grant BCS-2218803, a grant from the Sloan Foundation, and a Simons Investigator Award.
-Work began while a Visiting Researcher at the Bocconi University Department of Computing Sciences, supported by Luca Trevisan’s ERC Project GA-834861.
-Appendix A Overview and Modification of [18]
-In this section, we overview and modify the results of [18].
-The original proof relies on the existence of a constant $T$ that upper bounds the number of queries adversaries ask; however, our modified construction does not.
-Moreover, an algorithmic construction in [18]’s proof needs the answer set to be finite.
-Using mathematical tools, we formalize their construction and weaken this assumption to having discrete answer distributions.
-We show:
-Lemma A.1.
+In Algorithm 4, the summation $q(c+h)+\mathrm{Lap}(3/\epsilon)$ can be rewritten as an instance of Laplace mechanism with privacy parameter $\epsilon/3$ and input $q(c+h)$.
+Consequently, the mechanism HSS executes:
+  1. (1)
+    one instance of the $d$-dimensional counter d-counter, which consists of $d$ instances of a scalar counter,
+  2. (2)
+    an arbitrary number of SVT instances, and
+  3. (3)
+    an arbitrary number of Laplace mechanisms.
+By Corollary 7.5, d-counter is $\epsilon/3$-DP (or $(\epsilon/3,\delta)$-DP when approximate counters are used).
+By Lemma 7.4, each SVT instance is $\epsilon/3$-DP.
+Finally, the Laplace mechanism with parameter $\epsilon/3$ is a non-interactive mechanism (NIM) that is well known to satisfy $\epsilon/3$-DP on neighboring real inputs differing by at most $1$.
+The idea is to show that the concurrent composition of multiple instances of these mechanisms created by HSS is $\epsilon$-DP (or $(\epsilon,\delta)$-DP when d-counter is approx DP).
+Then, we will then show that HSS satisfies the required properties in Corollary 6.5, and thus it has the same privacy guarantees as the concurrent composition.
+As concurrent compositions theorems in this paper require the composed mechanisms to have discrete answer distributions, we first show that we can assume the output space of d-counter and Laplace mechanism is discrete.
+Since SVT has a finite output space, its answer distributions are also discrete.
+Discretization of outputs.
+As argued earlier, the counters executed by d-counter outputs integer values, and thus d-counter has discrete answer distributions.
+Likewise, we can assume that the Laplace mechanism in HSS rounds its noisy real-valued output to the nearest integer.
+By the post-processing lemma, this rounding does not degrade privacy, and by definition, it increases the additive error by at most $1$, which is unimportant for the accuracy guarantees of [13].
+Hence, both d-counter and the Laplace mechanism can be treated as having discrete output distributions.
+Pseudocode privacy analysis.
+Recall from Section 6.3 that decision variables of a mechanism are either (i) outputs of submechanisms, or (ii) variables computed solely as functions of other decision variables.
+In Algorithm 4, the variables $h$, outout, $j$, $t$, and Thresh are decision variables, while $c$ is not: Variable $h$ denotes the privatized outcomes of d-counter.
+Variable outout equals $q(h)$, where $h$ is a decision variable.
+Index $j$ is initially $1$ and is increased by $1$ when the privatized output of SVT is $\top$.
+Time step $t$ is initially 0 and is increase by $1$ at each step, independent of the input and other variables.
+Finally, variable Thresh is initialized and updated to values determined by $\epsilon,\beta,j,t$, where $\epsilon,\beta$ are fixed and $j,t$ are decision variables.
+It is important to note that the decision of updating variables $h$ (and consequently outout) and Thresh depends only on the privatized output of SVT.
+We first show that the destination property holds.
+For that we have to show that the decision of whether to produce an output, create a new submechanism (and if so, which one) or interact with an existing one (and if so, which one) are solely based on decision variables.
+All submechanisms of HSS are instantiated either in the initial step (and, thus, whether it receives input is independent of non-decision variables and the inputs $x$) or based on outputs of the private submechanism SVT.
+The active SVT instance receives input at every step, and thus whether it receives input is independent of non-decision variables and the inputs $x$.
+Also, whether the mechanism d-counter gets updated depends only on the privatized output of SVT.
+Finally, HSS returns an output after the final update of Thresh, without any condition on the inputs $x$ or the non-decision variables.
+Thus, the destination property in Corollary 6.5 holds.
+Moreover, the final output of HSS is outout, which is a decision variable, so the response property in that corollary also holds.
+It remains to show HSS satisfies the $f\to f^{\prime}$ property, where $f$ is the verification function corresponding to the neighbor relation $\sim$ in Definition 7.2 and $f^{\prime}$ is a verification satisfying that the $f^{\prime}$-concurrent composition of the continual mechanisms is $(\epsilon,\delta)$-DP.
+Fix some values for all decision variables at every time step.
+Conditioned on these values, when HSS processes two neighboring input sequences, the resulting inputs to its submechanisms behave as follows:
+  • d-counter receives neighboring sequences,
+  • all but one instance of SVT receive identical inputs, with at most one receiving neighboring inputs, and
+  • all but one Laplace mechanism receive identical inputs, with at most one receiving neighboring inputs.
+Hence, the concurrent composition of all submechanisms of HSS can be decomposed into three concurrent compositions: (1) all Laplace mechanisms, (2) all SVT instances, and (3) the $d$ continual counters forming d-counter.
+The first two are the parallel compositions of purely differentially private mechanisms.
+By Corollary 5.5, both are $\epsilon/3$-DP.
+The third one is, by construction, the concurrent composition of $d$ fixed continual counters.
+In Corollary 7.5, we showed that this composition, i.e., d-counter, is $\epsilon/3$-DP (or $(\epsilon/3,\delta)$-DP when approx DP counters are used).
+By basic composition and Theorem 4.9, the overall concurrent composition of all three groups of mechanisms is $\epsilon$-DP when d-counter uses pure counters, and $(\epsilon,\delta)$-DP when approximate counters are used.
+(More precisely, it is DP w.r.t.
+a verification function $f^{\prime}$ that ensures all submechanisms but one SVT instance, one Laplace mechanism, and the $d$ counters of d-counter are assigned pairs of identical inputs, while the first and second inputs for each of these mechanisms form neighboring sequences.) We note that the concurrent compositions are themselves continual mechanisms and can be composed; an explicit example of composing two compositions in the message-passing model is given in the proof of Lemma 5.11.
+By Corollary 6.5, since HSS satisfies the destination, response, and neighboring-mapping properties, and since the concurrent composition of its submechanisms is $\epsilon$-DP (or $(\epsilon,\delta)$-DP), the mechanism HSS satisfies the same privacy guarantee.
+We recall that every privacy guarantees in this section holds against adaptive adversaries.
+8. Reduction to Interactive Randomized Response
+Lyu [19] shows that for every $(\epsilon,\delta)$-DP IM $\mathcal{M}$ and every pair of neighboring initial states $s_{0},s_{1}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ are equivalent.
+Meaning that, $\mathcal{P}$ can simulate $\mathcal{M}(s_{b})$ given access only to the output of a randomized response mechanism $\mathrm{RR}_{\epsilon,\delta}(b)$.
+In this section, inspired by this work, we construct a new IPM $\mathcal{P}$ that, rather than simulating $\mathcal{M}$ based on the outcome of a single instance of $\mathrm{RR}_{\epsilon,\delta}$, operates on the output of $\mathrm{RR}_{0,\delta}$ and only when necessary uses the output of $\mathrm{RR}_{\epsilon,0}$.
+In other words, instead of interacting with the NIM $\mathrm{RR}_{\epsilon,\delta}$, $\mathcal{P}$ interacts with the IM $\mathrm{IRR}_{\epsilon,\delta}$, and under certain conditions, it queries $\mathrm{IRR}_{\epsilon,\delta}$ at most once.
+In addition, our construction removes or relaxes the following implicit assumptions in Lyu’s proof:
+  • Lyu’s result relies on the existence of a constant $T$ that upper bounds the number of queries adversaries ask.
+This restriction comes from a backward definition of a so-called control function.
+We define a function which is the limit over this control function and, using monotone convergence, show that this “limit function” satisfies all the necessary conditions to be used in the rest of Lyu’s proof.
+This extends the result to an unbounded number of rounds of interaction.
+  • Lyu’s proof uses an algorithmic construction iterating over the answer set, implying that the answer set must be finite.
+Instead with a mathematical approach, we extend the proof to work with a countably infinite answer set.
+Our result is formalized in Lemma 5.6, which is restated below:
+Lemma (Restatement of Lemma 5.6).
 For $\epsilon>0$ and $0\leq\delta\leq 1$, let $\mathcal{M}:S_{\mathcal{M}}\times Q_{\mathcal{M}}\to S_{\mathcal{M}}\times A_{\mathcal{M}}$ be an $(\epsilon,\delta)$-DP IM w.r.t.
-a neighboring relation $\sim$.
+a neighbor relation $\sim$.
 Suppose that $\mathcal{M}$ has discrete answer distributions.
-Then, for any two neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,\delta}(b)$ are identical.
-Notation A.2.
+Then, for every two neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$:
+  • The mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\!\circ^{*}\!\mathrm{IRR}_{\epsilon,\delta}(b)$ are equivalent.
+  • For every $k\in\mathbb{N}$ and every query sequence $(q_{1},\dots,q_{k})\in Q_{\mathcal{M}}^{k}$, if the distributions of answers produced by $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to $(q_{1},\dots,q_{k})$ are identical, then when $\mathcal{P}\!\circ^{*}\!\mathrm{IRR}_{\epsilon,\delta}(b)$ receives the queries $q_{1},\dots,q_{k}$, the IPM $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}(b)$ only once.
+Otherwise, $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}(b)$ at most twice.
+Notation 8.1.
 For short, let $Q=Q_{\mathcal{M}}$ and $A=A_{\mathcal{M}}$.
-For any $T\in\mathbb{N}$, let $[T]=\{1,\dots,T\}$.
 For each $t\in\mathbb{N}$ and $b\in\{0,1\}$, we define the function $\mu_{t}^{b}:(Q\times A)^{t}\to[0,1]$ such that $\mu_{t}^{b}(q_{1},\dots,a_{t})$ denotes the probability of $\mathcal{M}(s_{b})$ returning answers $a_{1},\dots,a_{t}$ to the queries $q_{1},\dots,q_{t}$.
-Fixing an upper bound $T$ on the number of queries, Lyu [18] reduces the construction of $\mathcal{P}$ to defining a sequence of functions $\phi_{t}^{0}$ and $\phi_{t}^{1}$ for each $t\in[T]$, satisfying three conditions.
-They then inductively construct each $\phi_{t}^{b}$ and verify that it satisfies the required conditions.
+Fixing an upper bound $T$ on the number of queries, Lyu [19] reduces the construction of $\mathcal{P}$ to defining a sequence of functions $\phi_{t}^{0}$ and $\phi_{t}^{1}$ for $t\in[T]$ that satisfy three conditions.
+They then inductively construct $\phi_{t}^{b}$ and verify that it satisfies the required conditions.
 The definition of these conditions is based on a sequence of control function $\mathrm{L}_{t,T}^{0}$ and $\mathrm{L}_{t,T}^{1}$, which have backward recursive definitions for $t$ from $T$ to $1$.
-In Section A.1.1, we overview the definition of $\mathrm{L}_{t,T}^{b}$ and its properties.
-In Section A.1.2, we introduce a new control function $\mathrm{L}_{t}^{b}$ with no dependency on $T$.
-The goal is to use $\mathrm{L}_{t}^{b}$ instead of $\mathrm{L}_{t,T}^{b}$ in [18]’s proofs.
-So, in Section A.1.3, we show some properties for $\mathrm{L}_{t}^{b}$, required in later proofs.
-In Section A.2, we overview the reduction presented in [18], with the control functions $\mathrm{L}_{t,T}^{b}$ replaced by $\mathrm{L}_{t}^{b}$.
-In this section, we prove the reduction still holds with the new control functions.
-Finally, in Section A.3, we present a construction for $\phi_{t}$ and show it satisfies all three conditions.
-We note that this construction is identical in essence to one given in [18], but we extend the proof to work with weaker assumptions.
-A.1. Preliminaries
-A.1.1. Control Functions $\mathrm{Lower}_{t,T}$ and $\mathrm{L}_{t,T}$.
+In Section 8.1, we overview the definition of $\mathrm{L}_{t,T}^{b}$ and its properties.
+In Section 8.1.1, we introduce a new control function $\mathrm{L}_{t}^{b}$ with no dependency on $T$.
+The goal is to use $\mathrm{L}_{t}^{b}$ instead of $\mathrm{L}_{t,T}^{b}$.
+In Section 8.1.2, we show some properties for $\mathrm{L}_{t}^{b}$, required in later proofs.
+In Section 8.2, we summarize the reduction introduced in [19], with the control functions $\mathrm{L}_{t,T}^{b}$ replaced by $\mathrm{L}_{t}^{b}$.
+We also modify this reduction by adding a new condition to the original three that the functions $\phi_{t}$ must satisfy.
+This additional condition is used to limit the number of interactions between $\mathcal{P}$ and $\mathrm{IRR}_{\epsilon,\delta}(b)$.
+Finally, in Section 8.3, we provide a new construction of $\phi_{t}$ and show that it satisfies all four conditions.
+8.1. Control Functions $\mathrm{Lower}_{t,T}$ and $\mathrm{L}_{t,T}$.
 Lyu [19] introduces the following control function:
-Definition A.3 ($\mathrm{Lower}_{t,T}$).
+Definition 8.2 ($\mathrm{Lower}_{t,T}$).
 For $T\in\mathbb{N}$, $b\in\{0,1\}$, and $t\in[T]$, the function $\mathrm{Lower}_{t,T}^{b}:(Q\times A)^{t-1}\times Q\to[0,1]$ is defined recursively as follows: for each transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
@@
 We introduce another control function $\mathrm{L}_{t,T}$ based on $\mathrm{Lower}_{t,T}$:
-Definition A.4 ($\mathrm{L}_{t,T}^{b}$).
+Definition 8.3 ($\mathrm{L}_{t,T}^{b}$).
 For $T\in\mathbb{N}$, $b\in\{0,1\}$, and $t\in[T]$, the function $\mathrm{L}_{t,T}^{b}:(Q\times A)^{t}\to[0,1]$ is defined recursively as follows: for each transcript $(q_{1},\dots a_{t})\in(Q\times A)^{t}$,
@@
 The following corollary follows directly from the definitions of $\mathrm{Lower}_{t,T}^{b}$ and $\mathrm{L}_{t,T}^{b}$:
-Corollary A.5.
+Corollary 8.4.
 For every $T\in\mathbb{N}$, $b\in\{0,1\}$, $t\in[T]$, and $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
@@
 Lyu [19] shows the following two claims about $\mathrm{Lower}_{t,T}^{b}$:
-Claim A.6 ([18]).
+Claim 8.5 ([19]).
 For each $b\in\{0,1\}$, $T\in\mathbb{N}$, and $q_{1}\in Q$ $\mathrm{Lower}_{1,T}^{b}(q_{1})\leq\delta$.
-Claim A.7 ([18]).
-For each $b\in\{0,1\}$, $T\in\mathbb{N}$, $t\in[T]$ [^3] In the original claim in [18], $t$ is assumed to be at most $T-1$, which is a typo., and transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$, we have
+Claim 8.6 ([19]).
+For each $b\in\{0,1\}$, $T\in\mathbb{N}$, $t\in[T]$, [^7] In the original claim in [19], $t$ is assumed to be at most T1T-1, which is a typo.
+and transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$, we have
 $\mathrm{Lower}_{t,T}^{b}(q_{1},a_{1},\dots,q_{t})$ $-e^{-\epsilon}\mathrm{Lower}_{t,T}^{1-b}(q_{1},a_{1},\dots,q_{t})\leq\mu_{t-1}^{b}(q_{1},\dots,a_{t-1})-e^{-\epsilon}\mu_{t-1}^{1-b}(q_{1},\dots,a_{t-1}).$
-Claim A.6 combined with the definition of $\mathrm{L}_{t,T}^{b}$ imply:
-Corollary A.8.
+Claim 8.5 combined with the definition of $\mathrm{L}_{t,T}^{b}$ imply:
+Corollary 8.7.
 For each $b\in\{0,1\}$, and $T\in\mathbb{N}$, $\mathrm{L}_{0,T}^{b}()\leq\delta$.
-As we point out later, in some part of the proof, Lyu [18] wrongly refers to Claim A.7, while the following lemma is required. For completeness, we prove this lemma and use it in the next section.
-Lemma A.9.
+In some part of the proof, Lyu [19] wrongly refers to Claim 8.6, while the following lemma is required. For completeness, we prove this lemma and use it in the next section.
+Lemma 8.8.
 For each $b\in\{0,1\}$, $T\in\mathbb{N}$, $t\in[T]\cup\{0\}$, and transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, we have
@@
 For short, we refer to the LHS and RHS as $A$ and $B$, respectively. There are three cases:
-  • If $\mu_{t}^{b}(q_{1},\dots,a_{t})\leq e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$, then $A=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}),$ which is at most $B$.
-  • If $\mu_{t}^{b}(q_{1},\dots,a_{t})\geq e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$, then, $A=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}),$ which is at most $B$ as $e^{\epsilon}\geq e^{-\epsilon}$.
+  • If $\mu_{t}^{b}(q_{1},\dots,a_{t})\leq e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$, then $A=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}),$ which equals $B$.
+  • If $\mu_{t}^{b}(q_{1},\dots,a_{t})\geq e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$, then, $A=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}),$ which is at most $B$ as $e^{\epsilon}\geq e^{-\epsilon}$.
   • Otherwise, if $e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})<\mu_{t}^{b}(q_{1},\dots,a_{t})<e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$, $A=0$ is at most $B$ since by the condition of the case distinction, $B=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$ is positive.
@@
 $\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{b}(q_{1},\dots,a_{t}),$
-where the last inequality holds because for every $q_{t+1}\in Q$, by Claim A.7
+where the last inequality holds because for every $q_{t+1}\in Q$, by Claim 8.6
 $\mathrm{Lower}_{t+1,T}^{b}(q_{1},a_{1},\dots,q_{t+1})$ $-e^{-\epsilon}\mathrm{Lower}_{t+1,T}^{1-b}(q_{1},a_{1},\dots,q_{t+1})\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{b}(q_{1},\dots,a_{t}).$
-A.1.2. Control Function $\mathrm{L}_{t}^{b}$ with is the Limit of $\mathrm{L}_{t,T}$ (Definition).
+8.1.1. Control Function $\mathrm{L}_{t}^{b}=\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}$ (Definition).
 This section introduces our main new technical contribution.
-We introduce $\mathrm{L}_{t}^{b}$ as limit of $\mathrm{L}_{t,T}^{b}$ and show that it satisfies all properties needed by the proof of [18].
-Thus, differently from [18], we will use $\mathrm{L}_{t}^{b}$ to define the functions $\phi_{t}^{0}$ and $\phi_{t}^{1}$.
-However, we first need to argue that the limit of $\mathrm{L}_{t,T}^{b}$ exists.
-the value of $\mathrm{L}_{t,T}^{b}$ is upper bounded by $1$.
-We also show that this function is non-decreasing in $T$.
-As a result, for each each $t\in\mathbb{N}\cup\{0\}$ and $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the limit $\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$ exists.
-To remove the assumption of having an upper bound $T$ on the number of queries, we introduce the following control function:
-Definition A.10 ($\mathrm{L}_{t}^{b}$).
+To remove the assumption of having an upper bound $T$ on the number of queries, we introduce $\mathrm{L}_{t}^{b}$ as limit of $\mathrm{L}_{t,T}^{b}$ and show that it satisfies all properties needed by the proof of [19].
+Thus, differently from [19], we will use $\mathrm{L}_{t}^{0}$ and $\mathrm{L}_{t}^{1}$ to define the function $\phi_{t}$.
+Definition 8.9 ($\mathrm{L}_{t}^{b}$).
 For each $b\in\{0,1\}$ and $t\in\mathbb{N}$, the function $\mathrm{L}_{t}^{b}:(Q\times A)^{t}\to[0,1]$ is defined as follows: for each transcript $(q_{1},\dots a_{t})\in(Q\times A)^{t}$,
 $\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t})=\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t}).$
+To ensure $\mathrm{L}_{t}^{b}$ is well-defined, we need to show that the limit of $\mathrm{L}_{t,T}^{b}$ exists.
+We will upper bound the values of $\mathrm{L}_{t,T}^{b}$ by $1$ and show that this function is non-decreasing in $T$.
+As a result, for each each $t\in\mathbb{N}\cup\{0\}$ and $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the limit $\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$ exists.
 We need the following lemma throughout the section:
-Lemma A.11.
+Lemma 8.10.
 For every $t\in\mathbb{N}$, transcript $(q_{1},\dots a_{t-1})\in(Q\times A)^{t-1}$, and query $q_{t}\in Q$, we have
@@
 To upper bound $\mathrm{L}_{t,T}$, we first need to upper bound $\mathrm{Lower}_{t,T}$:
-Lemma A.12.
+Lemma 8.11.
 For every $T\in\mathbb{N}$, $b\in\{0,1\}$, $t\in[T]$, and transcript $(q_{1},\dots a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
@@
 Proof.
-Fix $T\geq t$. The proof is by induction on $t$ from $T$ to $1$. For $t=T$, by definition and Lemma A.11, we have
+Fix $T\geq t$. The proof is by induction on $t$ from $T$ to $1$. For $t=T$, by definition and Lemma 8.10, we have
 $\mathrm{Lower}_{t,T}^{b}(q_{1},\dots a_{t-1},q_{t})$ $=\sum_{a_{t}\in A}\max\{0,\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})\}$
@@
 The following lemma implies that the values of $\mathrm{L}_{t,T}$ are upper bounded by $1$.
-Lemma A.13.
+Lemma 8.12.
 For every $T\in\mathbb{N}$, $b\in\{0,1\}$, $t\in[T]\cup\{0\}$, and transcript $(q_{1},\dots a_{t})\in(Q\times A)^{t}$,
@@
 Proof.
-For $t=T$, by definition and Lemma A.11, we have
+For $t=T$, by definition and Lemma 8.10, we have
 $\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$ $=\max\{0,\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})\}$
 $\leq\max\{0,\mu_{t}^{b}(q_{1},\dots,a_{t})\}=\mu_{t}^{b}(q_{1},\dots,a_{t})$
-For $t<T$, by Lemma A.12,
+For $t<T$, by Lemma 8.11,
 $\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$ $=\sup_{q_{t+1}\in Q}\mathrm{Lower}_{t+1,T}^{b}(q_{1},\dots,a_{t},q_{t+1})\leq\sup_{q_{t}\in Q}\mu_{t}^{b}(q_{1},\dots,a_{t})=\mu_{t}^{b}(q_{1},\dots,a_{t})$
 To ensure the function $\mathrm{L}_{t}^{b}$ is well-defined, it remains to show that $\mathrm{L}_{t,T}^{b}$ is non-decreasing in $T$. To prove so, we first need to show:
-Lemma A.14.
+Lemma 8.13.
 For each $b\in\{0,1\}$ and $t\in\mathbb{N}$, the function $\mathrm{Lower}_{t,T}^{b}$ is monotonically non-decreasing in $T$.
@@
 $\mathrm{Lower}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t})\leq\mathrm{Lower}_{t,T+1}^{b}(q_{1},\dots,a_{t-1},q_{t}).$
-Fix $T\in\mathbb{N}$ and $b\in\{0,1\}$. We show this by reverse induction on $t$.
-The base case is $t=T$. By Lemma A.11, for every $(q_{1},\dots,a_{t-1},q_{t})$, we have
+Fix $T\in\mathbb{N}$ and $b\in\{0,1\}$. We show this by reverse induction on $t$: The base case is $t=T$. By Lemma 8.10, for every $(q_{1},\dots,a_{t-1},q_{t})$, we have
 $\mathrm{Lower}_{T,T}^{b}(q_{1},\dots,a_{T-1},q_{T})=\sum\limits_{a_{T}\in A}\max\{0,\mu_{T}^{b}(q_{1},\dots,a_{T})-e^{\epsilon}\mu_{T}^{1-b}(q_{1},\dots,a_{T})\}$
@@
 $=\mathrm{Lower}_{T,T+1}^{b}(q_{1},\dots,a_{T-1},q_{T})$
-For the induction step, assume that $t<T$ and that for every $(q_{1},\dots,a_{t},q_{t+1})$, we have $\mathrm{Lower}_{t+1,T}^{b}(q_{1},\dots,a_{t},q_{t+1})\leq\mathrm{Lower}_{t+1,T+1}^{b}(q_{1},\dots,a_{t},q_{t+1})$.
-We need to show $\mathrm{Lower}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t})\leq\mathrm{Lower}_{t,T+1}^{b}(q_{1},\dots,a_{t-1},q_{t})$ for every $(q_{1},\dots,a_{t-1},q_{t})$.
-By definition,
+For the induction step, assume that $t<T$ and that for every $(q_{1},\dots,a_{t},q_{t+1})$, we have
+$\mathrm{Lower}_{t+1,T}^{b}(q_{1},\dots,a_{t},q_{t+1})\leq\mathrm{Lower}_{t+1,T+1}^{b}(q_{1},\dots,a_{t},q_{t+1}).$
+We need to show $\mathrm{Lower}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t})\leq\mathrm{Lower}_{t,T+1}^{b}(q_{1},\dots,a_{t-1},q_{t})$ for every $(q_{1},\dots,a_{t-1},q_{t})$. By definition,
 $\mathrm{Lower}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t})$ $=\sum\limits_{a_{t}\in A}\sup_{q_{t+1}\in Q}\mathrm{Lower}_{t+1,T}^{b}(q_{1},\dots a_{t},q_{t+1})$
@@
 $=\mathrm{Lower}_{t,T+1}^{b}(q_{1},\dots,a_{t-1},q_{t})$
-Lemma A.15.
+Lemma 8.14.
 For each $b\in\{0,1\}$ and $t\in\mathbb{N}\cup\{0\}$, the function $\mathrm{L}_{t,T}^{b}$ is monotonically non-decreasing in $T$.
@@
 $\mathrm{L}_{t,T}(q_{1},\dots,a_{t})\leq\mathrm{L}_{t,T+1}(q_{1},\dots,a_{t})$
-Fix $T\in\mathbb{N}$ and $b\in\{0,1\}$. We show this by reverse induction on $t$ from $T$ to 00.
-The base case is $t=T$. By Lemma A.11, for every $(q_{1},\dots,a_{T})$ we have
+Fix $T\in\mathbb{N}$ and $b\in\{0,1\}$. We show this by reverse induction on $t$ from $T$ to 0.
+The base case is $t=T$. By Lemma 8.10, for every $(q_{1},\dots,a_{T})$ we have
 $\mathrm{L}_{T,T}^{b}(q_{1},\dots,a_{T})=\max\{0,\mu_{T}^{b}(q_{1},\dots,a_{T})-e^{\epsilon}\mu_{T}^{1-b}(q_{1},\dots,a_{T})\}$
@@
 We need to show $\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})\leq\mathrm{L}_{t,T+1}^{b}(q_{1},\dots,a_{t})$ for every $(q_{1},\dots,a_{t})$.
-By definition of $\mathrm{L}_{t,T+1}$ and Lemma A.14, we have
+By definition of $\mathrm{L}_{t,T+1}$ and Lemma 8.13, we have
 $\mathrm{L}_{t,T}^{b}(q_{1},\dots,q_{t},a_{t})$ $=\sup_{q_{t+1}\in Q}\mathrm{Lower}_{t+1,T}^{b}(q_{1},\dots a_{t},q_{t+1})$
@@
 $=\mathrm{L}_{t,T+1}^{b}(q_{1},\dots,q_{t},a_{t})$
-A.1.3. Control Function $\mathrm{L}_{t}^{b}$ (Properties).
+8.1.2. Control Function $\mathrm{L}_{t}^{b}=\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}$ (Properties).
 In this section, we present the properties of $\mathrm{L}_{t}^{b}$, required in the next sections.
-The definition of $\mathrm{L}_{t}^{b}$ combined with Corollary A.8, Lemma A.9, and Lemma A.13 give the following corollaries, respectively:
-Corollary A.16.
+The definition of $\mathrm{L}_{t}^{b}$ combined with Corollary 8.7, Lemma 8.8, and Lemma 8.12 give the following corollaries, respectively:
+Corollary 8.15.
 For each $b\in\{0,1\}$, $\mathrm{L}_{0}^{b}()\leq\delta$.
-Corollary A.17.
+Corollary 8.16.
 For each $b\in\{0,1\}$, $t\in\mathbb{N}\cup\{0\}$, and transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, we have
 $\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t})$ $-e^{-\epsilon}\mathrm{L}_{t}^{1-b}(q_{1},\dots,a_{t})\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}).$
-Lemma A.18.
+Lemma 8.17.
 For each $b\in\{0,1\}$, $t\in\mathbb{N}$, and transcript $(q_{1},\dots a_{t})\in(Q\times A)^{t}$,
@@
 To show the next property, we need the following known fact:
-Lemma A.19 (Monotone Convergence).
+Lemma 8.18 (Monotone Convergence).
 Let $f:\mathbb{N}^{2}\to\mathbb{R}$ be a function such that $\sum_{m}f(m,1)>$-$\infty$ and that is non-decreasing in the second parameter. Then
 $\lim_{n\to\infty}\sum_{m=1}^{\infty}f(m,n)=\sum_{m=1}^{\infty}\lim_{n\to\infty}f(m,n)$
-Lemma A.19 combined with Lemma A.15 implies:
-Corollary A.20.
-For any $b\in\{0,1\}$, $t\in\mathbb{N}$, and $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
+Lemma 8.18 combined with Lemma 8.14 implies:
+Corollary 8.19.
+For every $b\in\{0,1\}$, $t\in\mathbb{N}$, and $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
 $\sum_{a_{t}\in A}\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})=\lim_{T\geq t,T\to\infty}\sum_{a_{t}\in A}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$
-Lemma A.21.
+Lemma 8.20.
 For every $b\in\{0,1\}$, $t\in\mathbb{N}$, transcript $(q_{1},\dots a_{t-1})\in(Q\times A)^{t-1}$, and query $q_{t}\in Q$,
@@
 Proof.
-First, we will show that for any $T>t-1$,
+First, we will show that for every T>t1T>t-1,
 $\mathrm{L}_{t-1,T}^{b}(q_{1},\dots a_{t-1})\geq\sum_{a_{t}\in A}\mathrm{L}_{t,T}^{b}(q_{1},\dots a_{t})$
-Since $t\neq T$, by definition of $\mathrm{L}_{t-1,T}^{b}$ and Corollary A.5, we have
+Since $t-1\neq T$, by definition of $\mathrm{L}_{t-1,T}^{b}$ and Corollary 8.4, we have
 $\mathrm{L}_{t-1,T}^{b}(q_{1},\dots a_{t-1})$ $=\sup_{q\in Q}\mathrm{Lower}_{t,T}(q_{1},\dots a_{t-1},q)\geq\mathrm{Lower}_{t,T}(q_{1},\dots a_{t-1},q_{t})=\sum_{a_{t}\in A}\mathrm{L}_{t,T}^{b}(q_{1},\dots a_{t}).$
-Thus, by Corollary A.20, we have
+Thus, by Corollary 8.19, we have
 $\mathrm{L}_{t-1}^{b}(q_{1},\dots a_{t-1})$ $=\lim_{T\geq t-1,T\to\infty}\mathrm{L}_{t-1,T}^{b}(q_{1},\dots a_{t-1})=\lim_{T\geq t,T\to\infty}\mathrm{L}_{t-1,T}^{b}(q_{1},\dots a_{t-1})$
 $\geq\lim_{T\geq t,T\to\infty}\sum_{a_{t}\in A}\mathrm{L}_{t,T}^{b}(q_{1},\dots a_{t})=\sum_{a_{t}\in A}\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots a_{t})=\sum_{a_{t}\in A}\mathrm{L}_{t}^{b}(q_{1},\dots a_{t})$
-A.2. Reduction.
-To define $\mathcal{P}$, Lyu [18] constructs four families of PMFs $\mathcal{F}_{(0,\bot)}$, $\mathcal{F}_{(1,\bot)}$, $\mathcal{F}_{(0,\top)}$, and $\mathcal{F}_{(1,\top)}$, corresponding to each outcome of $\mathrm{RR}_{\epsilon,\delta}$.
+8.2. Reduction.
+To define $\mathcal{P}$ and simulate $\mathcal{M}$, Lyu [19] constructs four families of probability mass functions (PMFs), corresponding to each outcome of $\mathrm{RR}_{\epsilon,\delta}$.
 Intuitively, each family describes an IM.
-Specifically, each family consists of PMFs over the answer set $A$ for every $t\in[T]$ and every history $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t}\times Q$, representing the probability of returning each answer in response to $q_{t}$, conditioned on $q_{1},\dots,a_{t-1}$ being the history of queries and answers.
-A.2.1. Constructing $\mathcal{P}$ using $\mathcal{F}_{r}$s.
-Roughly speaking, in the interaction between $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,\delta}(b)$ and an adversary $\mathcal{A}$, $\mathcal{P}$ interacts with $\mathrm{RR}_{\epsilon,\delta}(b)$ once and stores its output $r$ as thee first parameter in its state.
-It then answers the adversary’s queries by itself using the PMFs provided by $\mathcal{F}_{r}$.
-Specifically, $\mathcal{P}$ has a single initial state $\left(\text{null},()\right)$, where null is an empty field and ()()( ) represents an empty sequence.
-Let $h$ denote the sequence of interactions between $\mathcal{P}$ and $\mathcal{A}$, which is stored by $\mathcal{P}$ as second parameter of its state.
-After interacting with $\mathrm{RR}_{\epsilon,\delta}(b)$, $\mathcal{P}$ updates null to $r$.
-Moreover, upon receiving a query $q$ and before returning an answer $a$, $\mathcal{P}$ appends $q$ and $a$ to the history sequence $h$.
-To generate a response $a$ to a query $q$, $\mathcal{P}$ samples $a$ from the PMF corresponding to $(h,q)$ in $\mathcal{F}_{r}$, where $h$ is the history of all previous queries and answers.
-A.2.2. Constructing $\mathcal{F}_{(b,\bot)}$s.
-Lyu [18] constructs a function $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for each $t\in[T]\cup\{0\}$.
-Let $\phi_{t}^{0}$ and $\phi_{t}^{1}$ denote the functions returning the first and second elements of the output pairs of $\phi_{t}$.
-For each $b\in\{0,1\}$, the family $\mathcal{F}_{(b,\bot)}$ is defined based on $\phi_{t}^{b}$s.
-Intuitively, $\phi_{t}^{b}(q_{1},\dots,a_{t})$ represents the probability that $\mathcal{M}(s_{b})$ returns the answers $a_{1},\dots,a_{t}$ to the fixed queries $q_{1},\dots,q_{t}$.
-That is, for each query sequence $q_{1},\dots,q_{t}$, $\phi_{t}^{b}(q_{1},\cdot,q_{2},\cdot,\dots,q_{t},\cdot)$ is a PMF over $A^{t}$.
-We do not provide the construction of $\phi_{t}$ in this section.
-However, assuming the existence of $\phi_{t}$ satisfying the following three conditions, we show that for every $b\in\{0,1\}$, the mechanisms $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}$ and $\mathcal{M}(s_{b})$ are equivalent.
-These three conditions were originally stated in [18], with the control function $\mathrm{L}_{t,T}^{b}$ instead of $\mathrm{L}_{t}^{b}$.
-For $t=0$, let $\phi_{t}()=(1,1)$. Suppose for every $t\in[T]$, there exists a function $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ satisfying the following conditions for each $b\in\{0,1\}$:
+Specifically, each family describes an IM and consists of PMFs over the answer set $A$ for every $t\in[T]$ and every history $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t}\times Q$.
+These PMFs represent the probability of returning each answer in response to $q_{t}$, conditioned on $q_{1},\dots,a_{t-1}$ being the history of queries and answers.
+Similarly, we define $\mathcal{P}$ based on four families of PMFs: $\mathcal{F}_{(\bot,0)}$, $\mathcal{F}_{(\bot,1)}$, $\mathcal{F}_{(\top,0)}$, and $\mathcal{F}_{(\top,1)}$, corresponding to the four possible output sequences of $\mathrm{IRR}_{\epsilon,\delta}$.
+We note that in our construction, each family includes a PMF over $A$ for every $t\in\mathbb{N}$ and every history $(q_{1},\dots,a_{t-1},q_{t})$.
+Moreover, the definitions of these PMFs are not the same as the ones in [19].
+8.2.1. Constructing $\mathcal{P}$.
+In [19], the post-processing mechanism initially interacts with $\mathrm{RR}_{\epsilon,\delta}$ and stores its output $r=(\tau,c)\in\{\top,\bot\}\times\{0,1\}$.
+It then answers the adversary’s queries by itself using the family of PMFs corresponding to $r$.
+Our high-level idea is to construct the families $\mathcal{F}_{(\bot,0)}$, $\mathcal{F}_{(\bot,1)}$, $\mathcal{F}_{(\top,0)}$, and $\mathcal{F}_{(\top,1)}$ in such a way that, under certain conditions, knowledge of $\tau$ alone suffices to choose the appropriate PMF.
+Recall that the IPM $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$ instead of $\mathrm{RR}_{\epsilon,\delta}$.
+We design $\mathcal{P}$ to initially interact with $\mathrm{IRR}_{\epsilon,\delta}$ and store its response.
+$\mathcal{P}$ then interacts with $\mathrm{IRR}_{\epsilon,\delta}$ for the second time when it cannot choose a PMF using only the first response.
+[…]
+To generate a response $a$ to query $q$, $\mathcal{P}$ proceeds as follows: If $q$ is the first query (i.e., $r=()$), then $\mathcal{P}$ initially interacts with $\mathrm{IRR}_{\epsilon,\delta}$ and appends its response to $r$.
+After this, whether $q$ is the first query or not, when the following condition does not hold for the first time, $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$ for the second time and appends its answer to $r$: the PMFs corresponding to $(h,q)$ in $\mathcal{F}_{(\top,0)}$ and $\mathcal{F}_{(\top,1)}$ are identical, and the PMFs corresponding to $(h,q)$ in $\mathcal{F}_{(\bot,0)}$ and $\mathcal{F}_{(\bot,1)}$ are identical.
+After that, $\mathcal{P}$ never interacts with $\mathrm{IRR}_{\epsilon,\delta}$ again.
+After the potential interactions with $\mathrm{IRR}_{\epsilon,\delta}$, $\mathcal{P}$ generates the answer $a$ as follows:
+  • If $r\in\{\top,\bot\}\times\{0,1\}$, then $\mathcal{P}$ samples $a$ from the PMF corresponding to $(h,q)$ in $\mathcal{F}_{r}$.
+  • Otherwise, by design, $r\in\{\top,\bot\}$, and the PMFs in $\mathcal{F}_{(r,0)}$ and $\mathcal{F}_{(r,1)}$ corresponding to $(h,q)$ are identical.
+In this case, $\mathcal{P}$ samples $a$ from this common PMF.
+8.2.2. Constructing $\mathcal{F}_{(\bot,b)}$s.
+Using the control function $\mathrm{Lower}_{t,T}$, Lyu [19] introduces three conditions and constructs a sequence of functions $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for $t\in\{1,\dots,T\}$ that satisfy these conditions.
+The families of PMFs are defined then based on these functions.
+We remove the dependence on the bound $T$ by replacing $\mathrm{Lower}_{t,T}$ with $\mathrm{L}_{t}$ in the conditions, and extend the construction of $\phi_{t}$ to all $t\in\mathbb{N}$.
+Furthermore, we introduce an additional condition to those given in [19], which we will later use to bound the number of times $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$.
+Let $\phi_{t}^{0}$ and $\phi_{t}^{1}$ denote the functions that return the first and second components, respectively, of the output of $\phi_{t}$.
+Intuitively, each $\phi_{t}^{b}$ represents an IM: for every query sequence $q_{1},\dots,q_{t}$, it assigns a probability to each answer sequence $a_{1},\dots,a_{t}$.
+That is, for any query sequence $q_{1},\dots,q_{t}$, the function $\phi_{t}^{b}(q_{1},\cdot,q_{2},\cdot,\dots,q_{t},\cdot)$ defines a PMF over $A^{t}$.
+For $t=0$, we define $\phi_{t}()=(1,1)$.
+We defer the detailed construction of the functions $\phi_{t}$ to Section 8.3.
+However, we assume the existence of $\phi_{t}$ satisfying the following conditions for each $b\in\{0,1\}$:
   1. (I)
-    For any transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$,
+    For every transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$,
     $\delta\phi_{t}^{b}(q_{1},\dots,a_{t})\geq\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t}).$
   2. (II)
-    For any transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$,
+    For every transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$,
     $\delta\phi_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\delta\phi_{t}^{1-b}(q_{1},\dots,a_{t})\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}).$
   3. (III)
-    For any transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
+    For every transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
     $\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)=\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
-For each $b\in\{0,1\}$, the family $\mathcal{F}_{(b,\bot)}$ consists of functions $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}:A\to[0,1]$ for each $(q_{1},\dots,a_{t-1},q_{t})\in\bigcup_{t=1}^{T}(Q\times A)^{t-1}\times Q$, where $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ is defined as follows:
+  4. (IV)
+    For each query sequence $(q_{1},\dots,q_{t})\in Q^{t}$, if $\mu_{t}^{0}(q_{1},\dots,a_{t})=\mu_{t}^{1}(q_{1},\dots,a_{t})$ for every answer sequence $(a_{1},\dots,a_{t})\in A^{t}$, then
+    $\phi_{t}^{0}(q_{1},\dots,a_{t})=\phi_{t}^{1}(q_{1},\dots,a_{t})\geq\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}$
+    for every $(a_{1},\dots,a_{t})\in A^{t}$.
+For each $b\in\{0,1\}$, the family $\mathcal{F}_{(\bot,b)}$ consists of functions $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}:A\to[0,1]$ defined for every $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{*}\times Q$, where $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ is defined as follows:
   • If $\phi_{t-1}(q_{1},\dots,a_{t-1})>0$, then for each answer $a_{t}\in A$,
     $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=\frac{\phi_{t}(q_{1},\dots,a_{t})}{\phi_{t-1}(q_{1},\dots,a_{t-1})}.$
-  • Otherwise, if $\phi_{t-1}(q_{1},\dots,a_{t-1})=0$, $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=0$ for every answer, except for one fixed arbitrary answer $a^{*}\in A$, where $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a^{*})=1$.
-This case is included only for formality, and the corresponding PMFs are never used by $\mathcal{P}$.
+  • Otherwise, if $\phi_{t-1}(q_{1},\dots,a_{t-1})=0$, $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=0$ for all $a_{t}\in A$, except for one fixed arbitrary answer $a^{*}\in A$, for which $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a^{*})=1$.
+This case is included only for formality, as the corresponding PMFs are never used by $\mathcal{P}$.
 Since $\phi_{t-1}$ and $\phi_{t}$ are non-negative, $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})$ is also non-negative.
 Furthermore, by (III), we have $\sum_{a_{t}\in A}f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a)=1$, implying that $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})$ is a PMF.
-A.2.3. Constructing $\mathcal{F}_{(b,\top)}$s.
-Similar to the previous section, $\mathcal{F}_{(0,\top)}$ and $\mathcal{F}_{(1,\top)}$ are defined based on a sequence of functions $\psi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for each $t\in[T]\cup\{0\}$.
+8.2.3. Constructing $\mathcal{F}_{(\top,b)}$s.
+Similar to the previous section, $\mathcal{F}_{(0,\top)}$ and $\mathcal{F}_{(1,\top)}$ are defined based on a sequence of functions $\psi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for each $t\in\mathbb{N}\cup\{0\}$.
 For $t=0$, $\psi_{t}()=(1,1)$.
-For $t\in[T]$, $\psi_{t}$ is defined as follows:
+For $t\in\mathbb{N}$, we define $\psi_{t}$ as follows:
   • If $\delta=1$, then $\psi_{t}$ is identical to $\phi_{t}$.
@@
     $\frac{1}{(1-\delta)(e^{\epsilon}-1)}\Big(e^{\epsilon}\mu_{t}^{b}(q_{1},\dots,a_{t})-\delta e^{\epsilon}\phi_{t}^{b}(q_{1},\dots,a_{t})-\mu_{t}^{1-b}(q_{1},\dots,a_{t})+\delta\phi_{t}^{1-b}(q_{1},\dots,a_{t})\Big).$
-For each $b\in\{0,1\}$, the family $\mathcal{F}_{(b,\top)}$ consists of functions $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}:A\to[0,1]$ for each $(q_{1},\dots,a_{t-1},q_{t})\in\bigcup_{t=1}^{T}(Q\times A)^{t-1}\times Q$, where $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ is defined as follows:
+For each $b\in\{0,1\}$, the family $\mathcal{F}_{(\top,b)}$ consists of functions $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}:A\to[0,1]$ for each $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{*}\times Q$, where $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ is defined as follows:
   • If $\psi_{t-1}(q_{1},\dots,a_{t-1})>0$, then for each answer $a_{t}\in A$,
     $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=\frac{\psi_{t}(q_{1},\dots,a_{t})}{\psi_{t-1}(q_{1},\dots,a_{t-1})}.$
-  • Otherwise, $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=0$ for every answer, except for one fixed arbitrary answer $a^{*}\in A$, where $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a^{*})=1$.
-By condition (II), each $\psi_{t}^{b}$ is non-negative. Also, by Lemma A.11 and condition (III), for every $(q_{1},\dots,a_{t},q_{t})\in(Q\times A)^{t-1}\times Q$,
+  • Otherwise, $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=0$ for all $a_{t}\in A$, except for one fixed arbitrary answer $a^{*}\in A$, for which $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a^{*})=1$.
+By condition (II), $\psi_{t}^{b}$ is non-negative. Also, by Lemma 8.10 and condition (III),
 $\sum_{a\in A}\psi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)=\psi_{t-1}^{b}(q_{1},\dots,a_{t-1}),$
-implying that each $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ forms a valid PMF.
-A.2.4. Proof of Lemma A.1.
-By definition of $\mathrm{RR}$ and of $\mathcal{P}$, for each $t\in\mathbb{N}$ and transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the probability of $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,\delta}(b)$ returning the answers $a_{1},\dots,a_{t}$ to the queries $q_{1},\dots,q_{t}$ is
+implying that $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ is a PMF.
+8.2.4. Proof of Lemma 5.6.
+By definition of $\mathrm{IRR}_{\epsilon,\delta}$ and $\mathcal{P}$, for every $t\in\mathbb{N}$ and every transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the probability of $\mathcal{P}\!\circ^{*}\!\mathrm{IRR}_{\epsilon,\delta}(b)$ returning the answers $a_{1},\dots,a_{t}$ to the queries $q_{1},\dots,q_{t}$ is
 $\delta\phi_{t}^{b}(q_{1},\dots,a_{t})+(1-\delta)\frac{e^{\epsilon}}{1+e^{\epsilon}}\psi_{t}^{b}(q_{1},\dots,a_{t})+(1-\delta)\frac{1}{1+e^{\epsilon}}\psi_{t}^{1-b}(q_{1},\dots,a_{t}).$
@@
 $=\mu_{t}^{b}(q_{1},\dots,a_{t}).$
-Thus, the views of any adversary interacting with $\mathcal{M}(s_{b})$ and $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,\delta}(b)$ are identically distributed, implying that these mechanisms are equivalent.
-A.3. Construction of $\phi_{t}$
-In this section, we formalize the construction of [18] differently to remove the finiteness assumption on $A$.
-Roughly speaking, Lyu [18] initializes the values of $\phi_{t}$ with the ones from $\mathrm{L}_{t}$.
-Then, for each $q_{1},\dots,a_{t-1},q_{t}$, they iterate over all possible answers $a_{t}\in A$ in a fixed order and gradually increase the value of $\phi_{t}(q_{1},\dots,a_{t})$ to satisfy conditions (II) and (III), while ensuring that condition (I) remains valid.
+Thus, the views of any adversary interacting with $\mathcal{M}(s_{b})$ and $\mathcal{P}\!\circ^{*}\!\mathrm{IRR}_{\epsilon,\delta}(b)$ are identically distributed, implying that these mechanisms are equivalent.
+Furthermore, for every query sequence $(q_{1},\dots,q_{k})$, if the answer distributions of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to $(q_{1},\dots,q_{k})$ are identical, then by Lemma 8.10, for every $t\leq k$ and every answer sequence $(a_{1},\dots,a_{t})\in A^{t}$, we have
+$\mu_{t}^{0}(q_{1},\dots,a_{t})=\mu_{t}^{1}(q_{1},\dots,a_{t}).$
+Therefore, by condition (IV), for every $t\leq k$ and every $(a_{1},\dots,a_{t})\in A^{t}$,
+$\phi_{t}^{0}(q_{1},\dots,a_{t})=\phi_{t}^{1}(q_{1},\dots,a_{t}).$
+Consequently,
+$\psi_{t}^{0}(q_{1},\dots,a_{t})=\psi_{t}^{1}(q_{1},\dots,a_{t}).$
+By the definitions of $\mathcal{F}_{(\top,0)}$, $\mathcal{F}_{(\top,1)}$, $\mathcal{F}_{(\bot,0)}$, and $\mathcal{F}_{(\bot,1)}$, this implies that for every $t\leq k$ and answer sequence $(a_{1},\dots,a_{t-1})$, the PMF associated with $(q_{1},\dots,a_{t-1},q_{t})$ is identical in both $\mathcal{F}_{(\bot,0)}$ and $\mathcal{F}_{(\bot,1)}$, and likewise in $\mathcal{F}_{(\top,0)}$ and $\mathcal{F}_{(\top,1)}$.
+Hence, by the construction of $\mathcal{P}$, this mechanism interacts with $\mathrm{IRR}_{\epsilon,\delta}$ only once.
+Moreover, by design, $\mathcal{P}$ never interacts with $\mathrm{IRR}_{\epsilon,\delta}$ more than twice.
+8.3. Construction of $\phi_{t}$
+For each $q_{1},\dots,a_{t-1},q_{t}$, Lyu [19] initializes the values of $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ with the ones from $\mathrm{L}_{t,T}(q_{1},\dots,a_{t-1},q_{t},\cdot)$.
+(Removing the upper bound assumption, they are initialized by the values of $\mathrm{L}_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$.) Then, they iterate over all possible answers $a_{t}\in A$ in a fixed order and gradually increase the value of $\phi_{t}(q_{1},\dots,a_{t})$ to satisfy conditions (II) and (III), while ensuring that condition (I) remains valid.
 This iterative procedure requires $A$ to be finite; otherwise, the construction algorithm does not terminate.
-To extend this approach to infinite $A$, we define $\phi_{t}$ mathematically and apply the transfinite recursion theorem to prove its existence.
-For each $t\in\mathbb{N}$, knowing that $\phi_{t-1}$ exists and satisfies condition (I), we construct $\phi_{t}$ and show it satisfies conditions (I), (II), and (III).
-For $t=0$, we set $\phi_{0}()=(1,1)$.
-By Corollary A.16, $\phi_{0}$ satisfies condition (I).
-Remark A.22.
-In the rest of this section, we assume $\phi_{t-1}$ exists and satisfies condition (I).
-In Section A.3.1, we introduce the necessary preliminaries for formally defining $\phi_{t}$ when $A$ is infinite.
-In Section A.3.2, we construct $\phi_{t}$, and in Section A.3.3, we prove that $\phi_{t}$ satisfies conditions (I) to (III).
-A.3.1. Preliminaries.
+Moreover, the fixed iteration order can lead to an imbalance in which earlier values grow significantly larger than those that come later in the order.
+In our new construction, we control and balance this increase.
+Intuitively, we introduce a temporary upper bound and increase the values of the function $\mathrm{L}_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ to reach the minimum of the original upper bounds and the new temporary ones.
+This results in an intermediate function, denoted by $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$.
+Then, we remove the temporary upper bound and continue increasing the values of $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ (if needed) to meet the original upper bounds.
+The final function is the desired $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$.
+We will show that if $\mu_{t}^{0}(q_{1},\dots,a_{t})=\mu_{t}^{1}(q_{1},\dots,a_{t})$ for every answer sequence $(a_{1},\dots,a_{t})\in A^{t}$, then the functions $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ and $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ are identical.
+We set the temporary upper bound in a way that this implies the condition (IV) holds.
+Formally, we inductively construct $\phi_{t}$ for each $t\in\mathbb{N}$, assuming that $\phi_{t-1}$ exists and satisfies conditions (I), (II) and (IV).
+We then show that $\phi_{t}$ satisfies all conditions from (I) to (IV).
+For the base case $t=0$, we define $\phi_{0}()=(1,1)$.
+By Corollary 8.15, this choice satisfies condition (I).
+Moreover, for $\phi_{0}^{0}()=\phi_{0}^{1}()=1$ and $\mu_{0}^{0}()=\mu_{0}^{1}()=1$, condition (II) is $\delta\cdot 1-e^{-\epsilon}\cdot\delta\cdot 1\leq 1-e^{-\epsilon}\cdot 1$, which is true because $\delta\leq 1$.
+Furthermore, since $\phi_{0}^{0}()=\phi_{0}^{1}()=1\geq\min\{1,\delta+\delta\}$, condition (IV) is also satisfied.
+Remark 8.21.
+In the rest of this section, we assume $\phi_{t-1}$ exists and satisfies conditions (I), (II) and (IV).
+Section 8.3.1 outlines the preliminaries required for defining $\xi_{t}$ and $\phi_{t}$.
+In Section 8.3.2, we define the intermediate function $\xi_{t}$.
+Section 8.3.3 presents the new construction of $\phi_{t}$, and finally, in Section 8.3.4, we prove that $\phi_{t}$ satisfies conditions (I) to (IV).
+8.3.1. Preliminaries.
 In this section, we introduce the fundamental concepts and notations needed for defining $\phi_{t}$ when $A$ is not finite.
-Definition A.23 (Well-Order).
+Definition 8.22 (Well-Order).
 A well-order on a set $X$ is a total order (i.e., for every $x,y\in X$, either $x<y$, $x>y$, or $x=y$) such that every non-empty subset of $X$ has a least element.
-Since $A$ is countable, we can inject $A$ into $\mathbb{N}$.
-By listing elements of $A$ in the order given by their images under this injection, we get a well‐order $<_{A}$ on $A$.
-Note that for any non-empty subset $B\subseteq A$, the image of $B$ in $\mathbb{N}$ has a least element, whose preimage is the least element of $B$ under $<_{A}$.
-We will later use the following remark when constructing $\xi_{t}$ and $\phi_{t}$.
-Remark A.24.
-By the definition of $<_{A}$, each element of $A$ is either the least element of $A$ or has an immediate predecessor in $A$.
-The query set $Q$ is not necessarily countable.
-However, by the well‐ordering theorem (which is equivalent to the axiom of choice), there exists a well‐order $<_{Q}$ on $Q$, regardless of its cardinality.
-Definition A.25 (Lexicographic Order $<_{t}$ on $(Q\times A)^{t}$).
-For $t\in\mathbb{N}\cup\{0\}$, we define an order $<_{t}$ on $(Q\times A)^{t}$ by lexicographic comparison: For any two distinct tuples $(q_{1},\dots,a_{t})$ and $(q_{1}^{\prime},\dots,a_{t}^{\prime})$ in $(Q\times A)^{t}$, let $i$ be the smallest index such that $(q_{i},a_{i})\neq(q_{i}^{\prime},a_{i}^{\prime})$.
-Then:
-$(q_{1},\dots,a_{t})<_{t}(q_{1}^{\prime},\dots,a_{t}^{\prime})\text{ if and only if }\left[q_{i}<_{Q}q_{i}^{\prime}\right]\text{ or }\left[q_{i}=q_{i}^{\prime}\text{ and }a_{i}<_{A}a_{i}^{\prime}\right]$
-Since $<_{A}$ and $<_{Q}$ are well-orders, $<_{t}$ is also a well-order: $<_{t}$ is clearly a total order.
-We will show that every non-empty subset $B\subseteq(Q\times A)^{t}$ has a least element.
-Define
-$B_{1}^{Q}$ =$=\{q\in Q\mid\exists a_{1},\dots,a_{t}\in A,\exists q_{2},\dots,q_{t}\in Q\text{ such that }(q,a_{1},q_{2},\dots,a_{t})\in B\}\text{ and }$
-$B_{1}^{A}$ =$=\{a\in A\mid\exists a_{2},\dots,a_{t}\in A,\exists q_{2},\dots,q_{t}\in Q\text{ such that }(q_{1}^{*},a,q_{2},\dots,a_{t})\in B\}.$
-Since $B$ is non-empty, $B_{1}^{Q}$ is also non-empty.
-By the well-ordering of $Q$, $B_{1}^{Q}$ has a least element $q_{1}^{*}$.
-Similarly, since $B_{1}^{A}$ is non-empty, it has a least element $a_{1}^{*}$.
-Again, since $B_{2}^{Q}$ is non-empty, it has a least element $q_{2}^{*}$. Repeating this process $t$ times, we obtain the least element $(q_{1}^{*},\dots,a_{t}^{*})$ of $B$.
-Lemma A.26 (Transfinite Recursion).
-Let $f$ be a function recursively defined over $(Q\times A)^{t}$ according to the well-order $<_{t}$.
-If for each sequence $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the value of $f(q_{1},\dots,a_{t})$ is uniquely determined by the values $f(q_{1}^{\prime},\dots,a_{t}^{\prime})$ for every $(q_{1}^{\prime},\dots,a_{t}^{\prime})<_{t}(q_{1},\dots,a_{t})$, then $f$ exists and is unique.
-In later sections, we define functions $f$ recursively over $(Q\times A)^{t}$ using the well-order $<_{t}$.
-Specifically, we determine $f(q_{1},\dots,a_{t})$ based on the values $f(q_{1},\dots,a_{t-1},q_{t},a)$ for every $a<_{A}a_{t}$.
-By the definition of $<_{t}$, we have:
-$(q_{1},\dots,a_{t-1},q_{t},a)<_{t}(q_{1},\dots,a_{t}).$
-Finally, we define two key sets that will be used in our constructions: for any $a_{t}\in A$,
-$\Gamma_{<a_{t}}=\{a\in A\mid a<_{A}a_{t}\},$
-and
-$\Gamma_{>a_{t}}=\{a\in A\mid a_{t}<_{A}a\}.$
-A.3.2. Defining $\phi_{t}$ and showing its existence.
-We define the function $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ recursively for each $(q_{1},\dots,a_{t})$ based on the $<_{t}$ order.
-For each $(q_{1},\dots,a_{t})$, we determine $\phi_{t}(q_{1},\dots,a_{t})$ by solving an optimization problem with the variables $\phi_{t}^{0}(q_{1},\dots,a_{t})$ and $\phi_{t}^{1}(q_{1},\dots,a_{t})$.
-The constraints for this problem involve the known functions $\mu_{t}$, $\phi_{t-1}$, $\mathrm{L}_{t}$, and the known values $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a)$ for every $a\in\Gamma_{<a_{t}}$.
-We define $\phi_{t}^{0}(q_{1},\dots,a_{t})$ and $\phi_{t}^{1}(q_{1},\dots,a_{t})$ as the largest possible real numbers in $[0,1]$ satisfying the following constraints:
-  1. ($\phi$1)
+Since $A$ is countable, $A$ can be injected into $\mathbb{N}$.
+Listing elements of $A$ in the order given by their images under this injection gives a well‐order $_{A}$ on $A$.
+Note that for any non-empty subset $B\subseteq A$, the image of $B$ in $\mathbb{N}$ has a least element, whose preimage is the least element of $B$ under $_{A}$.
+We will later use the following remark when constructing $\phi_{t}$.
+Remark 8.23.
+By the definition of $_{A}$, each element of $A$ is either the least element of $A$ or has an immediate predecessor in $A$.
+Definition 8.24 ($\Gamma_{<a}$ and $\Gamma_{>a}$).
+For any answer $a\in A$, $\Gamma_{<a}=\{a^{\prime}\in A\mid a^{\prime}<_{A}a\}$ and $\Gamma_{>a}=\{a^{\prime}\in A\mid a<_{A}a^{\prime}\}$.
+Lemma 8.25 (Transfinite Recursion [22]).
+Let $f$ be a function recursively defined over $A$ according to the well-order $_{A}$.
+If for each answer $a\in A$, the value of $f(a)$ is uniquely determined by the values $f(a^{\prime})$ for every $a^{\prime}\in\Gamma_{<a}$, then $f$ exists and is unique.
+In the next section, we construct functions $f:A\to[0,1]^{2}$ and apply Lemma 8.25 to prove that they are well-defined.
+To apply this lemma, we must ensure that each value $f(a)$ is uniquely determined based on the previously defined values $f(a^{\prime})$ for all $a^{\prime}\in\Gamma_{<a}$.
+Thus, to define $f(a)$, we first select a subset $\mathcal{C}\subseteq[0,1]^{2}$ of desired values satisfying certain properties derived from the values of $f(a^{\prime})$ for $a^{\prime}\in\Gamma_{<a}$.
+We then set $f(a)$ to the lexicographically maximal element of $\mathcal{C}$.
+The following definitions and lemma formalizes this and guarantee that $f(a)$ is uniquely defined in this way.
+Definition 8.26 (Maximal Pair).
+Let $\mathcal{C}\subseteq[0,1]^{2}$.
+A pair $(x_{0},x_{1})\in\mathcal{C}$ is said to be maximal in $\mathcal{C}$ if there does not exist a pair $(y_{0},y_{1})\in\mathcal{C}$ such that $(x_{0},x_{1})\neq(y_{0},y_{1})$, $x_{0}\leq y_{0}$, and $x_{1}\leq y_{1}$.
+Definition 8.27 (Lexicographic Order $_{lex}$).
+The relation $_{lex}$ is a binary ordering on $0,1]^{2}$ defined as follows: for any $(r_{1},r_{2}),(r_{3},r_{4})\in[0,1]^{2}$, we have $(r_{1},r_{2})<_{lex}(r_{3},r_{4})$ if and only if $r_{1}\leq r_{3}$, or $r_{1}=r_{3}$ and $r_{2}\leq r_{4}$.
+Lemma 8.28.
+Let $\mathcal{C}$ be a non-empty compact subset of $0,1]^{2}$.
+  1. (a)
+    $(x_{0},x_{1})<_{lex}(x_{0}^{*},x_{1}^{*})$ for all $(x_{0},x_{1})\in\mathcal{C}$.
+  2. (b)
+    $(x_{0}^{*},x_{1}^{*})$ is maximal in $\mathcal{C}$.
+Proof.
+(a): Define $x_{0}^{*}=\sup\{x_{0}\mid(x_{0},x_{1})\in\mathcal{C}\}$.
+Since $\mathcal{C}$ is non-empty, the set $\{x_{0}\mid(x_{0},x_{1})\in\mathcal{C}\}$ is non-empty.
+Also, by definition, $x_{0}^{*}\leq 1$.
+As we are taking a continuous projection $(x_{0},x_{1})\to x_{0}$ of a compact set $\mathcal{C}$, this supremum is actually a maximum, so there exists $(x_{0},x_{1})\in\mathcal{C}$ such that $x_{0}=x_{0}^{*}$.
+Define $x_{1}^{*}=\sup\{x_{1}\mid(x_{0}^{*},x_{1})\in\mathcal{C}\}$.
+Again, the set $\{x_{1}\mid(x_{0}^{*},x_{1})\in\mathcal{C}\}$ is non-empty since there exists $(x_{0},x_{1})\in\mathcal{C}$ with $x_{0}=x_{0}^{*}$.
+By the same compactness argument, this supremum is also attained.
+Hence, $(x_{0}^{*},x_{1}^{*})\in\mathcal{C}$.
+By definition, $(x_{0},x_{1})<_{lex}(x_{0}^{*},x_{1}^{*})$ for every $(x_{0},x_{1})\in\mathcal{C}_{q_{1},\dots,a_{t}}$, which trivially implies that $(x_{0}^{*},x_{1}^{*})$ is the only pair in $\mathcal{C}$ satisfying this property.
+(b): Let $(x_{0},x_{1})$ be a pair in $\mathcal{C}$ satisfying $x_{0}\geq x_{0}^{*}$ and $x_{1}\geq x_{1}^{*}$.
+Since $x_{0}\geq x_{0}^{*}$ and $(x_{0},x_{1})\in\mathcal{C}$, by definition of $x_{0}^{*}$, we have $x_{0}^{*}=x_{0}$.
+Since $x_{1}\geq x_{1}^{*}$, $x_{0}^{*}=x_{0}$, and $(x_{0},x_{1})\in\mathcal{C}$, by definition of $x_{1}^{*}$, we have $x_{1}^{*}=x_{1}$.
+Therefore, $(x_{0},x_{1})=(x_{0}^{*},x_{1}^{*})$, implying that $(x_{0}^{*},x_{1}^{*})$ is maximal.
+∎
+8.3.2. Defining $\xi_{t}$ and showing its existence.
+To define the function $\xi_{t}:(Q\times A)^{t}\to[0,1]^{2}$, we fix $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$ and recursively define $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ for each answer $a_{t}\in A$ according to the order $_{A}$.
+We define $\xi_{t}(q_{1},\dots,a_{t})=(x_{0}^{*},x_{1}^{*})$, where $x_{0}^{*}$ and $x_{1}^{*}$ are uniquely chosen from the maximal feasible solutions of a constraint satisfaction problem with two variables $x_{0}$ and $x_{1}$.
+The constraints for this problem involve the known functions $\mu_{t}$, $\phi_{t-1}$, $\mathrm{L}_{t}$, and the known values $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},a)$ for every $a\in\Gamma_{<a_{t}}$.
+Consider the following constraints for the variables $x_{0}$ and $x_{1}$:
+  1. ($\xi$.1)
     For each $b\in\{0,1\}$,
-    $\delta\phi_{t}^{b}(q_{1},\dots,a_{t})\geq\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t}).$
-  2. ($\phi$2)
+    $x_{b}\geq\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t}).$
+  2. ($\xi$.2)
     For each $b\in\{0,1\}$,
-    $\sum_{a\in\Gamma_{<a_{t}}}$ $\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)+\phi_{t}^{b}(q_{1},\dots,a_{t})+\sum_{a\in\Gamma_{>a_{t}}}{\frac{1}{\delta}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$
-  3. ($\phi$3)
+    $\sum_{a\in\Gamma_{<a_{t}}}$ $\xi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)+x_{b}+\sum_{a\in\Gamma_{>a_{t}}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$
+  3. ($\xi$.3)
     For each $b\in\{0,1\}$,
-    $\delta\phi_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\delta\phi_{t}^{1-b}(q_{1},\dots,a_{t})\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$
-Although $\Gamma_{<a_{t}}$ is always finite, $\Gamma_{>a_{t}}$ could be infinite. Thus, to ensure ($\phi$2) is well-defined, we need to show:
-Claim A.27.
-For any $t\in[T]$, $b\in\{0,1\}$, and $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$, the summation $\sum_{a\in\Gamma_{>a_{t}}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ exists.
+    $x_{b}-e^{-\epsilon}x_{1-b}\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$
+  4. ($\xi$.4)
+    For each $b\in\{0,1\}$,
+    $x_{b}\leq\min\{\mu_{t}^{b}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}$
+Let the set $\mathcal{C}_{q_{1},\dots,a_{t}}$ denote the set of feasible solutions $(x_{0},x_{1})$ that satisfy all the constraints above.
+Each constraint is a non-strict inequality of the form $h_{i}(x_{0},x_{1})\geq 0$, where each $h_{i}$ is a continuous function.
+Since the solution set of a single non-strict inequality defined by a continuous function is closed, and the intersection of finitely many closed sets is also closed, the set $\mathcal{C}_{q_{1},\dots,a_{t}}$ is closed.
+Furthermore, as it is a subset of the compact set $0,1]^{2}$, it is itself compact.
+If $\mathcal{C}_{q_{1},\dots,a_{t}}$ is nonempty, then by Lemma 8.28, the lexicographically maximum pair $(x_{0}^{*},x_{1}^{*})$ of $\mathcal{C}_{q_{1},\dots,a_{t}}$ exists, belongs to $\mathcal{C}_{q_{1},\dots,a_{t}}$, and is maximal in $\mathcal{C}_{q_{1},\dots,a_{t}}$.
+We define $\xi_{t}(q_{1},\dots,a_{t})=(x_{0}^{*},x_{1}^{*})$.
+Claim 8.29.
+The function $\xi_{t}$ is well-defined and exists.
 Proof.
-By Lemma A.18,
-$\sum\limits_{a\in\Gamma_{>a_{t}}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\sum\limits_{a\in A}\mu_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)=\mu_{t-1}^{b}(q_{1},\dots,a_{t-1})\leq 1.$
-Moreover each summand $\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ is non-negative, implying that the summation is within the range $[0,1]$, and thus it exists. ∎
-Claim A.28.
+The proof consists of two steps: (i) proving that $\mathcal{C}_{q_{1},\dots,a_{t}}$ is non-empty, and (ii) applying the transfinite recursion theorem (Lemma 8.25) and concluding that $\xi_{t}$ exists.
+(i): Fix any $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$.
+By Remark 8.23, each answer in $A$ is either the smallest element of $A$ or has an immediate predecessor.
+We will inductively show that for every $a_{t}\in A$, the values $\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})$ and $\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})$ satisfy the constraints for defining $\xi_{t}(q_{1},\dots,a_{t})$, and consequently, $\mathcal{C}_{q_{1},\dots,a_{t}}$ is not empty.
+Trivially, $\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t})$ satisfies Constraint ($\xi$.1).
+Moreover, by Corollary 8.16, it satisfies Constraint ($\xi$.3).
+To prove Constraint ($\xi$.2) is satisfied, we need to show that
+(2) $\sum_{a\in\Gamma_{<a_{t}}}\xi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ $+\sum_{a\in\Gamma_{>a_{t}}\cup\{a_{t}\}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\delta\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}),$
+We prove this by induction on $a_{t}\in A$:
+Base Case: Suppose $a_{t}$ is the least element of $A$. Then, $\Gamma_{<a_{t}}$ is empty, reducing our goal to showing that
+$\sum_{\begin{subarray}{c}a\in A\end{subarray}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\delta\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
+We know that $\phi_{t-1}(q_{1},\dots,a_{t-1})$ satisfies condition (I). (see Remark 8.21.) Thus, by Lemma 8.20,
+$\delta\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$ $\geq\mathrm{L}_{t-1}^{b}(q_{1},\dots,a_{t-1})\geq\sum_{\begin{subarray}{c}a\in A\end{subarray}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a),$
+finishing the proof for the base case.
+Inductive Step: Let $a^{*}$ denote the immediate predecessor of $a_{t}$.
+By inductive hypothesis, $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},a^{*})$ exists and satisfies Constraint ($\xi$.3).
+Since $\Gamma_{<a_{t}}=\Gamma_{<a^{*}}\cup\{a^{*}\}$, Constraints ($\xi$.2) for defining $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},a^{*})$ is identical to (2).
+Thus, by inductive hypothesis, inequality (2) holds, completing the induction.
+It remains to prove that Constraint ($\xi$.4) is satisfied.
+By Lemma 8.17, $\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t})\leq\mu_{t}^{b}(q_{1},\dots,a_{t})$.
+Also, as $\mathrm{L}_{t}^{1-b}$ is non-negative, $\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t})\leq\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})$.
+Thus,
+$\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t})\leq\min\{\mu_{t}^{b}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\},$
+and Constraint ($\xi$.4) is satisfied.
+(ii): For every $(q_{1},\dots,a_{t-1},q_{t})$, the values of the function $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ is uniquely determined by the preceding values of this function in the well-order $_{A}$.
+Thus, by the transfinite recursion theorem, the function $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ exists and is uniquely defined.
+Consequently, the function $\xi_{t}$ exists and is uniquely defined.
+∎
+8.3.3. Defining $\phi_{t}$ and showing its existence.
+To define the function $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$, we fix $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$ and recursively define $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ for each answer $a_{t}\in A$ based on the $_{A}$ order.
+Similar to the definition of $\xi_{t}$, for each $(q_{1},\dots,a_{t})$, we determine $\phi_{t}(q_{1},\dots,a_{t})$ using a set of constraints.
+These constraints involve the known functions $\xi_{t}$, $\mu_{t}$, $\phi_{t-1}$, $\mathrm{L}_{t}$, and the known values $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a)$ for every $a\in\Gamma_{<a_{t}}$.
+Consider the following set of constraints for the variables $x_{0}$ and $x_{1}$:
+  1. ($\phi$.1)
+    For each $b\in\{0,1\}$,
+    $\delta x_{b}\geq\xi_{t}^{b}(q_{1},\dots,a_{t}).$
+  2. ($\phi$.2)
+    For each $b\in\{0,1\}$,
+    $\sum_{a\in\Gamma_{<a_{t}}}$ $\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)+x_{b}+\sum_{a\in\Gamma_{>a_{t}}}{\frac{1}{\delta}}\xi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$
+  3. ($\phi$.3)
+    For each $b\in\{0,1\}$,
+    $\delta x_{b}-e^{-\epsilon}\delta x_{1-b}\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$
+Let $\mathcal{C}_{q_{1},\dots,a_{t}}$ denote the set of feasible solutions $(x_{0},x_{1})$ satisfying the constraints above.
+As in Section 8.3.2, one can show that $\mathcal{C}_{q_{1},\dots,a_{t}}$ is compact.
+Therefore, if $\mathcal{C}_{q_{1},\dots,a_{t}}$ is nonempty, then by Lemma 8.28, the lexicographically maximum pair $(x_{0}^{*},x_{1}^{*})$ of $\mathcal{C}_{q_{1},\dots,a_{t}}$ exists, belongs to $\mathcal{C}_{q_{1},\dots,a_{t}}$, and is maximal in $\mathcal{C}_{q_{1},\dots,a_{t}}$.
+Define $\mathcal{G}_{q_{1},\dots,a_{t}}\subseteq\mathcal{C}_{q_{1},\dots,a_{t}}$ to be the set of all maximal elements in $\mathcal{C}_{q_{1},\dots,a_{t}}$.
+By the definition of maximality, this set either contains no pair of the form $(z,z)$ with identical components, or it contains exactly one such pair, denoted $(z^{*},z^{*})$.
+We define $\phi_{t}(q_{1},\dots,a_{t})=(z^{*},z^{*})$ if there exists $(z^{*},z^{*})\in\mathcal{G}_{q_{1},\dots,a_{t}}$ and $\phi_{t}(q_{1},\dots,a_{t})=(x_{0}^{*},x_{1}^{*})$ otherwise.
+In either case, $\phi_{t}(q_{1},\dots,a_{t})$ is a maximal solution for conditions ($\phi$.1) to ($\phi$.3).
+Claim 8.30.
 The function $\phi_{t}$ is well-defined and exists.
 Proof.
-The proof consists of two steps: (i) ensuring that the optimization problems are feasible, and (ii) applying the transfinite recursion theorem (Lemma A.26) and proving that $\phi_{t}$ exists.
-Fix any $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$.
-By Remark A.24, each answer in $A$ is either the smallest element of $A$ or has an immediate predecessor.
-We will inductively show that for every $a_{t}\in A$, the values $\frac{1}{\delta}\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})$ and $\frac{1}{\delta}\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})$ satisfy the constraints of the optimization problem defining $\phi_{t}^{0}(q_{1},\dots,a_{t})$ and $\phi_{t}^{1}(q_{1},\dots,a_{t})$.
-$\frac{1}{\delta}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t})$ trivially satisfies constraint ($\phi$1).
-Moreover, by Corollary A.17, it satisfies constraint ($\phi$3).
-To prove constraints ($\phi$2) is satisfied, we need to show that
-(2) $\sum_{a\in\Gamma_{<a_{t}}}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ $+\sum_{a\in\Gamma_{>a_{t}}\cup\{a_{t}\}}\frac{1}{\delta}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}),$
-We prove this for the base case (i.e., when $a_{t}$ is the smallest element in $A$) and the inductive step (i.e., when $a_{t}$ has a predecessor $a^{*}$) separately:
+The proof consists of two steps: (i) proving that $\mathcal{C}_{q_{1},\dots,a_{t}}$ is non-empty, and (ii) applying the transfinite recursion theorem (Lemma 8.25) and concluding that $\phi_{t}$ exists.
+(i): Fix any $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$.
+By Remark 8.23, each answer in $A$ is either the smallest element of $A$ or has an immediate predecessor.
+We will inductively show that for every $a_{t}\in A$, the values $\frac{1}{\delta}\xi_{t}^{0}(q_{1},\dots,a_{t})$ and $\frac{1}{\delta}\xi_{t}^{1}(q_{1},\dots,a_{t})$ satisfy all the constraints for the optimization problem defining $\phi_{t}(q_{1},\dots,a_{t})$, and consequently, $\mathcal{C}_{q_{1},\dots,a_{t}}$ is not empty.
+First, note that $\frac{1}{\delta}\xi_{t}^{b}(q_{1},\dots,a_{t})$ trivially satisfies Constraint ($\phi$.1).
+Moreover, by Constraint ($\xi$.3), it satisfies Constraint ($\phi$.3).
+To prove Constraint ($\phi$.2) is satisfied,it suffices to show that
+(3) $\sum_{a\in\Gamma_{<a_{t}}}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ $+\sum_{a\in\Gamma_{>a_{t}}\cup\{a_{t}\}}\frac{1}{\delta}\xi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}),$
+We prove this by induction on $a_{t}\in A$:
 Base Case: Suppose $a_{t}$ is the least element of $A$. Then, $\Gamma_{<a_{t}}$ is empty, reducing our goal to showing that
-$\sum_{\begin{subarray}{c}a\in A\end{subarray}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\delta\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
-We know that $\phi_{t-1}(q_{1},\dots,a_{t-1})$ satisfies condition (I). Thus, by Lemma A.21,
-$\delta\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$ $\geq\mathrm{L}_{t-1}^{b}(q_{1},\dots,a_{t-1})\geq\sum_{\begin{subarray}{c}a\in A\end{subarray}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a),$
-finishing the proof for the base case.
-Inductive Step: Let $a_{*}$ be the immediate predecessor of $a_{t}$.
-Since $\Gamma_{<a_{t}}=\Gamma_{<a^{*}}\cup\{a^{*}\}$, constraints ($\phi$2) for defining $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a^{*})$ is identical to (2).
-By inductive hypothesis, $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a^{*})$ exists and satisfies these constraints.
+$\sum_{\begin{subarray}{c}a\in A\end{subarray}}\xi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\delta\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
+Let $a+t=a_{1}^{\prime}<_{A}a_{2}^{\prime}<_{A}\dots$ denote the ordered elements of $A$. As $\mathrm{L}_{t}$ is non-negative, by Constraint ($\xi$.2), for every $n\in\mathbb{N}$, we have
+$\sum_{\begin{subarray}{c}a\in\Gamma_{<a_{n}^{\prime}}\cup\{a_{n}^{\prime}\}\end{subarray}}\xi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\delta\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
+Taking $\lim_{n\to\infty}$ when $A$ is infinite or setting $n=|A|$ when $A$ is finite gives inequality (3).
+Inductive Step: Let $a^{*}$ be the immediate predecessor of $a_{t}$.
+Since $\Gamma_{<a_{t}}=\Gamma_{<a^{*}}\cup\{a^{*}\}$, Constraints ($\phi$.2) for defining $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a^{*})$ is identical to (3).
+By inductive hypothesis, $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a^{*})$ exists and satisfies ($\phi$.2).
 Thus, inequality (3) holds, completing the induction.
-Since $<_{t}$ is a well-order and the value of $\phi_{t}(q_{1},\dots,a_{t})$ depends only on preceding values in this order, by the transfinite recursion theorem, the function $\phi_{t}$ is well-defined and exists.
+(ii): For every $(q_{1},\dots,a_{t-1},q_{t})$, the values of the function $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ is uniquely determined by the preceding values of this function in the well-order $_{A}$.
+Thus, by the transfinite recursion theorem, the function $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ exists and is uniquely defined.
+Consequently, the function $\phi_{t}$ exists and is uniquely defined.
 ∎
-A.3.3. Satisfying Conditions (I) to (III)
-Our main objective is to construct $\phi_{t}^{b}$ so that it fulfills conditions (I), (II), and (III).
-Constraints ($\phi$1) and ($\phi$3) directly imply conditions (I) and (II), respectively.
-It remains to show $\phi_{t}$ satisfies conditions (III).
-First, we claim for each $b\in\{0,1\}$ and every transcript $(q_{1},\dots,a_{t-1},q_{t})$, $\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ is at most $\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$: By constraint ($\phi$2) and non-negativity of $\mathrm{L}^{b}$, for every $a_{t}$,
-$\sum_{a\in\Gamma_{<a_{t}}\cup\{a_{t}\}}\phi_{t}^{b}(q_{1},\dots,a_{t})\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
-Thus, when $A$ is finite and $a^{\prime}$ is the maximum element of $A$, we have
-$\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)=\sum_{\mathclap{a\in\Gamma_{<a^{\prime}}\cup\{a^{\prime}\}}}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a^{\prime})\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
-Moreover, when $A$ is infinite and $a_{1}^{\prime}<_{A}a_{2}^{\prime}<_{A}\dots$ are the sorted elements of $A$, we have
-$\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ $=\lim_{k\to\mathbb{N}}\sum_{a\in\Gamma_{<a_{k}^{\prime}}\cup\{a_{k}^{\prime}\}}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a_{k}^{\prime})\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
-The rest of proof is identical to the proof given in [18]. We repeat it for completeness. For the sake of contradiction, suppose there exists $b\in\{0,1\}$ and $(q_{1},\dots,a_{t-1},q_{t})$ such that
+8.3.4. Satisfying Conditions (I) to (IV)
+Our objective was to construct $\phi_{t}^{b}$ so that it fulfills conditions (I) to (IV).
+Combination of Constraints ($\phi$.1) and ($\xi$.1) gives condition (I).
+Moreover, Constraint ($\phi$.3) directly imply condition (II).
+The proof of the following claim is identical to a proof in [19]:
+Claim 8.31.
+$\phi_{t}$ satisfies conditions (III).
+Proof.
+For each $b\in\{0,1\}$ and every transcript $(q_{1},\dots,a_{t-1},q_{t})$, by Constraint ($\phi$.2) and non-negativity of $\xi^{b}$, for every $a_{t}$,
+$\sum_{a\in\Gamma_{<a_{t}}\cup\{a_{t}\}}\phi_{t}^{b}(q_{1},\dots,a_{t})\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}),$
+which in both cases where $A$ is finite and countable implies that
+$\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
+For the sake of contradiction, suppose there exists $b\in\{0,1\}$ and $(q_{1},\dots,a_{t-1},q_{t})$ such that
 (4) $\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)<\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
 Fix $b$.
-As $\phi_{t}^{b}$ is lower bounded by $\mathrm{L}_{t}^{b}/\delta$, this implies that for every $a_{t}\in A$, in the optimization problem determining the value of $\phi_{t}(q_{1},\dots,a_{t})$, constraint ($\phi$2) for $b$ is not tight.
+As $\phi_{t}^{b}$ is lower bounded by $\xi_{t}^{b}/\delta$ and the values of $\phi_{t}$ are maximal solutions, this strict inequality implies that for every $a_{t}\in A$, Constraint ($\phi$.2) for $b$ is not tight.
 Consequently, for every $a_{t}\in A$, we must have had
-$\phi_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\phi_{t}^{1-b}(q_{1},\dots,a_{t})=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$
-otherwise, the value of $\phi_{t}^{b}(q_{1},\dots,a_{t})$ could have increased, which contradicts the maximality of $\phi^{t}_{b}$.
-Taking a sum over all $a_{t}\in A$, we get
+$\delta\phi_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\delta\phi_{t}^{1-b}(q_{1},\dots,a_{t})=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$
+otherwise, the value of $\phi_{t}^{b}(q_{1},\dots,a_{t})$ could have increased, which contradicts the maximality. Taking a sum over all $a_{t}\in A$, we get
 (5) $\sum_{a_{t}\in A}\delta\phi_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\delta\sum_{a_{t}\in A}\phi_{t}^{1-b}(q_{1},\dots,a_{t})=\mu_{t-1}^{b}(q_{1},\dots,a_{t-1})-e^{-\epsilon}\mu_{t-1}^{1-b}(q_{1},\dots,a_{t-1})$
-By (II) and the inductive hypothesis, we know that the RHS is at least
+By (II) and Remark 8.21, we know that the RHS is at least
 $\delta\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})-e^{-\epsilon}\delta\phi_{t-1}^{1-b}(q_{1},\dots,a_{t-1})$
 Plugging this to (5) and reformulating it, we get
-$\sum_{a_{t}\in A}\phi_{t}^{b}(q_{1},\dots,a_{t})$ $-\phi_{t}^{b}(q_{1},\dots,a_{t-1})\geq e^{-\epsilon}(\sum_{a_{t}\in A}\phi_{t}^{1-b}(q_{1},\dots,a_{t})-\phi_{t}^{1-b}(q_{1},\dots,a_{t-1}))$
+$\sum_{a_{t}\in A}\phi_{t}^{b}(q_{1},\dots,a_{t})$ $-\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})\geq e^{-\epsilon}(\sum_{a_{t}\in A}\phi_{t}^{1-b}(q_{1},\dots,a_{t})-\phi_{t-1}^{1-b}(q_{1},\dots,a_{t-1}))$
 By (4), the LHS is negative, implying that the RHS is negative too.
@@
 Hence, by an identical argument, we have
-$\sum_{a_{t}\in A}\phi_{t}^{1-b}(q_{1},\dots,a_{t})$ $-\phi_{t}^{1-b}(q_{1},\dots,a_{t-1})\geq e^{-\epsilon}(\sum_{a_{t}\in A}\phi_{t}^{b}(q_{1},\dots,a_{t})-\phi_{t}^{b}(q_{1},\dots,a_{t-1})),$
-which is a contradiction when $\epsilon>0$, which implies that condition (III) holds.
-Appendix B Overview of [20]
-For $\epsilon\geq 0$, let $\mathcal{M}:S_{\mathcal{M}}\times Q_{\mathcal{M}}\to S_{\mathcal{M}}\times A_{\mathcal{M}}$ be an $\epsilon$-DP IM w.r.t.
-a neighboring relation $\sim$.
-Suppose that $\mathcal{M}$ has discrete answer distributions.
-Vadhan and Wang [20] shows that for any two neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,0}(b)$ are identical.
-In this section, we overview the construction of $\mathcal{P}$ and show it satisfies the following property:
-Claim B.1.
-For any $t\in\mathbb{N}$ and any query sequence $(q_{1},\dots,q_{t})\in Q_{\mathcal{M}}^{t}$, if the distributions of the first $t$ answers of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to the queries $(q_{1},\dots,q_{t})$ are identical, then for every $b\in\{0,1\}$, when $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,0}(b)$ receives the queries $q_{1},\dots,q_{t}$, $\mathcal{P}$ generates answers itself, without interacting with $\mathrm{RR}_{\epsilon,0}(b)$.
-Since $\delta=0$, $\mathrm{RR}_{\epsilon,0}(b)$ returns $(b,\bot)$ with probability $\frac{e^{\epsilon}}{1+e^{\epsilon}}$ and $(1-b,\bot)$ otherwise.
-For each $b$ and query sequence $(q_{1},\dots,q_{t})\in Q_{\mathcal{M}}^{t}$, let $\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{b})}$ denote the distribution of the first $t$ answers of $\mathcal{M}(s_{b})$ in response to $q_{1},\dots,q_{t}$.
-If $\epsilon=0$, then define $\nu_{q_{1},\dots,q_{t}}^{b}$ to be identical to $\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{b})}$.
-Otherwise, set
-$\nu_{q_{1},\dots,q_{t}}^{b}=\frac{e^{\epsilon}\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{b})}-\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{1-b})}}{e^{\epsilon}-1}$
-For $t\in\mathbb{N}$, $b\in\{0,1\}$, and $(q_{1},\dots,a_{t-1},q_{t})\in(Q_{\mathcal{M}}\times A_{\mathcal{M}})^{t-1}\times Q$, if $\nu_{q_{1},\dots,q_{t-1}}^{b}(q_{1},\dots,a_{t-1})>0$, the function $T_{q_{1},\dots,a_{t-1},q_{t}}:A_{\mathcal{M}}\to[0,1]$ is defined as
-$T_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=\frac{\nu_{q_{1},\dots,q_{t}}^{b}(q_{1},\dots,a_{t})}{\nu_{q_{1},\dots,q_{t-1}}^{b}(q_{1},\dots,a_{t-1})}$
-Vadhan and Wang [20] define $\mathcal{P}$ as follows: The initial state of $\mathcal{P}$ is $\left(\text{null},()\right)$, where null is an empty field and ()()( ) represents an empty sequence.
-Let $h$ denote the sequence.
-Let $(r,\bot)$ denote the outcome of $\mathrm{RR}_{\epsilon,0}(b)$.
-After interacting with $\mathrm{RR}_{\epsilon,0}(b)$, $\mathcal{P}$ updates null to $r$.
-Moreover, upon receiving a query $q$ and before returning an answer $a$, $\mathcal{P}$ appends $q$ and $a$ to the history sequence $h$.
-To generate a response $a$ to a query $q$, $\mathcal{P}$ samples $a$ from the distribution $T_{h,q}^{r}$, where $h$ is the history of all previous queries and answers.
-We observe that as long as $T_{h,q}^{0}$ and $T_{h,q}^{1}$ are identical, $\mathcal{P}$ does not need to know the value of $r$.
-Thus, to defer interaction with $\mathrm{RR}_{\epsilon,0}(b)$, we define $\mathcal{P}$ to interact with $\mathrm{RR}_{\epsilon,0}(b)$ only when these distributions are different for the first time.
-Proof of Claim B.1.
-That the distributions $\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{0})}$ and $\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{1})}$ are identical implies that for every $k\leq t$, $\mu_{q_{1},\dots,q_{k}}^{\mathcal{M}(s_{0})}$ and $\mu_{q_{1},\dots,q_{k}}^{\mathcal{M}(s_{1})}$ are identical.
-Consequently, by definition, both distributions $\nu_{q_{1},\dots,q_{k}}^{0}$ and $\nu_{q_{1},\dots,q_{k}}^{1}$ are identical to $\mu_{q_{1},\dots,q_{k}}^{0}$.
-Therefore, the conditional distributions $T_{q_{1},\dots,a_{t-1},q_{t}}^{0}$ and $T_{q_{1},\dots,a_{t-1},q_{t}}^{1}$ used by $\mathcal{P}$ are also identical, implying that $\mathcal{P}$ does not interact with $\mathrm{RR}_{\epsilon,0}$ to generate the first $t$ answers.
+$\sum_{a_{t}\in A}\phi_{t}^{1-b}(q_{1},\dots,a_{t})$ $-\phi_{t-1}^{1-b}(q_{1},\dots,a_{t-1})\geq e^{-\epsilon}(\sum_{a_{t}\in A}\phi_{t}^{b}(q_{1},\dots,a_{t})-\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})),$
+which is a contradiction when $\epsilon>0$, which implies that condition (III) holds. ∎
+To show $\phi_{t}$ satisfies conditions (IV), we need the following lemma:
+Lemma 8.32.
+For any query sequence $(q_{1},\dots,q_{t})\in Q^{t}$, if $\mu_{t}^{0}(q_{1},\dots,a_{t})=\mu_{t}^{1}(q_{1},\dots,a_{t})$ for every answer sequence $(a_{1},\dots,a_{t})\in A^{t}$, then
+$\xi_{t}^{0}(q_{1},\dots,a_{t})=\xi_{t}^{1}(q_{1},\dots,a_{t})=\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}$
+for every answer sequence $(a_{1},\dots,a_{t})\in A^{t}$.
+Proof.
+Fix $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$. We will inductively show that for every $a_{t}\in A$,
+$\xi_{t}^{0}(q_{1},\dots,a_{t})=\xi_{t}^{1}(q_{1},\dots,a_{t})=\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}.$
+Recall $\xi_{t}(q_{1},\dots,a_{t})$ is defined as the lexicographically maximum feasible solution $(x_{0}^{*},x_{1}^{*})$. By Constraint ($\xi$.4), for each $b\in\{0,1\}$, we have
+$x_{b}^{*}\leq\min\{\mu_{t}^{b}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}.$
+Therefore, to prove the lemma, it suffices to show
+$x_{0}=x_{1}=\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}$
+is a feasible solution:
+  • ($\xi$.1): By Lemma 8.17 and non-negativity of $\xi_{t}$, this constraint is satisfied.
+  • ($\xi$.2): To show Constraint ($\xi$.2) is satisfied, we need to show for each $b\in\{0,1\}$,
+    $\sum_{a\in\Gamma_{<a_{t}}\cup\{a_{t}\}}$ $\min\{\mu_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t-1},q_{t},a)+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t-1},q_{t},a)\}$
+    $+\sum_{a\in\Gamma_{>a_{t}}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$
+    Since
+    $\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\min\{\mu_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t-1},q_{t},a)+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t-1},q_{t},a)\},$
+    it suffices to prove
+    $\sum_{a\in A}$ $\min\{\mu_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t-1},q_{t},a)+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t-1},q_{t},a)\}\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$
+    By condition (IV) and Remark 8.21,
+    $\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})\geq\min\{\mu_{t-1}^{b}(q_{1},\dots,a_{t-1}),\mathrm{L}_{t-1}^{0}(q_{1},\dots,a_{t-1})+\mathrm{L}_{t-1}^{1}(q_{1},\dots,a_{t-1})\}$
+    Therefore, by Lemma 8.10 and Lemma 8.20,
+    $\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})\geq\min\{\mu_{t-1}^{b}(q_{1},\dots,a_{t-1}),\mathrm{L}_{t-1}^{0}(q_{1},\dots,a_{t-1})+\mathrm{L}_{t-1}^{1}(q_{1},\dots,a_{t-1})\}$
+    $\geq\min\left\{\sum_{a\in A}\mu_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a),\sum_{a\in A}\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t-1},q_{t},a)+\sum_{a\in A}\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t-1},q_{t},a)\right\}$
+    $\geq\sum_{a\in A}\min\{\mu_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t-1},q_{t},a)+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t-1},q_{t},a)\},$
+    completing the proof.
+  • ($\xi$.3): Recall that $\mu_{t}^{0}(q_{1},\dots,a_{t})=\mu_{t}^{1}(q_{1},\dots,a_{t})$ for every answer sequence $(a_{1},\dots,a_{t})\in A^{t}$.
+To prove this constraint is satisfied, we need to show
+    $(1-e^{-\epsilon})\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}\leq(1-e^{-\epsilon})\mu_{t}^{0}(q_{1},\dots,a_{t}),$
+    which is trivially true.
+  • ($\xi$.4): This constraint is trivially satisfied as equality holds.
+Claim 8.33.
+$\phi_{t}$ satisfies conditions (IV).
+Proof.
+For any query sequence $(q_{1},\dots,q_{t})\in Q^{t}$, if
+($\star$) $\mu_{t}^{0}(q_{1},\dots,a_{t})=\mu_{t}^{1}(q_{1},\dots,a_{t})$
+for every answer sequence $(a_{1},\dots,a_{t})\in A^{t}$, then by Lemma 8.32,
+($\star\star$) $\xi_{t}^{0}(q_{1},\dots,a_{t})=\xi_{t}^{1}(q_{1},\dots,a_{t})=\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}$
+for all $(a_{1},\dots,a_{t})\in A^{t}$.
+Moreover, by Lemma 8.10 and $\star$8.3.4, $\mu_{t-1}^{0}(q_{1},\dots,a_{t-1})=\mu_{t-1}^{1}(q_{1},\dots,a_{t-1})$ for all $(a_{1},\dots,a_{t})\in A^{t}$.
+Thus, by Remark 8.21 and condition (IV),
+($\star\star\star$) $\phi_{t-1}^{0}(q_{1},\dots,a_{t-1})=\phi_{t-1}^{1}(q_{1},\dots,a_{t-1}).$
+Therefore, by equations ($\star$8.3.4), ($\star\star$8.3.4) and ($\star\star\star$8.3.4), Constraints ($\phi$.1), ($\phi$.2), and ($\phi$.3) are symmetric for $x_{0}$ and $x_{1}$.
+Hence, there exists a maximal feasible solution $(z^{*},z^{*})$ satisfying all the constraints.
+By definition of $\phi_{t}(q_{1},\dots,a_{t})$, we have $\phi_{t}(q_{1},\dots,a_{t})=(z^{*},z^{*})$.
+Also, by Constraint ($\phi$.1), $z^{*}$ is at least $\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}$.
+9. $f^{\delta}_{\mathrm{RR}}$-concurrent composition
+In this section, we prove:
+Lemma (Restatement of Lemma 5.9).
+For every $0\leq\delta\leq 1$, the $f^{\delta}_{\mathrm{RR}}$-concurrent composition of continual mechanisms is $(0,\delta)$-differentially private.
+In Section 4, we assumed that all composed CMs have discrete answer distributions.
+Accordingly, we also prove Lemma 5.9 under this assumption.
+During this section, we assume the domain of all variables is finite or countable.
+For every random variable $Y$, we denote its support by $\mathrm{supp}(Y)$.
+We begin by proving some intermediate results needed for the final proof.
+Vadhan et al.
+[24] show that to guarantee an IM is differentially private, it suffices to consider only deterministic adversaries.
+Specifically, the following holds:
+Lemma 9.1 ([24]).
+An IM $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t.
+a neighboring relation $\sim$ if and and only if for every deterministic adversary $\mathcal{A}$ and every pair of neighboring initial states $s\sim s^{\prime}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}(s^{\prime}))$ are $(\epsilon,\delta)$-indistinguishable.
+By Definition 3.7, this result extends naturally to CMs:
+Corollary 9.2.
+A CM $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t.
+a verification function $f$ if and and only if for every deterministic adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
+The definition of $(0,\delta)$-DP CMs can be equivalently expressed using the notion of total variation:
+Definition 9.3 (Total Variation).
+Let $Y^{0}$ and $Y^{1}$ be two variables with the same domain $\mathcal{Y}$. The total variation distance between $Y^{0}$ and $Y^{1}$, denoted $\mathrm{TV}(Y^{0},Y^{1})$, is defined as
+$\mathrm{TV}(Y^{0},Y^{1})=\sup_{S\subseteq\mathcal{Y}}\Pr[Y^{0}\in S]-\Pr[Y^{1}\in S].$
+Lemma 9.4.
+Let $Y^{0}$ and $Y^{1}$ be two random variables over the same domain $\mathcal{Y}$.
+For every $0\leq\delta\leq 1$, $Y^{0}$ and $Y^{1}$ are $(0,\delta)$-indistinguishable if and only if $\mathrm{TV}(Y^{0},Y^{1})\leq\delta$.
+Proof.
+By definition, $Y^{0}$ and $Y^{1}$ are $(0,\delta)$-indistinguishable if and only if for every $S\subseteq\mathcal{Y}$, $\Pr[Y^{0}\in S]-\Pr[Y^{1}\in S]\leq\delta$.
+Therefore, $Y^{0}$ and $Y^{1}$ are $(0,\delta)$-indistinguishable if and only if $\mathrm{TV}(Y^{0},Y^{1})=\sup_{S\subseteq\mathcal{Y}}\Pr[Y^{0}\in S]-\Pr[Y^{1}\in S]\leq\delta$.
 ∎
+Lemma 9.5.
+[…]
+$\forall k>0,\quad\mathrm{TV}\left((W_{1}^{0},\dots,W_{k}^{0}),(W_{1}^{1},\dots,W_{k}^{1})\right)\leq\delta.$
+Then,
+$\mathrm{TV}(V^{0},V^{1})\leq\delta.$
+Proof.
+We have
+$\mathrm{TV}(V^{0},V^{1})$ $=\frac{1}{2}\sum_{\sigma\in\Sigma^{*}}\big|\Pr[V^{0}=\sigma]-\Pr[V^{1}=\sigma]\big|$
+$=\frac{1}{2}\sum_{\sigma\in\Sigma^{*}}\big|\Pr[W^{0}=\sigma\cdot\mathdollar^{\omega}]-\Pr[W^{1}=\sigma\cdot\mathdollar^{\omega}]\big|$
+$=\frac{1}{2}\sum_{\sigma^{\prime}\in\Sigma^{*}\cdot\mathdollar^{\omega}}\big|\Pr[W^{0}=\sigma^{\prime}]-\Pr[W^{1}=\sigma^{\prime}]\big|=\mathrm{TV}(W^{0},W^{1})$
+Thus, we need to show $\mathrm{TV}(W^{0},W^{1})\leq\delta$.
+For each $i\in\mathbb{N}$ and $b\in\{0,1\}$, let $W_{1:i}^{b}=(W_{1}^{b},\dots,W_{i}^{b})$ denote the prefix of length $i$, and let $W_{i:}^{b}=(W_{i}^{b},W_{i+1}^{b},\dots)$ denote the infinite suffix of $W^{b}$.
+Define:
+$T=\mathrm{TV}(W^{0},W^{1}),\quad T_{i}=\mathrm{TV}(W_{1:i}^{0},W_{1:i}^{1}),\quad\text{and}\quad\alpha(i)=\Pr[W_{i}^{0}\neq\mathdollar]+\Pr[W_{i}^{1}\neq\mathdollar$].
+We will show that
+  1. (1)
+    $T_{i}$ is non-decreasing in $i$.
+  2. (2)
+    For all $i\in\mathbb{N}$, we have $T-T_{i}\leq\alpha(i+1)$.
+Since each $T_{i}\leq 1$ and the sequence is non-decreasing, the limit $T^{*}=\lim_{i\to\infty}T_{i}$ exists.
+Moreover, since $W_{i}^{0}=\mathdollar$ if and only if $|V^{b}|<i$ and $V^{b}$ is finite with probability $1$, the probability $\Pr[W_{i}^{b}\neq\mathdollar$] converges to zero as $i\to\infty$.
+Therefore,
+$T-T^{*}=\lim_{i\to\infty}T-T_{i}\leq\lim_{i\to\infty}\alpha(i)=0.$
+Since $T_{i}\leq\delta$ for all $i$, we conclude that $T^{*}\leq\delta$. Thus, $T=\mathrm{TV}(W^{0},W^{1})\leq\delta$.
+It remains to prove (1) and (2).
+Since $W_{1:i}^{0}$ and $W_{1:i}^{1}$ in $T_{i}=\mathrm{TV}(W_{1:i}^{0},W_{1:i}^{1})$ equal $W_{1:i+1}^{0}$ and $W_{1:i+1}^{1}$ in $T_{i}=\mathrm{TV}(W_{1:i+1}^{0},W_{1:i+1}^{1})$ with their last entries removed, the data-processing inequality implies $T_{i}\leq T_{i+1}$.
+For (2), define $X_{i}^{b}=W_{1:i}^{b}\mathdollar^{\omega}$ for each $b\in\{0,1\}$.
+By definition, $T_{i}=\mathrm{TV}(X_{i}^{0},X_{i}^{1})$.
+We have
+$T-T_{i}$ $=\mathrm{TV}(W^{0},W^{1})-\mathrm{TV}(X_{i}^{0},X_{i}^{1})$
+$=\sup_{S}\left(\Pr[W^{0}\in S]-\Pr[W^{1}\in S]\right)-\sup_{S}\left(\Pr[X^{0}\in S]-\Pr[X^{1}\in S]\right)$
+$\leq\sup_{S}\left(\Pr[W^{0}\in S]-\Pr[W^{1}\in S]-\Pr[X^{0}\in S]+\Pr[X^{1}\in S]\right)$
+$\leq\sup_{S}\left(\Pr[W^{0}\in S]-\Pr[X^{0}\in S]\right)+\sup_{S}\left(\Pr[X^{1}\in S]-\Pr[W^{1}\in S]\right)$
+$=\Pr[W^{0}\neq X_{i}^{0}]+\Pr[W^{1}\neq X_{i}^{1}]$
+[…]
+$T-T_{i}\leq\Pr[W_{i+1}^{0}\neq\mathdollar]+\Pr[W_{i+1}^{1}\neq\mathdollar]=\alpha(i+1),$
+finishing the proof. ∎
+Lemma 9.6 (The Maximal Coupling Lemma).
+Let $Y^{0}$ and $Y^{1}$ be random variables with domain $\mathcal{Y}$.
+There exists a joint distribution for a pair of random variables $(Z^{0},Z^{1})$ on $\mathcal{Y}\times\mathcal{Y}$, called a coupling, such that:
+  1. (i)
+    For every $b\in\{0,1\}$, the marginal distribution of $Z^{b}$ is identical to the distribution of $Y^{b}$.
+  2. (ii)
+    Let $S^{0}=\mathrm{supp}(Y^{0})$ and $S^{1}=\mathrm{supp}(Y^{1})$. The probability of the event $Z^{0}=Z^{1}$ is maximized and satisfies
+    $\Pr[Z^{0}=Z^{1}]=\sum_{y\in\mathcal{Y}}\min\left\{\Pr[Z^{0}=y],\Pr[Z^{1}=y]\right\}=\sum_{y\in S^{0}\cap S^{1}}\min\left\{\Pr[Z^{0}=y],\Pr[Z^{1}=y]\right\}.$
+Lemma 9.7 (The Coupling Lemma).
+Let $Y^{0}$ and $Y^{1}$ be random variables with domain $\mathcal{Y}$.
+For every (not necessarily maximal) coupling $(Z^{0},Z^{1})$ of $(Y^{0},Y^{1})$, where the marginals of $Z^{0}$ and $Z^{1}$ are identically distributed as of $Y^{0}$ and $Y^{1}$, we have
+$\mathrm{TV}(Y^{0},Y^{1})\leq\Pr[Z^{0}\neq Z^{1}].$
+Lemma 9.8.
+Let $k\in\mathbb{N}$, and let $Y^{0}=(Y_{1}^{0},\dots,Y_{k}^{0})$ and $Y^{1}=(Y_{1}^{1},\dots,Y_{k}^{1})$ be random variables.
+For $i\in[k-1]$ and $b\in\{0,1\}$, set $S_{i}^{b}=\mathrm{supp}(W_{1:i}^{b})$ and $S_{i}=S_{i}^{0}\cap S_{i}^{1}$, with $S_{0}=S_{0}^{b}=\emptyset$.
+For $i\in[k]$, let $\Delta_{i}:S_{i-1}\to[0,1]$ be a function, and fix $\gamma\in[0,1]$.
+Suppose the following conditions hold:
+  • For every $i\in[k]$ and every $y_{1:i-1}\in S_{i-1}$,
+    $\mathrm{TV}\left((Y_{i}^{0}|Y_{1:i-1}^{0}=y_{1:i-1}),(Y_{i}^{1}|Y_{1:i-1}^{1}=y_{1:i-1})\right)\leq\Delta_{i}(w_{1:i-1}).$
+  • For every $y_{1:k-1}\in S_{k-1}$, $1-\prod_{j=1}^{i}\left(1-\Delta_{j}(w_{1:j})\right)\leq\gamma$.
+Then,
+$\mathrm{TV}(Y^{0},Y^{1})\leq\gamma.$
+Proof.
+We construct a coupling $\left(Z^{0}=(Z_{1}^{0},\dots,Z_{k}^{0}),Z^{1}=(Z_{1}^{1},\dots,Z_{k}^{1})\right)$ for $(Y^{0},Y^{1})$ and show that $\Pr[Z^{0}\neq Z^{1}]\leq\gamma$.
+By Lemma 9.7, this implies $\mathrm{TV}(Y^{0},Y^{1})\leq\gamma$.
+Throughout the proof, let $\mathcal{Y}$ denote the union of all possible outcomes of $Y_{i}^{b}$ over $i\in[k]$ and $b\in\{0,1\}$.
+Define $(Z_{1}^{0},Z_{1}^{1})$ to be the maximal coupling of $Y_{1}^{0}$ and $Y_{1}^{1}$ given by Lemma 9.6.
+For $i=2$ to $k$, given samples $(Z_{1:i-1}^{0},Z_{1:i-1}^{1})$, we define $(Z_{i}^{0},Z_{i}^{1})$ as follows:
+  • If the histories are identical, i.e.
+$Z_{1:i-1}^{0}=Z_{1:i-1}^{1}=y_{1:i-1}$, then $(Z_{i}^{0},Z_{i}^{1})$ is drawn from the maximal coupling of $Y_{i}^{0}\mid Y_{1:i-1}^{0}=y_{1:i-1}$} and $Y_{i}^{1}\mid Y_{1:i-1}^{1}=y_{1:i-1}$}, given by Lemma 9.6.
+  • Otherwise, if $Z_{1:i-1}^{0}=y_{1:i-1}^{0}$ and $Z_{1:i-1}^{1}=y_{1:i-1}^{1}$ differ, then $Z_{i}^{0}$ and $Z_{i}^{1}$ are drawn independently from their respective conditional distributions $Y_{i}^{0}\mid Y_{1:i-1}^{0}=y_{1:i-1}^{0}$} and $Y_{i}^{1}\mid Y_{1:i-1}^{1}=y_{1:i-1}^{1}$}.
+By the chain rule, the probability $\Pr[Z^{0}=Z^{1}]=\sum_{y_{1:k}\in S_{k}}\Pr[Z^{0}=Z^{1}\!=\!y_{1:k}]$ equals
+$\sum_{y_{1:k}\in S_{k}}\prod_{i=1}^{k}\Pr\!\left[Z_{i}^{0}\!=\!Z_{i}^{1}\!=\!y_{i}\,\middle|\,Z_{1:i-1}^{0}\!=\!Z_{1:i-1}^{1}\!=\!y_{1:i-1}\right]=$
+$\sum_{y_{1:k-1}\in S_{k-1}}\prod_{i=1}^{k-1}\Pr\!\left[Z_{i}^{0}\!=\!Z_{i}^{1}\!=\!y_{i}\,\middle|\,Z_{1:i-1}^{0}\!=\!Z_{1:i-1}^{1}\!=\!y_{1:i-1}\right]\!\!\!\!\sum_{\begin{subarray}{c}y_{k}\in\mathcal{Y}\\ \text{s.t.
+}y_{1:k}\in S_{k}\end{subarray}}\!\!\!\!\Pr\!\left[Z_{k}^{0}\!=\!Z_{k}^{1}\!=\!y_{k}\,\middle|\,Z_{1:k-1}^{0}\!=\!Z_{1:k-1}^{1}\!=\!y_{1:k-1}\right].$
+Conditioning on $Z_{1:k-1}^{0}=Z_{1:k-1}^{1}=y_{1:k-1}$, the pair $(Z_{k}^{0},Z_{k}^{1})$ is the maximal coupling of $Y_{i}^{0}\mid Y_{1:i-1}^{0}=y_{1:i-1}^{0}$} and $Y_{i}^{1}\mid Y_{1:i-1}^{1}=y_{1:i-1}^{1}$}.
+The innermost sum is taken over the common support of these conditional random variables.
+Thus, by Lemma 9.6, this summation equals
+$1-\mathrm{TV}\left((Y_{k}^{0}\mid Y_{1:k-1}^{0}=y_{1:k-1}),(Y_{k}^{1}\mid Y_{1:k-1}^{1}=y_{1:k-1})\right).$
+Hence, by the first assumption,
+(6) $\sum_{\begin{subarray}{c}y_{k}\in\mathcal{Y}\\ \text{s.t.
+}y_{1:k}\in S_{k}\end{subarray}}\!\!\!\!\Pr\!\left[Z_{k}^{0}\!=\!Z_{k}^{1}\!=\!y_{k}\,\middle|\,Z_{1:k-1}^{0}\!=\!Z_{1:k-1}^{1}\!=\!y_{1:k-1}\right]\geq 1-\Delta_{k}(y_{1:k-1})$
+By the second assumption, we know
+$1-\Delta_{k}(y_{1:k-1})\geq\frac{1-\gamma}{\prod_{i=1}^{k-1}(1-\Delta_{i}(y_{1:i-1}))}.$
+Therefore,
+$\Pr[Z^{0}=Z^{1}]$ $\geq(1-\gamma)\sum_{y_{1:k-1}\in S_{k-1}}\prod_{i=1}^{k-1}\frac{\Pr\!\left[Z_{i}^{0}\!=\!Z_{i}^{1}\!=\!y_{i}\,\middle|\,Z_{1:i-1}^{0}\!=\!Z_{1:i-1}^{1}\!=\!y_{1:i-1}\right]}{1-\Delta_{i}(y_{1:i-1})}$
+$=(1-\gamma)\sum_{y_{1:k-2}\in S_{k-2}}\prod_{i=1}^{k-2}\frac{\Pr\!\left[Z_{i}^{0}\!=\!Z_{i}^{1}\!=\!y_{i}\,\middle|\,Z_{1:i-1}^{0}\!=\!Z_{1:i-1}^{1}\!=\!y_{1:i-1}\right]}{1-\Delta_{i}(y_{1:i-1})}$
+$\qquad\qquad\qquad\qquad\quad\sum_{\begin{subarray}{c}y_{k-1}\in\mathcal{Y}\\ \text{s.t.
+}y_{1:k-1}\in S_{k-1}\end{subarray}}\!\!\!\!\frac{\Pr\!\left[Z_{k-1}^{0}\!=\!Z_{k-1}^{1}\!=\!y_{k-1}\,\middle|\,Z_{1:k-2}^{0}\!=\!Z_{1:k-2}^{1}\!=\!y_{1:k-1}\right]}{1-\Delta_{k-1}(y_{1:k-2})}.$
+The term $\Delta_{k-1}(y_{1:k-2})$ is independent of $y_{k-1}$. Moreover, by an argument identical to the one for Inequality 6, we have
+$\sum_{\begin{subarray}{c}y_{k-1}\in\mathcal{Y}\\ \text{s.t.
+}y_{1:k-1}\in S_{k-1}\end{subarray}}\!\!\!\!\Pr\!\left[Z_{k-1}^{0}\!=\!Z_{k-1}^{1}\!=\!y_{k-1}\,\middle|\,Z_{1:k-2}^{0}\!=\!Z_{1:k-2}^{1}\!=\!y_{1:k-1}\right]\geq 1-\Delta_{k-1}(y_{1:k-2}).$
+Thus,
+$\Pr[Z^{0}=Z^{1}]\geq(1-\gamma)\sum_{y_{1:k-2}\in S_{k-2}}\prod_{i=1}^{k-2}\frac{\Pr\!\left[Z_{i}^{0}\!=\!Z_{i}^{1}\!=\!y_{i}\,\middle|\,Z_{1:i-1}^{0}\!=\!Z_{1:i-1}^{1}\!=\!y_{1:i-1}\right]}{1-\Delta_{i}(y_{1:i-1})}$
+A backward iteration with the same proof implies that $\Pr[Z^{0}=Z^{1}]\geq 1-\gamma$. Consequently, $\Pr[Z^{0}\neq Z^{1}]\leq\gamma$, and by Lemma 9.7, $\mathrm{TV}(Y^{0},Y^{1})\leq\gamma$. ∎
+Proof of Lemma 5.9.
+By Corollary 9.2 and the definition of concurrent composition using the continual mechanism ExtConComp, it suffices if we show that for every deterministic adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{V}[f^{\delta}_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}})$ and $\mathrm{View}(\mathcal{A},\mathcal{V}[f^{\delta}_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}})$ are $(0,\delta)$-indistinguishable.
+Fix a deterministic adversary $\mathcal{A}$.
+Since $\mathcal{A}$ is deterministic, its sequence of random coins is empty (or predetermined), and its $i$-th query message to the verifier $\mathcal{V}[f^{\delta}_{\mathrm{RR}}]$ can be determined by the first i1i-1 responses from the verifier.
+Therefore, there exists a post-processing function $g$ such that for every $b\in\{0,1\}$, $g(V^{b})$ is equivalently distributed as $\mathrm{View}(\mathcal{A},\mathcal{V}[f^{\delta}_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}})$, where $V^{b}$ is the (finite) sequence of responses $V^{b}_{i}$ sent by $\mathcal{V}[f^{\delta}_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ to $\mathcal{A}$.
+By the (non-interactive) post-processing lemma, it suffices to prove that the random variables $V^{0}$ and $V^{1}$ are $(0,\delta)$-indistinguishable.
+[…]
+As $\mathcal{A}$ is deterministic, for every $i\in\mathbb{N}$ and $w_{1:i-1}=(w_{1},\dots,w_{i-1})\in S_{i-1}^{b}$, the random variable $W_{i}^{b}$ conditioned on $W_{1:i-1}=w_{1:i-1}$ is distributed as follows:
+  • If $w_{i-1}=\mathdollar$ (i.e., the communication has halted before) or the design of $\mathcal{A}$ is such that it halts the communication after receiving the i1i-1-st response $w_{i-1}$ given the history $w_{1:i-1}$, then $W_{1:i-1}=\mathdollar$ with probability $1$.
+  • Otherwise, $W_{1:i-1}$ is equivalently distributed to $\mathrm{RR}_{0,\delta_{i}^{\prime}}(b)$, where $\delta_{i}^{\prime}=\delta_{i}^{\prime}(w_{1:i-1},\mathcal{A})$ is determined by the history $w_{1:i-1}$ and the design of $\mathcal{A}$.
+Define $\Delta_{i}(w_{1:i-1})=0$ in the former case and $\Delta_{i}(w_{1:i-1})=\delta_{i}^{\prime}(w_{1:i-1},\mathcal{A})$ in the latter.
+By Lemma 9.4 and Lemma 2.4, for every $i\in\mathbb{N}$ and every $w_{1:i-1}\in S_{i-1}^{b}$,
+$\mathrm{TV}\left((W_{i}^{0}|W_{1:i-1}^{0}=w_{1:i-1}),(W_{i}^{1}|W_{1:i-1}^{1}=w_{1:i-1})\right)\leq\Delta_{i}(w_{1:i-1}).$
+Fix $w_{1:k-1}\in S_{k-1}^{0}$.
+Let $j$ be the smallest index such that $w_{j}=\mathdollar$.
+If there is no such index, set $j=k$.
+By Definition 5.8 (iii), $1-\prod_{i=1}^{j}(1-\Delta_{i}(w_{1:i-1}))\leq\delta$ for every $w_{1:i-1}\in S_{i-1}^{0}\cap S_{i-1}^{1}$.
+Moreover, by the definition of the functions $\Delta_{i}$, we have $\prod_{i=j+1}^{k}(1-\Delta_{i}(w_{1:i-1}))=1$.
+Therefore, $1-\prod_{i=1}^{k}(1-\Delta_{i}(w_{1:i-1}))\leq\delta$.
+Thus, by Lemma 9.8,
+$\mathrm{TV}(W_{1:k}^{0},W_{1:k}^{1})\leq\delta,$
+finishing the proof. ∎
+10. Concurrent Filter Composition
+In the parallel composition of continual mechanisms, the adversary sends pairs of identical queries to all but $k$ mechanisms with predetermined privacy parameters.
+In this section, we study a concurrent composition where the adversary can ask pairs of non-identical queries from all created mechanisms and choose the privacy parameters of all mechanisms adaptively.
+To limit the adversary’s power in this more general setting, we impose restrictions on the sequence of adaptively chosen privacy parameters.
+These restrictions are captured using a class of functions known as filters:
+Definition 10.1.
+A filter is a function $\mathrm{Filt}:\left([0,\infty)\times[0,1]\right)^{*}\to\{\top,\bot\}$ that takes a finite sequence of privacy parameters $(\epsilon_{i},\delta_{i})$ and returns either $\top$ or $\bot$.
+The output $\bot$ indicates that the privacy budget has been exhausted, and the interaction must be halted.
+Given a filter $\mathrm{Filt}$, we define a verification function $f[\mathrm{Filt}]$ as follows and refer to the $f[\mathrm{Filt}]$-concurrent composition of CMs as the concurrent $\mathrm{Filt}$-filter composition of CMs.
+The verification function $f[\mathrm{Filt}]$ verifies the validity of message formats, extracts the privacy parameters from the creation queries, and apply $\mathrm{Filt}$ on them.
+Definition 10.2.
+Let $\mathrm{Filt}$ be a filter.
+For every $t\in\mathbb{N}$ and every input sequence of messages $m_{1},\dots,m_{t}$, the verification function $f^{\mathrm{Filt}}$ returns $\top$ if and only if all of the following conditions hold:
+  1. (i)
+    The sequence $m_{1},\dots,m_{t}$ is suitable.
+  2. (ii)
+    Let $\ell$ denote the number of pairs of identical creation queries in the sequence $m_{1},\dots,m_{t}$, and let $\sigma=\left((\epsilon_{j}^{\prime},\delta_{j}^{\prime})\right)_{j=1}^{\ell}$ be the corresponding sequence of privacy parameters for the mechanisms created in those pairs.
+It must hold that $\mathrm{Filt}(\sigma)=\top$.
+Rogers et al.
+[21] studied the $\mathrm{Filt}$-filter composition of NIMs for the first time.
+More recently, Haney et al.
+[11] showed that if the $\mathrm{Filt}$-filter composition of NIMs is $(\epsilon,\delta)$-DP, then the same guarantee extends to the $\mathrm{Filt}$-filter composition of IMs.
+We further generalize this result to the $\mathrm{Filt}$-filter composition of CMs.
+To state our result, we need to define the $\mathrm{Filt}$-filter composition of randomized response mechanisms:
+Definition 10.3 ($f^{\mathrm{Filt}}_{\mathrm{RR}}$).
+Let $\mathrm{Filt}$ be a filter.
+For every $t\in\mathbb{N}$ and every input sequence of messages $m_{1},\dots,m_{t}$, the function $f^{\mathrm{Filt}}_{\mathrm{RR}}$ returns $\top$ if and only if all of the following conditions hold:
+  1. (i)
+    $f^{\mathrm{Filt}}(m_{1},\dots,m_{t})=\top$.
+  2. (ii)
+    All mechanisms created in $m_{1},\dots,m_{t}$ are randomized response mechanisms.
+We refer to the $f^{\mathrm{Filt}}_{\mathrm{RR}}$-concurrent composition of CMs as the $\mathrm{Filt}$-filter composition of randomized response mechanisms.
+Theorem 10.4.
+For every filter $\mathrm{Filt}$, if the $\mathrm{Filt}$-filter composition of randomized response mechanisms is $(\epsilon,\delta)$-DP, then the $\mathrm{Filt}$-filter composition of CMs is also $(\epsilon,\delta)$-DP.
+Proof.
+Let $f=f^{\mathrm{Filt}}$ and $f_{\mathrm{RR}}=f^{\mathrm{Filt}}_{\mathrm{RR}}$.
+This proof is identical to the proof of Theorem 4.9, except the following modifications: First, the right mechanism of the IPM $\mathcal{P}$ is $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ instead of ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((b)_{i=1}^{k}\right)$.
+Second, the variable $\sigma_{1}$ is set differently.
+In Theorem 4.9, $\mathcal{P}$ initially interacted with its right mechanism, ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((b)_{i=1}^{k}\right)$, and set $\sigma_{1}$ to its output sequence $(r_{1},\dots,r_{k})$, without ever changing it.
+Here, $\mathcal{P}$ interacts with its right mechanism, $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, each time it receives a pair of creation queries as a left message and appends a randomized response $r_{i}$ to $\sigma_{1}$.
+Specifically, the same as before, the state of $\mathcal{P}$ consists of two sequences $\sigma_{1}$ and $\sigma_{2}$, both initialized empty.
+Upon receiving a left message of the form $(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})^{2}$, $\mathcal{P}$ takes the following steps: First, $\mathcal{P}$ appends the pair $(\mathcal{T}_{i},t_{i})$ to the sequence $\sigma_{2}$, where $\mathcal{T}_{i}$ is the IPM corresponding to $\mathcal{M}_{i}$ from Lemma 4.6 and $t_{i}$ is its initial state.
+Then, it sends the message $(\beta_{\epsilon_{i},\delta_{i}},\beta_{\epsilon_{i},\delta_{i}})$ to its right mechanism $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, where $\beta_{\epsilon_{i},\delta_{i}}$ is the creation query corresponding to $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$.
+Upon receiving the acknowledgment response $\top$ from $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, $\mathcal{P}$ discards it and sends the message $\left((0,i),(1,i)\right)$ to its right mechanism again.
+This message will be transformed by $\mathcal{I}(b)$ to $(b,i)$.
+Receiving $(b,i)$, the mechanism ExtConComp sends $b$ to $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$ and returns its response $r_{i}$, which will be forwarded unchanged to $\mathcal{P}$.
+$\mathcal{P}$ then appends $r_{i}$ to $\sigma_{1}$.
+Finally, $\mathcal{P}$ returns the acknowledgment message $\top$ to its left mechanism.
+Upon receiving a left message $\left((q_{0},i),(q_{1},i)\right)$, exactly the same as in the proof of Theorem 4.9, $\mathcal{P}$ uses $r_{i}$ and $(\mathcal{T}_{i},t_{i})$ to generate a left response.
+To ensure that the privacy argument from Theorem 4.9 still holds under this modified construction, we must verify that $\mathcal{P}$ can always provide each $\mathcal{T}_{i}$ with a fresh output of $\mathrm{RR}_{\epsilon_{i},\delta_{i}}(b)$.
+This follows directly from the definitions of $f$ and $f_{\mathrm{RR}}$ and the design of $\mathcal{P}$.
+Specifically, the verifier $\mathcal{V}[f_{\mathrm{RR}}]$ never halts the communication and always returns the requested randomized response.
+∎
+Acknowledgments
+1Monika Henzinger and Roodabeh Safavi were supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No.
+101019564), and the Austrian Science Fund (FWF) under grants DOI 10.55776/Z422, DOI 10.55776/I5982, and DOI 10.55776/P33775, with additional funding from the netidee SCIENCE Stiftung, 2020–2024.
+2Salil Vadhan was supported by NSF grant BCS-2218803, a grant from the Sloan Foundation, and a Simons Investigator Award.
+Work began while a Visiting Researcher at the Bocconi University Department of Computing Sciences, supported by Luca Trevisan’s ERC Project GA-834861.

### v4

@@
 In addition, we prove concurrent filter compositions theorems for the scenarios in which the privacy parameters are adaptively chosen.
+We extend these results to other measures of differential privacy, including Rényi DP and $f$-DP.
 We then formalize a set of general conditions on a continual mechanism $\mathcal{M}$ that runs multiple continual sub-mechanisms such that the privacy guarantees of $\mathcal{M}$ follow directly using the above concurrent composition theorems on the sub-mechanisms, without further privacy loss.
-This enables us to give a simpler and more modular privacy analysis of a recent continual histogram mechanism of Henzinger, Sricharan, and Steiner.
+This enables us to give a simpler and modular privacy analysis of a recent continual histogram mechanism of Henzinger, Sricharan, and Steiner.
 In the case of approximate DP, ours is the first proof that shows that its privacy holds against adaptive adversaries.
-11footnotetext: Institute of Science and Technology, Klosterneuburg, Austria (email: monika.henzinger@ista.ac.at).22footnotetext: Institute of Science and Technology, Klosterneuburg, Austria (email: roodabeh.safavi@ista.ac.at).33footnotetext: Harvard University (email: salil_vadhan@harvard.edu).
+We also provide a framework that simplifies the analysis of local differential privacy when the protocol includes multi-round server-user interactions.
+Using this result, we simplify the privacy analysis of the core decomposition protocol of Dhulipala, Henzinger, Li, Liu, Sricharan, and Zhu (dhulipala2025near, ).
+copyright: none
 1. Introduction
-Differential privacy [6] is the now-standard framework for protecting privacy when performing statistical analysis or machine learning on sensitive datasets about individuals.
+Differential privacy (dwork2006calibrating, ) is the now-standard framework for protecting privacy when performing statistical analysis or machine learning on sensitive datasets about individuals.
 In addition to having a rich and rapidly growing scientific literature, differential privacy has also seen large-scale adoption by technology companies and government agencies.
@@
 a collection of statistics or the parameters of a machine learning model) $y\in\mathcal{Y}$.
-Differential privacy requires that every two datasets $x$ and $x^{\prime}$ that “differ on one individual’s data,” the output distributions $\mathcal{M}(x)$ and $\mathcal{M}(x^{\prime})$ are “close” to each other.
+Differential privacy requires that for every two datasets $x$ and $x^{\prime}$ that “differ on one individual’s data,” the output distributions $\mathcal{M}(x)$ and $\mathcal{M}(x^{\prime})$ are “close” to each other.
 Formally, we have a neighboring relation $\sim$ on $\mathcal{X}$, which specifies which datasets “differ on one individual’s data,” or more generally what information requires privacy protection.
@@
 This simple definition is very useful and easy to work with, but does not capture many scenarios in which we want to apply differential privacy.
-1.1. Differential Privacy Over Time
+Differential Privacy Over Time.
 Many intended uses of differential privacy involve a system that is set up to run continuously over a long period of time, making more statistical releases as either queries come in or the dataset is updated, for example with people leaving and entering.
-(The latter case, with dataset updates, is known as differential privacy under continual observation, and was introduced in the concurrent works of Dwork, Naor, Pitassi, and Rothblum [7] and Chan, Shi, and Song [2].) These settings introduce a new attack surface, where an adversary may be adaptive, issuing queries or influencing dataset updates based on the responses it receives.
+(The latter case, with dataset updates, is known as differential privacy under continual observation, and was introduced in the concurrent works of Dwork, Naor, Pitassi, and Rothblum (Dwork2010, ) and Chan, Shi, and Song (DBLP:journals/tissec/ChanSS11, ) – see (DBLP:conf/esa/HenzingerS25, ) for a survey.) These settings introduce a new attack surface, where an adversary may be adaptive, issuing queries or influencing dataset updates based on the responses it receives.
 When $\delta>0$, there are mechanisms that are differentially private against oblivious adversaries but not against adaptive ones.
 [^1] A simple artificial example is the following: the mechanism’s first release can include a uniformly random number $r$ from $\{1,2,\ldots,\lceil 1/\delta\rceil\}$, and then if the adversary’s next query is $r$, the mechanism publishes the entire dataset in the clear.
-One natural example is the “advanced composition theorem” for differentially privacy [9], which is known to fail if the adversary can select the privacy-loss parameters $(\epsilon_{i},\delta_{i})$ of the composed mechanisms adaptively [21].
+One natural example is the “advanced composition theorem” for differential privacy (dwork2010boosting, ), which is known to fail if the adversary can select the privacy-loss parameters $(\epsilon_{i},\delta_{i})$ of the composed mechanisms adaptively (rogers2016privacy, ).
 Until recently, adaptivity has been treated in an ad hoc manner in the differential privacy literature.
-For example, the aforementioned advanced composition theorem of [9] was formalized in a way that allows an adversary to adaptively choose the mechanisms and pairs of adjacent datasets, just not the privacy-loss parameters, and new composition theorems that allow for adaptive choices of privacy-loss parameters (known as privacy filters and privacy odometers) were given by [21, 25].
-Other early examples where adaptive queries were analyzed include the Sparse Vector Technique (SVT) [8, 12, 18] and the Private Multiplicative Weights (PMW) algorithm [12].
-A few years ago, Vadhan and Wang [23] initiated a systematic study of differentially private mechanisms $\mathcal{M}$ that support adaptive queries.
+For example, the aforementioned advanced composition theorem of (dwork2010boosting, ) was formalized in a way that allows an adversary to adaptively choose the mechanisms and pairs of adjacent datasets, just not the privacy-loss parameters, and new composition theorems that allow for adaptive choices of privacy-loss parameters (known as privacy filters and privacy odometers) were given by (rogers2016privacy, ; whitehouse2023fully, ).
+Other early examples where adaptive queries were analyzed include the Sparse Vector Technique (SVT) (dwork2009complexity, ; HardtRo10, ; lyu2016understanding, ) and the Private Multiplicative Weights (PMW) algorithm (HardtRo10, ).
+A few years ago, Vadhan and Wang (vadhan2021concurrent, ) initiated a systematic study of differentially private mechanisms $\mathcal{M}$ that support adaptive queries.
 Specifically, they formalized the notion of an interactive mechanism $\mathcal{M}$, which begins with a dataset as input and maintains a potentially secret state as it answers queries that come in over time.
@@
 They defined $\mathcal{M}$ to be an $(\epsilon,\delta)$-DP interactive mechanism if for every pair of adjacent datasets $x\sim x^{\prime}$ and every adaptive adversary $\mathcal{A}$, the view of $\mathcal{A}$ when interacting with $\mathcal{M}(x)$ is $(\epsilon,\delta)$-indistinguishable (in the sense of Inequality (1)) from its view when interacting with $\mathcal{M}(x^{\prime})$, where the view of $\mathcal{A}$ consists of the random coin tosses of $\mathcal{A}$ and the transcript of the interaction.
-Vadhan and Wang [23] asked whether our existing composition theorems for differential privacy, where the mechanisms being composed are noninteractive mechanisms, extend to the concurrent composition of interactive mechanisms.
+Vadhan and Wang (vadhan2021concurrent, ) asked whether our existing composition theorems for differential privacy, where the mechanisms being composed are noninteractive mechanisms, extend to the concurrent composition of interactive mechanisms.
 In concurrent composition of interactive mechanisms, an adaptive adversary can interleave its queries to the different interactive mechanisms being composed.
 This is a realistic attack scenario when multiple interactive differentially private mechanisms are deployed in practice, and is also a useful primitive for analyzing differentially private algorithms that use several interactive mechanisms as building blocks.
-The challenge in analyzing concurrent composition comes from the fact that the adversary can base its queries to $\mathcal{M}_{1}$ on answers given by $\mathcal{M}_{2}$ and vice-versa.
-The privacy guarantees on an interactive mechanism $\mathcal{M}$ only promise privacy against adversaries that do not get additional information about the sensitive dataset in the middle of its interaction with $\mathcal{M}$.
-Reasoning about the effect of such interleaved interactions is particularly tricky when the mechanisms $\mathcal{M}_{1}$ and/or $\mathcal{M}_{2}$ are not pure DP and can maintain secret state (like in the Sparse Vector Technique).
-Thus, it took several years before we had concurrent composition theorems that matched our composition theorems for noninteractive mechanisms.
-Note that the extension of composition theorems for noninteractive mechanism to concurrent composition theorems for interactive mechanisms is non-trivial: Let us call an interactive mechanism $M$ $M$-adaptive if it is differentially private against an adversary that was only interactive with $M$.
+Vadhan and Wang showed that the basic composition theorem for pure differential privacy (dwork2006calibrating, ) extends to concurrent composition.
+This result was later generalized by Lyu (lyu2022composition, ) to all composition theorems for approximate differential privacy.
+Subsequent works gave concurrent composition theorems for other flavors of differential privacy, such as Rényi DP (lyu2022composition, ) and $f$-DP (vadhan2022concurrent, ).
+Note that the extension of composition theorems for noninteractive mechanisms to concurrent composition theorems for interactive mechanisms is non-trivial: Let us call an interactive mechanism $M$ $M$-adaptive private if it is differentially private against an adversary that was only interactive with $M$.
 Consider an adversary that first interacts adaptively with $M_{1}$, then $M_{2}$ and then with $M_{1}$ again.
 Note that for the second sequence of interactions with $M_{1}$ the adversary takes into account its prior interaction with $M_{1}$ as well as with $M_{2}$.
-Thus, the adversary has more information about the data than if it had only interacted with $M_{1}$ before and the fact that $M_{1}$ is $M_{1}$-adaptive does not mean that it is adaptive against this interleaved composition of $M_{1}$ and $M_{2}$.
-Vadhan and Wang showed that the basic composition theorem for pure differential privacy [6] does extend to concurrent composition, but for approximate differential privacy, they obtained a much weaker bound for the concurrent case.
-A tight bound for concurrent composition of approximate differential privacy was given by Lyu [19].
-A series of works gave concurrent composition theorems for other flavors of differential privacy and composition, such as Rényi DP [19], $f$-DP [24], and adaptive choices of privacy parameters [11].
-Haney, Shoemate, Tian, Vadhan, Vyrros, Wang, and Xu [11] study the concurrent extension of composition theorems for non-interactive mechanisms where the adversary instantiates arbitrary mechanisms with adaptive privacy parameters.
+Thus, the adversary has more information about the data than if it had only interacted with $M_{1}$ before and the fact that $M_{1}$ is $M_{1}$-adaptive private does not mean that it is private against this interleaved composition of $M_{1}$ and $M_{2}$.
+Haney, Shoemate, Tian, Vadhan, Vyrros, Wang, and Xu (haney2023concurrent, ) study the concurrent extension of composition theorems for noninteractive mechanisms where the adversary instantiates arbitrary mechanisms with adaptive privacy parameters.
 This setting is referred to as filter composition, where a filter is a function $\mathrm{Filt}$ that maps the privacy parameters $(\epsilon_{i},\delta_{i})$ of the created mechanisms to an element of $\{\top,\bot\}$.
-Suppose an adversary adaptively creates noninteractive mechanisms such that $g$ applied to the privacy parameters of the created mechanisms equals $\top$ at every point in time.
+Suppose an adversary adaptively creates noninteractive mechanisms such that $\mathrm{Filt}$ applied to the privacy parameters of the created mechanisms equals $\top$ at every point in time.
 The $\mathrm{Filt}$-filter composition of noninteractive mechanisms is said to satisfy $(\epsilon,\delta)$-DP if the view of every such adversary is $(\epsilon,\delta)$-indistinguishable when all created mechanisms are run on dataset $x$ or its neighboring dataset $x^{\prime}$.
@@
 extend this definition to the concurrent $\mathrm{Filt}$-filter composition of interactive mechanisms, where the adversary may not only create new mechanisms but also interact with existing interactive ones.
-They prove that if the $\mathrm{Filt}$-filter composition of non-interactive mechanisms is $(\epsilon,\delta)$-DP, then the concurrent $\mathrm{Filt}$-filter composition of interactive mechanisms is also $(\epsilon,\delta)$-DP.
+They prove that if the $\mathrm{Filt}$-filter composition of noninteractive mechanisms is $(\epsilon,\delta)$-DP, then the concurrent $\mathrm{Filt}$-filter composition of interactive mechanisms is also $(\epsilon,\delta)$-DP.
 All of the above works are for the case where the queries to the mechanisms are adaptive, but the datasets are static, given as input to the mechanisms at the start of the interaction.
-[^2] One can extend the concurrent composition theorems to the case where, similarly to the advanced composition theorem [9], an adversary can adaptively select a pair of adjacent datasets $(x_{i},x_{i}^{\prime})$ for the $i$’th mechanism $\mathcal{M}_{i}$, but once the interactive mechanism $\mathcal{M}_{i}$ is started, its dataset is fixed and no further updates are allowed.
+[^2] One can extend the concurrent composition theorems to the case where, similarly to the advanced composition theorem (dwork2010boosting, ), an adversary can adaptively select a pair of adjacent datasets $(x_{i},x_{i}^{\prime})$ for the $i$’th mechanism $\mathcal{M}_{i}$, but once the interactive mechanism $\mathcal{M}_{i}$ is started, its dataset is fixed and no further updates are allowed.
 Thus, whenever $(x_{i+1},x_{i+1}^{\prime})$ are given, a new interactive mechanism $\mathcal{M}_{i+1}$ has to be started that does not know the secret state of $\mathcal{M}_{i}$.
@@
 For instance, Ghazi et al.
-[10] study an advertising problem in which the server receives both private data and queries over time.
-They design algorithms that accept dataset updates while applying non-interactive mechanisms—such as the discrete Laplace mechanism—to answer queries and analyze the privacy guarantees for their algorithm using existing results on privacy filters for composing non-interactive mechanisms with adaptive privacy-loss parameters.
-This independent work [^3] The first version of this work was available on arXiv prior to [10].
+(ghazi2025privacy, ) study an advertising problem in which the server receives both private data and queries over time.
+They design algorithms that accept dataset updates while applying noninteractive mechanisms—such as the discrete Laplace mechanism—to answer queries and analyze the privacy guarantees for their algorithm using existing results on privacy filters for composing noninteractive mechanisms with adaptive privacy-loss parameters.
+This independent work [^3] The first version of this work was available on arXiv prior to (ghazi2025privacy, ).
 shows that if a mechanism receiving both queries and dataset updates produces a differentially private output at each time step $i$ (with adaptive parameters $\epsilon_{i},\delta_{i}$ at step $i$) such that $\sum_{i}\epsilon_{i}\leq\epsilon$ and $\sum_{i}\delta_{i}\leq\delta$, then the overall mechanism is $(\epsilon,\delta)$-DP.
-This guarantee can be seen as a special case of our filter composition theorem, where the composition rule is simple summation and each composed mechanism is restricted to a non-interactive mechanism corresponding to a single step.
-While some continual mechanisms do indeed employ independent non-interactive mechanisms at each step, many important examples—such as the Sparse Vector Technique (SVT)—generate outputs that are correlated across time.
+This guarantee can be seen as a special case of our filter composition theorem, where the composition rule is simple summation and each composed mechanism is restricted to a noninteractive mechanism corresponding to a single step.
+While some continual mechanisms do indeed employ independent noninteractive mechanisms at each step, many important examples—such as the Sparse Vector Technique (SVT)—generate outputs that are correlated across time.
 For these mechanisms, privacy must be analyzed over the entire output sequence rather than on a step-by-step basis to get the desired guarantees.
-(If the $\epsilon$-DP SVT mechanism were analyzed on a step-by-step basis, then the privacy guarantee would degrade to $\Theta(q\cdot\epsilon)$-DP, where $q$ is the number of queries issued, defeating the entire point of SVT, which is to have a privacy guarantee and noise scale that is independent of the number of queries.) As a consequence, in the concurrent composition of continual mechanisms (or even interactive mechanisms), one cannot simply reduce the analysis to the composition of per-step non-interactive mechanisms.
-Indeed, this is why concurrent composition theorems for interactive mechanisms took several years to establish [23, 19, 24, 11].
+(If the $\epsilon$-DP SVT mechanism were analyzed on a step-by-step basis, then the privacy guarantee would degrade to $\Theta(q\cdot\epsilon)$-DP, where $q$ is the number of queries issued, defeating the entire point of SVT, which is to have a privacy guarantee and noise scale that is independent of the number of queries.) As a consequence, in the concurrent composition of continual mechanisms (or even interactive mechanisms), one cannot simply reduce the analysis to the composition of per-step noninteractive mechanisms.
+Indeed, this is why concurrent composition theorems for interactive mechanisms took several years to establish (vadhan2021concurrent, ; lyu2022composition, ; vadhan2022concurrent, ; haney2023concurrent, ).
 Thus, the results of Ghazi et al.
-[10] do not apply to the more general setting of (filter) concurrent composition of continual mechanisms.
+(ghazi2025privacy, ) do not apply to the more general setting of (filter) concurrent composition of continual mechanisms.
 [^4] It is Ghazi et al.’s definition of Individual DP (IDP, their Definition 4.1) and their composition theorem that is framed in terms of IDP that restricts to step-by-step privacy analyses corresponding to composition of noninteractive mechanisms.
 Their definition of $(\epsilon,\delta)$-DP interactive mechanisms (their Definition 3.6) is somewhat more general than their definition of IDP, but it still does not allow for exploiting correlated randomness across queries to reduce privacy loss, as in SVT.
-(This is because their Definition 3.4 requires the next state of the mechanism to be a function of only the dataset and the query, and not of the randomness used to generate the answer.
-Jain, Raskhodnikova, Sivakumar, and Smith [14] recently gave a formalization of adaptive adversaries under continual observation (where there are dataset updates but the queries are fixed) for the special case of event-level privacy, where two streams of updates are considered adjacent if they differ on only one update.
-The lack of a more general formalism and toolkit for reasoning about adaptivity in dataset updates has led to complicated and ad hoc proofs of privacy of continual mechanisms (see e.g.
-[13]).
+(This is because their Definition 3.4 requires the next state of the mechanism to be a function of only the dataset and the query, and not of the randomness used to generate the answer.)
+Jain, Raskhodnikova, Sivakumar, and Smith (pricedpco, ) recently gave a formalization of adaptive adversaries under continual observation (where there are dataset updates but the queries are fixed) for the special case of event-level privacy, where two streams of updates are considered adjacent if they differ on only one update.
+Denisov, McMahan, Rush, Smith and Thakurta (denisov2022improved, ) extended this definition to arbitrary neighbor relations on update sequences.
+The lack of a general toolkit for reasoning about adaptivity in dataset updates has led to complicated and ad hoc proofs of privacy of continual mechanisms (see e.g.
+(henzingerSS2023, )).
 Rectifying this situation is what motivated our work.
@@
 We note that the different flavors of mechanisms discussed above are differential privacy analogues of familiar variants of randomized algorithms and data structures.
-A non-interactive mechanism is simply a static (or batch) algorithm.
-An interactive mechanism corresponds to static data structure.
+A noninteractive mechanism is simply a static (or batch) algorithm.
+An interactive mechanism corresponds to a static data structure.
 Our focus in this paper is on continual mechanisms, which allow both queries and updates, and these can be viewed as dynamic data structures.
 Adaptivity is a common concern for the correctness properties of randomized data structures; here we are concerned instead with its effect on privacy properties.
-1.2. Our contributions
-(A) We give a formalism that captures privacy against adaptive adversaries for continual mechanisms, which support both queries and dataset updates.
-This generalizes the previously studied cases of interactive mechanisms (which have static datasets) and continual observation (which answer fixed queries).
-See Figure 1 for an illustration.
-We note that specific mechanisms, like variants of SVT, have been proposed to handle both dataset updates and queries [15], but without offering a definition of privacy against adversaries that can adaptively select both the updates and queries.
+Our contributions (A) We give a formalism that captures privacy against adaptive adversaries for continual mechanisms, which support both queries and dataset updates (Definition 3.7).
+This generalizes the previously studied cases of interactive mechanisms (which have static datasets) and continual observation (where mechanisms answer fixed queries).
+See Figure 4 in Appendix A for an illustration.
+We note that specific mechanisms, like variants of SVT, have been proposed to handle both dataset updates and queries (JainSmWa24, ), but without offering a definition of privacy against adversaries that can adaptively select both the updates and queries.
+(B) We prove concurrent composition theorems for continual mechanisms, where an adaptive adversary may interleave queries and updates across mechanisms based on the entire history of observed outputs.
+As discussed above for the concurrent composition of interactive mechanisms, the concurrent setting involves additional adaptivity that is not supported by composition theorems for noninteractive mechanisms.
+[^5] Existing results on the concurrent composition of interactive mechanisms do not apply to the concurrent composition of continual mechanisms, since in our model the adversary adaptively selects both (a) the mechanisms receiving neighboring inputs (including their privacy parameters and neighboring definition) and (b) the neighboring update sequences they receive over time, which is not the case for interactive mechanisms.
+We show tight privacy bounds for the following useful forms of concurrent composition of continual mechanisms.
+Further explanations are provided in Appendix B.
+  1. (1)
+    We show that when the privacy-loss parameters are fixed, the standard composition theorems for noninteractive approximate differential privacy extend to the concurrent composition of continual mechanisms (Theorem 4.12).
+This result also holds for Rényi differential privacy (Corollary 13.4) and for $f$-DP [^6] The $f$-DP results in this paper assume finite communication: for every created continual mechanism, there exists a constant $C$ bounding the sizes of its query and answer spaces and the number of queries it can answer before halting.
+This assumption was also made for interactive mechanisms in (vadhan2022concurrent, ).
+(Corollary 14.7).
+  1. (2)
+    We extend the classical “parallel composition theorem” for noninteractive mechanisms to the concurrent parallel composition of an unbounded [^7] Throughout this paper, we use the term “unbounded number” to refer to a quantity of unknown (potentially arbitrary) size.
+number of adaptively chosen interactive mechanisms with an adaptively chosen pair of neighboring datasets (Corollary 5.15).
+We show similar extensions for Rényi differential privacy (Corollary 13.7) and for $f$-DP (Corollary 14.9).
+  2. (3)
+    We extend the above concurrent parallel composition of interactive mechanisms to continual mechanisms when the composed mechanisms satisfy pure differential privacy (Corollary 5.5).
+For approximate DP, we construct a counterexample demonstrating that the concurrent parallel composition of an unbounded number of continual mechanisms with adaptively chosen dataset updates is not private (Theorem 5.2).
+Since $(\epsilon,\delta)$-DP is a special case of $f$-DP, this lower bound also applies to $f$-DP.
+We further construct an analogous counterexample for the case where the composed mechanisms satisfy Rényi DP instead of approximate DP (Theorem 13.5).
+  3. (4)
+    By imposing an upper bound on $\prod_{i}(1-\delta_{i})$, where $\delta_{i}$ is the approximate privacy parameter of the $i$-th created mechanism, we derive a new tight composition theorem for concurrent parallel composition of continual mechanisms (Theorem 5.4).
+  4. (5)
+    As an immediate consequence, our new theorem extends the classical parallel composition theorem for noninteractive mechanisms to the concurrent composition of an unbounded number of adaptively chosen continual mechanisms satisfying pure differential privacy (Corollary 5.5).
+  5. (6)
+    We show that, for every filter $\mathrm{Filt}$, the concurrent $\mathrm{Filt}$-filter composition of continual mechanisms enjoys the same privacy guarantees as the $\mathrm{Filt}$-filter composition of noninteractive mechanisms (Theorem 12.4).
+This result is also proved for Rényi DP (Corollary 13.8) and $f$-DP (Corollary 14.10).
+A summary of the positive results for the aforementioned types of concurrent composition of $(\epsilon,\delta)$-DP interactive and continual mechanisms is provided in Table 1.
+Concurrent Composition with Fixed Parameters Concurrent Parallel Composition Concurrent Filter Composition
+CMs
+Theorem 4.12 for $(\epsilon,\delta)$-DP
+Corollary 5.5 for $\epsilon$-DP
+Theorem 5.4 for $(\epsilon,\delta)$-DP
+Theorem 12.4 for $(\epsilon,\delta)$-DP
+IMs
+(vadhan2021concurrent, ) for $\epsilon$-DP
+(vadhan2022concurrent, ; lyu2022composition, ) for $(\epsilon,\delta)$-DP
+Corollary 5.15 for $(\epsilon,\delta)$-DP
+(haney2023concurrent, ) for $(\epsilon,\delta)$-DP
+Table 1. Concurrent composition results for $(\epsilon,\delta)$-DP interactive mechanisms (IMs) and continual mechanisms (CMs).
 […]
-Figure 1. Comparison between interactive mechanism (IMs), continual observation (CO), and continual mechanisms (CMs)
-(B) We prove concurrent composition theorems for continual mechanisms, where an adaptive adversary may interleave queries and updates across mechanisms based on the entire history of observed outputs.
-As discussed for the concurrent composition of interactive mechanisms, the concurrent setting involves additional adaptivity that is not supported by composition theorems for noninteractive mechanisms.
-Existing results on the concurrent composition of interactive mechanisms do not apply to the concurrent composition of continual mechanisms, since in our model the adversary adaptively selects both (a) the mechanisms receiving neighboring inputs (including their privacy parameters and neighboring definition) and (b) the neighboring update sequences they receive over time, which is not the case for interactive mechanisms.
-We show tight privacy bounds for the following useful forms of concurrent composition of continual mechanisms:
-(1) The first shows that, when the privacy-loss parameters are fixed, the standard composition theorems for noninteractive approximate differential privacy extend to the concurrent composition of continual mechanisms.
-(2) We extend the classical “parallel composition theorem” for noninteractive mechanisms to the concurrent parallel composition of an unbounded number of adaptively chosen interactive mechanisms with an adaptively chosen pair of neighboring datasets.
-(3) We construct a counterexample demonstrating that the concurrent parallel composition of an unbounded number of continual mechanisms with adaptively chosen dataset updates is not private.
-Then imposing an upper bound on the sum of the privacy parameter $\delta_{i}$ of the created mechanisms, we introduce a new composition theorem for concurrent parallel composition with this upper bound.
-(4) As a special case, our new composition theorem extends the “parallel composition theorem” for noninteractive mechanisms to the concurrent composition of an unbounded number of adaptively chosen continual mechanisms when the composed mechanisms satisfy pure differential privacy.
-(C) We formalize a set of general conditions on a continual mechanism $\mathcal{M}$ that runs multiple continual sub-mechanisms such that the privacy guarantees of $\mathcal{M}$ follow directly using the above concurrent composition theorems on the sub-mechanisms, without further privacy loss.
-These conditions are also reformulated as a set of properties for pseudocode describing continual mechanisms.
-(D) We use our concurrent composition theorems for continual mechanisms, with the aforementioned general conditions, to give a simpler and more modular privacy analysis of a recent continual histogram mechanism [13].
+(C) We formalize a set of general conditions on a continual mechanism $\mathcal{M}$ that runs multiple continual sub-mechanisms such that the privacy guarantees of $\mathcal{M}$ follow directly by using the above concurrent composition theorems on the sub-mechanisms, without further privacy loss (Theorem 6.1).
+These conditions are also reformulated as a set of properties for pseudocode describing continual mechanisms (Corollary 6.6).
+(D) We use our concurrent composition theorems for continual mechanisms, with the aforementioned general conditions, to give a simpler and more modular privacy analysis of a recent continual histogram mechanism (henzingerSS2023, ).
 In the case of approximate DP, ours is the first proof that the privacy of the mechanism holds against adaptive adversaries.
-(E) We how that, for every filter $\mathrm{Filt}$, the concurrent $\mathrm{Filt}$-filter composition of continual mechanisms enjoys the same privacy guarantees as the $\mathrm{Filt}$-filter composition of non-interactive mechanisms.
+(E) We model local protocols in the local differential privacy model as continual mechanisms with a modular structure, and demonstrate how our concurrent composition theorems for continual mechanisms—particularly the concurrent parallel composition theorem for interactive mechanisms—can simplify their privacy analysis (Theorem 8.7 and Corollary 8.8).
+As an example, we give a simple privacy analysis for the core decomposition local protocol proposed by Dhulipala, Henzinger, Li, Liu, Sricharan, and Zhu (dhulipala2025near, ).
 We elaborate on all of these contributions below.
@@
 To model static datasets (as in standard interactive differential privacy), the entire dataset can be given to $\mathcal{M}$ as its first message.
-Towards defining privacy, we consider adaptive adversaries $\mathcal{A}$ that send messages $m_{i}$ for $i=1,2,3,\dots$ that always are parsed as pairs $m_{i}=(m_{i}[0],m_{i}[1])$.
+Towards defining privacy, we consider adaptive adversaries $\mathcal{A}$ that send messages $m_{i}$ for $i=1,2,3,\dots$ that are always parsed as pairs $m_{i}=(m_{i}[0],m_{i}[1])$.
 Privacy is defined with respect to a verification function $f$ that takes a sequence of such pairs $(m_{1},m_{2},\ldots,m_{t})$ and outputs either $\top$ or $\bot$.
@@
 For continual mechanisms, we can interleave adaptive queries among the dataset updates by having $f$ require that all $m_{i}[0]=m_{i}[1]$ whenever either one is a query (rather than a dataset update).
-To go from a verification function $f$ to a definition of privacy against adaptive adversaries, we utilize the notion of an interactive post-processing mechanism (IPM) [11], which is a stateful procedure that can stand between an interactive mechanism and an adversary, or two interactive or continual mechanisms, or even two IPMs, and interact with both of them adaptively, sending and receiving messages to the mechanisms on its left and right.
-Specifically, we place two IPMs between the adversary $\mathcal{A}$ and the continual mechanism $\mathcal{M}$.
+To go from a verification function $f$ to a definition of privacy against adaptive adversaries, we utilize the notion of an interactive post-processing mechanism (IPM) (haney2023concurrent, ), which is a stateful procedure that can stand between an interactive mechanism and an adversary, or two interactive or continual mechanisms, or even two IPMs, and interact with both of them adaptively, sending and receiving messages to the mechanisms on its left and right.
+Specifically, as illustrated in Figure 1, we place two IPMs between the adversary $\mathcal{A}$ and the continual mechanism $\mathcal{M}$.
 The first is a verifier $\mathcal{V}[f]$ that, every time it receives a new message pair $m_{k}$ from $\mathcal{A}$, applies $f$ to the entire history of message pairs $(m_{1},\ldots,m_{k})$ sent by $\mathcal{A}$, and forwards $m_{k}$ onward only if $f$ returns $\top$.
@@
 The second IPM is the identifier $\mathcal{I}$, which starts out with a secret bit $b\in\{0,1\}$, and every time it receives a message pair $(m_{i}[0],m_{i}[1])$ from $\mathcal{V}[f]$, it forwards $m_{i}[b]$ on to the mechanism $\mathcal{M}$.
+The identifier $\mathcal{I}$ with bit $b$ is shown by $\mathcal{I}(b)$.
 When $\mathcal{M}$ returns an answer $a_{i}$, $\mathcal{I}$ and $\mathcal{V}[f]$ just forward it back to $\mathcal{A}$.
-See Figure 2 for an illustration of an interaction between $\mathcal{A}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}[1]\!\circ^{*}\!\mathcal{M}$ consisting of three messages from $\mathcal{A}$ to $\mathcal{V}[f]$.
+We denote the interactive mechanism representing the combination of $\mathcal{M}$, $I$, and $\mathcal{V}[f]$ by $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$.
 […]
 Figure 1. Illustration of the interactions between $\mathcal{A}$, $\mathcal{V}[f]$, $\mathcal{I}(b)$, and $\mathcal{M}$
+Illustration of the interactions between $\mathcal${A}\$, $\mathcal${V}[f]\$, $\mathcal{I}$(b)\$, and $\mathcal{M}$$
 This leads to the following definition of differential privacy for continual mechanisms:
@@
 Let $f$ be a verification function.
-A continual mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP with respect to $f$ if for every adaptive adversary $\mathcal{A}$, the view of $\mathcal{A}$ when interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}[0]\!\circ^{*}\!\mathcal{M}$ is $(\epsilon,\delta)$-indistinguishable from its view when interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}[1]\!\circ^{*}\!\mathcal{M}$.
-The idea of defining security against adaptive adversaries by having the adversary send pairs of messages and trying to distinguish whether the first message in each pair or the second message in each pair is being used goes back to the notion of left-or-right indistinguishability in crytography [1].
+A continual mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP with respect to $f$ if for every adaptive adversary $\mathcal{A}$, the view of $\mathcal{A}$ when interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}$ is $(\epsilon,\delta)$-indistinguishable from its view when interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M}$.
+The idea of defining security against adaptive adversaries by having the adversary send pairs of messages and trying to distinguish whether the first message in each pair or the second message in each pair is being used goes back to the notion of left-or-right indistinguishability in cryptography (BellareDJR97, ).
 What is different here is its combination with a verification function $f$ to enforce the neighboring condition, which is necessary to capture differential privacy.
@@
 Recall that with concurrent composition of interactive mechanisms, an adaptive adversary (or honest analyst) can interleave its queries to the different interactive mechanisms being composed; for concurrent composition of continual mechanisms, we also allow the dataset updates to be interleaved with the queries and each other.
-(See Figure 3.)
+See Figure 2 for illustration.
+[…]
+Figure 2. Illustration of the interactions between an honest adversary $\mathcal{A}$ and the concurrent composition of $k$ continual mechanisms
+Illustration of the interactions between an honest adversary $\mathcal${A}\$ and the concurrent composition of \$k\$ continual mechanisms
 Our first concurrent composition theorem is as follows:
@@
 $f_{i}$ is $(\epsilon,\delta)$-DP for the same $(\epsilon,\delta)$ that holds for composing noninteractive mechanisms that are $(\epsilon_{i},\delta_{i})$-DP.
-[…]
-Figure 3. Illustration of the interactions between an honest adversary $\mathcal{A}$ and the concurrent composition of $k$ continual mechanisms
 Notably, this result allows the adversary to adaptively and concurrently choose the neighboring sequences fed to each of the $\mathcal{M}_{i}$’s based on the responses received from the other mechanisms.
+Analogous results are shown for Rényi DP in Corollary 13.4 and, under the finite-communication assumption, for $f$-DP in Corollary 14.7.
 A key step in the proof of Theorem 1.2 is to show that for any $(\epsilon_{i},\delta_{i})$-DP continual mechanism $\mathcal{M}$, the interactive mechanism $\mathcal{M}^{\prime}(\cdot)=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(\cdot)\!\circ^{*}\!\mathcal{M}$ can be simulated by post-processing the randomized response mechanism $\mathrm{RR}_{\epsilon_{i},\delta_{i}}(\cdot)$.
-Lyu [19] proves such a result under the assumptions that the answer space of $\mathcal{M}$ is finite and that there exists a predetermined upper bound $T$ on the number of interactions.
+Lyu (lyu2022composition, ) proves such a result under the assumptions that the answer space of $\mathcal{M}$ is finite and that there exists a predetermined upper bound $T$ on the number of interactions.
+Vadhan and Zhang (vadhan2022concurrent, ) additionally assume that the query set is finite.
 We remove the latter assumption and relax the former to having an infinite (but discrete) answer space.
-While we focus on composing approximate DP with fixed privacy loss parameters $(\epsilon_{i},\delta_{i})$ in Theorem 1.2, we expect that there should also be analogues of it for other settings in which concurrent composition of interactive mechanisms have been studied, such as Rényi DP [19], $f$-DP [24], and adaptively chosen privacy-loss parameters [11].
 Next, we explore the concurrent and continual analogue of parallel composition of differential privacy.
-To review parallel composition, suppose that $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$ are each $(\epsilon_{0},\delta_{0})$-DP noninteractive mechanisms with respect to $\sim$, where each $\mathcal{M}_{i}$ expects a dataset $x_{i}$ as input, and for a parameter $k\in\{0,1,\ldots,\ell\}$, we define a neighboring relation $\sim_{k}$ on $\ell$-tuples as follows: $(x_{1}[0],\ldots,x_{\ell}[0])\sim_{k}(x_{1}[1],\ldots,x_{\ell}[1])$ if:
-  1. (1)
-    For every $i\in[\ell]$, $x_{i}[0]\sim x_{i}[1]$ or $x_{i}[0]=x_{i}[1]$.
-  2. (2)
-    For at most $k$ values of $i\in[\ell]$, $x_{i}[0]\neq x_{i}[1]$.
-We refer to the mechanism $\mathcal{M}^{\prime}(\cdot)=(\mathcal{M}_{1}(\cdot),\ldots,\mathcal{M}_{\ell}(\cdot))$ with neighbor relation $\sim_{k}$ as the $k$-sparse parallel composition of $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$.
-If $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$ are noninteractive mechanism, then it is known that $\mathcal{M}^{\prime}$ satisfies the same level of $(\epsilon,\delta)$-DP as the (standard) composition of $k$ (rather than $\ell$) $(\epsilon_{0},\delta_{0})$-DP mechanisms.
-We study the extent to which parallel composition extends to the concurrent composition of continual mechanisms, meaning the neighboring relation is as above, but the adversary can interact concurrently with the mechanisms, and in particular decide adaptively on which $k$ of them to have different update sequences.
-Moreover, we also consider the case where $\ell$ is unbounded, meaning the adversary (or honest analyst) can adaptively decide how many mechanisms to interact with.
-We refer to this as concurrent $k$-sparse parallel composition.
-A more restricted version of this composition, which is limited to noninteractive mechanisms, was previously studied in [20].
+To review parallel composition, suppose that $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$ are each $(\epsilon_{0},\delta_{0})$-DP noninteractive mechanisms with respect to a neighbor relation $\sim^{*}$, where each $\mathcal{M}_{i}$ expects a dataset $x_{i}$ as input, and for a parameter $k\in\{1,\ldots,\ell\}$, we define a neighboring relation $\sim^{(k)}$ on $\ell$-tuples as follows: $(x_{1},\ldots,x_{\ell})\sim^{(k)}(x_{1}^{\prime},\ldots,x_{\ell}^{\prime})$ if and only if $x_{i}=x_{i}^{\prime}$ for all but at most $k$ indices $i\in[\ell]$, for which $x_{i}\sim^{*}x_{i}^{\prime}$.
+We refer to the mechanism $\mathcal{M}^{\prime}(\cdot)=(\mathcal{M}_{1}(\cdot),\ldots,\mathcal{M}_{\ell}(\cdot))$ with neighbor relation $\sim^{(k)}$ as the $k$-sparse parallel composition of $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$.
+If $\mathcal{M}_{1},\ldots,\mathcal{M}_{\ell}$ are noninteractive mechanisms, then it is known that $\mathcal{M}^{\prime}$ satisfies the same level of $(\epsilon,\delta)$-DP as the (standard) composition of $k$ (rather than $\ell$) $(\epsilon_{0},\delta_{0})$-DP mechanisms.
+We extend this notion to the concurrent $k$-sparse parallel composition of continual mechanisms, where an adversary (or an honest analyst) may adaptively create continual mechanisms (with no bound $\ell$ on the number of mechanism), where each mechanism satisfies $(\epsilon_{0},\delta_{0})$-DP with respect to a verification function $f^{*}$, and interact with them concurrently.
+Formally, instead of fixing one dataset pair $(x_{i},x_{i}^{\prime})$ for each mechanism in advance, the adversary issues pairs of messages $m=(m^{0},m^{1})$ to mechanisms over time.
+All but at most $k$ mechanisms receive identical message pairs throughout the interaction, while for each of the $k$ exceptional mechanisms, the sequence of message pairs $\sigma$ satisfies $f^{*}(\sigma)=\top$.
+If messages are data records to be inserted in the dataset, then this concurrent composition captures adaptive partitioning of data among mechanisms over time.
+Note that an adversary can decide to designate a mechanism $\mathcal{M}_{i}$ as one of the $k$ “exceptional” mechanisms, i.e., issue at least one pair of non-identical messages to $\mathcal{M}_{i}$, after potentially interacting with other mechanisms and sending multiple identical message pairs to $\mathcal{M}_{i}$ itself.
 First, we extend the parallel composition of noninteractive mechanisms to concurrent composition of interactive mechanisms (where there are adaptive queries but datasets $x_{i}$ are static, rather than dynamically updated).
+[^8] As an intermediate step, Qiu et al.(qiu2022differential, ) analyzes the composition of infinitely many SVT instances (sharing a secret dataset) that are executed sequentially, where each instance begins only after the previous one halts, and where a differing element between two neighboring datasets can affect the output of at most one instance.
+Their result is weaker than ours as it does not allow for concurrent interaction of the adversary with the interactive mechanisms.
+We recall that an interactive mechanism is modeled as a continual mechanism receiving a dataset as its first message and queries as subsequent messages.
+(The verification function $f^{*}$ for this mechanism ensures that the first message is a pair of neighbor datasets, while the subsequent pairs are identical queries.) A formal statement of the following theorem is given in Corollary 5.15, which more generally composes interactive mechanisms $\mathcal{M}_{i}$ with differing $(\epsilon_{i},\delta_{i})$ parameters.
 Theorem 1.3 (concurrent parallel composition of interactive mechanisms).
 For every $\epsilon_{0}>0$ and $0\leq\delta_{0}\leq 1$ and $k\in\mathbb{N}$, and every $(\epsilon_{0},\delta_{0})$-DP interactive mechanism $\mathcal{M}_{0}$, the concurrent $k$-sparse parallel composition of an unbounded number of copies of $\mathcal{M}_{0}$ satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},\delta_{0})$-DP mechanisms.
-As discussed later in the paper, we also prove a generalization of Theorem 1.3 for the case of composing interactive mechanisms $\mathcal{M}_{i}$ with differing $(\epsilon_{i},\delta_{i})$ parameters.
-Then we turn to the case of continual mechanisms, where there can also be adaptive dataset updates. We prove a negative result for approximate DP ($\delta_{0}>0$):
+We prove analogous results for Rényi DP in Corollary 13.7 and, under the finite-communication assumption, for $f$-DP in Corollary 14.9.
+Then we turn to the case of continual mechanisms, where there can also be adaptive dataset updates.
+We prove a negative result for approximate DP ($\delta_{0}>0$):
 Theorem 1.4 (counterexample for the concurrent parallel composition of approximate DP continual mechanisms, informally stated).
 For every $\delta_{0}>0$, there is an $(0,\delta_{0})$-DP continual mechanism $\mathcal{M}_{0}$ such that for every $\ell\in\mathbb{N}$, the concurrent 1-sparse parallel composition of $\ell$ copies of $\mathcal{M}_{0}$ is not $(0,\delta)$-DP for any $\delta<1-(1-\delta_{0})^{\ell}$.
-This theorem says that, at least in terms of the $\delta$ parameter, concurrent parallel composition of continual mechanisms does not provide any benefit over non-parallel composition, since the bound of $1-(1-\delta_{0})^{\ell}$, is the bound that we would get by just applying composition over all $\ell$ mechanisms.
-Inspired by this counterexample, we prove a weaker version of Theorem 1.3 for the concurrent parallel composition of continual mechanisms.
+This theorem says that, at least in terms of the $\delta$ parameter, concurrent parallel composition of continual mechanisms does not provide any benefit over non-parallel composition, since the bound of $1-(1-\delta_{0})^{\ell}$ is the bound that we would get by just applying composition over all $\ell$ mechanisms.
+The high-level idea of this counterexample is to define a continual mechanism $\mathcal{M}_{0}$ that, upon receiving its first input message (regardless of the value), outputs $\bot$ with probability $\delta$ and $\top$ otherwise.
+If the first output is $\bot$, then on its second input the mechanism simply returns that input, revealing it entirely.
+Consider an adversary $\mathcal{A}$ that creates $\ell$ instances of $\mathcal{M}_{0}$, sends the pair $(0,0)$ to each instance, and then sends the pair $(0,1)$ to the first mechanism whose response was $\bot$.
+With probability $1-(1-\delta)^{\ell}$, such a mechanism exists and its second response reveals its input bit, which equals the secret bit used by the identifier to filter the adversary’s message pairs.
+An analogous counterexample for Rényi DP is provided in Theorem 13.5.
+Inspired by this counterexample, we prove a weaker version of Theorem 1.3 for the concurrent parallel composition of continual mechanisms, stated below.
+The formal version of this result is given in Theorem 5.4 and applies to $(\epsilon_{i},\delta_{i})$-DP continual mechanisms with possibly different $(\epsilon_{i},\delta_{i})$ parameters.
 Theorem 1.5 (concurrent parallel composition of approx DP continual mechanisms, informally stated).
-For every $\epsilon_{0}>0$, $0\leq\delta^{\prime}\leq 1$, and $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},\delta_{i})$-DP continual mechanisms, where the parameters $\delta_{i}$ are chosen adaptively such that $\sum_{i}\delta_{i}\leq\delta^{\prime}$, satisfies $(\epsilon,\delta+\delta^{\prime})$-DP if the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms is $(\epsilon,\delta)$-DP.
-We note that the formal version of this result applies to $(\epsilon_{i},\delta_{i})$-DP continual mechanisms with possibly different $(\epsilon_{i},\delta_{i})$ parameters.
+For every $\epsilon_{0}>0$, $0\leq\delta^{\prime}\leq 1$, and $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},\delta_{i})$-DP continual mechanisms, where the parameters $\delta_{i}$ are chosen adaptively such that $\prod_{i}(1-\delta_{i})\geq 1-\delta^{\prime}$, satisfies $(\epsilon,1-(1-\delta)(1-\delta^{\prime}))$-DP if the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms is $(\epsilon,\delta)$-DP.
 The proof of Theorem 1.5 relies on a key property of our new post-processing construction (recall that $\mathrm{RR}_{\epsilon,\delta}(b)$ denotes the randomized response mechanism with parameters $\epsilon$ and $\delta$ and input $b\in\{0,1\}$):
-Lemma 1.6 (post-processing of randomized response mechanisms, informal statement).
+Proposition 1.6 (post-processing of randomized response mechanisms, informal statement).
 For every $(\epsilon,\delta)$-DP interactive mechanism $\mathcal{M}$ and every pair of neighboring datasets $x_{0},x_{1}$, there exists an interactive post-processing mechanism $\mathcal{P}$ that receives the outputs of $\mathrm{RR}_{\epsilon,0}(b)$ and $\mathrm{RR}_{0,\delta}(b)$ and simulates $\mathcal{M}(x_{b})$ identically.
 As long as the answer distributions of $\mathcal{M}(x_{0})$ and $\mathcal{M}(x_{1})$ to the queries asked so far are identical, $\mathcal{P}$ does not access the outcome of $\mathrm{RR}_{\epsilon,0}(b)$.
-Setting $\delta^{\prime}=0$, Theorem 1.5 implies:
+The first sentence of this proposition was proven by Lyu (lyu2022composition, ) and Vadhan and Zhang (vadhan2022concurrent, ) with finiteness condition on the communication.
+The novelty in our result is that $\mathcal{P}$ does not access the outcome of $\mathrm{RR}_{\epsilon,0}(b)$ until $\mathcal{M}(x_{0})$ and $\mathcal{M}(x_{1})$ differ.
+Our construction builds on Lyu’s proof, where an IPM is initially given the outcome of $\mathrm{RR}_{\epsilon,\delta}(b)$.
+Let $\mathcal{M}^{\prime}$ be a continual mechanism that is $(\epsilon,\delta)$-DP with respect to a verification function $f$.
+Consider the interaction between $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}^{\prime}$ and an adversary.
+As long as the adversary sends pairs of identical messages, the answer distributions of $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}^{\prime}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M}^{\prime}$ will be identical.
+Thus, by replacing $\mathcal{M}(x_{b})$ in Proposition 1.6 with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}^{\prime}$, we conclude that there exists an IPM $\mathcal{P}$ that simulates $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}^{\prime}$ using only the output of $\mathrm{RR}_{0,\delta}(b)$ until the adversary issues a pair of non-identical messages.
+At that point, $\mathcal{P}$ requires the output of $\mathrm{RR}_{\epsilon,0}(b)$ to continue its simulation.
+This result is the core idea to show that in the concurrent $k$-sparse parallel composition of continual mechanisms, the $\epsilon$ parameters of the (potentially unboundedly many) composed mechanisms that are always assigned pairs of identical messages do not contribute to the overall privacy loss.
+In the concurrent parallel composition of interactive mechanisms, the adversary sends to each mechanism a pair of identical or neighboring datasets as first message and identical query pairs as subsequent messages.
+The key idea in the proof of Theorem 1.3 is that we can determine whether a mechanism is one of the $k$ exceptional ones solely from its first pair, before the mechanism produces any output.
+Recall that all messages of the adversary are filtered by a secret bit $b\in\{0,1\}$.
+If a mechanism is non-exceptional, whether the bit $b$ is zero or one makes no difference, and the first element of every pair can always be forwarded.
+In this proof, we simulate a mechanism by post-processing of a randomized-response mechanism with input $b$ only if it is exceptional, and show that we incur no privacy cost—neither in $\epsilon$ nor in $\delta$—from the non-exceptional mechanisms.
+Setting $\delta^{\prime}=0$ in Theorem 1.5 yields the following corollary, stated formally (and for differing $\epsilon_{i}$) in Corollary 1.7.
 Corollary 1.7 (concurrent parallel composition of pure DP continual mechanisms, informally stated).
-For every $\epsilon_{0}>0$ and integer $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},0)$-DP mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms.
-Like Theorem 1.5, we also prove a generalization of Corollary 1.7 for the case of composing mechanisms $\mathcal{M}_{i}$ with differing $\epsilon_{i}$ parameters.
+For every $\epsilon_{0}>0$ and integer $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},0)$-DP continual mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms.
+[^9] In some composition theorems (e.g., advanced composition), composing pure DP mechanisms yields $(\epsilon,\delta)$-DP with $\delta>0$.
 Concurrent filter composition theorems for continual mechanisms.
@@
 Let $\mathrm{Filt}:\left([0,\infty)\times[0,1]\right)^{*}\to\{\top,\bot\}$ be a filter.
-The concurrent $\mathrm{Filt}$-filter composition of continual mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantee as the $\mathrm{Filt}$-filter composition of non-interactive mechanisms.
+The concurrent $\mathrm{Filt}$-filter composition of continual mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantees as the $\mathrm{Filt}$-filter composition of noninteractive mechanisms.
+This result is also shown for Rényi DP in Corollary 13.8 and for $f$-DP, under the finite-communication assumption, in Corollary 14.10.
 General Conditions for Privacy Analysis of Continual Mechanisms.
@@
 One is to bound the accumulated privacy loss when DP systems are used repeatedly over time.
-Another is to facilitate the design of complex DP algorithms from simpler ones.
-The concurrent composition theorems for interactive mechanisms have already found an application in the privacy analysis of matching algorithms in local edge differential privacy model [4].
-(This work also has a result in the continual observation model, but the privacy is only proven for oblivious adversaries, which allows it to be proven without using a composition theorem for continual mechanisms.) Our concurrent composition theorem allows for a simpler and more modular privacy analysis of continual mechanisms built out of other continual mechanisms.
-There are various applications, where a continual mechanism executes multiple differentially private mechanisms, such as the Sparse Vector Technique (SVT) and continual counters, as sub-mechanisms and concurrently interacts with them.
+Another is to facilitate the design of complex DP algorithms from simpler ones, for example in modular software designs.
+The concurrent composition theorems for interactive mechanisms have already found an application in the privacy analysis of matching algorithms in local edge differential privacy model (DinitzLiLiZh25, ).
+(That work also has a result in the continual observation model, but the privacy is only proven for oblivious adversaries, which allows it to be proven without using a composition theorem for continual mechanisms.) Our concurrent composition theorem allows for a simpler and more modular privacy analysis of continual mechanisms built out of other continual mechanisms.
+There are various applications where a continual mechanism executes multiple differentially private mechanisms, such as the Sparse Vector Technique (SVT) and continual counters, as sub-mechanisms and concurrently interacts with them.
 Let $\mathcal{M}$ denote such a composite mechanism.
 Upon receiving an input, $\mathcal{M}$ may create new sub-mechanisms, interact multiple times with the existing ones, and then return an output.
-$\mathcal{M}$ generates inputs for its sub-mechanisms based on its own inputs and the previous outputs of all sub-mechanisms.(See Figure 4 for an illustration.) Analyzing the privacy of such mechanism designs is challenging because the creation and inputs of sub-mechanisms are interdependent and adaptive.
+$\mathcal{M}$ generates inputs for its sub-mechanisms based on its own inputs and the previous outputs of all sub-mechanisms.
+(See Figure 3 for an illustration.) Analyzing the privacy of such mechanism designs is challenging because the creation and inputs of sub-mechanisms are interdependent and adaptive.
 Even if the inputs of $\mathcal{M}$ are non-adaptive, the inputs to its sub-mechanisms may become adaptive: Let $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ denote the sub-mechanisms of $\mathcal{M}$.
@@
 In that design, the inputs of the counters depend on the output of the (active instance of the) SVT and vice versa.
+[…]
+Figure 3. Illustration of the interactions between an honest adversary $\mathcal{A}$ and a mechanism with $k$ sub-mechanisms
+Illustration of the interactions between an honest adversary $\mathcal${A}\$ and a mechanism with \$k\$ sub-mechanisms
 In such modular designs, we model $\mathcal{M}$ as the post-processing of a specific concurrent composition of its (continual) sub-mechanisms.
@@
 We therefore reformulate our conditions in this setting.
-We call a variable a decision variable if it is either the output of a private sub-mechanism or computed solely based on other decision variables.
+We call a variable an untainted variable if it is either the output of a private sub-mechanism or computed solely based on other untainted variables.
 Our conditions are now as follows:
-  • Destination condition: Decisions about whether to produce an output, instantiate a new sub-mechanism (and of which type), or interact with an existing one (and which one) must be based only on decision variables.
-  • Response condition: The value of $\mathcal{M}$’s output must be determined solely by decision variables.
-  • Mapping condition: Fixing the values of decision variables, neighboring inputs to $\mathcal{M}$ must be mapped to input sequences for the sub-mechanisms with a structure for which concurrent composition privacy is private (This condition is referred to as $f\to f^{\prime}$ property in Section 6).
+  • Destination condition: Decisions about whether to produce an output, instantiate a new sub-mechanism (and of which type), or interact with an existing one (and which one) must be based only on untainted variables.
+  • Response condition: The values of $\mathcal{M}$’s outputs must be determined solely by untainted variables.
+  • Mapping condition: Fixing the values of untainted variables, neighboring inputs to $\mathcal{M}$ must be mapped to input sequences for the sub-mechanisms with a structure for which concurrent composition is private (This condition is referred to as the $f\to f^{\prime}$ property in Section 6.)
 We show that if these conditions are satisfied, then the privacy guarantee of $\mathcal{M}$ is the same as that of the underlying concurrent composition.
@@
 The same conclusion holds for the concurrent $k$-sparse parallel composition of sub-mechanisms, as established in Theorem 1.3 and Corollary 1.7.
-[…]
-Figure 4. Illustration of the interactions between an honest adversary $\mathcal{A}$ and a mechanism with $k$ sub-mechanisms
 An Algorithmic Application.
-We give a new privacy proof for the continual monotone histogram mechanism of Henzinger, Sricharan, and Steiner (HSS) [13].
+We give a new privacy proof for the continual monotone histogram mechanism of Henzinger, Sricharan, and Steiner (HSS) (henzingerSS2023, ).
 This mechanism involves concurrent executions of a simpler continual histogram mechanism and multiple instantiations of the Sparse Vector Technique mechanism, and the original privacy analysis was rather intricate even for oblivious adversaries (because the mechanism itself is adaptive in its queries and dataset updates to the building-block mechanisms).
@@
 Moreover, for the case of approximate DP, the original proof was only for oblivious adversaries, whereas our proof also applies to adaptive adversaries.
-Our Variants of Concurrent Composition.
-Here we elaborate on the different types of concurrent composition that we study in this paper, with more general and more precise formulations than described above.
-(1) Concurrent Composition of Continual Mechanisms: In the concurrent composition of fixed continual mechanism, $k$ continual mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ are fixed at the beginning, and an adversary adaptively issues pairs of messages for them such that for each $\mathcal{M}_{i}$, the sequences formed by the first messages of all message pairs and the sequence formed by the second messages of all message pairs are neighboring.
-A secret bit $b\in\{0,1\}$ determines whether each mechanism receives the first or second sequence from each pair.
-The adversary’s goal is to guess $b$ using the answers of all $k$ mechanisms.
-Theorem 4.9, which is the formal statement of Theorem 1.2 above, analyzes the privacy of this composition.
-A natural extension is the concurrent composition of continual mechanism with fixed parameters, where instead of fixing the mechanisms themselves, $k$ privacy parameter pairs $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed in advance.
-The adversary then adaptively selects mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ over time, ensuring that, for each $1\leq i\leq k$, the mechanism $\mathcal{M}_{i}$ is $(\epsilon_{i},\delta_{i})$-DP.
-The privacy guarantee for this variant is the same as Theorem 4.9 and is stated in Theorem 4.11.
-(2) Concurrent $k$-Sparse Parallel Composition of Continual Mechanisms: This composition generalizes the concurrent composition of continual mechanisms with $k$ fixed parameters by permitting the creation of an arbitrary number of continual mechanisms.
-As before, $k$ privacy parameters $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed at the beginning.
-The adversary adaptively creates a potentially unlimited number of continual mechanisms and issues pairs of messages to each mechanism such that for all but at most $k$ of these mechanisms, each message pair consists of identical messages.
-Let $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ denote the exceptional mechanisms, which may receive differing message sequences.
-For each $1\leq i\leq k$, the mechanism $\mathcal{M}_{i}$ must be $(\epsilon_{i},\delta_{i})$-DP, and the sequences of the first and second messages for $\mathcal{M}_{i}$ must be neighboring.
-We note the choice of mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ is made adaptively over time.
-Specifically, the adversary decides whether to issue a pair of differing messages for a mechanism based on the previous answers of all mechanisms, including the outputs of that mechanism.
-As in previous compositions, a secret bit $b$ determines whether each mechanism receives the first or second message from each pair, and the adversary seeks to guess $b$ based on the outputs.
-Let $\mathcal{M}_{j}^{\prime}$ denote the $j$-th created mechanism, and let $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ be its privacy parameter.
-Theorem 5.4, informally stated in Theorem 1.5, analyzes the privacy of the concurrent $k$-sparse parallel composition of continual mechanisms when $\sum\delta_{j}^{\prime}$ is upper bounded by a predetermined value.
-Corollary 5.5, informally stated in Corollary 1.7, sets this upper bound to 0 and provides privacy guarantees for the concurrent $k$-sparse parallel composition of purely differentially private continual mechanisms.
-(3) Concurrent $k$-Sparse Parallel Composition of Interactive Mechanisms: In this composition, the adversary creates an unbounded number of interactive mechanisms over time, instead of continual ones.
-As before, $k$ privacy parameter pairs $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed in advance.
-Each time the adversary creates an interactive mechanism, they also select a pair of neighboring datasets for it.
-Instead of message pairs, the adversary adaptively sends (single) queries to each mechanism.
-A secret bit $b$ determines whether the mechanism is run on the first or second dataset of its respective pair.
-The adversary then tries to infer $b$ from the responses of all mechanisms.
-Theorem 1.3 analyzes the privacy of this composition when all $\mathcal{M}_{i}$ are $(\epsilon_{0},\delta_{0})$-DP.
-The more general case for mechanisms with possibly different $(\epsilon_{i},\delta_{i})$ parameters is stated in Theorem 5.15.
+An Application in Local Differential Privacy (LDP)
+Consider an untrusted server that concurrently interacts with multiple users, each holding a private dataset, with the goal of computing a function on the users’ aggregated data.
+A local protocol specifies the interaction rules between the server and the users, where both the server and the users are modeled as interactive mechanisms.
+Differential privacy for a local protocol is defined in terms of the distribution of the server’s view, consisting of its internal randomness and the transcript of all exchanged messages.
+Specifically, a protocol $P$ is said to satisfy $(\epsilon,\delta)$-LDP if the server’s view is $(\epsilon,\delta)$-indistinguishable for every pair of neighboring sequences of user datasets.
+A useful notion of neighboring in local differential privacy (LDP) considers two assignments of datasets to users that are identical for all but at most $k$ users, where the exceptional users datasets are neighboring with respect to a given binary relation on user datasets.
+An example of this is a setting in which each user represents a hospital and each patient may have records in at most $k$ hospitals.
+Analyzing the privacy of local protocols in which the server interacts concurrently with users over multiple rounds can be challenging.
+We observe that such interactions can be modeled as a concurrent parallel composition of interactive mechanisms, and our new privacy results for this concurrent composition directly yield a simplified privacy analysis for local protocols:
+Theorem 1.9 (informal statement).
+Let $P$ be a local protocol with $n$ users, and let $\sim^{*}$ be a neighbor relation on (individual) user datasets.
+Define two sequences of $n$ datasets to be neighbors if they are identical for all but at most $k\in\mathbb{N}$ users, where each differing user’s datasets are neighbors with respect to $\sim^{*}$.
+Suppose the interactive mechanism corresponding to every user is $(\epsilon_{0},\delta_{0})$-DP with respect to $\sim^{*}$ against adaptive adversaries.
+If the composition of $k$ noninteractive $(\epsilon_{0},\delta_{0})$-DP mechanisms is $(\epsilon,\delta)$-DP, then the protocol $P$ is $(\epsilon,\delta)$-LDP.
+In the formal version of Theorem 1.9 (see Corollary 8.8), we incorporate the fact that the server in $P$ restricts the queries sent to each user and enforces that they satisfy certain properties.
+Consequently, in Theorem 1.9, each user’s privacy guarantee only needs to hold for adversaries whose queries obey these properties.
+Furthermore, we provide a more general statement of Theorem 1.9 for an arbitrary choice of neighbor relation on sequences of user datasets in Theorem 8.7.
+The concurrent composition theorem used in the general result must be selected based on the given neighbor relation.
+We illustrate an application of Theorem 1.9 by simplifying the privacy analysis of the local protocol proposed by Dhulipala et al.
+(dhulipala2025near, ) for the core decomposition problem.
+In that work, a sequence of datasets $x=(x_{1},\dots,x_{n})$ represents a graph with $n$ vertices, where $x_{i}$ denotes the set of vertices adjacent to vertex $i$.
+(Users are vertices.) Differential privacy is defined with respect to edge-neighboring: two dataset sequences $x$ and $x^{\prime}$ are considered neighbors if their corresponding graphs differ by a single edge.
+The edge-neighboring relation can be generalized to a relation $\sim$, where $x\sim x^{\prime}$ if and only if the adjacency sets $x_{i}$ and $x_{i}^{\prime}$ are identical for all vertices $i$ but at most $2$ of them, and for those vertices, $x_{i}$ and $x_{i}^{\prime}$ are add-remove neighbors, meaning that they differ in the presence or absence of a single (adjacent) vertex.
+By defining $\sim^{*}$ in Theorem 1.9 to be the add-remove neighbor relation and setting $k=2$, we obtain a simpler privacy analysis for the protocol in (dhulipala2025near, ).
+We note that Theorem 1.9 and its generalizations are stated for protocols in which each user holds a static dataset.
+In many real-world settings, however, users may join or leave, and their data may evolve over time.
+In such cases, users must be modeled more generally as continual mechanisms rather than interactive mechanisms.
+The privacy guarantees of these protocols can still be computed using our concurrent composition theorems for continual mechanisms.
+We also note that in the above definition, all components (i.e., the server and the users) are assumed to be honest, meaning that they follow the protocol, and non-colluding, meaning that they do not share their internal states outside the protocol.
+In the presence of possible dishonesty or collusion, all users except the target users who hold the private data, together with the server, must be treated as a single adversary that interacts concurrently with the target users.
+In this case, privacy of the local protocol follows from the concurrent composition of the interactive and continual mechanisms, depending on whether the datasets evolve over time.
 The paper is organized as follows.
-Section 2 introduces all necessary notation, including non-interactive mechanisms, interactive mechanisms, and interactive post-processing mechanism.
+Section 2 introduces all necessary notation, including noninteractive mechanisms, interactive mechanisms, and interactive post-processing mechanism.
 Section 3 defines continual mechanisms, identifiers $\mathcal{I}$, verification functions $f$, verifiers $\mathcal{V}$ and differential privacy for continual mechanisms.
@@
 Section 5 introduces concurrent $k$-sparse parallel composition of continual mechanisms and interactive mechanisms, proves the theorems for this type of composition and shows the counterexample.
-In Section 6, we state the conditions and the theorem for analyzing the privacy of complex continual mechanisms that use non-interactive, interactive and continual sub-mechanisms.
-In Section 7, we apply our theorem to analyze the privacy of the continual histogram mechanism of [13].
-Finally, in Section 8 and Section 9, we prove two technical intermediate lemmas, required for the previous results.
+In Section 6, we state the conditions and the theorem for analyzing the privacy of complex continual mechanisms that use noninteractive, interactive and continual sub-mechanisms.
+In Section 7, we apply our theorem to analyze the privacy of the continual histogram mechanism of (henzingerSS2023, ).
+In Section 8, we show how local differential privacy can be analyzed using our concurrent composition theorems.
+In Section 9, we simplify the privacy analysis of the core decomposition local protocol by (dhulipala2025near, ).
+In Section 10 and Section 11, we prove two technical intermediate lemmas, required for the previous results.
+In Section 12, we introduce concurrent filter composition of continual mechanisms and reduce it to the filter composition for noninteractive mechanisms.
+Finally, in Section 13 and Section 14, we provide analogous results for Rényi DP and $f$-DP, respectively.
 2. Preliminaries
@@
 To define differential privacy, we require the definition of indistinguishability for random variables:
-Definition 2.1 ($(\epsilon,\delta)$-indistinguishability [17]).
-For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, two random variables $X_{1}$ and $X_{2}$ over the same domain $\mathcal{Y}$ are $(\epsilon,\delta)$-indistinguishable if for any subset $Y\subseteq\mathcal{Y}$,
+Definition 2.1 ($(\epsilon,\delta)$-indistinguishability (kasiviswanathan2014semantics, )).
+For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, two random variables $X_{1}$ and $X_{2}$ over the same domain $\mathcal{Y}$ are $(\epsilon,\delta)$-indistinguishable if for every measurable subset $Y\subseteq\mathcal{Y}$,
 $\Pr[X_{1}\in Y]\leq e^{\epsilon}\cdot\Pr[X_{2}\in Y]+\delta,$
@@
 Differential privacy for NIMs is defined as follows:
-Definition 2.2 ($(\epsilon,\delta)$-DP for NIMs [6]).
+Definition 2.2 ($(\epsilon,\delta)$-DP for NIMs (dwork2006calibrating, )).
 For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, a randomized mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-differentially private (or $(\epsilon,\delta)$-DP for short) with respect to (w.r.t.) a neighbor relation $\sim$ if for every two datasets $x_{0},x_{1}\in\mathcal{X}$ satisfying $x_{0}\sim x_{1}$, the random variables $\mathcal{M}(x_{0})$ and $\mathcal{M}(x_{1})$ are $(\epsilon,\delta)$-indistinguishable.
@@
 $(1-\delta)\frac{1}{1+e^{\epsilon}}$}\end{cases}$
-Lemma 2.4 ([16]).
+Lemma 2.4 ((KairouzOV15, )).
 For every $\epsilon\geq 0$ and $0\leq\delta\leq 1$, the approx randomized response $\mathrm{RR}_{\epsilon,\delta}$ is $(\epsilon,\delta)$-DP.
-The reason [16] introduced $\mathrm{RR}_{\epsilon,\delta}$ is that it is ”universal” among $(\epsilon,\delta)$-DP NIMs in the following sense:
-Theorem 2.5 ([16]).
+The reason (KairouzOV15, ) introduced $\mathrm{RR}_{\epsilon,\delta}$ is that it is ”universal” among $(\epsilon,\delta)$-DP NIMs in the following sense:
+Theorem 2.5 ((KairouzOV15, )).
 A mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-DP w.r.t.
-a neighbor relation $\sim$ if and only if for any $x_{0}\sim x_{1}$, there exists a randomized ”post-processor” $\mathcal{P}:\mathcal{Y}\to\{0,1\}\times\{\top,\bot\}$ such that for both $b\in\{0,1\}$, $\mathcal{M}(x_{b})$ is identically distributed to $\mathcal{P}(\mathrm{RR}_{\epsilon,\delta}(b))$.
+a neighbor relation $\sim$ if and only if for any $x_{0}\sim x_{1}$, there exists a randomized “post-processor” $\mathcal{P}:\mathcal{Y}\to\{0,1\}\times\{\top,\bot\}$ such that for both $b\in\{0,1\}$, $\mathcal{M}(x_{b})$ is identically distributed to $\mathcal{P}(\mathrm{RR}_{\epsilon,\delta}(b))$.
 2.2. Interactive Mechanisms (IM)
@@
 In the differential privacy literature, a query typically refers to a question about the dataset held by a mechanism.
-However, in this paper, we use query as a shorthand for query message, meaning an input message to a mechanism that requires a response.
+However, in this paper, we use query as a shorthand for query message, meaning any input message to a mechanism that requires a response.
 The state of an IM $\mathcal{M}$ serves as its memory.
@@
 Ghazi et al.
-[10] define this mechanism to initially flip a coin and determine whether to answer all subsequent queries truthfully or in a privacy-preserving manner.
-Our definition of interactive randomized response mechanism is different but shares the same intuition.
+(ghazi2025privacy, ) define this mechanism to initially flip a coin and determine whether to answer all subsequent queries truthfully or in a privacy-preserving manner.
+Our definition of the interactive randomized response mechanism is different but shares the same intuition.
 Recall that the standard randomized response mechanism $\mathrm{RR}_{\epsilon,\delta}$ is an NIM that outputs a pair from the space $\{\top,\bot\}\times\{0,1\}$.
@@
 Let $s$ and $s^{\prime}$ be initial states for $\mathcal{M}$ and $\mathcal{M}^{\prime}$, respectively.
-The mechanisms $\mathcal{M}(s)$ and $\mathcal{M}^{\prime}(s^{\prime})$ are equivalent if for any adversary $\mathcal{A}$ with $A_{\mathcal{A}}=Q_{\mathcal{M}}$ and $Q_{\mathcal{A}}=A_{\mathcal{M}}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}^{\prime}(s^{\prime}))$ are identically distributed.
+The mechanisms $\mathcal{M}(s)$ and $\mathcal{M}^{\prime}(s^{\prime})$ are equivalent if for every adversary $\mathcal{A}$ with $A_{\mathcal{A}}=Q_{\mathcal{M}}$ and $Q_{\mathcal{A}}=A_{\mathcal{M}}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}^{\prime}(s^{\prime}))$ are identically distributed.
 2.2.2. Differential Privacy for IMs.
@@
 Moreover, since adversaries are stateful algorithms, they can choose their messages adaptively based on the history of messages exchanged, meaning that all adversaries in this paper are adaptive.
-Existing works [23, 24, 19, 11] view the initial state of $\mathcal{M}$ is a dataset, i.e., $S_{\mathcal{M}}^{\text{init}}=\mathcal{X}$, and (as is usual) they define neighbor relations as binary relations on $\mathcal{X}$.
+Existing works (vadhan2021concurrent, ; vadhan2022concurrent, ; lyu2022composition, ; haney2023concurrent, ) view the initial state of $\mathcal{M}$ as a dataset, i.e., $S_{\mathcal{M}}^{\text{init}}=\mathcal{X}$, and (as is usual) they define neighbor relations as binary relations on $\mathcal{X}$.
 However, for us it will be convenient to conceptually separate $S_{\mathcal{M}}^{\text{init}}$ from the data (which is changing continuously over time), but still define DP w.r.t.
@@
 Differential privacy for NIMs is known to be preserved under post-processing.
-Haney et al. [11] generalize post-processing functions to interactive post-processing mechanisms (IPMs) as follows and show that interactive post-processing preserves privacy guarantees for IMs.
-Definition 2.14 (Interactive Post-Processing Mechanism (IPM) [11]).
+Haney et al.
+(haney2023concurrent, ) generalize post-processing functions to interactive post-processing mechanisms (IPMs) as follows and show that interactive post-processing preserves privacy guarantees for IMs.
+Definition 2.14 (Interactive Post-Processing Mechanism (IPM) (haney2023concurrent, )).
 An interactive post-processing mechanism (IPM) is a stateful randomized algorithm $\mathcal{P}:S_{\mathcal{P}}\times\left((\{{\texttt{L}}\}\times Q_{\mathcal{P}}^{L})\cup(\{{\texttt{R}}\}\times Q_{\mathcal{P}}^{R})\right)\rightarrow S_{\mathcal{P}}\times\left((\{{\texttt{L}}\}\times A_{\mathcal{P}}^{L})\cup(\{{\texttt{R}}\}\times A_{\mathcal{P}}^{R})\right)$ that stands between two interactive mechanisms, referred to as the left and right mechanisms, communicating with them via message passing.
@@
 The formal definition of $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ is given below:
-Definition 2.15 (Post-Processing of an IM $\mathcal{M}$ by an IPM $\mathcal{P}$ ($\mathcal{P}\!\circ^{*}\!\mathcal{M}$)[11]).
+Definition 2.15 (Post-Processing of an IM $\mathcal{M}$ by an IPM $\mathcal{P}$ ($\mathcal{P}\!\circ^{*}\!\mathcal{M}$)(haney2023concurrent, )).
 Let $\mathcal{M}$ be an IM and $\mathcal{P}$ an IPM such that $Q_{\mathcal{P}}^{R}=A_{\mathcal{M}}$ and $A_{\mathcal{P}}^{R}=Q_{\mathcal{M}}$.
@@
 For readability, when $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ is initialized with the state $(s^{\mathcal{M}},s^{\mathcal{P}})\in S_{\mathcal{P}\!\circ^{*}\!\mathcal{M}}^{\text{init}}$, we denote it as $\mathcal{P}(s^{\mathcal{P}})\!\circ^{*}\!\mathcal{M}(s^{\mathcal{M}})$ instead of $\left(\mathcal{P}\!\circ^{*}\!\mathcal{M}\right)\left((s^{\mathcal{M}},s^{\mathcal{P}})\right)$.
-Algorithm 1 Post-Processing of $\mathcal{M}$ by $\mathcal{P}$, denoted by $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ [11].
+Algorithm 1 Post-Processing of $\mathcal{M}$ by $\mathcal{P}$, denoted by $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ (haney2023concurrent, ).
 procedure $\left(\mathcal{P}\!\circ^{*}\!\mathcal{M}\right)$($s,m$):
@@
 Haney et al.
-[11] shows that if an IM $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private, then for any IPM $\mathcal{P}$ with a singleton initial state space, $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ is also $(\epsilon,\delta)$-differentially private.
+(haney2023concurrent, ) shows that if an IM $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private, then for any IPM $\mathcal{P}$ with a singleton initial state space, $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ is also $(\epsilon,\delta)$-differentially private.
 This result is formally stated below.
-Lemma 2.16 (Post-Processing Lemma for IMs [11]).
-Let $\mathcal{M}$ be an IM $\mathcal{M}$ with an initial state space $S_{\mathcal{M}}^{\text{init}}$ and $\mathcal{P}$ be an IPM with a singleton initial state space $S_{\mathcal{P}}^{\text{init}}=\{s\}$.
+Lemma 2.16 (Post-Processing Lemma for IMs (haney2023concurrent, )).
+Let $\mathcal{M}$ be an IM with an initial state space $S_{\mathcal{M}}^{\text{init}}$ and $\mathcal{P}$ be an IPM with a singleton initial state space $S_{\mathcal{P}}^{\text{init}}=\{s\}$.
 For every pair of initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$, the following holds: If for every adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s_{0}))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}(s_{1}))$ are $(\epsilon,\delta)$-indistinguishable, then for every adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{P}\!\circ^{*}\!\mathcal{M}(s_{0}))$ and $\mathrm{View}(\mathcal{A},\mathcal{P}\!\circ^{*}\!\mathcal{M}(s_{1}))$ are $(\epsilon,\delta)$-indistinguishable.
-[^5] In [11], the definitions of IMs and IPMs are more restrictive.
-Specifically, for any IM $\mathcal{M}$, $S_{\mathcal{M}}=Q_{\mathcal{M}}=A_{\mathcal{M}}=\{0,1\}^{*}$, and for any IPM $\mathcal{P}$, $S_{\mathcal{P}}=M_{\mathcal{P}}=\{0,1\}^{*}$; However, their results do not rely on these restrictions, and the interactive post-processing lemma holds for arbitrary sets $S_{\mathcal{M}}$, $Q_{\mathcal{M}}$, $A_{\mathcal{M}}$, $S_{\mathcal{P}}$, and $M_{\mathcal{P}}$.
+[^10] In (haney2023concurrent, ), the definitions of IMs and IPMs are more restrictive.
+Specifically, for any IM $\mathcal{M}$, $S_{\mathcal{M}}=Q_{\mathcal{M}}=A_{\mathcal{M}}=\{0,1\}^{*}$, and for any IPM $\mathcal{P}$, $S_{\mathcal{P}}=M_{\mathcal{P}}=\{0,1\}^{*}$; however, their results do not rely on these restrictions, and the interactive post-processing lemma holds for arbitrary sets $S_{\mathcal{M}}$, $Q_{\mathcal{M}}$, $A_{\mathcal{M}}$, $S_{\mathcal{P}}$, and $M_{\mathcal{P}}$, which are not necessarily countable.
 IPMs are not limited to interacting with IMs. Two IPMs can interact with each other as well, as follows:
@@
 This section is based on two crucial observations: (i) A message, as sent or received by IMs and IPMs, is a general and powerful concept capable of encoding various types of instructions, and (ii) the essence of DP is to protect the privacy of a single bit that distinguishes neighboring states, sequences, etc.
-The second observation has been used many times in the DP literature and was specifically used in [14] to model adaptivity in continual observation.
+The second observation has been used many times in the DP literature and was specifically used in (pricedpco, ) to model adaptivity in continual observation.
 Using these key observations, we first reformulate the definition of DP for IMs and continual observation.
@@
 By Definition 2.13, an interactive mechanism $\mathcal{M}$ with initial state space $S_{\mathcal{M}}^{\text{init}}$ is $(\epsilon,\delta)$-DP w.r.t.
-a neighboring relation $\sim$, if for every two initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$ such that $s_{0}\sim s_{1}$ and every adversary $\mathcal{A}$, the view of $\mathcal{A}$ in the interaction with $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ is $(\epsilon,\delta)$-indistinguishable.
+a neighboring relation $\sim$, if for every two initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$ such that $s_{0}\sim s_{1}$ and every adversary $\mathcal{A}$, the views of $\mathcal{A}$ in the interaction with $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ are $(\epsilon,\delta)$-indistinguishable.
 This means that for the worst-case (i.e., adversarial) choice of neighboring initial states, an adversary by adaptively asking queries should not be able to determine whether $\mathcal{M}$ is running on $s_{0}$ or $s_{1}$.
@@
 We also define an IPM $\mathcal{I}$ with the initial state space $S_{\mathcal{I}}^{\text{init}}=\{0,1\}$ to stand between the adversary and $\mathcal{M}$, filtering the pairs based on its, i.e., $\mathcal{I}$’s, initial state.
-Specifically, for $b\in\{0,1\}$, in an interaction between an adversary $\mathcal{A}$ and $\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$, upon receiving a message $(m_{0},m_{1})$ from $\mathcal{A}$, $\mathcal{I}(b)$ forwards $m_{b}$ to $\mathcal{M}$ and returns $\mathcal{M}$’s answer unchanged to $\mathcal{A}$.
-With this formulation, an interactive mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP if for every adversary of the above form, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}[0]\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}[1]\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
-Actually, we will enforce the restrictions on the adversary also by an IPM, which we call a verifier $\mathcal{V}$.
-The verifier checks whether the initial message is a pair of neighboring datasets and the subsequent messages of the adversary are of the form $(q,q)$; only if these checks pass, are the messages forwarded to $\mathcal{I}$.
+Specifically, for $b\in\{0,1\}$, in the interaction between an adversary $\mathcal{A}$ and $\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$, upon receiving a message $(m_{0},m_{1})$ from $\mathcal{A}$, $\mathcal{I}(b)$ forwards $m_{b}$ to $\mathcal{M}$ and returns $\mathcal{M}$’s answer unchanged to $\mathcal{A}$.
+With this formulation, an interactive mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP if for every adversary of the above form, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
+We will enforce the restrictions on the adversary by an IPM, which we call a verifier $\mathcal{V}$.
+The verifier checks whether the initial message is a pair of neighboring datasets and the subsequent messages of the adversary are of the form $(q,q)$; only if these checks pass are the messages forwarded to $\mathcal{I}$.
 3.2. An Alternative Definition of DP under Continual Observation.
-Dwork, Naor, Pitassi, and Rothblum [7] introduced differential privacy under continual observation, where the dataset is not given at the initialization time and instead evolves over time.
+Dwork, Naor, Pitassi, and Rothblum (Dwork2010, ) introduced differential privacy under continual observation, where the dataset is not given at the initialization time and instead evolves over time.
 In this model, the mechanism receives either an empty record or a data record at each step.
@@
 Then, whether the input record is empty or not, the mechanism answers a predetermined question about its current dataset.
-To capture ”event-level privacy” (where the privacy unit is a single record), Dwork et al.
-[7] define neighboring as a binary relation on sequences of data and empty records: two sequences are neighboring if they are identical except for the presence or absence of a record in a single step.
+To capture “event-level privacy” (where the privacy unit is a single record), Dwork et al.
+(Dwork2010, ) define neighboring as a binary relation on sequences of data and empty records: two sequences are neighboring if they are identical except for the presence or absence of a record in at most one step.
 They called a mechanism $(\epsilon,\delta)$-DP under continual observation if the distributions of answer sequences are $(\epsilon,\delta)$-indistinguishable for every two neighboring input sequences.
 This definition does not allow for data sequences that are adaptively chosen based on the answers released.
-Jain, Raskhodnikova, Sivakumar, and Smith [14] formalized the definition of DP against adaptive adversaries in the continual observation model as follows: An adaptive adversary $\mathcal{A}$ chooses a single (data or empty) record at every step, except for one challenge step, selected by the adversary, where $\mathcal{A}$ sends a pair of records.
-They called a mechanism $\mathcal{M}$ $(\epsilon,\delta)$-differentially private against adaptive adversaries in the continual observation model if for any such adaptive adversary $\mathcal{A}$, the distributions of $\mathcal{M}$’s answer sequences are $(\epsilon,\delta)$-indistinguishable when the first or the second element of the pair in the challenge step is handed to $\mathcal{M}$.
+Jain, Raskhodnikova, Sivakumar, and Smith (pricedpco, ) formalized the definition of DP against adaptive adversaries in the continual observation model as follows: An adaptive adversary $\mathcal{A}$ chooses a single (data or empty) record at every step, except for one challenge step, selected by the adversary, where $\mathcal{A}$ sends a pair of records.
+They called a mechanism $\mathcal{M}$ $(\epsilon,\delta)$-differentially private against adaptive adversaries in the continual observation model if for any such adaptive adversary $\mathcal{A}$, the views of $\mathcal{A}$—consisting of $\mathcal{A}$’s internal randomness and the transcript of messages sent and received by $\mathcal{A}$—are $(\epsilon,\delta)$-indistinguishable when the first or the second element of the pair in the challenge step is handed to $\mathcal{M}$.
+Denisov, McMahan, Rush, Smith and Thakurta (denisov2022improved, ) generalize this definition by allowing the adversary to send a pair of (data or empty) records at every step such that the sequence of first records and the sequence of second records are neighbors under an arbitrary binary relation $\sim$ on sequences of data and empty records.
+A mechanism $\mathcal{M}$ is said to satisfy $(\epsilon,\delta)$-DP against an adaptive adversary if, for every adversary $\mathcal{A}_{\sim}$ that adaptively sends pairs of records forming $\sim$-neighbor sequences, the adversary’s views are $(\epsilon,\delta)$-indistinguishable when $\mathcal{M}$ receives the first elements of each pair and when it receives the second elements.
 A mechanism $\mathcal{M}$ in the continual observation model can be viewed as an IM with a single possible initial state, representing an empty dataset.
 The input message set of $\mathcal{M}$ consists of the empty record and all possible data records.
-Instead of having a challenge step, we can alternatively restrict to adversaries that send pairs of identical records at every step except for one.
-Under this formulation, the definition of differential privacy proposed by [14] can be restated as follows: A mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP differential privacy against adaptive adversaries if for any such adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}[0]\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}[1]\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable, where $\mathcal{I}$ is the interactive post-processing mechanism discussed previously.
+Under this formulation, the above definition of differential privacy can be restated as follows: For every adversary $\mathcal{A}_{\sim}$, the random variables $\mathrm{View}(\mathcal{A}_{\sim},\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A}_{\sim},\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable, where $\mathcal{I}$ is the interactive post-processing mechanism discussed previously.
 Again, below we will enforce the restriction on the adversary via another IPM $\mathcal{V}$.
-In this alternative definition, all adversary messages are pairs of identical data or empty records, except for a single message that is a pair of different records.
-Recall that in the alternative definition of DP for IMs, the first message sent by an adversary must be a pair of neighboring initial datasets, while all subsequent adversary messages must be pairs of identical queries.
+In this alternative definition, all adversary messages are pairs of data or empty records, and the sequences of first and second elements are neighbors.
+Recall that in the alternative definition of DP for IMs, the first adversary message must be a pair of neighboring initial datasets, while all subsequent adversary messages must be pairs of identical queries.
 Next, we generalize the definition of differential privacy by allowing arbitrary conditions on the sequence of pairs sent by adversaries, without restricting the content of the messages.
@@
 To the best of our knowledge, all mechanisms in DP literature can be represented as CMs:
-  • Every non-interactive mechanism $\mathcal{N}$, such as the Laplace mechanism, can be modeled as a CM that receives a dataset $x$ as its first message and returns $\mathcal{N}(x)$ as a response.
+  • Every non-interactive mechanism $\mathcal{N}$, such as the Laplace mechanism, can be modeled as a CM that receives a dataset $x$ as its first message and returns $\mathcal{N}(x)$ in response.
 Upon receiving any subsequent input message, the CM representing $\mathcal{N}$ halts the communication.
@@
 This CM then answers the next input messages the same as $\mathcal{M}(s)$.
-  • Mechanisms such as the Sparse Vector Technique (SVT) [8, 12, 18] and the continual counter [6, 2, 3, 5] in the streaming settings can also be represented as CMs receiving two types of messages: question queries and data-update queries.
+  • Mechanisms such as the Sparse Vector Technique (SVT) (dwork2009complexity, ; HardtRo10, ; lyu2016understanding, ) and the continual counter (dwork2006calibrating, ; DBLP:journals/tissec/ChanSS11, ; mcmahan2022private, ; dvijotham2024efficient, ) in the streaming settings can also be represented as CMs receiving two types of messages: question queries and data-update queries.
 Question queries request information about the current dataset, while data-update queries consist of instructions for modifying the dataset.
@@
 Comparing the sequences formed by the first and second elements of the pairs, $f$ can be used to encode a neighboring property.
-If the pair sequence satisfies the properties, $\mathcal{V}[f]$ forwards $\mathcal{A}$’s message to $\mathcal{I}\!\circ^{*}\!\mathcal{M}$; otherwise, it halts the communication.
+If the sequences satisfy the property, $\mathcal{V}[f]$ forwards $\mathcal{A}$’s message to $\mathcal{I}\!\circ^{*}\!\mathcal{M}$; otherwise, it halts the communication.
 The formal definitions of identifiers, verification functions, and verifiers are given below.
@@
 Since the set of all possible messages does not exist, we assume that all message spaces considered in this paper are subsets of some universal set $\mathcal{U}$.
-For instance, in [11], the universal set is taken to be $\{0,1\}^{*}$.
+For instance, in (haney2023concurrent, ), the universal set is taken to be $\{0,1\}^{*}$.
 The domain of verification functions is $\mathcal{U}^{*}$.
@@
 Since the initial state space of $\mathcal{V}[f]$ is a singleton, we denote an instance of $\mathcal{V}[f]$ with the initial state $()$ as $\mathcal{V}[f]$.
-$\mathcal{V}[f]$ stands between two IMs, referred to as the left and right mechanisms.
+The IPM $\mathcal{V}[f]$ stands between two IMs, referred to as the left and right mechanisms.
 The state of $\mathcal{V}[f]$ consists of the sequence of messages it has received from the left mechanism so far.
@@
 Thus, the initial state space of $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ is $\left\{((),0,s^{\text{init}}),((),1,s^{\text{init}})\right\}$.
-Define $\sim_{01}$ as a neighboring on this set, where only $((),0,s^{\text{init}})\sim_{01}((),1,s^{\text{init}})$.
+Define $\sim_{01}$ as the universal relation on this set, where every two elements are neighboring.
 Comparing Definition 3.7 and Definition 2.13 leads to the following corollary:
@@
 Concurrent Composition of IMs.
-Vadhan and Wang [23] extended the definition of the composition of NIMs to the concurrent composition of IMs, where an adversary interacts with $k$ interactive mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ at the same time.
+Vadhan and Wang (vadhan2021concurrent, ) extended the definition of the composition of NIMs to the concurrent composition of IMs, where an adversary interacts with $k$ interactive mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ at the same time.
 In concurrent composition, the adversary generates a message and determines the receiver of this message (one of the $\mathcal{M}_{i}$s) adaptively based on all the previous messages exchanged between the adversary and the mechanisms $M_{1},\dots,M_{k}$.
-To formalize this, Vadhan and Wang [23] introduced an interactive mechanism ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$, which internally runs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ and, upon receiving a message $(q,j)$, forwards the query $q$ to $\mathcal{M}_{j}$ and returns its response.
+To formalize this, Vadhan and Wang (vadhan2021concurrent, ) introduced an interactive mechanism ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$, which internally runs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ and, upon receiving a message $(q,j)$, forwards the query $q$ to $\mathcal{M}_{j}$ and returns its response.
 The initial state space of ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ is $S_{\mathcal{M}_{1}}^{\text{init}}\times\dots\times S_{\mathcal{M}_{k}}^{\text{init}}$, where $S_{\mathcal{M}_{i}}^{\text{init}}$ is the initial state space of $\mathcal{M}_{i}$.
 Let $\sim_{i}$ be a neighboring relation on each $S_{\mathcal{M}_{i}}^{\text{init}}$, and define the neighboring relation $\sim$ on $S_{\mathcal{M}_{1}}^{\text{init}}\times\dots\times S_{\mathcal{M}_{k}}^{\text{init}}$ as follows: for every two initial states $s=(s_{1},\dots,s_{k})$ and $s^{\prime}=(s_{1}^{\prime},\dots,s_{k}^{\prime})$, we have $s\sim s^{\prime}$ if and only if $s_{i}\sim_{i}s_{i}^{\prime}$ for all $i\in[k]$.
-Vadhan and Wang [23] show that if each IM $\mathcal{M}_{i}$ is $\epsilon_{i}$-DP w.r.t $\sim_{i}$, and if the composition of NIMs $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ is $(\epsilon,\delta)$-DP, then ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ is also $(\epsilon,\delta)$-DP against an adaptive adversary.
-Lyu [19] extends this result for $(\epsilon_{i},\delta_{i})$-DP IMs.
+Vadhan and Wang (vadhan2021concurrent, ) show that if each IM $\mathcal{M}_{i}$ is $\epsilon_{i}$-DP w.r.t $\sim_{i}$, and if the composition of NIMs $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ is $(\epsilon,\delta)$-DP, then ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ is also $(\epsilon,\delta)$-DP against an adaptive adversary.
+Lyu (lyu2022composition, ) and Vadhan and Zhang (vadhan2022concurrent, ) extend this result for $(\epsilon_{i},\delta_{i})$-DP IMs.
 In the concurrent composition of IMs, queries asked to each mechanism are chosen adaptively, not only based on that mechanism’s previous responses but also using the responses of the other mechanisms up to that point.
-The results of [23] and [19] show that this concurrency and additional adaptivity do not compromise privacy guarantees.
-As a result, all $(\epsilon,\delta)$-DP composition theorems for NIMs, such as the basic [6], advanced [9], and optimal [16] composition theorems, can also be applied to the concurrent composition of IMs.
+The results of (vadhan2022concurrent, ) and (lyu2022composition, ) show that this concurrency and additional adaptivity do not compromise privacy guarantees.
+As a result, all $(\epsilon,\delta)$-DP composition theorems for NIMs, such as the basic (dwork2006calibrating, ), advanced (dwork2010boosting, ), and optimal (KairouzOV15, ) composition theorems, can also be applied to the concurrent composition of IMs.
 Concurrent Composition of CMs.
@@
 In Section 4.1, we generalize ConComp to ExtConComp and give a formal definition for the concurrent composition of CMs.
-To extend Lyu’s result on the privacy guarantee of the concurrent composition of IMs to the concurrent composition of CMs, we want to use the following intermediate lemma from their paper: For every $(\epsilon,\delta)$-DP interactive mechanism $\mathcal{M}$ and every pair of neighboring initial states $s$ and $s^{\prime}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the IMs $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(b)$ are equivalent.
-However, as we discuss in Section 4.2, there are hidden assumptions in the proof of this lemma.
-Additionally, we show there that the construction of $\mathcal{P}$ in [23], for $\epsilon_{i}$-DP IMs, satisfies a property, required for a later proof.
+To extend the proofs of (lyu2022composition, ; vadhan2022concurrent, ) on the privacy guarantee of the concurrent composition of IMs to the concurrent composition of CMs, we want to use the following intermediate lemma from their paper: For every $(\epsilon,\delta)$-DP interactive mechanism $\mathcal{M}$ and every pair of neighboring initial states $s$ and $s^{\prime}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the IMs $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(b)$ are equivalent.
+However, as we discuss in Section 4.2, there are hidden assumptions in Lyu’s proof of this lemma.
 In Section 4.3, we analyze the privacy of the concurrent composition of $k$ fixed CMs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$.
@@
 4.1. Formalization.
-We extend the definition of ConComp in [23] and introduce ExtConComp that allows the creation of new mechanisms in addition to queries to existing ones.
+We extend the definition of ConComp in (vadhan2021concurrent, ) and introduce ExtConComp that allows the creation of new mechanisms in addition to queries to existing ones.
 A creation query $(\mathcal{M},s,\epsilon,\delta,f)$ is a message that requests the creation of a continual mechanism $\mathcal{M}$ with the initial state $s$ and asserts that $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t.
@@
 Both query set and answer set of ExtConComp are the universal message set $\mathcal{U}$ in Remark 3.4.
+Remark 4.2.
+By Russell’s paradox, a set of all continual mechanisms cannot exist.
+Therefore, to ensure that Definition 4.1 is mathematically correct and the query and answer sets of ExtConComp are well defined, we restrict creation queries to select mechanisms from a well-defined set and verification functions from a well-defined set.
+These sets may be infinite.
+The mechanism ExtConComp itself may belong to the mechanism set and, consequently, may be recursively created by itself.
 To define DP for the extended concurrent composition, we must impose constraints on the created mechanisms and their query pairs to limit the adversary’s power.
@@
 Later, when defining specific verification functions $f$ for different concurrent compositions, we set $f(m_{1},\dots,m_{k})=\bot$ whenever the sequence of messages $m_{1},\dots,m_{k}$ is not suitable.
+For convenience, in the rest of the paper, we represent a pair of identical creation queries $(\mathcal{M},s,\epsilon,\delta,f)$ by $(\mathcal{M},s,\epsilon,\delta,f)^{2}$.
 Definition 4.5 (Suitable Sequence of Messages).
@@
 4.2. $(\epsilon,\delta)$-DP IMs and $\mathrm{RR}_{\epsilon,\delta}$
-Lyu [19] shows that for every $(\epsilon,\delta)$-DP IM $\mathcal{M}$ and every pair of neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{T}$ (with a single initial state) such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(s_{b})$ are identical.
-Although not explicitly stated, the construction of $\mathcal{T}$ proposed in [19] requires the answer distributions of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to be discrete.
-This assumption is formally defined as:
+Lyu (lyu2022composition, ) and Vadhan and Zhang (vadhan2022concurrent, ) show that for every $(\epsilon,\delta)$-DP IM $\mathcal{M}$ and every pair of neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{T}$ (with a single initial state) such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(s_{b})$ are equivalent.
+Although not explicitly stated, the construction of $\mathcal{T}$ proposed in (lyu2022composition, ) requires the answer sets of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to be finite.
+Vadhan and Zhang (vadhan2022concurrent, ) make an even stronger assumption that the query set must be finite.
+We instead assume that the answer distribution is discrete, which is formally defined as follows:
 Definition 4.6.
-An IM $\mathcal{M}$ has discrete answer distributions if $A_{\mathcal{M}}$ is countable, and for every initial state $s$, integer $t\in\mathbb{N}$, history $(q_{1},a_{1},\dots,q_{t-1},a_{t-1})\in(Q_{\mathcal{M}}\times A_{\mathcal{M}})^{t-1}$, and new query $q_{t}\in Q_{\mathcal{M}}$, the distribution of $\mathcal{M}(s)$’s answer in respond to $q_{t}$ conditioned on having the history $(q_{1},\dots,a_{t-1})$ has a probability mass function (PMF).
+An IM $\mathcal{M}$ has discrete answer distributions if $A_{\mathcal{M}}$ is countable, and for every initial state $s$, integer $t\in\mathbb{N}$, history $(q_{1},a_{1},\dots,q_{t-1},a_{t-1})\in(Q_{\mathcal{M}}\times A_{\mathcal{M}})^{t-1}$, and new query $q_{t}\in Q_{\mathcal{M}}$, the distribution of $\mathcal{M}(s)$’s answer in response to $q_{t}$ conditioned on having the history $(q_{1},\dots,a_{t-1})$ has a probability mass function (PMF).
 This is generally not a restrictive assumption in differential privacy.
@@
 An example of this approach will be provided in Section 7.
-Lyu [19] additionally assumes that there exists a pre-fixed upper bound $c$ on the number of queries $\mathcal{M}$ answers.
+Lyu (lyu2022composition, ) additionally assumes that there exists a pre-fixed upper bound $c$ on the number of queries $\mathcal{M}$ answers.
 This assumption stems from their use of a backward-defined control function.
@@
 We have:
-Lemma 4.6 ([19]).
+Lemma 4.7 (slight extension of (lyu2022composition, ; vadhan2022concurrent, )).
 For $\epsilon>0$ and $0\leq\delta\leq 1$, let $\mathcal{M}$ be a CM that is $(\epsilon,\delta)$-DP w.r.t.
@@
 By definition, the choice of neighboring initial states for the interactive mechanisms is non-adaptive.
-However, in the concurrent composition of CMs, the choice of differences between query sequences, i.e., the neighboring query sequences is adaptive, based on not only the responses of the corresponding mechanism but also the responses of other mechanisms.
-(Note that, unlike IMs, CMs inherently have a single fixed initial state; hence, no choice of initial state is involved.) Despite this additional adaptivity, in the following theorem, we show that the privacy results of [19] for ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ also hold for $\mathcal{V}[f^{\alpha_{1},\dots,\alpha_{k}}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$.
+However, in the concurrent composition of CMs, the choice of differences between query sequences, i.e., the neighboring query sequences is adaptive, based not only on the responses of the corresponding mechanism but also on the responses of other mechanisms.
+(Note that, unlike IMs, CMs inherently have a single fixed initial state; hence, no choice of initial state is involved.) Despite this additional adaptivity, in the following theorem, we show that the privacy results of (lyu2022composition, ; vadhan2021concurrent, ) for ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ also hold for $\mathcal{V}[f^{\alpha_{1},\dots,\alpha_{k}}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$.
 Theorem 4.10.
@@
 A slight adaptation of the proof of Theorem 4.10 implies the following theorem.
-For convenience, in the rest of the paper, we represent a pair of identical creation queries $(\mathcal{M},s,\epsilon,\delta,f)$ by $(\mathcal{M},s,\epsilon,\delta,f)^{2}$.
 Theorem 4.12.
@@
 That is, the concurrent $k$-sparse parallel composition of purely differentially private CMs with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$ is $(\epsilon,\delta)$-DP against adaptive adversaries.
-In the proofs of the previous composition theorems (namely, Theorems 4.9 and 4.11), each created CM $\mathcal{M}_{j}^{\prime}$ with privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ is simulated by the IPM $\mathcal{T}_{j}$ from Lemma 4.6.
+In the proofs of the previous composition theorems (namely, Theorems 4.10 and 4.12), each created CM $\mathcal{M}_{j}^{\prime}$ with privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ is simulated by the IPM $\mathcal{T}_{j}$ from Lemma 4.7.
 For each $b\in\{0,1\}$, the mechanism $\mathcal{T}_{j}$ interacts with $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ once and simulates $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ identically.
@@
 The high-level idea behind the proof of Theorem 5.4 is to separately compose $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ and $\mathrm{RR}_{0,\delta_{1}^{\prime}},\mathrm{RR}_{0,\delta_{2}^{\prime}},\dots$, and then apply basic composition to these two sets of composed mechanisms to conclude the final result.
-Since the parameters $\delta_{j}^{\prime}$ are chosen adaptively, the composition of $\mathrm{RR}_{0,\delta_{1}^{\prime}},\mathrm{RR}_{0,\delta_{2}^{\prime}},\dots$ is an instance of the filter composition of NIMs [21].
+Since the parameters $\delta_{j}^{\prime}$ are chosen adaptively, the composition of $\mathrm{RR}_{0,\delta_{1}^{\prime}},\mathrm{RR}_{0,\delta_{2}^{\prime}},\dots$ is an instance of the filter composition of NIMs (rogers2016privacy, ).
 We do not state this composition in terms of filters and define it using a verification function $f^{\delta^{\prime}}_{\mathrm{RR}}$ that only allows the creation of randomized response mechanisms $\mathrm{RR}_{0,\delta_{j}^{\prime}}$ and returns $\bot$ if $1-\prod(1-\delta_{j}^{\prime})>\delta^{\prime}$.
@@
 The following lemma is proved in Section 11.
-[^6] By setting $\epsilon_{1}=\epsilon_{2}=\dots=0$ in Theorem 2 of [25], it follows that for every $\epsilon>0$, $\delta^{\prime}>0$, and $0\leq\delta^{\prime\prime}\leq 1$, the composition of randomized response mechanisms $\mathrm{RR}_{0,\delta_{1}},\mathrm{RR}_{0,\delta_{2}},\dots$ with adaptively chosen privacy parameters $\delta_{1},\delta_{2},\dots$ satisfying $\sum_{i}\delta_{i}\leq\delta^{\prime\prime}$ guarantees $(\epsilon,\delta^{\prime}+\delta^{\prime\prime})$-DP.
+[^11] By setting $\epsilon_{1}=\epsilon_{2}=\dots=0$ in Theorem 2 of (whitehouse2023fully, ), it follows that for every $\epsilon>0$, $\delta^{\prime}>0$, and $0\leq\delta^{\prime\prime}\leq 1$, the composition of randomized response mechanisms $\mathrm{RR}_{0,\delta_{1}},\mathrm{RR}_{0,\delta_{2}},\dots$ with adaptively chosen privacy parameters $\delta_{1},\delta_{2},\dots$ satisfying $\sum_{i}\delta_{i}\leq\delta^{\prime\prime}$ guarantees $(\epsilon,\delta^{\prime}+\delta^{\prime\prime})$-DP.
 Taking $\epsilon$ and $\delta^{\prime}$ to zero, we can conclude that this composition is $(0,\delta^{\prime\prime})$-DP.
@@
 Lemma 5.11.
-For every $\epsilon\geq 0$ and $0\leq\delta,\delta^{\prime}\leq 1$, if ${\textrm{Comp}}(\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0})$ is $(\epsilon,\delta)$-DP, then the $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta^{\prime},RR}$-concurrent composition of CMs is $(\epsilon,\delta+\delta^{\prime})$-DP.
+For every $\epsilon\geq 0$ and $0\leq\delta,\delta^{\prime}\leq 1$, if ${\textrm{Comp}}(\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0})$ is $(\epsilon,\delta)$-DP, then the $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta^{\prime},RR}$-concurrent composition of CMs is $(\epsilon,1-(1-\delta)(1-\delta^{\prime}))$-DP.
 Proof.
@@
 $f^{\mathrm{RR}}$.
-Therefore, by Theorem 4.9 and the fact that the composition of $\mathrm{RR}_{\epsilon,\delta}$ and $\mathrm{RR}_{0,\delta^{\prime}}$ is $(\epsilon,\delta+\delta^{\prime})$-DP, ExtConComp is $(\epsilon,\delta+\delta^{\prime})$-DP w.r.t.
+By the optimal composition theorem of (murtagh2015complexity, ), we have that the composition of $\mathrm{RR}_{\epsilon,\delta}$ and $\mathrm{RR}_{0,\delta^{\prime}}$ is $(\epsilon,1-(1-\delta)(1-\delta^{\prime}))$-DP.
+[^12] Replacing $k=2$, $\epsilon_{1}=\epsilon$, $\delta_{1}=\delta$, $\epsilon_{2}=0$, $\delta_{2}=\delta^{\prime}$, and $\epsilon_{g}=\epsilon$ in Theorem 1.5 of (murtagh2015complexity, ) implies that for every $\delta_{g}$ such that $0\leq 1-\frac{1-\delta_{g}}{(1-\delta)(1-\delta^{\prime})}$, the composition of $\mathrm{RR}_{\epsilon,\delta}$ and $\mathrm{RR}_{0,\delta^{\prime}}$ is $(\epsilon,\delta_{g})$-DP.
+$\delta_{g}=1-(1-\delta)(1-\delta^{\prime})$ satisfies this inequality.
+Therefore, by Theorem 4.10, ExtConComp is $(\epsilon,1-(1-\delta)(1-\delta^{\prime}))$-DP w.r.t.
 the verification function $f^{*}=f^{\gamma_{1},\gamma_{2}}$, finishing the proof.
 ∎
-Remark 5.12 (Improved Basic Composition).
-An immediate consequence of Lemma 5.11 is an improvement in the well-known basic composition theorem, stating that the composition of $k$ randomized response mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $\left(\sum_{i=1}^{k}\epsilon_{i},\sum_{i=1}^{k}\delta_{i}\right)$-DP.
-By separately composing $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ using the original basic composition and composing $\mathrm{RR}_{0,\delta_{1}},\dots,\mathrm{RR}_{0,\delta_{k}}$ using Lemma 5.9, we can conclude that the composition of $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $\left(\sum_{i=1}^{k}\epsilon_{i},1-\prod_{i=1}^{k}(1-\delta_{i})\right)$-DP, where $1-\prod_{i=1}^{k}(1-\delta_{i})\leq\sum_{i=1}^{k}\delta_{i}$ always holds.
 Proof of Theorem 5.4.
@@
 Based on a secret bit $b$, either all created mechanisms are initialized with the first initial state in their respective pairs, or all are initialized with the second.
-The goal of the adversary is to guess $b$.
+The goal of the adversary is to distinguish whether $b=0$ or $b=1$.
 This scenario is an special case of FPC concurrent $k$-sparse parallel composition of CMs, where, instead of creating $\mathcal{M}_{i}$ and assigning a pair of initial states to it, the adversary sends $(\mathcal{M}_{i}^{\prime},\text{null},\epsilon_{i},\delta_{i},f_{\sim_{i}})^{2}$, where $\mathcal{M}_{i}^{\prime}$ is the CM simulating $\mathcal{M}_{i}$.
 Moreover, to ask a query $q$ from the $i$-th created IM, the adversary sends $\left((q,i),(q,i)\right)$.
-Since each $f_{\sim_{i}}$ is first-pair consistent, Theorem 5.15 applies to the concurrent $k$-sparse parallel composition of IM with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$ (that are differentially private w.r.t.
+Since each $f_{\sim_{i}}$ is first-pair consistent, Theorem 5.14 applies to the concurrent $k$-sparse parallel composition of IM (that are differentially private w.r.t.
 neighbor relations).
+Corollary 5.15.
+For $\epsilon_{1},\dots,\epsilon_{k}>0$, $\epsilon\geq 0$ and $0\leq\delta,\delta_{1},\dots,\delta_{k}\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the concurrent $k$-sparse parallel composition of IM with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$ is also $(\epsilon,\delta)$-DP.
 NIMs are a special subclass of IMs that halt the communication upon receiving a second input message.
 Since Theorem 5.14 applies to all IMs, it also holds for the concurrent $k$-sparse parallel composition of NIMs with privacy parameters $\left((\epsilon_{1},\delta_{i}\right)_{i=1}^{k}$.
-We note that the restricted version of Theorem 5.15 for the $k$-sparse parallel composition of (unbounded number of) NIMs was previously proved in [20] as an intermediate lemma.
 In the next section, we compare the $k$-sparse parallel composition of IMs to the classical parallel composition theorem for NIMs.
@@
 The privacy analysis of such mechanisms $\mathcal{M}$ is challenging as the creation of sub-mechanisms and their inputs is adaptive and dependent.
-The goal of this section is to give a set of conditions for $\mathcal{M}$ and prove a theorem that shows that under these conditions the privacy of $\mathcal{M}$ can directly be computed by applying the known concurrent composition theorems for continual mechanisms to the sub-mechanisms.
+The goal of this section is to give a set of conditions for $\mathcal{M}$ and prove a theorem that shows that under these conditions the privacy of $\mathcal{M}$ can directly be bounded by applying our concurrent composition theorems for continual mechanisms to the sub-mechanisms.
 Specifically, given that the $f^{\prime}$-concurrent composition of CMs is $(\epsilon,\delta)$-DP w.r.t.
@@
 To prove that $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t.
-$f$, we need to show that the view of every adversary $\mathcal{A}$ interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M}=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable.
-Intuitively, the messages that $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ receives from $\mathcal{I}(b)$ contains information about $b$.
+$f$, we need to show that the views of every adversary $\mathcal{A}$ interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M}=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ are $(\epsilon,\delta)$-indistinguishable.
+Intuitively, the messages that $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ receives from $\mathcal{I}(b)$ contain information about $b$.
 Therefore, to deduce that $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-DP w.r.t.
@@
 $f^{\prime}$, the IPM $\mathcal{P}$ must not directly use the messages from $\mathcal{I}$ to respond.
-Instead, it should only use these messages to determine which queries it wants to send to its its sub-mechanisms via ExtConComp.
-More precisely, $\mathcal{P}$ must generate an answer for $\mathcal{I}$ only based on the privatized answers provided by ExtConComp and not dependent on the un-privatized input sequence (Response Property).
+Instead, it should only use these messages to determine which queries it wants to send to its sub-mechanisms via ExtConComp.
+More precisely, $\mathcal{P}$ must generate an answer for $\mathcal{I}$ only based on the privatized answers provided by ExtConComp and independent of the un-privatized input sequence (Response Property).
 Moreover, $\mathcal{P}$ must decide whether to continue interacting with ExtConComp or to return a response based on these privatized answers (Destination Property).
@@
 In contrast, mechanisms are usually described more simply using pseudocode rather than message-passing state machines.
-In Section 6.3, we reformulate our results in terms of pseudocode, enabling new mechanisms to be analyzed directly without the need for reformulating to the setting of in this paper.
+In Section 6.3, we reformulate our results in terms of pseudocode, enabling new mechanisms to be analyzed directly without the need for reformulating to the setting in this paper.
 6.1. Properties.
@@
 Recall that the goal of this section is to show $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-DP w.r.t.
-$f^{\prime}$ if ExtConComp is $(\epsilon,\delta)$-DP w.r.t.
-$f$.
-That is, we need to show that the view of every adversary interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable.
+$f$ if ExtConComp is $(\epsilon,\delta)$-DP w.r.t.
+$f^{\prime}$.
+That is, we need to show that the views of every adversary interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ are $(\epsilon,\delta)$-indistinguishable.
 The messages received by $\mathcal{P}$ from $\mathcal{I}$ contain information about the bit used by $\mathcal{I}$ to filter the adversary’s pairs.
-Therefore, to directly use the existing compositions theorems for ExtConComp to analyze the privacy of $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$, we need to ensure that $\mathcal{P}$ only utilizes the messages from $\mathcal{I}$ to generate queries for the existing private sub-mechanisms executed by ExtConComp.
+Therefore, to directly use the existing composition theorems for ExtConComp to analyze the privacy of $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$, we need to ensure that $\mathcal{P}$ only utilizes the messages from $\mathcal{I}$ to generate queries for the existing private sub-mechanisms executed by ExtConComp.
 Moreover, we need to guarantee that the sequence $(m_{1}^{0\prime},m_{1}^{1\prime}),(m_{2}^{0\prime},m_{2}^{1\prime}),\dots$ is $f^{\prime}$-valid, where $m_{1}^{b\prime},m_{2}^{b\prime},\dots$ is the sequence of messages that $\mathcal{P}$ sends to ExtConComp when it receives messages from $\mathcal{I}(b)\!\circ^{*}\!\mathcal{V}[f]$.
@@
   • $f\to f^{\prime}$ Property.
-For $v\in\{{\texttt{L}},{\texttt{R}}\}$, let the pair $(m,v)$ denote a message $m$ sent to or received from the left or right mechanism of $\mathcal{P}$.
+For $v\in\{{\texttt{L}},{\texttt{R}}\}$, let the pair $(m,v)$ denote a message $m$ sent to or received from the $v$’th mechanism (i.e., left or right) of $\mathcal{P}$.
 For $b\in\{0,1\}$, let $\sigma^{b}=\left((m_{1}^{b},v_{1}^{b}),\dots,(m_{t}^{b},v_{t}^{b})\right)$ be a sequence of $t$ input messages to $\mathcal{P}$, and let $\tau^{b}=\left((m_{1}^{b\prime},v_{1}^{b\prime}),\dots,(m_{t}^{b\prime},v_{t}^{b\prime})\right)$ be the corresponding sequence of $t$ output messages from $\mathcal{P}$.
@@
 $f$ against adaptive adversaries.
-In Theorem 6.1, we assume that the view of every adversary interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable.
+In Theorem 6.1, we assume that the views of every adversary interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ are $(\epsilon,\delta)$-indistinguishable.
 The goal is to show the same result for $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$.
-The key idea of the proof is to show that the IMs $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{T}\!\circ^{*}\!\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are identical, where $\mathcal{T}$ is an IPM with a singleton initial state space.
+The key idea of the proof is to show that the IMs $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{T}\!\circ^{*}\!\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are identical for some IPM $\mathcal{T}$ with a singleton initial state space.
 The proof consists of two steps.
@@
 When $\mathcal{P}^{2}$ receives a message $m\in Q_{\mathcal{P}}^{L}$ from the left mechanism, it forwards two copies of $m$ to both $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ as left messages.
-Specifically, it executes $(s_{0},v_{0},m_{0}^{\prime})\leftarrow\mathcal{P}_{0}(s_{0},{\texttt{L}},m)$ and $(s_{1},v_{1},m_{1}^{\prime})\leftarrow\mathcal{P}_{1}(s_{1},{\texttt{L}},m)$.
+Specifically, it executes $(s_{0},v_{0}^{\prime},m_{0}^{\prime})\leftarrow\mathcal{P}_{0}(s_{0},{\texttt{L}},m)$ and $(s_{1},v_{1}^{\prime},m_{1}^{\prime})\leftarrow\mathcal{P}_{1}(s_{1},{\texttt{L}},m)$.
 When $\mathcal{P}^{2}$ receives a pair of messages $(m_{0},m_{1})\in Q_{\mathcal{P}}^{R}\times Q_{\mathcal{P}}^{R}$ from the right mechanism, it passes each $m_{b}$ to $\mathcal{P}_{b}$ as a right message.
-Specifically, it executes $(s_{0},v_{0},m_{0}^{\prime})\leftarrow\mathcal{P}_{0}(s_{0},{\texttt{R}},m_{0})$ and $(s_{1},v_{1},m_{1}^{\prime})\leftarrow\mathcal{P}_{1}(s_{1},{\texttt{R}},m_{0})$.
+Specifically, it executes $(s_{0},v_{0}^{\prime},m_{0}^{\prime})\leftarrow\mathcal{P}_{0}(s_{0},{\texttt{R}},m_{0})$ and $(s_{1},v_{1}^{\prime},m_{1}^{\prime})\leftarrow\mathcal{P}_{1}(s_{1},{\texttt{R}},m_{0})$.
 By the destination property of $\mathcal{P}$ and the fact that both $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ receive identical left messages, we have $v_{0}^{\prime}=v_{1}^{\prime}$.
@@
 If $v_{0}^{\prime}=v_{1}^{\prime}={\texttt{L}}$, then by the response property, $m_{0}^{\prime}=m_{1}^{\prime}$, and $\mathcal{P}^{2}$ returns $m_{0}^{\prime}$ to its left mechanism.
-The following claim follow directly from the definition of $\mathcal{P}^{2}$:
+The following claim follows directly from the definition of $\mathcal{P}^{2}$:
 Claim 6.3.
@@
 $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}^{2}$ in Claim 6.4 is an IPM with a singleton initial state space.
-Thus, by Lemma 2.16, that for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable implies that for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable, which finishes the proof.
+Thus, by Lemma 2.16, the fact that for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ are $(\epsilon,\delta)$-indistinguishable implies that for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable, which finishes the proof.
 ∎
@@
 For clarity and practicality, we restate the results of this section in the pseudocode framework.
-Decision vs. Auxiliary Variables.
-We partition the variables of a continual mechanism $\mathcal{M}$ into two categories:
-  • Decision variables: These are either outputs of submechanisms or variables that are computed solely based on decision variables.
-Intuitively, it follows that they are post-processing of the privatized outputs.
-  • Auxiliary variables: These are all remaining variables. The values of these variables may depend on the mechanism’s inputs as well as on both decision and auxiliary variables.
+Tainted vs. Untainted Variables.
+To rigorously track sensitive data, we rely on concepts from computer security.
+Information flow characterizes how data transfers between variables during a program’s execution (denning1976lattice, ; sabelfeld2003language, ).
+Taint analysis builds upon this to monitor the propagation of sensitive inputs, known as taint sources (newsome2005dynamic, ; schwartz2010all, ).
+Any variable computed using a taint source becomes a tainted variable.
+To safely release data, it must be processed by a sanitizer—in our context, a differentially private submechanism.
+The sanitizer mathematically interrupts the sensitive information flow, rendering its outputs and all subsequent downstream computations as untainted variables.
+To formalize this, we partition the variables of a mechanism into two categories using a recursive definition based on their inputs:
+  • Tainted variables: The mechanism’s inputs are strictly tainted.
+Furthermore, any intermediate variable computed via a function where at least one input argument is a tainted variable is also considered tainted.
+These variables retain a direct dependence on the raw inputs.
+  • Untainted variables: Variables computed solely from privatized data or fixed values (such as the privacy parameters or constants) are untainted.
+Crucially, the output of any differentially private submechanism (the sanitizer) is defined as untainted, regardless of whether its inputs were tainted.
+Any subsequent variable computed as a function of only untainted variables remains untainted.
+Consequently, untainted variables are purely a post-processing of privatized outputs.
 Note.
-For simplicity, we assume pseudocodes are deterministic as any randomness could be modeled as a submechanism call.
+For simplicity, we assume the pseudocode of $\mathcal{M}$ is deterministic as any randomness could be modeled as a submechanism call.
 For instance, sampling from the uniform distribution could be represented as invoking a 0-DP submechanism that outputs such a sample on a fixed input.
-Properties.
+Definition 6.5.
 For $t\in\mathbb{N}$, given two sequences $\sigma^{0}=(m_{1}^{0},\dots,m_{t}^{0})$ and $\sigma^{1}=(m_{1}^{1},\dots,m_{t}^{1})$, their combination is defined as $\big((m_{1}^{0},m_{1}^{1}),\dots,(m_{t}^{0},m_{t}^{1})\big)$.
@@
     Destination property.
-A pseudocode satisfies this property if (i) the decision of whether to produce an output, create a new submechanism, or interact with an existing one, (ii) the choice of the new submechanism in the second case, and (iii) the selection of the submechanism to interact with in the third case is made solely based on decision variables.
+A pseudocode satisfies this property if (i) the decision of whether to produce an output, create a new submechanism, or interact with an existing one, (ii) the choice of the new submechanism in the second case, and (iii) the selection of the submechanism to interact with in the third case is made solely based on untainted variables.
   2. (2)
-    Response property. A pseudocode satisfies this property if its output is determined solely by decision variables.
+    Response property. A pseudocode satisfies this property if its output is determined solely by untainted variables.
   3. (3)
@@
 Let $f$ and $f^{\prime}$ be verification functions.
-A pseudocode satisfies this property if, for every $t\in\mathbb{N}$ and every pair of input sequences $(\sigma^{0},\sigma^{1})$ of length $t$ forming an $f$-valid combination, the following holds: Assume that all decision variables of $\mathcal{M}$ are updated identically through the end of time step $t$.
+A pseudocode satisfies this property if, for every $t\in\mathbb{N}$ and every pair of input sequences $(\sigma^{0},\sigma^{1})$ of length $t$ forming an $f$-valid combination, the following holds: Assume that all untainted variables of $\mathcal{M}$ are updated identically through the end of time step $t$.
 For each $b\in\{0,1\}$, define $\sigma^{b}$ to denote the sequence of creation queries and inputs to submechanisms when processing $\sigma^{b}$.
-Here, a creation query is a tuple specifying a created submechanisms, its privacy parameters, and associated verification functions.
+Here, a creation query is a tuple specifying a created submechanism, its privacy parameters, and associated verification functions.
 Also, an input $m$ to the $j$-th submechanism is represented by the pair $(m,j)$.
@@
 the neighbor relation $\sim$ (or simply it is $(\epsilon,\delta)$-DP) if the mechanism is $(\epsilon,\delta)$-DP w.r.t.
-to the verification function $f_{\sim}$ against adaptive adversaries, where $f_{\sim}$ is defined as follows: For every message sequence $m_{1},\dots,m_{t}$, $f_{\sim}(m_{1}\dots,m_{t})=\top$ if and only if the message sequence is suitable (implying $m_{j}=(m_{j}^{0},m_{j}^{1})$ for every $j\in\{1,\dots,t\}$) and the sequences $\sigma_{0}=(m_{1}^{0},\dots,m_{t}^{0})$ and $\sigma_{1}=(m_{1}^{1},\dots,m_{t}^{1})$ satisfy $\sigma_{0}\sim\sigma_{1}$.
+the verification function $f_{\sim}$ against adaptive adversaries, where $f_{\sim}$ is defined as follows: For every message sequence $m_{1},\dots,m_{t}$, $f_{\sim}(m_{1}\dots,m_{t})=\top$ if and only if the message sequence is suitable in the sense of Definition 4.5 (implying $m_{j}=(m_{j}^{0},m_{j}^{1})$ for every $j\in\{1,\dots,t\}$) and the sequences $\sigma_{0}=(m_{1}^{0},\dots,m_{t}^{0})$ and $\sigma_{1}=(m_{1}^{1},\dots,m_{t}^{1})$ satisfy $\sigma_{0}\sim\sigma_{1}$.
 Definition 7.2.
 Let $d,T\in\mathbb{N}$ and let $q:\mathbb{N}^{d}\to\mathbb{N}$ be a function that is monotone in its input.
-In the (differentially private) continual monotone histogram query problem, a continual mechanism $\mathcal{M}$ receives $T$ inputs of row vectors $x^{1},\dots,x^{T}$, each in $\{0,1\}^{d}$.
-Define $\sim$ to be a neighbor relation on $\left(\{0,1\}^{d}\right)^{*}$ where two input streams are neighboring if they are identical except for a single step.
-Differential privacy for $\mathcal{M}$ is defined with respect to this neighbor relation (see Remark 7.1).
+In the (differentially private) continual monotone histogram query problem, a continual mechanism $\mathcal{M}$ receives $T$ inputs of row vectors $x=(x^{1},\dots,x^{T})$, each in $\{0,1\}^{d}$.
+Differential privacy for $\mathcal{M}$ is defined with respect to the neighbor relation $\sim_{H}$ defined as follows (see Remark 7.1).
+For $T\in\mathbb{N}$, two input streams in $\left(\{0,1\}^{d}\right)^{T}$ are neighboring if they are identical except for a single step.
 The output of $\mathcal{M}$ at each time step $t$ is (an additive approximation to) $q(t)=q((\sum_{\ell=1}^{t}x_{i}^{\ell})_{i\in[d]})$.
 Definition 7.3.
-Recall function $q:\{0,1\}^{d}\to\mathbb{R}$ and neighbor relation $\sim$ from Definition 7.2. $q$ is said to be $1$-sensitive if $\max_{x^{0}\sim x^{1}}|q(x^{0})-q(x^{1})|$ is at most $1$.
-Henzinger, Sricharan, and Steiner [13] presented the best known differentially private mechanism for the monotone histogram query problem, where the query is $1$-sensitive.
-For $\epsilon$-differential privacy it has, with probability at least $1-\beta$, an additive error of $O((d\log^{2}(dq^{*}/\beta)+\log T)\epsilon^{-1})$, and for $(\epsilon,\delta)$-differential privacy, it has an additive error of $O((d\log^{1.5}(dq^{*}/\beta)+\log T)\epsilon^{-1}\log\frac{1}{\delta})$, where $q^{*}=\max_{t\in[T]}q(t)$.
-The privacy of the mechanism is proven “from scratch” using the concept of a privacy game [14] and the latter bound only holds against an oblivious adversary.
-In Section 7.1, we give an overview of the mechanism in [13], which we will call HSS, and in Section 7.2, we give a simple privacy proof for it using our approach.
+Recall function $q:\{0,1\}^{d}\to\mathbb{R}$ and neighbor relation $\sim_{H}$ from Definition 7.2. $q$ is said to be $1$-sensitive if $\max_{x^{0}\sim_{H}x^{1}}|q(x^{0})-q(x^{1})|$ is at most $1$.
+Henzinger, Sricharan, and Steiner (henzingerSS2023, ) presented the best known differentially private mechanism for the monotone histogram query problem, where the query is $1$-sensitive.
+For $\epsilon$-differential privacy it has, with probability at least $1-\beta$, an additive error of $O((d\log^{2}(dq^{*}/\beta)+\log T)/\epsilon)$, and for $(\epsilon,\delta)$-differential privacy, it has an additive error of $O((d\log^{1.5}(dq^{*}/\beta)+\log T)\log\frac{1}{\delta}/\epsilon)$, where $q^{*}=\max_{t\in[T]}q(t)$.
+The privacy of the mechanism is proven “from scratch” using the concept of a privacy game (pricedpco, ) and the bound was only shown for $(\epsilon,\delta)$-DP against an oblivious adversary.
+In Section 7.1, we give an overview of the mechanism in (henzingerSS2023, ), which we will call HSS, and in Section 7.2, we give a simple privacy proof for it using our approach.
 With the new privacy analysis, the result for $(\epsilon,\delta)$-differential privacy is stronger and holds against an adaptive adversary.
@@
 When $\texttt{SVT-init}(\epsilon,q,h)$ is invoked, the continual mechanism $\textrm{SVT}_{h,\epsilon,q}$ with the single initial state $(\epsilon,q,h,\tau=\text{null})$ is created.
-In the pseudocode of SVT-init, the variable $\tau$ takes value.
+In the pseudocode of SVT-init, the variable $\tau$ takes a value.
 In our setting, such initialization is considered to happen when $\textrm{SVT}_{h,\epsilon,q}$ receives its first input.
-The SVT-update operation then describes how $\textrm{SVT}_{h,\epsilon,q}$ maps its current state (i.e., set of variables) and an input message $m=(x,\texttt{Thresh})\in\{0,1\}^{d}\times\mathbb{R}$ to a new states (with updated variables) and a response $m^{\prime}\in\{\top,\bot\}$.
+The SVT-update operation then describes how $\textrm{SVT}_{h,\epsilon,q}$ maps its current state (i.e., set of variables) and an input message $m=(x,\texttt{Thresh})\in\{0,1\}^{d}\times\mathbb{R}$ to a new state (with updated variables) and a response $m^{\prime}\in\{\top,\bot\}$.
 Algorithm 3 Sparse vector technique for histogram queries
@@
 end procedure
-As part of their privacy proof, the following is shown in [13]:
+As part of their privacy proof, the following is shown in (henzingerSS2023, ):
 Lemma 7.4.
-If $q$ is a $1$-sensitive function, then for every $\epsilon$ and $h$, the mechanism $\textrm{SVT}_{\epsilon/6,q,h}$ is $\epsilon/3$-DP w.r.t. the neighbor relation $\sim$ in Definition 7.2.
+If $q$ is a $1$-sensitive function, then for every $\epsilon$ and $h$, the mechanism $\textrm{SVT}_{\epsilon/6,q,h}$ is $\epsilon/3$-DP w.r.t.
+the neighbor relation $\sim_{S}$ defined as follows (see Remark 7.1).
+For $T\in\mathbb{N}$, two input streams $\left((x_{1},\tau_{1}),\dots,(x_{T},\tau_{T})\right)$ and $\left((y_{1},\nu_{1}),\dots,(y_{T},\nu_{T})\right)$ in $\left((\{0,1\}^{d}\times\mathbb{R})\right)^{T}$ are neighboring if the (threshold) sequences $(\tau_{1},\dots,\tau_{T})$ and $(\nu_{1},\dots,\nu_{T})$ are identical at every entry and the (binary) sequences $(x_{1},\dots,x_{T})$ and $(y_{1},\dots,y_{T})$ are identical in all but one entry.
 $d$-dimensional continual counter.
 The mechanism d-counter is described by two operations.
-The first, $\texttt{d-counter-init}(\epsilon)$, creates $d$ differentially private continual binary counters with privacy parameter $\epsilon/d$.
-The second, $\texttt{d-counter-update}(x)$, gives the elements of its $d$-dimensional input $x$ to their corresponding binary counters and returns the aggregated outputs as a $d$-dimensional vector.
+The first, $\texttt{d-counter-init}(\epsilon)$, creates $d$ continual counters that are $\epsilon/d$-DP w.r.t.
+the neighbor relation $\sim^{\prime}$, where two real-valued sequences of the same length are neighbors if they differ only on a single step by at most $1$.
+The second, $\texttt{d-counter-update}(c)$, gives the elements of its $d$-dimensional input $c\in\mathbb{R}^{d}$ to their corresponding counters and returns the aggregated outputs as a $d$-dimensional vector.
 We denote this continual mechanism by $\texttt{d-counter}_{\epsilon}$.
-In another version, $\texttt{d-counter-init}(\epsilon,\delta)$ creates $d$ differentially private continual binary counters with privacy parameters $\epsilon/d$ and $\delta/d$.
+In another version, $\texttt{d-counter-init}(\epsilon,\delta)$ creates $d$ continual counters, each satisfying $(\epsilon/d,\delta/d)$-DP w.r.t.
+$\sim^{\prime}$.
 We represent this continual mechanism by $\texttt{d-counter}_{\epsilon,\delta}$.
@@
 Corollary 7.5.
-Mechanisms $\texttt{d-counter}_{\epsilon}$ and $\texttt{d-counter}_{\epsilon,\delta}$ are $\epsilon$-DP and $(\epsilon,\delta)$-DP, respectively.
+Mechanisms $\texttt{d-counter}_{\epsilon}$ and $\texttt{d-counter}_{\epsilon,\delta}$ are $\epsilon$-DP and $(\epsilon,\delta)$-DP w.r.t.
+the neighbor relation $\sim_{C}$ defined as follows (see Remark 7.1).
+For $T\in\mathbb{N}$, two input streams $x=(x_{1},\dots,x_{T})$ and $y=(y_{1},\dots,y_{T})$ in $(\mathbb{R}^{d})^{T}$ are neighboring if there exists $i\in[T]$ such that $x_{j}=y_{j}$ for all $j\in[T]\setminus\{i\}$ and $|x_{i}-y_{i}|_{\infty}\leq 1$, i.e., each entry of $x_{i}$ and $y_{i}$ differs by at most $1$.
 Continual histogram query mechanism.
-The mechanism HSS, described in Algorithm 4, is initialized with a privacy parameter $\epsilon>0$ and an accuracy parameter $0<\beta\leq 1$ and receives a vector $x\in\{0,1\}^{d}$ as input at each time step.
+The mechanism HSS, described in Algorithm 4, is initialized with a privacy parameter $\epsilon>0$, a query $q:\mathbb{N}^{d}\to\mathbb{N}$, and an accuracy parameter $0<\beta\leq 1$ and receives a vector $x\in\{0,1\}^{d}$ as input at each time step.
 This mechanism partitions the input stream into subsequences, called intervals.
@@
 Upon receiving an input $x$ at step $t$, the mechanism updates $c$ and invokes $\texttt{SVT-update}(x,\texttt{Thresh})$.
-If the response is $\top$, the current interval ends: $c$ is passed to d-counter, $h$ is updated with the output of d-counter, and $c$ is reset to $0^{d}$.
-Since SVT halts after returning $\top$, a new instance $\textrm{SVT}_{\epsilon/6,q,h}$ is created.
-In this case, HSS privately evaluates $q(c+h)$ via the Laplace noise and compares the noisy result to $\texttt{Thresh}-\xi(t,j,\beta,\epsilon)$.
+If the response, denoted by $a$, is $\top$, the current interval ends and HSS takes the following actions: (threshold update) It first privately evaluates $q(c+h)$ via the Laplace noise and compares the noisy result to $\texttt{Thresh}-\xi(t,j,\beta,\epsilon)$.
 If the inequality holds, Thresh is suitably increased.
-Moreover, independent of the outcome of this comparison, Thresh is updated again to reflect the increase in the number of intervals.
-Then, regardless of the output of $\texttt{SVT-update}(x,\texttt{Thresh})$, HSS updates Thresh once more to account for the increased number of time steps.
-Finally, HSS outputs $h$.
+Then independent of the outcome of this comparison, Thresh is updated again to reflect the increase in the number of intervals.
+(summation update) Subsequently, HSS passes $c$ to d-counter, updates $h$ with the output of d-counter, and resets $c$ to $0^{d}$.
+Since SVT halts after returning $\top$, HSS also creates a new instance $\textrm{SVT}_{\epsilon/6,q,h}$.
+Finally, regardless of the output of $\texttt{SVT-update}(x,\texttt{Thresh})$, HSS updates Thresh once more to account for the increased number of time steps and outputs $h$.
 Algorithm 4 Mechanism for answering a monotone histogram query
@@
   $\texttt{SVT-init}(\epsilon/6,q,h)$
+  $a\leftarrow\bot$
   $\texttt{Thresh}=\gamma(1,j,\beta,\epsilon)$
@@
   $t\leftarrow t+1$
-  $c_{i}\leftarrow c_{i}+x_{i}$ for all $i\in[d]$,
-  if $\texttt{SVT-update}(x,\texttt{Thresh})==\top$ then
+  $c_{i}\leftarrow c_{i}+x_{i}$ for all $i\in[d]$
+  $a=\texttt{SVT-update}(x,\texttt{Thresh})$
+  if $a==\top$ then
+   if $q(c+h)+\mathrm{Lap}(3/\epsilon)>\texttt{Thresh}-\xi(t,j,\beta,\epsilon)$ then
+     $\texttt{Thresh}=\texttt{Thresh}+\gamma(t,j,\beta,\epsilon)$
+   end if
    $h=\texttt{d-counter-update}(c)$
@@
    $c_{i}\leftarrow 0$ for all $i\in[d]$
-   if $q(c+h)+\mathrm{Lap}(3/\epsilon)>\texttt{Thresh}-\xi(t,j,\beta,\epsilon)$ then
-     $\texttt{Thresh}=\texttt{Thresh}+\gamma(t,j,\beta,\epsilon)$
-   end if
    $j\leftarrow j+1$
@@
 In Algorithm 4, the summation $q(c+h)+\mathrm{Lap}(3/\epsilon)$ can be rewritten as an instance of Laplace mechanism with privacy parameter $\epsilon/3$ and input $q(c+h)$.
+The privacy of the Laplace mechanism is defined with respect to the neighbor relation $\sim_{L}$, where two real values $r_{1},r_{2}\in\mathbb{R}$ satisfy $r_{1}\sim_{L}r_{2}$ if and only if $|r_{1}-r_{2}|\leq 1$.
 Consequently, the mechanism HSS executes:
@@
     an arbitrary number of Laplace mechanisms.
-By Corollary 7.5, d-counter is $\epsilon/3$-DP (or $(\epsilon/3,\delta)$-DP when approximate counters are used).
-By Lemma 7.4, each SVT instance is $\epsilon/3$-DP.
-Finally, the Laplace mechanism with parameter $\epsilon/3$ is a non-interactive mechanism (NIM) that is well known to satisfy $\epsilon/3$-DP on neighboring real inputs differing by at most $1$.
+By Corollary 7.5, d-counter is $\epsilon/3$-DP (or $(\epsilon/3,\delta)$-DP when approximate counters are used) w.r.t.
+$\sim_{C}$.
+By Lemma 7.4, each SVT instance is $\epsilon/3$-DP w.r.t.
+$\sim_{S}$.
+Finally, the Laplace mechanism with parameter $\epsilon/3$ is a non-interactive mechanism (NIM) that is well known to satisfy $\epsilon/3$-DP on neighboring real inputs differing by at most $1$ (i.e., w.r.t.
+to neighbor relation $\sim_{L}$ defined above).
 The idea is to show that the concurrent composition of multiple instances of these mechanisms created by HSS is $\epsilon$-DP (or $(\epsilon,\delta)$-DP when d-counter is approx DP).
-Then, we will then show that HSS satisfies the required properties in Corollary 6.5, and thus it has the same privacy guarantees as the concurrent composition.
-As concurrent compositions theorems in this paper require the composed mechanisms to have discrete answer distributions, we first show that we can assume the output space of d-counter and Laplace mechanism is discrete.
+Then, we will show that HSS satisfies the required properties in Corollary 6.6, and thus it has the same privacy guarantees as the concurrent composition.
+As concurrent composition theorems in this paper require the composed mechanisms to have discrete answer distributions, we first show that we can assume the output space of d-counter and Laplace mechanism is discrete.
 Since SVT has a finite output space, its answer distributions are also discrete.
 Discretization of outputs.
-As argued earlier, the counters executed by d-counter outputs integer values, and thus d-counter has discrete answer distributions.
+As argued earlier, the counters executed by d-counter output integer values, and thus d-counter has discrete answer distributions.
 Likewise, we can assume that the Laplace mechanism in HSS rounds its noisy real-valued output to the nearest integer.
-By the post-processing lemma, this rounding does not degrade privacy, and by definition, it increases the additive error by at most $1$, which is unimportant for the accuracy guarantees of [13].
+By the post-processing lemma, this rounding does not degrade privacy, and by definition, it increases the additive error by at most $1$, which is unimportant for the accuracy guarantees of (henzingerSS2023, ).
 Hence, both d-counter and the Laplace mechanism can be treated as having discrete output distributions.
 Pseudocode privacy analysis.
-Recall from Section 6.3 that decision variables of a mechanism are either (i) outputs of submechanisms, or (ii) variables computed solely as functions of other decision variables.
-In Algorithm 4, the variables $h$, outout, $j$, $t$, and Thresh are decision variables, while $c$ is not: Variable $h$ denotes the privatized outcomes of d-counter.
-Variable outout equals $q(h)$, where $h$ is a decision variable.
-Index $j$ is initially $1$ and is increased by $1$ when the privatized output of SVT is $\top$.
-Time step $t$ is initially 0 and is increase by $1$ at each step, independent of the input and other variables.
-Finally, variable Thresh is initialized and updated to values determined by $\epsilon,\beta,j,t$, where $\epsilon,\beta$ are fixed and $j,t$ are decision variables.
-It is important to note that the decision of updating variables $h$ (and consequently outout) and Thresh depends only on the privatized output of SVT.
+Recall from Section 6.3 that untainted variables of a mechanism are (i) outputs of submechanisms, (ii) fixed values (e.g., privacy and accuracy parameters), or (iii) variables computed solely as functions of other untainted variables.
+In Algorithm 4, the variables $\epsilon$, $\beta$, $q$, $a$, $h$, outout, $j$, $t$, and Thresh are untainted, while $c$ is not: The values of the variables $\epsilon$, $\beta$, and $q$ are fixed and independent of the sensitive data.
+Variables $a$ and $h$ denote the privatized outcomes of SVT and d-counter, respectively.
+Variable outout equals $q(h)$, where $h$ is an untainted variable.
+Index $j$ is initially $1$ and is increased by $1$ when $a=\top$.
+Time step $t$ is initially 0 and is increased by $1$ at each step, independent of the input and other variables.
+Finally, variable Thresh is initialized and updated to values determined by the untainted variables $\epsilon,\beta,j,t$.
+It is important to note that the decision of updating variables $h$ (and consequently outout) and Thresh depends only on the privatized output of SVT, i.e., the untainted variable $a$.
 We first show that the destination property holds.
-For that we have to show that the decision of whether to produce an output, create a new submechanism (and if so, which one) or interact with an existing one (and if so, which one) are solely based on decision variables.
-All submechanisms of HSS are instantiated either in the initial step (and, thus, whether it receives input is independent of non-decision variables and the inputs $x$) or based on outputs of the private submechanism SVT.
-The active SVT instance receives input at every step, and thus whether it receives input is independent of non-decision variables and the inputs $x$.
+For that we have to show that the decision of whether to produce an output, create a new submechanism (and if so, which one) or interact with an existing one (and if so, which one) is solely based on untainted variables.
+All submechanisms of HSS are instantiated either in the initial step (and, thus, whether it receives input is independent of tainted variables and the inputs $x$) or based on outputs of the private submechanism SVT.
+The active SVT instance receives input at every step, and thus whether it receives input is independent of tainted variables and the inputs $x$.
 Also, whether the mechanism d-counter gets updated depends only on the privatized output of SVT.
-Finally, HSS returns an output after the final update of Thresh, without any condition on the inputs $x$ or the non-decision variables.
+Finally, HSS returns an output after the final update of Thresh, without any condition on the inputs $x$ or the tainted variables.
 Thus, the destination property in Corollary 6.6 holds.
-Moreover, the final output of HSS is outout, which is a decision variable, so the response property in that corollary also holds.
-It remains to show HSS satisfies the $f\to f^{\prime}$ property, where $f$ is the verification function corresponding to the neighbor relation $\sim$ in Definition 7.2 and $f^{\prime}$ is a verification satisfying that the $f^{\prime}$-concurrent composition of the continual mechanisms is $(\epsilon,\delta)$-DP.
-Fix some values for all decision variables at every time step.
-Conditioned on these values, when HSS processes two neighboring input sequences, the resulting inputs to its submechanisms behave as follows:
-  • d-counter receives neighboring sequences,
-  • all but one instance of SVT receive identical inputs, with at most one receiving neighboring inputs, and
-  • all but one Laplace mechanism receive identical inputs, with at most one receiving neighboring inputs.
+Moreover, the final output of HSS is outout, which is an untainted variable, so the response property in that corollary also holds.
+It remains to show HSS satisfies the $f\to f^{\prime}$ property, where $f$ is the verification function corresponding to the neighbor relation $\sim_{H}$ in Definition 7.2 and $f^{\prime}$ is a verification function satisfying that the $f^{\prime}$-concurrent composition of the continual mechanisms is $(\epsilon,\delta)$-DP.
+Let $T\in\mathbb{N}$, and let $x=(x_{1},\dots,x_{T})$ and $y=(y_{1},\dots,y_{T})$ be two sequences in $\left(\{0,1\}^{d}\right)^{T}$ satisfying $x\sim_{H}y$, i.e., there exists $i\in[T]$ such that $x_{j}=y_{j}$ for all $j\in[T]\setminus\{i\}$.
+Fix arbitrary values for all untainted variables at every time step.
+Conditioned on these values, we compare the input streams of HSS’s submechanisms when this mechanisms runs on $x$ and $y$:
+  • Inputs of d-counter: By construction, the input stream of HSS is partitioned into intervals determined by the time steps at which SVT outputs $\top$.
+Since these outputs are untainted, they are identical in both executions to the fixed values specified in the condition.
+Let $1\leq t_{1}<\dots<t_{k}\leq T$ be the steps SVT outputs $\top$, and set $t_{0}=0$.
+The submechanism d-counter receives the interval sums $\left(\sum_{\ell=t_{j-1}+1}^{t_{j}}x_{\ell}\right)_{j=1}^{k}$ and $\left(\sum_{\ell=t_{j-1}+1}^{t_{j}}y_{\ell}\right)_{j=1}^{k}$ as input sequence when HSS runs on $x$ and $y$, respectively.
+All intervals not containing $i$ have identical sums.
+If $i>t_{k}$, no (closed) interval includes $i$ and thus all sums coincide.
+If $i\leq t_{k}$, let $j^{*}\in[k]$ denote the index satisfying $t_{j^{*}-1}+1\leq i\leq t_{j^{*}}$.
+Then
+    $\sum_{\ell=t_{j^{*}-1}+1}^{t_{j^{*}}}x_{\ell}-\sum_{\ell=t_{j^{*}-1}+1}^{t_{j^{*}}}y_{\ell}=x_{i}-y_{i}.$
+    Since $x_{i},y_{i}\in\{0,1\}^{d}$, we have $x_{i}-y_{i}\in\{-1,0,1\}^{d}$ and thus $\|x_{i}-y_{i}\|_{\infty}\leq 1$.
+Hence the inputs to d-counter differ in at most one coordinate by at most one, i.e., they are $\sim_{C}$-neighbors (see Corollary 7.5).
+  • Inputs of SVT: The threshold variable Thresh is untainted.
+Let $\texttt{Thresh}_{\ell}$ denote its fixed value at time $\ell$.
+Let time steps $t_{1},\dots,t_{k}$ be defined as above.
+Each time SVT outputs $\top$, the mechanism HSS reinitializes it.
+Therefore, HSS initiates $k+1$ instances of SVT.
+Let $t_{0}=0$ and $t_{k+1}=T$.
+For each $j\in[k+1]$, the $j$-th instance of SVT receives $\left((x_{\ell},\texttt{Thresh}_{\ell})\right)_{\ell=t_{j-1}+1}^{t_{j}}$ and $\left((y_{\ell},\texttt{Thresh}_{\ell})\right)_{\ell=t_{j-1}+1}^{t_{j}}$ as input sequence when HSS runs on $x$ and $y$, respectively.
+All instances except possibly $j^{*}$ (the one containing $i$) receive identical inputs.
+Let $j^{*}\in[k+1]$ denote the index where $t_{j^{*}-1}+1\leq i\leq t_{j^{*}}$.
+For every $j\in[k+1]\setminus\{j^{*}\}$, the input sequences of the $j$-th SVT submechanism are identical when HSS runs on $x$ and $y$.
+Moreover, the input sequences of the $j^{*}$-th SVT submechanism are identical except for the inputs $(x_{i},\texttt{Thresh}_{i})$ and $(y_{i},\texttt{Thresh}_{i})$, which have the same threshold $\texttt{Thresh}_{i}$.
+Hence, these input sequences are $\sim_{S}$-neighbors (see Lemma 7.4).
+  • Inputs of the Laplace Mechanism: The Laplace mechanism is applied to the value $q(h+c)$, where $c$ is the interval sum.
+The variable $h$ is untainted and therefore fixed and identical in both executions of HSS.
+As shown above, all interval sums coincide except possibly one, and that one differs by at most $1$ in $\ell_{\infty}$-norm.
+Hence the corresponding values $h+c$ also differ by at most $1$ in $\ell_{\infty}$-norm.
+Since $q$ is $1$-sensitive (see Definition 7.3), the real-valued outputs of $q$ on these inputs differ by at most $1$.
+Therefore, when HSS runs on $x$ and $y$, all corresponding Laplace submechanisms receive identical inputs, except possibly one pair whose inputs differ by at most $1$, that is, they are $\sim_{L}$-neighbors.
 Hence, the concurrent composition of all submechanisms of HSS can be decomposed into three concurrent compositions: (1) all Laplace mechanisms, (2) all SVT instances, and (3) the $d$ continual counters forming d-counter.
@@
 (More precisely, it is DP w.r.t.
-a verification function $f^{\prime}$ that ensures all submechanisms but one SVT instance, one Laplace mechanism, and the $d$ counters of d-counter are assigned pairs of identical inputs, while the first and second inputs for each of these mechanisms form neighboring sequences.) We note that the concurrent compositions are themselves continual mechanisms and can be composed; an explicit example of composing two compositions in the message-passing model is given in the proof of Lemma 5.11.
+a verification function $f^{\prime}$ that ensures all submechanisms but one SVT instance, one Laplace mechanism, and the $d$ counters of d-counter are assigned pairs of identical inputs, while the first and second inputs for each of these mechanisms form neighboring sequences.)
 By Corollary 6.6, since HSS satisfies the destination, response, and neighboring-mapping properties, and since the concurrent composition of its submechanisms is $\epsilon$-DP (or $(\epsilon,\delta)$-DP), the mechanism HSS satisfies the same privacy guarantee.
-We recall that every privacy guarantees in this section holds against adaptive adversaries.
+We recall that all privacy guarantees in this section hold against adaptive adversaries.
+8. Analyzing Local Differential Privacy
+So far, we have discussed the application of concurrent composition theorems in designing continual mechanisms with modular structures.
+Another natural application of these theorems arises in the setting of local differential privacy (LDP), where multiple users interact with an untrusted server over (possibly) multiple rounds.
+Each user holds a private data record (or dataset) and, upon receiving a query from the server, produces a response that satisfies differential privacy.
+After several rounds of interaction, the server aggregates the received messages and outputs an estimate of a function over the users’ data.
+Formally, a local protocol consists of an interactive mechanism for the server and an interactive mechanism for each user.
+Local differential privacy is defined with respect to a neighbor relation on sequences of datasets held by the users.
+Definition 8.1.
+Let $P$ be a local protocol with $n$ users.
+Let $\mathcal{X}$ denote the domain of datasets held by the users, and let $\sim$ be a neighbor relation on $\mathcal{X}^{n}$.
+For example, two dataset sequences $x=(x_{1},\dots,x_{n})$ and $x^{\prime}=(x_{1}^{\prime},\dots,x_{n}^{\prime})$ may satisfy $x\sim x^{\prime}$ if they differ in exactly one coordinate $i$, where $x_{i}\sim^{*}x_{i}^{\prime}$ for a neighbor relation $\sim^{*}$ on $\mathcal{X}$.
+The protocol $P$ is said to satisfy $(\epsilon,\delta)$-local differential privacy (or $(\epsilon,\delta)$-LDP) w.r.t.
+$\sim$ if, for every pair of neighboring dataset sequences $x\sim x^{\prime}$, the distributions of the server’s view—consisting of its random coins and the transcript of all exchanged messages—are $(\epsilon,\delta)$-indistinguishable when users hold the datasets in $x$ and $x^{\prime}$.
+Equivalence of a local protocol $P$ being LDP to a continual mechanism $\mathcal{M}_{P}$ being DP.
+Let $P$ be a local protocol as defined above, and let $\sim$ denote the neighboring relation on the sequences of users’ datasets.
+Define $f_{\sim}$ to output $\top$ only when it receives a single message of the form $(x,x^{\prime})$ where $x\sim x^{\prime}$.
+We construct a continual mechanism $\mathcal{M}_{P}$ that receives a dataset sequence as (a single) input, simulates the interaction of the server with the users in the protocol $P$, and returns the view of the server as output.
+(Indeed, $\mathcal{M}_{P}$ is a NIM, which is a special case of CMs.) We will show that $P$ is $(\epsilon,\delta)$-LDP w.r.t.
+$\sim$ if and only if $\mathcal{M}_{P}$ is $(\epsilon,\delta)$-DP w.r.t.
+$f_{\sim}$.
+We define $\mathcal{M}_{P}$ as
+$\mathcal{M}_{P}=\mathcal{S}_{P}\!\circ^{*}\!{\textrm{ExtConComp}},$
+where $\mathcal{S}_{P}$ is an IPM modeling the server’s behavior.
+The mechanism $\mathcal{S}_{P}$ is specifically defined as follows: $\mathcal{S}_{P}$ accepts only one left message of the form $x=(x_{1},\dots,x_{n})$, where each $x_{i}$ is a dataset for user $i$.
+Upon receiving $x$, $\mathcal{S}_{P}$ takes the following actions: (1) It creates $n$ user mechanisms by sending appropriate creation queries to ExtConComp, and (2) it sends each $x_{i}$ as the first message to the corresponding user mechanism through ExtConComp.
+(3) It then interacts with the user mechanisms according to the server’s interaction rules defined in protocol $P$.
+(4) Once the server decides to halt, $\mathcal{S}_{P}$ outputs the server’s view as its (left) response and then terminates.
+This view consists of the server’s internal randomness (i.e., random coins used by $\mathcal{S}_{P}$) and the transcript of all messages sent and received by the server.
+Thus, it does not include the creation queries or the initial dataset messages sent by $\mathcal{S}_{P}$.
+Claim 8.2.
+The local protocol $P$ is $(\epsilon,\delta)$-LDP w.r.t. the neighboring relation $\sim$ if and only if the CM $\mathcal{M}_{P}$ is $(\epsilon,\delta)$-DP w.r.t. the verification function $f_{\sim}$.
+Proof.
+The claim holds by the construction of $\mathcal{M}_{P}$ and Definition 8.1. ∎
+This equivalence allows us to analyze the privacy of the protocol $P$ by examining the privacy of the corresponding CM $\mathcal{M}_{P}$, which has a modular design, using Theorem 6.1.
+Remark 8.3.
+While the server in $P$ may use randomness, Theorem 6.1 requires $\mathcal{S}_{P}$ to be deterministic.
+This can be handled by defining $\mathcal{S}_{P}$ to request ExtConComp to create a $(0,0)$-DP mechanism that generates random coins or random samples whenever queried.
+Claim 8.4.
+The IPM $\mathcal{S}_{P}$ satisfies the destination and response properties defined in Section 6.1 and required by Theorem 6.1.
+Proof.
+By construction, upon receiving the left message $x$, the IPM $\mathcal{S}_{P}$ in $\mathcal{M}_{P}$ first sends exactly $2n$ messages to ExtConComp: $n$ creation queries and $n$ dataset messages.
+Although the content of these dataset messages depends on $x$, the decision of $\mathcal{S}_{P}$ sending a dataset message to its right mechanism, ExtConComp, is completely independent of $x$.
+(If the server is randomized, $\mathcal{S}_{P}$ additionally creates an auxiliary $(0,0)$-DP mechanism.
+This creation query to ExtConComp is also independent of $x$.)
+After these initial interactions, $\mathcal{S}_{P}$ proceeds to interact with the user mechanisms according to the server’s specification in $P$.
+The decision of whether to continue the interaction or to return a left response and halt is solely determined by the control flow of the server, which depends only on the users’ (and the auxiliary mechanism’s) responses and not directly on $x$.
+(The server has no access to $x$.) Hence, $\mathcal{S}_{P}$ satisfies the destination property.
+Moreover, the (single) left response of $\mathcal{S}_{P}$ is the server’s view, consisting of the messages exchanged with (the auxiliary mechanism and) the user mechanisms after their dataset is set.
+By design, this depends only on the privatized outputs of the mechanisms executed by ExtConComp.
+Therefore, $\mathcal{S}_{P}$ also satisfies the response property.
+∎
+To apply Theorem 6.1, it remains to show that for a verification function $f$, the IPM $\mathcal{S}_{P}$ satisfies the $f_{\sim}\to f$ property (see Section 6.1) and the $f$-concurrent composition of CMs is DP.
+We now propose a verification function $f$ for which the $f_{\sim}\to f$ property holds.
+Intuitively, this verification function enforces that all user mechanisms are first created, and that each is assigned a pair of datasets $(x_{i},x_{i}^{\prime})$ satisfying $x\sim x^{\prime}$.
+After this, queries $q$ can be asked concurrently from the users.
+Local protocols are typically defined so that the queries asked from each mechanism satisfy certain properties $\mathrm{prop}$, which play a key role in the privacy analysis.
+To later be able to show privacy guarantees for the $f$-concurrent composition of CMs, we also define $f$ to enforce the queries asked from each user to satisfy $\mathrm{prop}$.
+The formal definition of $f$ is given below.
+For convenience, we refer to the $i$-th created mechanism as user $i$.
+Definition 8.5 ($\sim_{i}$, $f_{\sim_{i},\mathrm{prop}_{i}}$, and $f_{P,\sim}^{\epsilon_{1},\delta_{1},\dots,\epsilon_{n},\delta_{n}}$).
+Let $P$ be a local protocol with $n$ users, where the queries issued by the server to each user $i\in[n]$ satisfy a property $\mathrm{prop}_{i}$.
+Let $\sim$ denote a neighboring relation on the sequences of users’ datasets.
+For every $i\in[n]$, define the neighbor relation $\sim_{i}$ on datasets of user $i$ as follows: for datasets $x_{i},x_{i}^{\prime}$, we have $x_{i}\sim_{i}x_{i}^{\prime}$ if and only if there exist datasets $x_{j},x_{j}^{\prime}$ for all $j\in[n]\setminus\{i\}$ such that $x\sim x^{\prime}$.
+(For example, assume $x$ represents a graph with $n$ vertices, where $x_{i}$ denotes the set of vertices adjacent to vertex $i$.
+Suppose $x\sim x^{\prime}$ iff the corresponding graphs of $x$ and $x^{\prime}$ differ by an edge.
+Then $x_{i}\sim_{i}x_{i}^{\prime}$ iff the adjacency sets $x_{i}$ and $x_{i}^{\prime}$ of vertex $i$ differ in the presence or absence of a single vertex.)
+Define the verification function $f_{\sim_{i},\mathrm{prop}_{i}}$ to accept a message sequence if and only if (1) the first message is a dataset pair $(x_{i},x_{i}^{\prime})$ with $x_{i}\sim_{i}x_{i}^{\prime}$, and (2) all subsequent messages are pairs of identical queries $(q,q)$, where the sequence of $q$ values satisfies $\mathrm{prop}_{i}$.
+For every $i\in[n]$, let $\mathcal{M}_{i}$ denote the CM associated with user $i$ and let $s_{i}$ be its single initial state.
+Assume $\mathcal{M}_{i}$ is $(\epsilon_{i},\delta_{i})$-DP w.r.t.
+$f_{\sim_{i},\mathrm{prop}_{i}}$.
+Then the verification function $f_{P,\sim}^{\epsilon_{1},\delta_{1},\dots,\epsilon_{n},\delta_{n}}$ is defined as follows: for every $t\in\mathbb{N}$ and message sequence $(m_{1},\dots,m_{t})$, we have $f_{P,\sim}^{\epsilon_{1},\delta_{1},\dots,\epsilon_{n},\delta_{n}}(m_{1},\dots,m_{t})=\top$ if and only if
+  1. (i)
+    the sequence $(m_{1},\dots,m_{t})$ is suitable; and
+  2. (ii)
+    it is a prefix of a sequence consisting of the following messages, in order:
+    1. (a)
+      pairs of identical creation queries $(\alpha_{i},\alpha_{i})$ for each user $i\in[n]$, where $\alpha_{i}=(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{\sim_{i},\mathrm{prop}_{i}})$ (and possibly an additional pair for an auxiliary sample-generating mechanism if the server is randomized);
+    2. (b)
+      dataset-assignment messages $((x_{i},i),(x_{i}^{\prime},i))$ for each user $i$, where $x\sim x^{\prime}$; and
+    3. (c)
+      arbitrarily many identical query messages $((q,i),(q,i))$ directed to the same user $i$.
+Note that condition (i) ensures that the query sequence for each user $i$ remains valid with respect to $f_{\sim_{i},\mathrm{prop}_{i}}$.
+Claim 8.6.
+The IPM $\mathcal{S}_{P}$ satisfies the $f_{\sim}\to f_{P,\sim}^{\epsilon_{1},\delta_{1},\dots,\epsilon_{n},\delta_{n}}$ property.
+Proof.
+By definition, every $f_{\sim}$-valid message sequence consists of a single message of the form $(x,x^{\prime})$ with $x\sim x^{\prime}$.
+Consider two executions of $\mathcal{S}_{P}$ receiving $x$ and $x^{\prime}$ as their respective left messages.
+By definition, both IPMs first identically create user mechanisms.
+Then one instance assigns datasets $x_{i}$ to the users by sending the right messages $(x_{i},i)$, while the other assigns datasets $x_{i}^{\prime}$ by sending $(x_{i}^{\prime},i)$.
+Once all datasets have been assigned, the future behavior of both instances becomes independent of their left inputs—their actions are determined solely by the server description in $P$, which has no access to the datasets and depends only on the (privatized) responses returned by the user mechanisms.
+Hence, if both instances of $\mathcal{S}_{P}$ receive identical right messages (from user mechanisms), they generate identical right queries to the same users, i.e., identical messages of the form $(q,i)$.
+Consequently, the sequence $\sigma$ formed by pairing the right messages produced by these two instances satisfies condition (ii) in Definition 8.5.
+We will show that this sequence also satisfies condition (i) in this definition, and thus it is $f_{P,\sim}$-valid.
+Consequently, $\mathcal{S}_{P}$ satisfies the $f_{\sim}\to f_{P,\sim}$ property.
+For condition (i), we recall that a message sequence is suitable if each message is either (i) a pair of identical creation queries $(\mathcal{M},s,\epsilon,\delta,f)^{2}$, where the CM $\mathcal{M}$ with single initial state $s$ is $(\epsilon,\delta)$-DP w.r.t.
+$f$, or (ii) a message of the form $\big((q_{0},\ell),(q_{1},\ell)\big)$, where at least $\ell$ mechanisms have already been created and $q_{0},q_{1}$ are valid inputs for the $\ell$-th one.
+Let $\mathcal{M}_{\ell}$ denote the $\ell$-th created mechanism and let $f_{\ell}$ be its associated verification function.
+The suitability condition also requires that the sequence of query pairs $(q_{0},q_{1})$ extracted from messages of the form $\big((q_{0},\ell),(q_{1},\ell)\big)$ be $f_{\ell}$-valid.
+By construction, the paired sequence $\sigma$ satisfies these conditions and is therefore suitable, finishing the proof.
+∎
+Therefore, by Claim 8.4 and Claim 8.6, Theorem 6.1 is applicable. This theorem combined with Claim 8.2 implies:
+Theorem 8.7.
+Let $P$ be a local protocol with $n$ users, where the queries asked by the server of each user $i\in[n]$ satisfy a property $\mathrm{prop}_{i}$.
+Let $\sim$ be a neighbor relation on sequences of $n$ user datasets, and for each $i\in[n]$, define the verification function $f_{\sim_{i},\mathrm{prop}_{i}}$ as in Definition 8.5.
+Suppose that for parameters $\epsilon_{1},\dots,\epsilon_{n},\epsilon\geq 0$ and $0\leq\delta_{1},\dots,\delta_{n},\delta\leq 1$, the following hold:
+  1. (i)
+    for every $i\in[n]$, the CM corresponding to user $i$ is $(\epsilon_{i},\delta_{i})$-DP w.r.t. $f_{\sim_{i},\mathrm{prop}_{i}}$; and
+  2. (ii)
+    the $f_{P,\sim}^{\epsilon_{1},\delta_{1},\dots,\epsilon_{n},\delta_{n}}$-concurrent composition of CMs is $(\epsilon,\delta)$-DP.
+Then the protocol $P$ is $(\epsilon,\delta)$-LDP w.r.t. $\sim$.
+We next simplify this result for a choice of the neighboring relation $\sim$ based on a shared neighbor relation for the users.
+Let $\sim^{*}$ denote a neighbor relation on (individual) users’ datasets.
+For $k\in\mathbb{N}$, define $\sim$ on sequences of $n$ user datasets as follows: For $x,x^{\prime}$, we have $x\sim x^{\prime}$ if and only if they differ in at most $k$ indices, and for each differing index $i$, we have $x_{i}\sim^{*}x_{i}^{\prime}$.
+We refer to $\sim$ as the $k$-sparse parallel-composition neighbor relation with user-neighbor relation $\sim^{*}$.
+We note that the relation $\sim^{*}$ is implicitly assumed to be reflexive, meaning each dataset is related to itself.
+Hence, the neighbor relation $\sim_{i}$ associated to each user $i$ in Definition 8.5 equals $\sim^{*}$.
+Corollary 8.8.
+Let $P$ be a local protocol with $n$ users following identical procedures, and let $\sim^{*}$ be a neighbor relation on individual users’ datasets.
+For $k\in\mathbb{N}$, let $\sim_{k,\sim^{*}}$ denote the $k$-sparse parallel-composition relation with user-neighbor relation $\sim^{*}$, defined above.
+Assume that in protocol $P$, the queries asked by the server of each user satisfy a property $\mathrm{prop}^{*}$.
+Let $\epsilon^{\prime}>0$, $\epsilon\geq 0$, and $0\leq\delta^{\prime},\delta\leq 1$.
+Suppose the following conditions hold:
+  • The CM corresponding to each user is $(\epsilon^{\prime},\delta^{\prime})$-DP w.r.t. the verification function $f_{\sim^{*},\mathrm{prop}^{*}}$ from Definition 8.5.
+  • The composition of $k$ instances of the randomized response mechanism $\mathrm{RR}_{\epsilon^{\prime},\delta^{\prime}}$ is $(\epsilon,\delta)$-DP.
+Then the protocol $P$ is $(\epsilon,\delta)$-LDP w.r.t. $\sim$.
+To prove Corollary 8.8, we need the following observation:
+Observation 8.9.
+Let $f_{1}$ and $f_{2}$ be verification functions such that every message sequence that is $f_{1}$-valid is also $f_{2}$-valid.
+If a CM $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t.
+$f_{2}$, then $\mathcal{M}$ is also $(\epsilon,\delta)$-DP w.r.t.
+$f_{1}$.
+Proof.
+By assumption, the IPMs $\mathcal{V}[f_{1}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ and $\mathcal{V}[f_{1}]\!\circ^{*}\!\mathcal{V}[f_{2}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ are equivalent.
+Hence, by the post-processing lemma, the claim follows immediately.
+∎
+Proof of Corollary 8.8.
+Recall the verification function $f^{(\epsilon^{\prime},\delta^{\prime})_{i=1}^{k}}_{\infty,\mathit{FPC}}$ from Definition 5.13.
+This verification function permits the creation of an arbitrary number of CMs over time, but guarantees that at most $k$ of them—each with privacy parameters $(\epsilon^{\prime},\delta^{\prime})$—are assigned pairs of non-identical inputs.
+Moreover, it enforces that if the first message directed to a mechanism contains identical inputs, then all subsequent messages to that mechanism must also contain identical inputs.
+Comparing the verification functions $f_{P,\sim}^{(\epsilon^{\prime},\delta^{\prime})_{i=1}^{n}}$ (Definition 8.5) and $f^{(\epsilon^{\prime},\delta^{\prime})_{i=1}^{k}}_{\infty,\mathit{FPC}}$, we observe that every $f_{P,\sim}^{(\epsilon^{\prime},\delta^{\prime})_{i=1}^{n}}$-valid message sequence is also $f^{(\epsilon^{\prime},\delta^{\prime})_{i=1}^{k}}_{\infty,\mathit{FPC}}$-valid.
+Applying Observation 8.9 with $\mathcal{M}={\textrm{ExtConComp}}$, we conclude that if the $f^{(\epsilon^{\prime},\delta^{\prime})_{i=1}^{k}}_{\infty,\mathit{FPC}}$-concurrent composition of CMs is $(\epsilon,\delta)$-DP, then so is the $f_{P,\sim}^{(\epsilon^{\prime},\delta^{\prime})_{i=1}^{n}}$-concurrent composition of CMs.
+By Theorem 5.14, if the composition of $k$ randomized response mechanisms $\mathrm{RR}_{\epsilon^{\prime},\delta^{\prime}}$ is $(\epsilon,\delta)$-DP, then the $f^{(\epsilon^{\prime},\delta^{\prime})_{i=1}^{k}}_{\infty,\mathit{FPC}}$-concurrent composition of CMs is also $(\epsilon,\delta)$-DP.
+Hence, under the second assumption of Corollary 8.8, the $f_{P,\sim}^{(\epsilon^{\prime},\delta^{\prime})_{i=1}^{n}}$-concurrent composition is $(\epsilon,\delta)$-DP.
+Furthermore, by the definition of $\sim$, for each $i\in[n]$, the neighbor relation $\sim_{i}$ in Definition 8.5 coincides with $\sim^{*}$.
+Thus, the first assumption of the corollary is equivalent to requiring that each user’s CM is $(\epsilon_{i},\delta_{i})$-DP w.r.t.
+$f_{\sim_{i},\mathrm{prop}_{i}}$.
+Thus, applying Theorem 8.7 yields that the protocol $P$ is $(\epsilon,\delta)$-LDP w.r.t.
+$\sim$, completing the proof.
+∎
+Remark 8.10.
+In real-world deployments, users may join or leave the system dynamically, and data records may be shared across a subset of users.
+Furthermore, a user’s local dataset may evolve over time.
+To cover these scenarios, users are CMs rather than IMs, and in the above discussions, instead of $\mathcal{M}_{P}$ receiving only a single message $x=(x_{1},\dots,x_{n})$, this mechanism receives creation queries for creating a new user, deletion messages informing that a specific user is gone, and messages of the form $(u,j)$ requesting $\mathcal{M}_{P}$ to forward the dataset-update $u$ to user $j$.
+Without repeating the whole argument, we note that Theorem 8.7 and Corollary 8.8 can be adapted and applied in this general scenario and our concurrent composition theorems for CMs (rather than IMs) can be used to simplify the privacy analysis of these local protocols.
+Remark 8.11.
+In this section, we assumed that all components (i.e., the server and the users) are honest, meaning that they follow the local protocol.
+We also assumed that there is no collusion, that is, parties do not share their internal states with each other.
+Suppose now that $k$ users hold private data, while the remaining nkn-k users and the server may be dishonest and collude to learn this data.
+We can model the colluding parties as a single adversary that interacts concurrently with the $k$ target users.
+In this case, privacy of the local protocol follows from the concurrent composition of the IMs corresponding to these $k$ users.
+If the users’ datasets evolve over time, then the concurrent composition of CMs must be used instead.
+9. Application (Core Decomposition via Multi-Dimensional SVT)
+In this section, we review the local protocol of Dhulipala, Henzinger, Li, Liu, Sricharan, and Zhu (dhulipala2025near, ) and demonstrate how its privacy analysis can be simplified using our concurrent parallel composition theorem for IMs.
+They study the core decomposition problem in the LDP model, where each user corresponds to a (publicly known) vertex of a graph and holds the set of adjacent vertices as its private dataset.
+A core decomposition protocol is said to satisfy $(\epsilon,\delta)$-edge-LDP if it is $(\epsilon,\delta)$-LDP w.r.t.
+the neighbor relation $\sim$, where two sequences of user datasets are considered neighbors whenever the corresponding graphs differ by at most one edge.
+The core decomposition problem.
+Let $G=(V,E)$ be a graph, where $V(G)$ and $E(G)$ denote its vertex and edge sets, respectively.
+For a subset $U\subseteq V(G)$, the subgraph of $G$ induced by $U$ is denoted by $G[U]$.
+The degree of a vertex $v\in V(G)$ is written as $\deg_{G}(v)$.
+The core number of a vertex $v$, denoted $k_{G}(v)$, is the largest integer $k$ such that $v$ belongs to a subgraph $H\subseteq G$ in which every vertex $u\in V(H)$ satisfies $\deg_{H}(u)\geq k$.
+The goal of the core decomposition problem is to compute the coreness vector
+$\vec{k}_{G}=(k_{G}(v):v\in V),$
+which assigns to each vertex its coreness value.
+Overview of the Local Protocol by (dhulipala2025near, ).
+Dhulipala et al.
+(dhulipala2025near, ) propose a local protocol $P$ for the core decomposition problem, inspired by the classical peeling algorithm.
+Let $G=(V,E)$ be a graph stored by $n=|V|$ users: each user corresponds to exactly one vertex $v\in V$ and stores a private dataset $N_{v}$, defined as
+$N_{v}=\left\{u\in V\mid\{v,u\}\in E\right\}.$
+The server knows only the vertex set $V$, while each user $v$ knows both $V$ and its private set of neighbors $N_{v}$. Next, we will describe the server and user mechanisms in $P$.
+Server.
+The server maintains four variables: (1) the current core value $d$, initialized to $1$; (2) the active vertex set $A$, initialized to $V$; and (3) the estimated coreness vector $\vec{k}$, initialized to $\vec{0}\in\mathbb{R}^{n}$.
+The server repeats the following procedure $n$ times:
+  1. (1)
+    Send the current active vertex set $A$ and core value $d$ to all users $v\in A$ and ask whether their number of adjacent vertices in $A$ is less than $d$, i.e., $|N_{v}\cap A|<d$.
+  2. (2)
+    Let $S$ denote the subset of vertices in $A$ responding with $\top$, indicating that $|N_{v}\cap A|<d$. Update $\vec{k}$ by assigning $k_{v}=d$ for all $v\in S$.
+  3. (3)
+    Remove the vertices in $S$ from $A$.
+Users.
+For each $v\in V$, let $\mathcal{M}_{v}$ denote the user IM corresponding to $v$.
+This mechanism is a variant of the Sparse Vector Technique (SVT) that holds $N_{v}$ as its dataset and maps inputs (A,d)(A,d)—where $A\subseteq V$ and $d\in\mathbb{N}$—to either $\top$ or $\bot$, indicating whether $|N_{v}\cap A|<d$.
+Two datasets are neighbors if they differ in at most one vertex, i.e., the adjacency set of $v$ changes in at most one element.
+The integer $d$ is a threshold, and the set $A$ represents a function $f_{A}(\cdot)=|\cdot\cap A|$, which is $1$-sensitive because it changes by at most one on neighboring vertex sets.
+Modeling $\mathcal{M}_{v}$ as a continual mechanism, differential privacy for $\mathcal{M}_{v}$ is defined with respect to a verification function $g$ that requires the first message to be a pair of identical or neighboring vertex sets, and every subsequent message to be a pair of identical vertex sets (representing $1$-sensitive functions).
+We say that $\mathcal{M}_{v}$ is $(\epsilon,\delta)$-DP if it satisfies $(\epsilon,\delta)$-DP with respect to $g$ against adaptive adversaries.
+Lemma 9.1 ((dhulipala2025near, )).
+Consider the local protocol $P$ for the core decomposition problem described above.
+If every user (SVT) mechanism satisfies $(\epsilon/2,0)$-DP, then the protocol $P$ satisfies $(\epsilon,0)$-LDP.
+Moreover, for a graph $G=(V,E)$ with $n=|V|$ and maximum degree $\Delta$, the following holds with high probability: for every vertex $v\in V$, the coreness value estimated by the server is in the range $k_{G}(v)\pm O(\log n\log\Delta)$ where $k_{G}(v)$ is the true core number of $v$.
+To prove the privacy part of this lemma, Dhulipala et al.
+(dhulipala2025near, ) analyze the privacy of a mechanism called Multi-Dimensional SVT, which generalizes SVT by mapping, at each step, a vector of $n$ 11-sensitive functions and a vector of $n$ thresholds to an element of $\{\top,\bot\}^{n}$.
+Their analysis is from scratch and technically involved.
+We observe that multi-dimensional SVT is a special case of the concurrent parallel composition of $n$ SVT mechanisms, where the adversary must issue queries to the SVT instances in a round-robin pattern.
+So we can simplify their analysis using Corollary 5.15 and Theorem 6.1.
+However, to illustrate the use of Corollary 8.8, we instead analyze the entire local protocol directly.
+Simplified Privacy Analysis.
+Recall that in the local model, a graph with public vertex set $V$ of size $n=|V|$ is represented as a sequence of adjacency sets $(N_{v})_{v\in V}$, where each $N_{v}$ corresponds to the dataset of the user mechanism associated with vertex $v$.
+Let $\sim^{*}$ denote the neighbor relation for user SVT mechanisms.
+(For every $N,N^{\prime}\subseteq V$, we have $N\sim^{*}N^{\prime}$ iff the sets differ in at most one vertex.) Let $\sim_{2,\sim^{*}}$ denote the $2$-sparse parallel-composition neighbor relation with user-neighbor relation $\sim^{*}$.
+By definition, the local protocol $P$ of (dhulipala2025near, ) satisfies $(\epsilon,\delta)$-edge-LDP if, for every two sequences of $n$ (adjacency) vertex sets whose graphs differ in exactly one edge, the server’s views are $(\epsilon,\delta)$-indistinguishable.
+Since such sequences differ in at most two adjacency sets, it suffices to show that $P$ is $(\epsilon,\delta)$-LDP w.r.t.
+$\sim_{2,\sim^{*}}$.
+By construction, functions corresponding to the vertex sets $A$ sent by the server are all $1$-sensitive; let this property be $\mathrm{prop}^{*}$.
+Hence, each user’s SVT mechanism is $(\epsilon/2,0)$-DP w.r.t.
+the verification function $f_{\sim^{*},\mathrm{prop}^{*}}$ from Definition 8.5.
+By basic composition, the composition of two instances of $\mathrm{RR}_{\epsilon/2,0}$ is $(\epsilon,0)$-DP.
+Therefore, by Corollary 8.8, the protocol $P$ is $(\epsilon,0)$-LDP w.r.t.
+$\sim_{2,\sim^{*}}$, and therefore $(\epsilon,0)$-edge-LDP.
 10. Reduction to Interactive Randomized Response
-Lyu [19] shows that for every $(\epsilon,\delta)$-DP IM $\mathcal{M}$ and every pair of neighboring initial states $s_{0},s_{1}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ are equivalent.
-Meaning that, $\mathcal{P}$ can simulate $\mathcal{M}(s_{b})$ given access only to the output of a randomized response mechanism $\mathrm{RR}_{\epsilon,\delta}(b)$.
+Lyu (lyu2022composition, ) shows that for every $(\epsilon,\delta)$-DP IM $\mathcal{M}$ and every pair of neighboring initial states $s_{0},s_{1}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ are equivalent.
+This means that $\mathcal{P}$ can simulate $\mathcal{M}(s_{b})$ given access only to the output of a randomized response mechanism $\mathrm{RR}_{\epsilon,\delta}(b)$.
 In this section, inspired by this work, we construct a new IPM $\mathcal{P}$ that, rather than simulating $\mathcal{M}$ based on the outcome of a single instance of $\mathrm{RR}_{\epsilon,\delta}$, operates on the output of $\mathrm{RR}_{0,\delta}$ and only when necessary uses the output of $\mathrm{RR}_{\epsilon,0}$.
@@
 Our result is formalized in Lemma 5.6, which is restated below:
-Lemma (Restatement of Lemma 5.6).
+Lemma 0 (Restatement of Lemma 5.6).
 For $\epsilon>0$ and $0\leq\delta\leq 1$, let $\mathcal{M}:S_{\mathcal{M}}\times Q_{\mathcal{M}}\to S_{\mathcal{M}}\times A_{\mathcal{M}}$ be an $(\epsilon,\delta)$-DP IM w.r.t.
@@
 Otherwise, $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}(b)$ at most twice.
-Notation 8.1.
+Notation 10.1.
 For short, let $Q=Q_{\mathcal{M}}$ and $A=A_{\mathcal{M}}$.
 For each $t\in\mathbb{N}$ and $b\in\{0,1\}$, we define the function $\mu_{t}^{b}:(Q\times A)^{t}\to[0,1]$ such that $\mu_{t}^{b}(q_{1},\dots,a_{t})$ denotes the probability of $\mathcal{M}(s_{b})$ returning answers $a_{1},\dots,a_{t}$ to the queries $q_{1},\dots,q_{t}$.
-Fixing an upper bound $T$ on the number of queries, Lyu [19] reduces the construction of $\mathcal{P}$ to defining a sequence of functions $\phi_{t}^{0}$ and $\phi_{t}^{1}$ for $t\in[T]$ that satisfy three conditions.
+Fixing an upper bound $T$ on the number of queries, Lyu (lyu2022composition, ) reduces the construction of $\mathcal{P}$ to defining a sequence of functions $\phi_{t}^{0}$ and $\phi_{t}^{1}$ for $t\in[T]$ that satisfy three conditions.
 They then inductively construct $\phi_{t}^{b}$ and verify that it satisfies the required conditions.
@@
 In Section 10.1.2, we show some properties for $\mathrm{L}_{t}^{b}$, required in later proofs.
-In Section 8.2, we summarize the reduction introduced in [19], with the control functions $\mathrm{L}_{t,T}^{b}$ replaced by $\mathrm{L}_{t}^{b}$.
+In Section 10.2, we summarize the reduction introduced in (lyu2022composition, ), with the control functions $\mathrm{L}_{t,T}^{b}$ replaced by $\mathrm{L}_{t}^{b}$.
 We also modify this reduction by adding a new condition to the original three that the functions $\phi_{t}$ must satisfy.
@@
 10.1. Control Functions $\mathrm{Lower}_{t,T}$ and $\mathrm{L}_{t,T}$.
-Lyu [19] introduces the following control function:
+Lyu (lyu2022composition, ) introduces the following control function:
 Definition 10.2 ($\mathrm{Lower}_{t,T}$).
@@
 $\mathrm{Lower}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t})=\sum_{a_{t}\in A}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t}).$
-Lyu [19] shows the following two claims about $\mathrm{Lower}_{t,T}^{b}$:
-Claim 8.5 ([19]).
-For each $b\in\{0,1\}$, $T\in\mathbb{N}$, and $q_{1}\in Q$ $\mathrm{Lower}_{1,T}^{b}(q_{1})\leq\delta$.
-Claim 8.6 ([19]).
-For each $b\in\{0,1\}$, $T\in\mathbb{N}$, $t\in[T]$, [^7] In the original claim in [19], $t$ is assumed to be at most T1T-1, which is a typo.
+Lyu (lyu2022composition, ) shows the following two claims about $\mathrm{Lower}_{t,T}^{b}$:
+Claim 10.5 ((lyu2022composition, )).
+For each $b\in\{0,1\}$, $T\in\mathbb{N}$, and $q_{1}\in Q$, we have $\mathrm{Lower}_{1,T}^{b}(q_{1})\leq\delta$.
+Claim 10.6 ((lyu2022composition, )).
+For each $b\in\{0,1\}$, $T\in\mathbb{N}$, $t\in[T]$, [^13] In the original claim in (lyu2022composition, ), $t$ is assumed to be at most T1T-1, which is a typo.
 and transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$, we have
@@
 Corollary 10.7.
-For each $b\in\{0,1\}$, and $T\in\mathbb{N}$, $\mathrm{L}_{0,T}^{b}()\leq\delta$.
-In some part of the proof, Lyu [19] wrongly refers to Claim 8.6, while the following lemma is required. For completeness, we prove this lemma and use it in the next section.
+For each $b\in\{0,1\}$ and $T\in\mathbb{N}$, we have $\mathrm{L}_{0,T}^{b}()\leq\delta$.
+In some part of the proof, Lyu (lyu2022composition, ) wrongly refers to Claim 10.6, while the following lemma is required. For completeness, we prove this lemma and use it in the next section.
 Lemma 10.8.
@@
 This section introduces our main new technical contribution.
-To remove the assumption of having an upper bound $T$ on the number of queries, we introduce $\mathrm{L}_{t}^{b}$ as limit of $\mathrm{L}_{t,T}^{b}$ and show that it satisfies all properties needed by the proof of [19].
-Thus, differently from [19], we will use $\mathrm{L}_{t}^{0}$ and $\mathrm{L}_{t}^{1}$ to define the function $\phi_{t}$.
+To remove the assumption of having an upper bound $T$ on the number of queries, we introduce $\mathrm{L}_{t}^{b}$ as the limit of $\mathrm{L}_{t,T}^{b}$ and show that it satisfies all properties needed by the proof of (lyu2022composition, ).
+Thus, differently from (lyu2022composition, ), we will use $\mathrm{L}_{t}^{0}$ and $\mathrm{L}_{t}^{1}$ to define the function $\phi_{t}$.
 Definition 10.9 ($\mathrm{L}_{t}^{b}$).
@@
 We will upper bound the values of $\mathrm{L}_{t,T}^{b}$ by $1$ and show that this function is non-decreasing in $T$.
-As a result, for each each $t\in\mathbb{N}\cup\{0\}$ and $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the limit $\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$ exists.
+As a result, for each $t\in\mathbb{N}\cup\{0\}$ and $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the limit $\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$ exists.
 We need the following lemma throughout the section:
@@
 $\mathrm{Lower}_{T,T}^{b}(q_{1},\dots,a_{T-1},q_{T})=\sum\limits_{a_{T}\in A}\max\{0,\mu_{T}^{b}(q_{1},\dots,a_{T})-e^{\epsilon}\mu_{T}^{1-b}(q_{1},\dots,a_{T})\}$
-$=\sum\limits_{a_{T}\in A}\sup_{q_{T+1}\in Q}\max\{0,\sum\limits_{a_{T+1}\in A}\mu_{T+1}^{b}(q_{1},\dots,,a_{T+1})-e^{\epsilon}\mu_{T+1}^{1-b}(q_{1},\dots,a_{T+1})\}$
+$=\sum\limits_{a_{T}\in A}\sup_{q_{T+1}\in Q}\max\{0,\sum\limits_{a_{T+1}\in A}\mu_{T+1}^{b}(q_{1},\dots,a_{T+1})-e^{\epsilon}\mu_{T+1}^{1-b}(q_{1},\dots,a_{T+1})\}$
 $\leq\sum\limits_{a_{T}\in A}\sup_{q_{T+1}\in Q}\sum\limits_{a_{T+1}\in A}\max\{0,\mu_{T+1}^{b}(q_{1},\dots,a_{T+1})-e^{\epsilon}\mu_{T+1}^{1-b}(q_{1},\dots,a_{T+1})\}$
@@
 In this section, we present the properties of $\mathrm{L}_{t}^{b}$, required in the next sections.
-The definition of $\mathrm{L}_{t}^{b}$ combined with Corollary 8.7, Lemma 8.8, and Lemma 8.12 give the following corollaries, respectively:
+The definition of $\mathrm{L}_{t}^{b}$ combined with Corollary 10.7, Lemma 10.8, and Lemma 10.12 gives the following corollaries, respectively:
 Corollary 10.15.
-For each $b\in\{0,1\}$, $\mathrm{L}_{0}^{b}()\leq\delta$.
+For each $b\in\{0,1\}$, we have $\mathrm{L}_{0}^{b}()\leq\delta$.
 Corollary 10.16.
@@
 10.2. Reduction.
-To define $\mathcal{P}$ and simulate $\mathcal{M}$, Lyu [19] constructs four families of probability mass functions (PMFs), corresponding to each outcome of $\mathrm{RR}_{\epsilon,\delta}$.
-Intuitively, each family describes an IM.
-Specifically, each family describes an IM and consists of PMFs over the answer set $A$ for every $t\in[T]$ and every history $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t}\times Q$.
+To define $\mathcal{P}$ and simulate $\mathcal{M}$, Lyu (lyu2022composition, ) constructs four families of probability mass functions (PMFs), corresponding to each outcome of $\mathrm{RR}_{\epsilon,\delta}$.
+Intuitively, each family describes an IM and consists of PMFs over the answer set $A$ for every $t\in[T]$ and every history $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t}\times Q$.
 These PMFs represent the probability of returning each answer in response to $q_{t}$, conditioned on $q_{1},\dots,a_{t-1}$ being the history of queries and answers.
@@
 We note that in our construction, each family includes a PMF over $A$ for every $t\in\mathbb{N}$ and every history $(q_{1},\dots,a_{t-1},q_{t})$.
-Moreover, the definitions of these PMFs are not the same as the ones in [19].
+Moreover, the definitions of these PMFs are not the same as the ones in (lyu2022composition, ).
 10.2.1. Constructing $\mathcal{P}$.
-In [19], the post-processing mechanism initially interacts with $\mathrm{RR}_{\epsilon,\delta}$ and stores its output $r=(\tau,c)\in\{\top,\bot\}\times\{0,1\}$.
+In (lyu2022composition, ), the post-processing mechanism initially interacts with $\mathrm{RR}_{\epsilon,\delta}$ and stores its output $r=(\tau,c)\in\{\top,\bot\}\times\{0,1\}$.
 It then answers the adversary’s queries by itself using the family of PMFs corresponding to $r$.
@@
 To generate a response $a$ to query $q$, $\mathcal{P}$ proceeds as follows: If $q$ is the first query (i.e., $r=()$), then $\mathcal{P}$ initially interacts with $\mathrm{IRR}_{\epsilon,\delta}$ and appends its response to $r$.
-After this, whether $q$ is the first query or not, when the following condition does not hold for the first time, $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$ for the second time and appends its answer to $r$: the PMFs corresponding to $(h,q)$ in $\mathcal{F}_{(\top,0)}$ and $\mathcal{F}_{(\top,1)}$ are identical, and the PMFs corresponding to $(h,q)$ in $\mathcal{F}_{(\bot,0)}$ and $\mathcal{F}_{(\bot,1)}$ are identical.
+After this, whether $q$ is the first query or not, the first time the following condition does not hold, $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$ for the second time and appends its answer to $r$: the PMFs corresponding to $(h,q)$ in $\mathcal{F}_{(\top,0)}$ and $\mathcal{F}_{(\top,1)}$ are identical, and the PMFs corresponding to $(h,q)$ in $\mathcal{F}_{(\bot,0)}$ and $\mathcal{F}_{(\bot,1)}$ are identical.
 After that, $\mathcal{P}$ never interacts with $\mathrm{IRR}_{\epsilon,\delta}$ again.
@@
 10.2.2. Constructing $\mathcal{F}_{(\bot,b)}$s.
-Using the control function $\mathrm{Lower}_{t,T}$, Lyu [19] introduces three conditions and constructs a sequence of functions $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for $t\in\{1,\dots,T\}$ that satisfy these conditions.
+Using the control function $\mathrm{Lower}_{t,T}$, Lyu (lyu2022composition, ) introduces three conditions and constructs a sequence of functions $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for $t\in\{1,\dots,T\}$ that satisfy these conditions.
 The families of PMFs are defined then based on these functions.
 We remove the dependence on the bound $T$ by replacing $\mathrm{Lower}_{t,T}$ with $\mathrm{L}_{t}$ in the conditions, and extend the construction of $\phi_{t}$ to all $t\in\mathbb{N}$.
-Furthermore, we introduce an additional condition to those given in [19], which we will later use to bound the number of times $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$.
+Furthermore, we introduce an additional condition to those given in (lyu2022composition, ), which we will later use to bound the number of times $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$.
 Let $\phi_{t}^{0}$ and $\phi_{t}^{1}$ denote the functions that return the first and second components, respectively, of the output of $\phi_{t}$.
@@
 10.2.3. Constructing $\mathcal{F}_{(\top,b)}$s.
-Similar to the previous section, $\mathcal{F}_{(0,\top)}$ and $\mathcal{F}_{(1,\top)}$ are defined based on a sequence of functions $\psi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for each $t\in\mathbb{N}\cup\{0\}$.
+Similar to the previous section, $\mathcal{F}_{(\top,0)}$ and $\mathcal{F}_{(\top,1)}$ are defined based on a sequence of functions $\psi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for each $t\in\mathbb{N}\cup\{0\}$.
 For $t=0$, $\psi_{t}()=(1,1)$.
@@
 10.3. Construction of $\phi_{t}$
-For each $q_{1},\dots,a_{t-1},q_{t}$, Lyu [19] initializes the values of $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ with the ones from $\mathrm{L}_{t,T}(q_{1},\dots,a_{t-1},q_{t},\cdot)$.
+For each $q_{1},\dots,a_{t-1},q_{t}$, Lyu (lyu2022composition, ) initializes the values of $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ with the ones from $\mathrm{L}_{t,T}(q_{1},\dots,a_{t-1},q_{t},\cdot)$.
 (Removing the upper bound assumption, they are initialized by the values of $\mathrm{L}_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$.) Then, they iterate over all possible answers $a_{t}\in A$ in a fixed order and gradually increase the value of $\phi_{t}(q_{1},\dots,a_{t})$ to satisfy conditions (II) and (III), while ensuring that condition (I) remains valid.
@@
 For any answer $a\in A$, $\Gamma_{<a}=\{a^{\prime}\in A\mid a^{\prime}<_{A}a\}$ and $\Gamma_{>a}=\{a^{\prime}\in A\mid a<_{A}a^{\prime}\}$.
-Lemma 8.25 (Transfinite Recursion [22]).
+Lemma 10.25 (Transfinite Recursion (suppes2012axiomatic, )).
 Let $f$ be a function recursively defined over $A$ according to the well-order $_{A}$.
@@
 Combination of Constraints ($\phi$.1) and ($\xi$.1) gives condition (I).
-Moreover, Constraint ($\phi$.3) directly imply condition (II).
-The proof of the following claim is identical to a proof in [19]:
+Moreover, Constraint ($\phi$.3) directly implies condition (II).
+The proof of the following claim is identical to a proof in (lyu2022composition, ):
 Claim 10.31.
-$\phi_{t}$ satisfies conditions (III).
+$\phi_{t}$ satisfies condition (III).
 Proof.
@@
 $\sum_{a_{t}\in A}\phi_{t}^{1-b}(q_{1},\dots,a_{t})$ $-\phi_{t-1}^{1-b}(q_{1},\dots,a_{t-1})\geq e^{-\epsilon}(\sum_{a_{t}\in A}\phi_{t}^{b}(q_{1},\dots,a_{t})-\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})),$
-which is a contradiction when $\epsilon>0$, which implies that condition (III) holds. ∎
+which is a contradiction when $\epsilon>0$, implying that condition (III) holds. ∎
 To show $\phi_{t}$ satisfies conditions (IV), we need the following lemma:
@@
 for all $(a_{1},\dots,a_{t})\in A^{t}$.
-Moreover, by Lemma 8.10 and $\star$8.3.4, $\mu_{t-1}^{0}(q_{1},\dots,a_{t-1})=\mu_{t-1}^{1}(q_{1},\dots,a_{t-1})$ for all $(a_{1},\dots,a_{t})\in A^{t}$.
+Moreover, by Lemma 10.10 and $\star$10.3.4, $\mu_{t-1}^{0}(q_{1},\dots,a_{t-1})=\mu_{t-1}^{1}(q_{1},\dots,a_{t-1})$ for all $(a_{1},\dots,a_{t})\in A^{t}$.
 Thus, by Remark 10.21 and condition (IV),
 ($\star\star\star$) $\phi_{t-1}^{0}(q_{1},\dots,a_{t-1})=\phi_{t-1}^{1}(q_{1},\dots,a_{t-1}).$
-Therefore, by equations ($\star$8.3.4), ($\star\star$8.3.4) and ($\star\star\star$8.3.4), Constraints ($\phi$.1), ($\phi$.2), and ($\phi$.3) are symmetric for $x_{0}$ and $x_{1}$.
+Therefore, by equations ($\star$10.3.4), ($\star\star$10.3.4) and ($\star\star\star$10.3.4), Constraints ($\phi$.1), ($\phi$.2), and ($\phi$.3) are symmetric for $x_{0}$ and $x_{1}$.
 Hence, there exists a maximal feasible solution $(z^{*},z^{*})$ satisfying all the constraints.
@@
 In this section, we prove:
-Lemma (Restatement of Lemma 5.9).
+Lemma 0 (Restatement of Lemma 5.9).
 For every $0\leq\delta\leq 1$, the $f^{\delta}_{\mathrm{RR}}$-concurrent composition of continual mechanisms is $(0,\delta)$-differentially private.
@@
 Vadhan et al.
-[24] show that to guarantee an IM is differentially private, it suffices to consider only deterministic adversaries.
+(vadhan2022concurrent, ) show that to guarantee an IM is differentially private, it suffices to consider only deterministic adversaries.
 Specifically, the following holds:
-Lemma 9.1 ([24]).
+Lemma 11.1 ((vadhan2022concurrent, )).
 An IM $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t.
-a neighboring relation $\sim$ if and and only if for every deterministic adversary $\mathcal{A}$ and every pair of neighboring initial states $s\sim s^{\prime}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}(s^{\prime}))$ are $(\epsilon,\delta)$-indistinguishable.
+a neighboring relation $\sim$ if and only if for every deterministic adversary $\mathcal{A}$ and every pair of neighboring initial states $s\sim s^{\prime}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}(s^{\prime}))$ are $(\epsilon,\delta)$-indistinguishable.
 By Definition 3.7, this result extends naturally to CMs:
@@
 A CM $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t.
-a verification function $f$ if and and only if for every deterministic adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
+a verification function $f$ if and only if for every deterministic adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
 The definition of $(0,\delta)$-DP CMs can be equivalently expressed using the notion of total variation:
@@
 Let $Y^{0}$ and $Y^{1}$ be random variables with domain $\mathcal{Y}$.
-For every (not necessarily maximal) coupling $(Z^{0},Z^{1})$ of $(Y^{0},Y^{1})$, where the marginals of $Z^{0}$ and $Z^{1}$ are identically distributed as of $Y^{0}$ and $Y^{1}$, we have
+For every (not necessarily maximal) coupling $(Z^{0},Z^{1})$ of $Y^{0}$ and $Y^{1}$, where the marginals $Z^{0}$ and $Z^{1}$ are distributed as $Y^{0}$ and $Y^{1}$, we have
 $\mathrm{TV}(Y^{0},Y^{1})\leq\Pr[Z^{0}\neq Z^{1}].$
@@
   • If the histories are identical, i.e.
-$Z_{1:i-1}^{0}=Z_{1:i-1}^{1}=y_{1:i-1}$, then $(Z_{i}^{0},Z_{i}^{1})$ is drawn from the maximal coupling of $Y_{i}^{0}\mid Y_{1:i-1}^{0}=y_{1:i-1}$} and $Y_{i}^{1}\mid Y_{1:i-1}^{1}=y_{1:i-1}$}, given by Lemma 9.6.
+$Z_{1:i-1}^{0}=Z_{1:i-1}^{1}=y_{1:i-1}$, then $(Z_{i}^{0},Z_{i}^{1})$ is drawn from the maximal coupling of teh distributions of $Y_{i}^{0}\mid Y_{1:i-1}^{0}=y_{1:i-1}$} and $Y_{i}^{1}\mid Y_{1:i-1}^{1}=y_{1:i-1}$}, given by Lemma 11.6.
   • Otherwise, if $Z_{1:i-1}^{0}=y_{1:i-1}^{0}$ and $Z_{1:i-1}^{1}=y_{1:i-1}^{1}$ differ, then $Z_{i}^{0}$ and $Z_{i}^{1}$ are drawn independently from their respective conditional distributions $Y_{i}^{0}\mid Y_{1:i-1}^{0}=y_{1:i-1}^{0}$} and $Y_{i}^{1}\mid Y_{1:i-1}^{1}=y_{1:i-1}^{1}$}.
@@
 $\Pr[Z^{0}=Z^{1}]\geq(1-\gamma)\sum_{y_{1:k-2}\in S_{k-2}}\prod_{i=1}^{k-2}\frac{\Pr\!\left[Z_{i}^{0}\!=\!Z_{i}^{1}\!=\!y_{i}\,\middle|\,Z_{1:i-1}^{0}\!=\!Z_{1:i-1}^{1}\!=\!y_{1:i-1}\right]}{1-\Delta_{i}(y_{1:i-1})}$
-A backward iteration with the same proof implies that $\Pr[Z^{0}=Z^{1}]\geq 1-\gamma$. Consequently, $\Pr[Z^{0}\neq Z^{1}]\leq\gamma$, and by Lemma 9.7, $\mathrm{TV}(Y^{0},Y^{1})\leq\gamma$. ∎
+A backward iteration with the same proof shows that $\Pr[Z^{0}=Z^{1}]\geq 1-\gamma$. Consequently, $\Pr[Z^{0}\neq Z^{1}]\leq\gamma$, and by Lemma 11.7, $\mathrm{TV}(Y^{0},Y^{1})\leq\gamma$. ∎
 Proof of Lemma 5.9.
@@
 As $\mathcal{A}$ is deterministic, for every $i\in\mathbb{N}$ and $w_{1:i-1}=(w_{1},\dots,w_{i-1})\in S_{i-1}^{b}$, the random variable $W_{i}^{b}$ conditioned on $W_{1:i-1}=w_{1:i-1}$ is distributed as follows:
-  • If $w_{i-1}=\mathdollar$ (i.e., the communication has halted before) or the design of $\mathcal{A}$ is such that it halts the communication after receiving the i1i-1-st response $w_{i-1}$ given the history $w_{1:i-1}$, then $W_{1:i-1}=\mathdollar$ with probability $1$.
+  • If $w_{i-1}=\mathdollar$ (i.e., the communication has halted before) or the design of $\mathcal{A}$ is such that it halts the communication after receiving the (i1)(i-1)-th response $w_{i-1}$ given the history $w_{1:i-1}$, then $W_{1:i-1}=\mathdollar$ with probability $1$.
   • Otherwise, $W_{1:i-1}$ is equivalently distributed to $\mathrm{RR}_{0,\delta_{i}^{\prime}}(b)$, where $\delta_{i}^{\prime}=\delta_{i}^{\prime}(w_{1:i-1},\mathcal{A})$ is determined by the history $w_{1:i-1}$ and the design of $\mathcal{A}$.
 Define $\Delta_{i}(w_{1:i-1})=0$ in the former case and $\Delta_{i}(w_{1:i-1})=\delta_{i}^{\prime}(w_{1:i-1},\mathcal{A})$ in the latter.
-By Lemma 9.4 and Lemma 2.4, for every $i\in\mathbb{N}$ and every $w_{1:i-1}\in S_{i-1}^{b}$,
+By Lemmas 11.4 and 2.4, for every $i\in\mathbb{N}$ and every $w_{1:i-1}\in S_{i-1}^{b}$,
 $\mathrm{TV}\left((W_{i}^{0}|W_{1:i-1}^{0}=w_{1:i-1}),(W_{i}^{1}|W_{1:i-1}^{1}=w_{1:i-1})\right)\leq\Delta_{i}(w_{1:i-1}).$
@@
 The output $\bot$ indicates that the privacy budget has been exhausted, and the interaction must be halted.
-Given a filter $\mathrm{Filt}$, we define a verification function $f[\mathrm{Filt}]$ as follows and refer to the $f[\mathrm{Filt}]$-concurrent composition of CMs as the concurrent $\mathrm{Filt}$-filter composition of CMs.
-The verification function $f[\mathrm{Filt}]$ verifies the validity of message formats, extracts the privacy parameters from the creation queries, and apply $\mathrm{Filt}$ on them.
+Given a filter $\mathrm{Filt}$, we define a verification function $f^{\mathrm{Filt}}$ as follows and refer to the $f^{\mathrm{Filt}}$-concurrent composition of CMs as the concurrent $\mathrm{Filt}$-filter composition of CMs.
+The verification function $f^{\mathrm{Filt}}$ verifies the validity of message formats, extracts the privacy parameters from the creation queries, and applies $\mathrm{Filt}$ on them.
 Definition 12.2.
@@
 Rogers et al.
-[21] studied the $\mathrm{Filt}$-filter composition of NIMs for the first time.
+(rogers2016privacy, ) studied the $\mathrm{Filt}$-filter composition of NIMs for the first time.
 More recently, Haney et al.
-[11] showed that if the $\mathrm{Filt}$-filter composition of NIMs is $(\epsilon,\delta)$-DP, then the same guarantee extends to the $\mathrm{Filt}$-filter composition of IMs.
-We further generalize this result to the $\mathrm{Filt}$-filter composition of CMs.
+(haney2023concurrent, ) showed that if the $\mathrm{Filt}$-filter composition of NIMs is $(\epsilon,\delta)$-DP, then the same guarantee extends to the concurrent $\mathrm{Filt}$-filter composition of IMs.
+We further generalize this result to the concurrent $\mathrm{Filt}$-filter composition of CMs.
 To state our result, we need to define the $\mathrm{Filt}$-filter composition of randomized response mechanisms:
@@
 Theorem 12.4.
-For every filter $\mathrm{Filt}$, if the $\mathrm{Filt}$-filter composition of randomized response mechanisms is $(\epsilon,\delta)$-DP, then the $\mathrm{Filt}$-filter composition of CMs is also $(\epsilon,\delta)$-DP.
+For every filter $\mathrm{Filt}$, if the $\mathrm{Filt}$-filter composition of randomized response mechanisms is $(\epsilon,\delta)$-DP, then the concurrent $\mathrm{Filt}$-filter composition of CMs is also $(\epsilon,\delta)$-DP.
 Proof.
 Let $f=f^{\mathrm{Filt}}$ and $f_{\mathrm{RR}}=f^{\mathrm{Filt}}_{\mathrm{RR}}$.
-This proof is identical to the proof of Theorem 4.9, except the following modifications: First, the right mechanism of the IPM $\mathcal{P}$ is $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ instead of ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((b)_{i=1}^{k}\right)$.
+This proof is identical to the proof of Theorem 4.10, except for the following modifications: First, the right mechanism of the IPM $\mathcal{P}$ is $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ instead of ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((b)_{i=1}^{k}\right)$.
 Second, the variable $\sigma_{1}$ is set differently.
@@
 Here, $\mathcal{P}$ interacts with its right mechanism, $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, each time it receives a pair of creation queries as a left message and appends a randomized response $r_{i}$ to $\sigma_{1}$.
-Specifically, the same as before, the state of $\mathcal{P}$ consists of two sequences $\sigma_{1}$ and $\sigma_{2}$, both initialized empty.
+Specifically, as before, the state of $\mathcal{P}$ consists of two sequences $\sigma_{1}$ and $\sigma_{2}$, both initialized empty.
 Upon receiving a left message of the form $(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})^{2}$, $\mathcal{P}$ takes the following steps: First, $\mathcal{P}$ appends the pair $(\mathcal{T}_{i},t_{i})$ to the sequence $\sigma_{2}$, where $\mathcal{T}_{i}$ is the IPM corresponding to $\mathcal{M}_{i}$ from Lemma 4.7 and $t_{i}$ is its initial state.
@@
 ∎
+13. Rényi DP
+In this section, we show concurrent composition theorems for Rényi differential privacy (RDP), defined as follows.
+Definition 13.1 (Rényi Divergence).
+For $\alpha>1$, the $\alpha$-Rényi divergence between two distributions $P$ and $Q$ is defined as
+$D_{\alpha}(P\|Q)=\frac{1}{\alpha-1}\log\mathbb{E}_{y\sim Q}\left[\left(\frac{P(y)}{Q(y)}\right)^{\alpha}\right].$
+By taking the limit as $\alpha\to\infty$, we obtain the $\infty$-Rényi divergence,
+$D_{\infty}(P\|Q)=\lim_{\alpha\to\infty}D_{\alpha}(P\|Q)=\log\max_{y}\frac{P(y)}{Q(y)}.$
+Definition 13.2 (Rényi Differential Privacy (RDP) (mironov2017renyi, )).
+For $1<\alpha\leq\infty$ and $\epsilon\geq 0$, a NIM $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is said to be $(\alpha,\epsilon)$-Rényi differentially private (or $(\alpha,\epsilon)$-RDP) w.r.t.
+a neighbor relation $\sim$ if for all neighboring datasets $x_{0},x_{1}\in\mathcal{X}$, the $\alpha$-Rényi divergence between the output distributions satisfies
+[…]
+Note that the above definition includes pure DP as a special case.
+Lyu (lyu2022composition, ) and Vadhan et al.
+(vadhan2022concurrent, ) extend this definition to IMs, which we further generalize to CMs:
+Definition 13.3 (RDP for CMs).
+For $1<\alpha\leq\infty$ and $\epsilon\geq 0$, a continual mechanism $\mathcal{M}$ is said to be $(\alpha,\epsilon)$-Rényi differentially private (or $(\alpha,\epsilon)$-RDP) w.r.t.
+a verification function $f$ against adaptive adversaries if for every adversary $\mathcal{A}$, the random variables $V^{0}=\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $V^{1}=\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ satisfy
+[…]
+Recall the concurrent composition of $k$ CMs with fixed privacy parameters $\{(\epsilon_{i},\delta_{i})\}_{i=1}^{k}$, which was formally defined via a verification function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$ in Definition 4.11.
+This verification function ensures that every creation query $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})$ is truthful, meaning that the mechanism $\mathcal{M}_{j}^{\prime}$ is $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$-DP w.r.t.
+the verification function $f_{j}^{\prime}$, and the multiset of parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ for all created mechanisms forms a sub-multiset of the predetermined parameters $\{(\epsilon_{i},\delta_{i})\}_{i=1}^{k}$.
+For RDP, creation queries are of the form $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\alpha_{j}^{\prime},\epsilon_{j}^{\prime},f_{j}^{\prime})$, and the corresponding concurrent composition verification function must ensure that each $\mathcal{M}_{j}^{\prime}$ is $(\alpha_{j}^{\prime},\epsilon_{j}^{\prime})$-RDP with respect to $f_{j}^{\prime}$.
+Fixing $\alpha$ and $\epsilon_{1},\dots,\epsilon_{k}$, this verification function must check that all created mechanisms satisfy $\alpha_{j}^{\prime}=\alpha$, and that the multiset of parameters $\epsilon_{j}^{\prime}$ is a sub-multiset of the prefix multiset $\{\epsilon_{i}\}_{i=1}^{k}$.
+With these modifications to Definition 4.11, we denote the resulting verification function by $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\mathrm{RDP},\alpha}$ and refer to the $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\mathrm{RDP},\alpha}$-concurrent composition of CMs as the $\alpha$-RDP concurrent composition of CMs with parameters $\epsilon_{1},\dots,\epsilon_{k}$.
+Lyu (lyu2022composition, ) analyzed the concurrent composition of $k$ fixed IMs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$, where each $\mathcal{M}_{i}$ is $(\alpha,\epsilon_{i})$-RDP w.r.t.
+a neighbor relation $\sim_{i}$.
+The analysis considers an adversary that concurrently interacts with $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$, initialized with states (datasets) $s_{1},\dots,s_{k}$, and compares this interaction to one where the mechanisms are initialized with neighboring states $s_{1}^{\prime},\dots,s_{k}^{\prime}$, such that $s_{i}\sim_{i}s_{i}^{\prime}$ for all $i\in[k]$.
+Lyu shows that the $\alpha$-Rényi divergence between the resulting adversarial view distributions is at most $\epsilon=\sum_{i=1}^{k}\epsilon_{i}$.
+This result was originally proved to hold assuming that the number of queries made by the adversary is bounded by a fixed integer $T$.
+However, Haney et al.
+(haney2023concurrent, ) later demonstrated that this assumption can be removed.
+(See the application of the result of Van Erven and Harremoës (van2014Renyi, ) in their paper.)
+We can replace each IM $\mathcal{M}_{i}$ by a universal interactive mechanism $\mathcal{UM}_{i}$ whose initial state space is $\{0,1\}$.
+Given an initial state $b\in\{0,1\}$, the mechanism $\mathcal{UM}_{i}$ receives as its first message a creation query of the form $(\mathcal{M}_{i}^{\prime},s_{i}^{\prime},\alpha,\epsilon_{i},f_{i})$ and executes $\mathcal{V}[f_{i}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{i}^{\prime}(s_{i}^{\prime})$ from that point on.
+If the initial message is malformed or if $\mathcal{M}_{i}^{\prime}$ does not satisfy $(\alpha,\epsilon_{i})$-RDP w.r.t.
+$f_{i}$, then $\mathcal{UM}_{i}$ immediately halts the interaction.
+It is straightforward to verify that $\mathcal{UM}_{i}$ is $(\alpha,\epsilon_{i})$-RDP w.r.t.
+the neighbor relation $\sim_{01}$ defined by $0\sim_{01}1$.
+By construction, the concurrent composition of the universal mechanisms $\mathcal{UM}_{1},\dots,\mathcal{UM}_{k}$ is equivalent to the $\alpha$-RDP concurrent composition of CMs with parameters $\epsilon_{1},\dots,\epsilon_{k}$.
+Hence, the result of Lyu (lyu2022composition, ) for IMs implies the following.
+Corollary 13.4.
+For $1<\alpha\leq\infty$, $k\in\mathbb{N}$, and $\epsilon_{1},\dots,\epsilon_{k}\geq 0$, the $\alpha$-RDP concurrent composition of CMs with parameters $\epsilon_{1},\dots,\epsilon_{k}$ is $(\alpha,\sum_{i=1}^{k}\epsilon_{i})$-RDP.
+Previously, the concurrent parallel composition of CMs was defined via the verification function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$ in Definition 5.1.
+Analogously, we define the verification function $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\mathrm{RDP},\alpha}$ for RDP with the following modifications: (i) Creation queries are of the form $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\alpha_{j}^{\prime},\epsilon_{j}^{\prime},f_{j}^{\prime})$ rather than $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})$.
+(ii) The verification function $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\mathrm{RDP},\alpha}$ ensures that each $\mathcal{M}_{j}^{\prime}$ is $(\alpha_{j}^{\prime},\epsilon_{j}^{\prime})$-RDP w.r.t.
+$f_{j}^{\prime}$, and that all $\alpha_{j}^{\prime}$ equal the fixed $\alpha$.
+We refer to the $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,RDP,\alpha}$-concurrent composition as the $\alpha$-RDP concurrent $k$-sparse parallel composition of CMs with privacy parameters $\epsilon_{1},\dots,\epsilon_{k}$.
+The following theorem shows that this concurrent composition is not necessarily Rényi private:
+Theorem 13.5.
+For every $1<\alpha<\infty$ and $\epsilon\geq 0$, the $f^{\epsilon}_{\infty,RDP,\alpha}$-concurrent composition is not $(\alpha,\epsilon^{\prime})$-DP for every $\epsilon^{\prime}\geq 0$.
+Proof.
+Recall the proof of Theorem 5.2, in which an adversary $\mathcal{A}_{\delta}$ repeatedly created instances of a $(0,\delta)$-DP CM, denoted $\mathcal{M}_{\delta}$, with a pair of zeros as its first input pair until one mechanism returned $\bot$ as response.
+At that point, $\mathcal{A}_{\delta}$ sent the input pair $(0,1)$ to the last mechanism, and the mechanism responded with the bit held by the identifier, thereby revealing it.
+In that construction, $\mathcal{M}_{\delta}$ was $(0,\delta)$-DP w.r.t.
+a verification function $g$ that was defined to accept only message sequences of length at most two, where each message was a pair of bits and at most one pair included different bits.
+This proof is identical to the proof of Theorem 5.2, except that the adversary $\mathcal{A}_{\alpha,\epsilon,\epsilon^{\prime}}$ creates instances of a mechanism $\mathcal{M}_{\alpha,\epsilon,\epsilon^{\prime}}$ that is $(\alpha,\epsilon)$-RDP w.r.t.
+the (same) verification function $g$, and instead of a full disclosure of the identifier’s bit, we will show that the $f^{\epsilon}_{\infty,RDP,\alpha}$-concurrent composition fails to satisfy $(\alpha,\epsilon^{\prime})$-DP.
+Definition of $\mathcal{M}_{\alpha,\epsilon,\epsilon^{\prime}}$: Like $\mathcal{M}_{\delta}$, the mechanism $\mathcal{M}_{\alpha,\epsilon,\epsilon^{\prime}}$ receives single-bit inputs and outputs a response in $\{0,1,\top,\bot\}$.
+Upon receiving its first input bit, $\mathcal{M}_{\alpha,\epsilon,\epsilon^{\prime}}$ outputs $\bot$ with probability
+$p=\min\left\{\frac{1}{2},\epsilon\cdot(\alpha-1)\cdot 6^{-\alpha}\cdot 2^{-\epsilon^{\prime}\cdot(\alpha-1)}\right\},$
+and outputs $\top$ otherwise.
+If the first response was $\top$, then on its second input, $\mathcal{M}_{\alpha,\epsilon,\epsilon^{\prime}}$ again outputs $\bot$ with probability $p$ and $\top$ otherwise.
+However, if the first response was $\bot$, then on its next input it outputs a bit correlated with the input bit.
+In the $(0,\delta)$-DP construction, $\mathcal{M}_{\delta}$ revealed the input bit exactly, resulting in a full disclosure of the private bit.
+In contrast, $\mathcal{M}_{\alpha,\epsilon,\epsilon^{\prime}}$ outputs the true input bit with probability
+$q=1-\frac{3}{4}2^{-2-\epsilon^{\prime}\frac{\alpha-1}{\alpha}},$
+and outputs the opposite bit otherwise. Note that $q\geq 1-\frac{3}{4}2^{-2}>\frac{3}{4}$ and $q\leq 1$.
+$\mathcal{M}_{\alpha,\epsilon,\epsilon^{\prime}}$ is $(\alpha,\epsilon)$-RDP w.r.t.
+$g$: By the definition of the verification function $g$, the mechanism $\mathcal{V}[g]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}_{\alpha,\epsilon,\epsilon^{\prime}}$ interacts with an adversary at most twice and accepts a pair of bits each time, where at most one pair contains differing bits.
+By construction of $\mathcal{M}_{\alpha,\epsilon,\epsilon^{\prime}}$, the $\alpha$-Rényi divergence between the adversary’s views is maximized when the adversary first sends identical bits, and when observing the response $\bot$, it sends a pair of differing bits as the second query.
+Let $\mathcal{A}$ denote such an adversary.
+For $b\in\{0,1\}$, we denote the view of $\mathcal{A}$ interacting with $\mathcal{V}[g]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{\alpha,\epsilon,\epsilon^{\prime}}$ by $V_{\mathcal{A}}^{b}$.
+For every view $v$ where the first response of $\mathcal{M}_{\alpha,\epsilon,\epsilon^{\prime}}$ is $\top$, the probabilities $\Pr[V_{\mathcal{A}}^{0}=v]$ and $\Pr[V_{\mathcal{A}}^{1}=v]$ are equal; hence,
+$\left(\frac{\Pr[V_{\mathcal{A}}^{0}=v]}{\Pr[V_{\mathcal{A}}^{1}=v]}\right)^{\alpha}=1.$
+For every view $v$ whose first response is $\bot$, the term $\bigl(\frac{\Pr[V_{\mathcal{A}}^{0}=v]}{\Pr[V_{\mathcal{A}}^{1}=v]}\bigr)^{\alpha}$ equals $(\frac{q}{1-q})^{\alpha}$ if the second response is 0, and $(\frac{1-q}{q})^{\alpha}$ if the second response is $1$.
+Since $q\in[3/4,1]$, we have $\frac{1-q}{q}\leq\frac{1}{1-q}$.
+Thus, both $(\frac{q}{1-q})^{\alpha}$ and $(\frac{1-q}{q})^{\alpha}$ are upper bounded by $(\frac{1}{1-q})^{\alpha}$.
+The probability that the first response in the random variable in $V_{\mathcal{A}}^{0}$ equals $\bot$ is $p$.
+Thus, for the worst-case adversary $\mathcal{A}$ we have
+$D_{\alpha}(V_{\mathcal{A}}^{0}\|V_{\mathcal{A}}^{1})$ $\leq\frac{1}{\alpha-1}\log\left((1-p)\cdot 1+p\cdot\left(\frac{1}{1-q}\right)^{\alpha}\right)$
+$\leq\frac{1}{\alpha-1}\log\left(1+p\cdot\left(\frac{4}{3}2^{2+\epsilon^{\prime}\frac{\alpha-1}{\alpha}}\right)^{\alpha}\right)$
+$\leq\frac{1}{\alpha-1}\log\left(1+p\cdot 6^{\alpha}\cdot 2^{\epsilon^{\prime}(\alpha-1)}\right)$
+$\leq\frac{1}{\alpha-1}\log\left(1+\epsilon\cdot(\alpha-1)\right)$
+$\leq\epsilon$
+In the last inequality, we use the fact that $\log(1+x)\leq x$ for every $x\geq 0$.
+By symmetry, $D_{\alpha}(V_{\mathcal{A}}^{1}\|V_{\mathcal{A}}^{0})\leq\epsilon$ as well, and therefore $\mathcal{M}_{\alpha,\epsilon,\epsilon^{\prime}}$ satisfies $(\alpha,\epsilon)$-RDP w.r.t.
+$g$.
+The $f^{\epsilon}_{\infty,\mathrm{RDP},\alpha}$-concurrent composition is not $(\alpha,\epsilon^{\prime})$-DP: For $b\in\{0,1\}$, let $V^{b}$ denote the view of $\mathcal{A}_{\alpha,\epsilon,\epsilon^{\prime}}$ interacting with $\mathcal{V}[f^{\epsilon}_{\infty,\mathrm{RDP},\alpha}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$.
+We show that $D_{\alpha}(V^{0}\|V^{1})>\epsilon^{\prime}$.
+By definition, $\mathcal{A}_{\alpha,\epsilon,\epsilon^{\prime}}$ eventually receives a $\bot$ response from an instance of $\mathcal{M}_{\alpha,\epsilon,\epsilon^{\prime}}$.
+It then sends $(0,1)$ to that instance and receives an answer $a\in\{0,1\}$.
+For every view $v$ where $a=0$,
+$\left(\frac{\Pr[V^{0}=v]}{\Pr[V^{1}=v]}\right)^{\alpha}=\left(\frac{q}{1-q}\right)^{\alpha},$
+and for every view $v$ with $a=1$,
+$\left(\frac{\Pr[V^{0}=v]}{\Pr[V^{1}=v]}\right)^{\alpha}=\left(\frac{1-q}{q}\right)^{\alpha}.$
+When sampling $v$ from $V^{0}$, the event $a=0$ occurs with probability $q$. Hence,
+$D_{\alpha}(V^{0}\|V^{1})$ $=\frac{1}{\alpha-1}\log\left(q\cdot\left(\frac{q}{1-q}\right)^{\alpha}+(1-q)\cdot\left(\frac{q}{q}\right)^{\alpha}\right)$
+$\geq\frac{1}{\alpha-1}\log\left(q\left(\frac{q}{1-q}\right)^{\alpha}\right)$
+$\geq\frac{1}{\alpha-1}\log\left(\frac{3}{4}\cdot\left(\frac{3}{4(1-q)}\right)^{\alpha}\right)$
+$=\frac{1}{\alpha-1}\log\left(\frac{3}{4}\cdot\left(\frac{3}{4}\cdot\frac{1}{\frac{3}{4}\cdot 2^{-2-\epsilon^{\prime}\frac{\alpha-1}{\alpha}}}\right)^{\alpha}\right)$
+$=\frac{1}{\alpha-1}\log\left(\frac{3}{4}\cdot 2^{2+\epsilon^{\prime}(\alpha-1)}\right)$
+$>\epsilon^{\prime}.$
+As in Section 5.3, to overcome this lower bound, we modify the verification function $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\mathrm{RDP},\alpha}$ to only create mechanisms $\mathcal{M}_{j}^{\prime}$ whose associated verification function $f_{j}^{\prime}$ is first-pair consistent (see Definition 5.12).
+Let $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\mathit{FPC},\mathrm{RDP},\alpha}$ denote this verification function.
+Theorem 13.6.
+For $\alpha>1$, $k\in\mathbb{N}$, and $\epsilon_{1},\dots,\epsilon_{k}\geq 0$, the $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\mathit{FPC},\mathrm{RDP},\alpha}$-concurrent composition of CMs is $(\alpha,\sum_{i=1}^{k}\epsilon_{i})$-RDP.
+Proof.
+The proof is identical to the proof of Theorem 5.14, with a few modifications in the construction of $\mathcal{P}$.
+Recall that in the proof of Theorem 5.14, receiving the $j$-th pair of identical creation queries $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$, the IPM $\mathcal{P}$ temporarily stored the creation query until it received the first query pair for mechanism $\mathcal{M}_{j}$, i.e., a message of the form $\left((q_{0},j),(q_{1},j)\right)$.
+Omitting the implementation details, the idea was that $\mathcal{P}$ instantiated either an IM $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ if $q_{0}=q_{1}$ on the first query pair or an IPM $\mathcal{T}_{j}$ otherwise, where $\mathcal{T}_{j}$ was the IPM corresponding to $\mathcal{M}_{j}^{\prime}$ from Lemma 4.7.
+This lemma constructed $\mathcal{T}_{j}$ so that the IMs $\mathcal{T}_{j}\!\circ^{*}\!\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ and $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ are equivalent for every $b\in\{0,1\}$.
+We will call $\mathcal{M}_{j}^{\prime}$ a type-1 CM in the former case and a type-2 CM in the latter.
+Upon receiving a left message of the form $\left((q_{0},j),(q_{1},j)\right)$, the IPM $\mathcal{P}$ forwarded $(q_{0},q_{1})$ either to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ (type 1) or to $\mathcal{T}_{j}$ (type 2), and returned the response as a left message.
+The IPM $\mathcal{T}_{j}$ might want to interact with its right mechanism, requesting the outcome of $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$.
+To supply answers, $\mathcal{P}$ initially interacted with ${\textrm{Comp}}(\mathrm{RR}_{\epsilon_{1},\delta_{1}},\ldots,\mathrm{RR}_{\epsilon_{k},\delta_{k}})$, stored its response $(r_{1},\dots,r_{k})$, and later used these $r_{i}$’s as right responses to the IPMs $\mathcal{T}_{j}$.
+This reduced the $f_{\infty,\mathit{FPC}}^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$-concurrent composition of CMs to the composition of the noninteractive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$.
+In this proof we make the following changes:
+  1. (1)
+    Creation queries now have the form $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\alpha_{j}^{\prime},\epsilon_{j}^{\prime},f_{j}^{\prime})$ instead of $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})$.
+  2. (2)
+    The right mechanism of $\mathcal{P}$ is
+    $\mathcal{V}[f^{\epsilon_{1},\dots,\epsilon_{k}}_{\mathrm{RDP},\alpha}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}},$
+    instead of ${\textrm{Comp}}(\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}})$.
+  3. (3)
+    Rather than instantiating $\mathcal{T}_{j}$ via Lemma 4.7 when $q_{0}=q_{1}$ on the first query pair, $\mathcal{P}$ now forwards the pair of identical creation queries $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\alpha_{j}^{\prime},\epsilon_{j}^{\prime},f_{j}^{\prime})^{2}$ to its right mechanism and stores the index $i_{j}$ indicating that $\mathcal{M}_{j}^{\prime}$ corresponds to the $i_{j}$-th mechanism created by the right mechanism of $\mathcal{P}$.
+  4. (4)
+    Upon receiving a left message of the form $\left((q_{0},j),(q_{1},j)\right)$, if $\mathcal{M}_{j}^{\prime}$ is a type-2 CM, then $\mathcal{P}$ forwards the message $\left((q_{0},i_{j}),(q_{1},i_{j})\right)$ to its right mechanism and returns that response to the left.
+If $\mathcal{M}_{j}^{\prime}$ is a type-1 CM, then as before $\mathcal{P}$ forwards $(q_{0},q_{1})$ to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ and returns the resulting answer.
+The exact same argument as in the proof of Theorem 5.14 shows that the $f^{\epsilon_{1},\ldots,\epsilon_{k}}_{\infty,\mathit{FPC},\mathrm{RDP},\alpha}$-concurrent composition of CMs satisfies the privacy guarantees of the $f^{\epsilon_{1},\ldots,\epsilon_{k}}_{\mathrm{RDP},\alpha}$-concurrent composition of CMs.
+By Theorem 13.4, the $f^{\epsilon_{1},\ldots,\epsilon_{k}}_{\mathrm{RDP},\alpha}$-concurrent composition of CMs is $(\alpha,\sum_{i=1}^{k}\epsilon_{i})$-RDP.
+Therefore, the $f^{\epsilon_{1},\ldots,\epsilon_{k}}_{\infty,\mathit{FPC},\mathrm{RDP},\alpha}$-concurrent composition of CMs is also $(\alpha,\sum_{i=1}^{k}\epsilon_{i})$-RDP, completing the proof.
+∎
+As discussed in Section 5.3.1, IMs satisfy the first-pair consistent condition. Thus, by Theorem 13.6, we have:
+Corollary 13.7.
+For $\alpha>1$, $k\in\mathbb{N}$, and $\epsilon_{1},\dots,\epsilon_{k}\geq 0$, the $\alpha$-RDP concurrent $k$-sparse parallel composition of IMs with privacy parameters $\epsilon_{1},\dots,\epsilon_{k}$ is $(\alpha,\sum_{i=1}^{k}\epsilon_{i})$-RDP.
+Finally, to define concurrent filter composition of RDP CMs, we modify Definition 12.1 so that a filter $\mathrm{Filt}$ maps a finite sequence of nonnegative reals in $0,\infty)^{*}$ (instead of a sequence in $\left([0,\infty)\times[0,1]\right)^{*}$) to an element in $\{\top,\bot\}$.
+Fix $\alpha>1$.
+As before, we obtain an RDP version of the verification function $f^{\mathrm{Filt}}$ in Definition 12.2 by replacing creation queries to use RDP parameters, ensuring that each created mechanism with parameters $(\alpha_{j},\epsilon_{j})$ is $(\alpha_{j},\epsilon_{j})$-RDP, and checking whether all $\alpha_{j}=\alpha$.
+We denote this verification function by $f^{\mathrm{Filt}}_{\mathrm{RDP},\alpha}$.
+For parameters $\alpha>1$ and $\epsilon\geq 0$, we define the filter $\mathrm{Filt}_{\mathrm{RDP},\alpha,\epsilon}$ as follows.
+For every $t\in\mathbb{N}$ and every $\epsilon_{1},\dots,\epsilon_{t}\in[0,\infty)^{t}$,
+$\mathrm{Filt}_{\mathrm{RDP},\alpha,\epsilon}(\epsilon_{1},\dots,\epsilon_{t})=\begin{cases}\top,&\text{if }\sum_{i=1}^{t}\epsilon_{i}\leq\epsilon,\\[4.0pt] \bot,&\text{otherwise.}\end{cases}$
+We refer to the $f^{\mathrm{Filt}_{\mathrm{RDP},\alpha,\epsilon}}$-concurrent composition of CMs as the concurrent $\mathrm{Filt}_{\mathrm{RDP},\alpha,\epsilon}$-filter composition of CMs.
+Haney et al.
+(haney2023concurrent, ) show that the concurrent $\mathrm{Filt}_{\mathrm{RDP},\alpha,\epsilon}$-filter composition of IMs is $(\alpha,\epsilon)$-RDP.
+Following the same argument as for the concurrent composition of $k$ fixed RDP CMs, this result extends to CMs by replacing each IM with the universal IM with the same privacy parameters:
+Corollary 13.8.
+For every $\alpha>1$ and $\epsilon\geq 0$, the concurrent $\mathrm{Filt}_{\mathrm{RDP},\alpha,\epsilon}$-filter composition of CMs is $(\alpha,\epsilon)$-RDP against adaptive adversaries.
+14. $f$-DP
+In this section, we show concurrent composition theorems for $f$-differential privacy ($f$-DP), which generalizes $(\epsilon,\delta)$-DP.
+$f$-DP takes a statistical point of view and uses trade-off functions to measure how indistinguishable the output distributions of a mechanism are for two neighboring datasets.
+Remark 14.1.
+To avoid ambiguity, throughout this section we use the symbol $g$ to denote verification functions and the symbol $f$ to denote privacy functions.
+Consider a NIM $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ and two neighboring datasets $x_{0},x_{1}\in\mathcal{X}$.
+Observing the outcome of $\mathcal{M}$, an analyst/adversary $\mathcal{A}$ must decide whether the input dataset was $x_{0}$ (null hypothesis $H_{0}$) or $x_{1}$ (alternative hypothesis $H_{1}$).
+A rejection rule $\phi:\mathcal{Y}\to\{0,1\}$ is a function that takes the answer of $\mathcal{M}$ as input and decides whether to reject the null hypothesis or not.
+Thus, for $b\in\{0,1\}$, $\mathcal{A}$ chooses dataset $x_{b}$ when $\phi$ returns $b$.
+To quantify the indistinguishability between $\mathcal{M}(x_{0})$ and $\mathcal{M}(x_{1})$, $f$-DP evaluates the error made by the analyst in selecting the correct hypothesis.
+Specifically, given a rejection rule $\phi$, the type I error $\alpha_{\phi}=\mathbb{E}[\phi(\mathcal{M}(x_{0}))]$ is the probability of rejecting the null hypothesis $H_{0}$ while $H_{0}$ is true (i.e., guessing $x_{1}$ when the dataset is $x_{0}$).
+Similarly, the type II error $\beta_{\phi}=1-\mathbb{E}[\phi(\mathcal{M}(x_{1}))]$ is the probability of not rejecting $H_{0}$ while $H_{1}$ is true (i.e., guessing $x_{0}$ when the dataset is $x_{1}$).
+Although an analyst could achieve zero type II error by always rejecting the null hypothesis, when an upper bound on the type I error is set, the minimum achievable type II error reflects the difficulty of distinguishing between $\mathcal{M}(x_{0})$ and $\mathcal{M}(x_{1})$.
+The following function measures the optimal trade-off between the type I and type II errors.
+Definition 14.2 (Trade-off Function).
+Let $Y_{0}$ and $Y_{1}$ be two random variables on the same domain $\mathcal{Y}$. The trade-off function between $Y_{0}$ and $Y_{1}$, denoted $T(Y_{0},Y_{1}):[0,1]\to[0,1]$, is defined as:
+$T(Y_{0},Y_{1})(\alpha)=\inf_{\phi}\{\beta_{\phi}:\alpha_{\phi}\leq\alpha\},$
+where the infimum is taken over all possible rejection rules. Given a function $f:[0,1]\to[0,1]$, we write $T(Y_{0},Y_{1})\geq f$ if $T(Y_{0},Y_{1})(\alpha)\geq f(\alpha)$ for every $\alpha\in[0,1]$.
+The (meta-)function $T$ in Definition 14.2 is indeed a function that maps a pair of random variables to a trade-off function between these random variables.
+Definition 14.3 ($f$-DP for NIMs (dong2022gaussian, )).
+Let $f:[0,1]\to[0,1]$ be a function.
+A NIM $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $f$-differentially private (or $f$-DP) w.r.t.
+a neighbor relation $\sim$ if for every two neighboring datasets $x_{0},x_{1}\in\mathcal{X}$,
+$T(\mathcal{M}(x_{0}),\mathcal{M}(x_{1}))\geq f,$
+where $T(\mathcal{M}(x_{0}),\mathcal{M}(x_{1}))$ denotes the trade-off function defined in Definition 14.2.
+For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, define the function $f_{\epsilon,\delta}:[0,1]\to[0,1]$ as $f_{\epsilon,\delta}(\alpha)=\max\{0,1-\delta-\exp(\epsilon)\alpha,\exp(-\epsilon)(1-\delta-\alpha)\}$ for every $\alpha\in[0,1]$.
+Dong, Roth, and Su (dong2022gaussian, ) show that $f_{\epsilon,\delta}$-DP is equivalent to $(\epsilon,\delta)$-DP.
+They also show that if a NIM $\mathcal{N}$ satisfies $f$-DP w.r.t.
+a neighbor relation $\sim$, where $f$ is the trade-off function of two random variables $Z_{0}$ and $Z_{1}$, then for every pair of neighboring datasets $x_{0}\sim x_{1}$, there exists a randomized post-processing function Post such that, for each $b\in\{0,1\}$, the distributions of $\mathcal{N}(x_{b})$ and $\textsc{Post}(Z_{b})$ are identical.
+Vadhan and Zhang (vadhan2022concurrent, ) prove an analogue of Lemma 4.7 for $f$-DP under an additional finite communication assumption, defined below.
+Definition 14.4 (Finite Communication for IMs).
+An IM $\mathcal{M}$ is said to have finite communication if there exists a constant $c$ such that both its query space $Q_{\mathcal{M}}$ and answer space $A_{\mathcal{M}}$ have cardinality at most $c$, and $\mathcal{M}$ terminates after receiving more than $c$ queries.
+Lemma 14.5 ((vadhan2022concurrent, )).
+Let $f$ be a privacy function, and let $\mathcal{M}$ be an IM that is $f$-DP w.r.t.
+a neighbor relation $\sim$ and has finite communication.
+Then, for every pair of neighboring datasets $x_{0}\sim x_{1}$, there exists an IPM $\mathcal{P}$ with a single initial state and a NIM $\mathcal{N}$ taking a bit as input such that the IMs $\mathcal{P}\!\circ^{*}\!\mathcal{N}(b)$ and $\mathcal{M}(x_{b})$ are equivalent for every $b\in\{0,1\}$, and the mechanism $\mathcal{N}$ is $f$-DP w.r.t.
+the neighbor relation $\sim_{01}$ defined by $0\sim_{01}1$.
+We now extend the definition of $f$-DP from IMs to CMs:
+Definition 14.6 ($f$-DP for CMs).
+Let $f:[0,1]\to[0,1]$ be a function.
+A continual mechanism $\mathcal{M}$ is said to be $f$-differentially private (or $f$-DP) w.r.t.
+a verification function $g$ against adaptive adversaries if, for every adversary $\mathcal{A}$, the random variables $V^{0}=\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $V^{1}=\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ satisfy
+$T(V^{0},V^{1})\geq f,$
+where $T(V^{0},V^{1})$ denotes the trade-off function defined in Definition 14.2.
+In the DP setting, creation queries were of the form $(\mathcal{M},s,\epsilon,\delta,f)$, where $\mathcal{M}_{i}$ is a CM with the initial state $s$ that is $(\epsilon,\delta)$-DP w.r.t.
+the verification function $f$.
+In the $f$-DP setting, creation queries are of the form $(\mathcal{M},s,f,g)$, where $\mathcal{M}$ is $f$-DP w.r.t.
+the verification function $g$.
+We emphasize that the symbol $f$ in this section denotes a privacy function, whereas $f$ in the previous DP sections denoted a verification function.
+Recall that the concurrent composition of $k$ CMs with fixed privacy parameters $\{(\epsilon_{i},\delta_{i})\}_{i=1}^{k}$ was defined via a verification function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$ in Definition 4.11.
+For the $f$-DP setting, we modify this verification function to accommodate creation queries of the form $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},f_{j}^{\prime},g_{j}^{\prime})$, rather than $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})$.
+Specifically, the new verification function must ensure that each created mechanism $\mathcal{M}_{j}^{\prime}$ is $f_{j}^{\prime}$-DP w.r.t.
+$g_{j}^{\prime}$ and has finite communication.
+Given privacy functions $f_{1},\dots,f_{k}$, the verification function must verify that the multiset of the privacy functions $f_{j}^{\prime}$ is a sub-multiset of the prefix functions $\{f_{i}\}_{i=1}^{k}$.
+With these modifications to Definition 4.11, we denote the resulting verification function by $g^{f_{1},\dots,f_{k}}_{\mathrm{FDP}}$ and refer to the $g^{f_{1},\dots,f_{k}}_{\mathrm{FDP}}$-concurrent composition of CMs as the concurrent composition of CMs with privacy functions $f_{1},\dots,f_{k}$.
+An argument identical to that of Theorem 4.12, with Lemma 14.5 being used instead of Lemma 2.16 implies:
+Corollary 14.7.
+For $k\in\mathbb{N}$, let $f,f_{1},\dots,f_{k}:[0,1]\to[0,1]$ be functions.
+Suppose that for every $k$ NIMs $\mathcal{N}_{1},\dots,\mathcal{N}_{k}$, where each $\mathcal{N}_{i}$ is $f_{i}$-DP w.r.t.
+a neighbor relation $\sim_{i}$, the composition of $\mathcal{N}_{1},\dots,\mathcal{N}_{k}$ is $f$-DP.
+Then, the concurrent composition of CMs with privacy functions $f_{1},\dots,f_{k}$ is also $f$-DP.
+Proof.
+The proof is exactly the same as the proof of Theorem 4.12, with Lemma 14.5 applied instead of Lemma 2.16.
+For completeness, we outline how the additional finite communication requirement of Lemma 14.5 is satisfied.
+Recall that in Theorem 4.12, the $i$-th created mechanism, $\mathcal{M}_{i}$, was $(\epsilon_{i},\delta_{i})$-DP w.r.t.
+a verification function $f_{i}$, which we denote here by $g_{i}$ to avoid confusion.
+The proof of that theorem showed that for each $b\in\{0,1\}$, the IMs $\mathcal{V}[f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}]\!\circ^{*}\!{\textrm{IECC}}(b)$ are equivalent, where ${\textrm{IECC}}(b)$ was an IM internally executing the mechanism $\mathcal{V}[g_{i}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{i}$ for each created mechanism $\mathcal{M}_{i}$.
+The argument then replaced each instance $\mathcal{V}[g_{i}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{i}$ by $\mathcal{T}_{i}\!\circ^{*}\!\mathrm{RR}_{\epsilon_{i},\delta_{i}}(b)$, where $\mathcal{T}_{i}$ is the IPM given by Lemma 2.16.
+In the $f$-DP setting, by the definition of $g^{f_{1},\dots,f_{k}}_{\mathrm{FDP}}$, each created mechanism $\mathcal{M}_{i}$ is required to have finite communication.
+Hence, its query space $Q_{\mathcal{M}_{i}}$ and answer space $A_{\mathcal{M}_{i}}$ are both finite, and $\mathcal{M}_{i}$ halts before a certain number of interactions.
+Thus, the mechanism $\mathcal{V}[g_{i}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{i}$ also has a finite answer space and terminates before the same number of steps.
+Although the verifier $\mathcal{V}[g_{i}]$ may, in principle, have an infinite query space, the outer verifier $\mathcal{V}[g^{f_{1},\dots,f_{k}}_{\mathrm{FDP}}]$ in $\mathcal{V}[g^{f_{1},\dots,f_{k}}_{\mathrm{FDP}}]\!\circ^{*}\!{\textrm{IECC}}$ ensures that all messages sent to $\mathcal{M}_{i}$ are $g_{i}$-valid.
+Consequently, the set of admissible query messages to $\mathcal{V}[g_{i}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{i}$ is restricted to $Q_{\mathcal{M}_{i}}\times Q_{\mathcal{M}_{i}}$, which is finite.
+Therefore, we can assume that this mechanism has finite communication, satisfying the condition of Lemma 14.5, and the rest of the proof follows as the proof of Theorem 4.12.
+∎
+Similar to the concurrent composition of $k$ CMs with fixed privacy parameters, we define, for given privacy functions $f_{1},\dots,f_{k}$, the verification function $g^{f_{1},\dots,f_{k}}_{\mathrm{FDP},\infty}$ as the $f$-DP analogue of $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$ from Definition 5.1.
+Since $(\epsilon,\delta)$-DP is a special case of $f$-DP, Theorem 5.2 immediately yields a counterexample: there exists a privacy function $f_{1}$ and an adversary $\mathcal{A}$ such that the views of $\mathcal{A}$ when interacting with $\mathcal{V}[g^{f_{1}}_{\mathrm{FDP},\infty}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ have disjoint supports for $b=0$ and $b=1$.
+Equivalently, the $g^{f_{1}}_{\mathrm{FDP},\infty}$-concurrent composition of CMs is not private.
+To overcome the lower bound, as in Section 5.3, we can define the verification function $g^{f_{1},\dots,f_{k}}_{\mathrm{FDP},\infty,\mathit{FPC}}$ identical to $g^{f_{1},\dots,f_{k}}_{\mathrm{FDP},\infty}$ with an additional condition that checks whether the created mechanisms are first-pair consistent.
+A proof identical to the proof of Theorem 5.14 shows:
+Corollary 14.8.
+For $k\in\mathbb{N}$, let $f,f_{1},\dots,f_{k}:[0,1]\to[0,1]$ be functions.
+Suppose that for every $k$ NIMs $\mathcal{N}_{1},\dots,\mathcal{N}_{k}$, where each $\mathcal{N}_{i}$ is $f_{i}$-DP w.r.t.
+a neighbor relation $\sim_{i}$, the composition of $\mathcal{N}_{1},\dots,\mathcal{N}_{k}$ is $f$-DP.
+Then, the $g^{f_{1},\dots,f_{k}}_{\mathrm{FDP},\infty,\mathit{FPC}}$-concurrent composition of CMs is also $f$-DP.
+We know that IMs satisfy the first-pair consistent condition. Define the concurrent $k$-sparse parallel composition of IMs with privacy functions $f_{1},\dots,f_{k}$ as in Section 5.3.1. We have:
+Corollary 14.9.
+For $k\in\mathbb{N}$, let $f,f_{1},\dots,f_{k}:[0,1]\to[0,1]$ be functions.
+Suppose that for every $k$ NIMs $\mathcal{N}_{1},\dots,\mathcal{N}_{k}$, where each $\mathcal{N}_{i}$ is $f_{i}$-DP w.r.t.
+a neighbor relation $\sim_{i}$, the composition of $\mathcal{N}_{1},\dots,\mathcal{N}_{k}$ is $f$-DP.
+Then, the concurrent $k$-sparse parallel composition of IMs with privacy functions $f_{1},\dots,f_{k}$ is also $f$-DP.
+Finally, we extend the definition of concurrent filter composition of CMs to the $f$-DP setting.
+To this end, we modify Definition 12.1 so that a filter $\mathrm{Filt}$ maps a finite sequence of functions $f_{1},\dots,f_{t}:[0,1]\to[0,1]$ to an element of $\{\top,\bot\}$.
+As before, we construct the $f$-DP version of the verification function $f^{\mathrm{Filt}}$ in Definition 12.2 by replacing the DP parameters with privacy functions and ensuring that every created CM has finite communication.
+We denote this verification function by $g^{\mathrm{Filt}}_{\mathrm{FDP}}$.
+Corollary 14.10.
+Let $\mathrm{Filt}$ be a filter and $f:[0,1]\to[0,1]$ a function.
+If the $\mathrm{Filt}$-filter composition of NIMs is $f$-DP, then the $g^{\mathrm{Filt}}_{\mathrm{FDP}}$-concurrent composition of CMs is also $f$-DP.
+Equivalently, the concurrent $\mathrm{Filt}$-filter composition of CMs is $f$-DP.
 Acknowledgments
-1Monika Henzinger and Roodabeh Safavi were supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No.
-101019564), and the Austrian Science Fund (FWF) under grants DOI 10.55776/Z422, DOI 10.55776/I5982, and DOI 10.55776/P33775, with additional funding from the netidee SCIENCE Stiftung, 2020–2024.
-2Salil Vadhan was supported by NSF grant BCS-2218803, a grant from the Sloan Foundation, and a Simons Investigator Award.
+1Salil Vadhan was supported by NSF grant BCS-2218803, a grant from the Sloan Foundation, and a Simons Investigator Award.
 Work began while a Visiting Researcher at the Bocconi University Department of Computing Sciences, supported by Luca Trevisan’s ERC Project GA-834861.
+2Monika Henzinger and Roodabeh Safavi were supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No.
+101019564), and the Austrian Science Fund (FWF) under grants DOI 10.55776/Z422, DOI 10.55776/I5982, and DOI 10.55776/P33775.
+For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
+Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency.
+Neither the European Union nor the granting authority can be held responsible for them.
+Appendix A Figures
+[…]
+Figure 4. Comparison between interactive mechanism (IMs), continual observation (CO), and continual mechanisms (CMs)
+Comparison between interactive mechanism (IMs), continual observation (CO), and continual mechanisms (CMs)
+Appendix B Variants of Concurrent Composition
+Here, we elaborate on the different types of concurrent composition that we study in this paper:
+(1) Concurrent Composition of Continual Mechanisms: In the concurrent composition of fixed continual mechanism, $k$ continual mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ are fixed at the beginning, and an adversary adaptively issues pairs of messages for them such that for each $\mathcal{M}_{i}$, the sequences formed by the first messages of all message pairs and the sequence formed by the second messages of all message pairs are neighboring.
+A secret bit $b\in\{0,1\}$ determines whether each mechanism receives the first or second sequence from each pair.
+The adversary’s goal is to guess $b$ using the answers of all $k$ mechanisms.
+Theorem 4.10, which is the formal statement of Theorem 1.2 above, analyzes the privacy of this composition.
+A natural extension is the concurrent composition of continual mechanism with fixed parameters, where instead of fixing the mechanisms themselves, $k$ privacy parameter pairs $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed in advance.
+The adversary then adaptively selects mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ over time, ensuring that, for each $1\leq i\leq k$, the mechanism $\mathcal{M}_{i}$ is $(\epsilon_{i},\delta_{i})$-DP.
+The privacy guarantee for this variant is the same as Theorem 4.10 and is stated in Theorem 4.12.
+(2) Concurrent $k$-Sparse Parallel Composition of Continual Mechanisms: This composition generalizes the concurrent composition of continual mechanisms with $k$ fixed parameters by permitting the creation of an arbitrary number of continual mechanisms.
+As before, $k$ privacy parameters $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed at the beginning.
+The adversary adaptively creates a potentially unlimited number of continual mechanisms and issues pairs of messages to each mechanism such that for all but at most $k$ of these mechanisms, each message pair consists of identical messages.
+Let $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ denote the exceptional mechanisms, which may receive differing message sequences.
+For each $1\leq i\leq k$, the mechanism $\mathcal{M}_{i}$ must be $(\epsilon_{i},\delta_{i})$-DP, and the sequences of the first and second messages for $\mathcal{M}_{i}$ must be neighboring.
+We note the choice of mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ is made adaptively over time.
+Specifically, the adversary decides whether to issue a pair of differing messages for a mechanism based on the previous answers of all mechanisms, including the outputs of that mechanism.
+As in previous compositions, a secret bit $b$ determines whether each mechanism receives the first or second message from each pair, and the adversary seeks to guess $b$ based on the outputs.
+Let $\mathcal{M}_{j}^{\prime}$ denote the $j$-th created mechanism, and let $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ be its privacy parameter.
+Theorem 5.4, informally stated in Theorem 1.5, analyzes the privacy of the concurrent $k$-sparse parallel composition of continual mechanisms when $\sum\delta_{j}^{\prime}$ is upper bounded by a predetermined value.
+Corollary 5.5, informally stated in Corollary 1.7, sets this upper bound to 0 and provides privacy guarantees for the concurrent $k$-sparse parallel composition of purely differentially private continual mechanisms.
+(3) Concurrent $k$-Sparse Parallel Composition of Interactive Mechanisms: In this composition, the adversary creates an unbounded number of interactive mechanisms over time, instead of continual ones.
+As before, $k$ privacy parameter pairs $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed in advance.
+Each time the adversary creates an interactive mechanism, they also select a pair of neighboring datasets for it.
+Instead of message pairs, the adversary adaptively sends (single) queries to each mechanism.
+A secret bit $b$ determines whether the mechanism is run on the first or second dataset of its respective pair.
+The adversary then tries to infer $b$ from the responses of all mechanisms.
+Theorem 1.3 analyzes the privacy of this composition when all $\mathcal{M}_{i}$ are $(\epsilon_{0},\delta_{0})$-DP.
+The more general case for mechanisms with possibly different $(\epsilon_{i},\delta_{i})$ parameters is stated in Theorem 5.14.